### v1

# Geometric QCD II: The Confining Twistor String and Meson Spectrum

Abstract
We develop a continuum framework for confining planar QCD ($N_{c}\to\infty$) by quantizing the Fermi-string (1981) degrees of freedom on the rigid Hodge-dual minimal surface constructed in Geometric QCD I. The internal Majorana fermions on the worldsheet provide the algebraic mechanism required by the Makeenko–Migdal loop equations: by the Pauli principle, they enforce cancellations of non-planar self-intersection contributions, leaving the planar factorization structure. The Liouville instability of the original random-surface formulation is avoided because the bulk geometry is rigid and fixed holographically by the boundary loop, with no summation over fluctuating worldsheet metrics.
A key outcome is a formulation in momentum loop space, where coordinate-space cusp/contact singularities are integrated out, and the relevant quark-loop amplitudes admit a finite local limit. A central technical step is an explicit change of variables in the quark-loop phase-space path integral: after gauge fixing the reparametrization symmetry (Virasoro constraint), the reduced loop-space measure is parametrized by boundary twistors and an induced Jacobian, yielding a concrete twistor-string functional integral derived directly from the planar QCD amplitude rather than postulated as an abstract string model.
In the local limit, the theory becomes a confining (analytic) twistor string: a boundary sigma model $S^{1}\to(S^{3}\times S^{3})/S^{1}$ coupled to a holographically determined Liouville field $\rho(z,\bar{z})=\log\!\bigl(|\lambda(z)||\mu(z)|\bigr)$ obtained by analytic continuation of the boundary twistor data (Hilbert/Plemelj reconstruction).
The meson spectrum is encoded in a one-dimensional functional integral over the boundary twistor trajectory $\lambda(z),\mu(z),|z|=1$. While we evaluate this integral numerically using a reweighted Complex Langevin approach—successfully extracting approximately linear Regge trajectories as a proof of concept—our ultimate finding is analytical. By analyzing the monodromies of complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles inside the unit circle. The simplest sector with just one pole has the spectrum $m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$.
1 Introduction
The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the loop equations. At leading order in the $1/N_{c}$ expansion, the Makeenko–Migdal hierarchy exhibits the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
Historically, the main obstruction is the ultraviolet singular nature of the coordinate-space Wilson loop functional $W[C]$. Even after renormalization, $W[C]$ exhibits cusp and perimeter divergences, and the loop equation contains a contact term $\delta^{(4)}(x-y)$ at self-intersections. In the standard coordinate representation, this makes the local limit technically delicate and conceptually opaque: the loop equation is local, while the perturbative solution is nonlocal and develops logarithmic singularities tied to cusp geometry.
In this paper, we adopt a different viewpoint: rather than attempting to “repair” the coordinate-space Wilson loop as an observable object, we formulate the dynamics directly in momentum loop space. The momentum loop functional $W[P]$ was introduced in earlier work as the functional Fourier transform of $W[C]$. Its decisive advantage is structural: the planar loop equation in momentum space becomes an exact algebraic-differential functional equation with no coordinate-space delta functions. In this representation, the geometric cusp singularities of $W[C]$ are integrated out, and the continuum equations admit a local formulation that is finite in the sense relevant for physical quark-loop amplitudes.
The goal of this second paper in the Geometric QCD series is to provide the dynamics that complement the kinematics established in Part I [1]. Part I constructs the rigid vacuum geometry: the Hodge-dual additive minimal surface $S_{\chi}[C]$ whose (anti)self-dual area derivative yields a zero mode of the loop operator, producing the confining dressing factor $\exp\left(-\kappa S[C]\right)$. Here, we attach dynamical worldsheet degrees of freedom to this rigid background without summing over fluctuating metrics and show how the (regularized) planar loop equations are satisfied by this Ansatz for $W[C]$.
Our framework synthesizes three ingredients:
  1. 1.
    Fermi string (“Elfin theory”). We revisit the fermionic construction in which internal Majorana degrees of freedom on the worldsheet implement the required sign structure at intersections. In particular, the Pauli principle enforces cancellations between nonplanar intersection configurations, leaving the planar factorization pattern.
  2. 2.
    Rigid Hodge-dual minimal surface. The fermions are not placed on a random surface. Instead, they live on the rigid Hodge-dual minimal surface determined holographically by the boundary loop. This removes the dynamical Liouville instability that afflicted the original random-surface formulation: there is no summation over worldsheet metrics, and the geometry is fixed by the loop data.
  3. 3.
    Momentum loop dynamics and the twistor resolution. In momentum loop space, the loop equation becomes an exact algebraic-differential equation, stripped of coordinate-space contact singularities. We rigorously demonstrate that while this equation admits a functional Taylor-Magnus expansion at low orders, a purely one-dimensional algebraic Master Field solution violently breaks down at the eighth order ($\mathcal{W}^{(8)}$). This mathematical catastrophe proves that a 1D topological string fundamentally lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD. To resolve this geometric anomaly, the reduced loop-space measure must be parametrized by 4D continuous boundary twistor variables. This leads to an effective “analytic twistor string” description: a boundary sigma-model (a coset built from two $S^{3}$ spinors modulo a local $U(1)$) coupled to a holographic Liouville field whose bulk values are determined by the unique holomorphic extension of the boundary twistors.
Remark 1.1 (Equation solving vs. model building).
The starting point of this work is not a proposed “string model of QCD,” but the planar Makeenko–Migdal loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD. The aim is constructive: to exhibit explicit variables and an explicit functional measure in which the loop equations admit a finite continuum solution and yield physical observables. The principal ingredients of the present construction are not adjustable assumptions. The rigid Hodge-dual minimal surface enters because strict additivity and an (anti)self-dual area derivative are algebraically required for compatibility with the loop hierarchy, while the Majorana determinant enters because its Pauli-principle sign structure enforces planar factorization.
[…]
The culminating result of this paper is the exact reduction of the meson spectroscopy problem to a one-dimensional functional integral over boundary twistor data $(\lambda(z),\mu(z),\,|z|=1)$. While we initially evaluate this integral numerically via a reweighted Complex Langevin (CLE) approach—successfully extracting approximately linear Regge trajectories as a computational proof of concept—our ultimate finding is analytical. By evaluating the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory. The path integral exactly localizes at critical resonance energies where the Lefschetz thimbles become degenerate. Because the integration variables are coupled holomorphic twistors, this degeneracy intrinsically possesses a corank of 2, mathematically upgrading generic semiclassical branch cuts into the exact simple poles of hadronic states. Planar QCD thus ceases to be a theory of stochastic quantum fluctuations; the extraction of the exact spectrum reduces to a deterministic, classical generalized eigenvalue problem. While we initially evaluate the one-dimensional functional integral over the boundary twistor trajectory numerically using a reweighted Complex Langevin approach—successfully extracting approximately linear Regge trajectories as a proof of concept—our ultimate finding in this paper is analytical.
By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory. The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral. We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent. Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle. This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$. By analytically continuing the action to the boundary and summing over the exact monodromy phases, we obtain an exact Bohr-Sommerfeld quantization condition. For the simplest sector with just one pole, this yields the spectrum:
$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (1.1)
Remarkably, the universally correct Lüscher open-string intercept (+1/12+1/12) emerges directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations.
Organization of the paper and navigation points.
To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with four specific navigation points that address anticipated physical and computational questions.
Section 2 summarizes the loop-equation zero modes and recalls why the Hodge-dual minimal surface provides the appropriate rigid geometric background. For readers concerned about (i) the physical meaning of loop-equation zero modes, Sec. 2.3 explicitly details the underlying multi-instanton and vacuum ambiguity in loop space. Sections 3 to  9 then develop the fermionic determinant on this rigid surface and derive the corresponding loop equation, including its local limit.
Sections 10 to 12 introduce momentum loop space and the twistor string framework. To clarify (ii) the status of Wilson-loop “renormalization,” Sec. 10.1 sharply outlines our perspective: $W[C]$ is kept regularized at a finite Elf mass, physical observables are constructed natively in momentum loop space $W[P]$ (where coordinate-space singularities are integrated out), and only then is the local limit taken. Following this, Sections 11 to  12 execute the central change of variables. Addressing (iii) the parametrization of the quark loop in phase space by boundary values of holomorphic twistors, Sec. 11 details how the linear phase-space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
Section 13 discusses the semiclassical (WKB) limit and the emergence of linear Regge behavior. Section 14.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
Then, we address (iv) the discrete approximation and numerical implementation of the path integral. Section 15 summarizes the CLE evaluation, while E provides exhaustive details on the CLE implementation and topological analysis. Section 16 details the unsupervised phase-information clustering used to isolate the spectral branches.
Finally, we highlight (v) the exact analytic extraction of the mass spectrum. Section 17.1 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory. This last observation reveals the true geometric nature of Planar QCD: the main theoretical result of this work. Section 17.6 discusses the geometric origin of QCD, and Section 18 concludes. The remaining appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C.
2 Zero modes of Loop equation
Let us summarize the theory of Hodge-dual minimal surfaces and its relevance to the MM loop equation. For a detailed discussion and proofs of relevant theorems, the reader may look into the latest paper [1], and the details of the new loop calculus without divergences can be found in [2, 3].
2.1 Self-dual area derivative and loop equation
Let us recall why we are looking for the self-dual minimal surface, as suggested in [2]. In this section, we follow the arguments of our recent paper [1], but our goal is not to present the Hodge-dual surface per se, but rather to view it as a hologram of the QCD string, with Fermi degrees of freedom attached.
The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
$\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=\frac{\delta^{2}W}{\delta\dot{C}_{\mu}(\theta-0)\delta\dot{C}_{\nu}(\theta+0)}-\mu\leftrightarrow\nu$ (2.1)
This definition implements the geometric definition of [4, 5] and is formally equivalent to the functional derivative implementation by Polyakov [6]. The crucial difference is the lack of divergences in this new definition and its unbroken parametric invariance. In the old, singular implementations, the zero modes were obscured by kinematic divergencies, interfering with potential small-distance singularities of perturbative QCD.
The new loop equation for the Yang–Mills gradient flow was shown in [2] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
$\mathcal{L}_{\nu}(S)=0;$ (2.2)
$\mathcal{L}_{\nu}=\left(\frac{\delta}{\delta\dot{C}_{\mu}(\theta+0)}-\frac{\delta}{\delta\dot{C}_{\mu}(\theta-0)}\right)\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$ (2.3)
The basic observation made in that paper was that there is a Hodge-dual solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.3) is identically satisfied for any functional $S[C]$ with self-dual (SD) or anti-dual (ASD) area derivative
$\frac{\delta S_{\pm}}{\delta\sigma_{\mu\nu}(\theta)}=\pm\ast\frac{\delta S_{\pm}}{\delta\sigma_{\mu\nu}(\theta)};$ (2.4)
$e_{\alpha\mu\nu\lambda}\left(\frac{\delta}{\delta\dot{C}_{\alpha}(\theta+0)}-\frac{\delta}{\delta\dot{C}_{\alpha}(\theta-0)}\right)\frac{\delta S_{\pm}}{\delta\sigma_{\mu\nu}(\theta)}\equiv 0;$ (2.5)
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [2] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor. This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
The existence of this zero mode is based on the so-called Leibniz property of the loop operator and similar singular loop operators in loop calculus.
$\mathcal{L}_{\nu}f(\Phi[C])=f^{\prime}(\Phi[C])\mathcal{L}_{\nu}\Phi[C];$ (2.6)
$\mathcal{L}_{\nu}(A[C]B[C])=\mathcal{L}_{\nu}(A[C])B[C]+A[C]\mathcal{L}_{\nu}(B[C])$ (2.7)
It was heuristically claimed in the original MM papers [7, 5], and rigorously proven in the loop calculus [2], section 3, eqs (39), (40). Due to the Leibnitz property,
$\mathcal{L}_{\nu}\left(\exp\left(-\kappa S[C]\right)W_{0}[C]\right)=$
$-\kappa\mathcal{L}_{\nu}(S[C])\exp\left(-\kappa S[C]\right)W_{0}[C]$
$+\exp\left(-\kappa S[C]\right)\mathcal{L}_{\nu}(W_{0}[C])=\exp\left(-\kappa S[C]\right)\mathcal{L}_{\nu}(W_{0}[C])$ (2.8)
We are looking for such a self-dual minimal surface which is additive at the self-intersecting loops. This condition will be discussed in detail below. It is critical for the whole theory—due to this additivity, the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the whole chain of the loop equations.
2.2 Compatibility with the MM loop equations.
The full chain of loop equations for the Wilson loops with a finite number of colors $N_{c}$ [7, 5] has the same loop operator (2.3) as the Yang–Mills gradient flow, except there is no time derivative. In this paper, we are only considering the leading order of the $1/N_{c}$ expansion. We use the loop calculus developed in the paper [2], as it applies to the loop diffusion operator (2.3), which is the same in the multiloop MM chain and in the Yang–Mills gradient flow equation. The MM loop equation has the structure (see Fig.1)
Figure 1: The MM equation, with the loop diffusion operator on the left side, and the integral term on the right side. The double line on the right stands for the delta function $\delta^{4}(x-y)$ enforcing self-intersection.
[…]
$\mathcal{L}_{\nu}(W[C])=\lambda\int_{C}dy_{\nu}\delta^{4}(x-y)W[C_{xy}]W[C_{yx}]$ (2.10)
Here $\lambda=N_{c}g_{0}^{2}$ is the ’t Hooft’s coupling constant in the limit of $N_{c}\to\infty$. This equation is singular at self-intersections, so it must be understood in the sense of distributions in loop space. It is manifestly parameterization-invariant, and the confining factor $\exp\left(-\kappa S[C]\right)$ factors out from the right side by additivity of $S[C]$ and cancels with the same factor on the left side.
Let us stress that this factorization takes place independently of the splitting points $x,y\in C$. For every self-intersection $x,y\in C$ such that $C(x)=C(y)$, the product of two factors $\exp\left(-\kappa S[C_{xy}]\right)$ and $\exp\left(-\kappa S[C_{yx}]\right)$ exactly equals the factor $\exp\left(-\kappa S[C]\right)$ on the left side of the MM equation; Thus, it can be taken out of the integral, after which it cancels on both sides of the MM equation.
The physical hypothesis of that paper was that the actual gluonic vacuum of QCD is described by this dressed solution, choosing the "fluctuative" factor $W_{\text{fluct}}$ as the seed. While the symmetry is exact, the separation into a fluctuative factor and a purely geometric area law is ambiguous. It is physically most distinct in the large-loop limit, where the area term dominates over $W_{\text{fluct}}$, which is assumed only to contribute the term proportional to a perimeter of the loop.
In the present paper, we construct an explicit solution for this factor $W_{\text{fluct}}$, adapting the fermionic determinant from the ’81 paper [8] to the Hodge-dual minimal surface instead of a random surface, assumed in that work. The confining factor becomes the vacuum energy of that rigid Fermionic string, and planar graphs emerge as a strong coupling expansion of that 2D fermi system in terms of a sum over fermionic loops on a Hodge-dual minimal surface. The Pauli principle leads to repulsion of these loops, making them topologically equivalent to planar graphs of QCD.
The comparison of the contour equation for these fermionic determinants from the ’81 work with the MM equation establishes the correspondence between the ’t Hooft coupling constant $\lambda$ and the parameters of the Elfin theory (the bare fermionic mass and the vacuum energy).
2.3 The Physical Vacuum and Multi-Instanton Resummation
Before we formulate the dynamics of the Fermi string on the minimal surface, we must clarify the physical origin of this rigid geometric background. The Makeenko-Migdal (MM) loop equations serve as the exact quantum equations of motion for Planar QCD. Like any equations of motion, they possess an inherent ambiguity corresponding to the choice of the physical vacuum. The mathematical freedom to multiply the perturbative loop functional by a zero mode exactly reflects this physical vacuum ambiguity.
In the traditional gauge-field picture, the nonperturbative QCD vacuum is a highly complex mixed state, often modeled as a nontrivial distribution in the multi-instanton space. Because any single multi-instanton solution explicitly breaks translation invariance by localizing topological charge at specific spacetime points, restoring Poincaré symmetry requires an exceedingly complicated integration over the entire multi-instanton moduli space (positions, sizes, and color orientations).
The loop-space formulation achieves this summation geometrically. Just as instantons are characterized by the self-duality of the gauge field strength ($F_{\mu\nu}=\pm*F_{\mu\nu}$), our exact zero mode is characterized by the self-duality of the loop-space area derivative.
Thus, the complex mixed state of multi-instantons in coordinate space simplifies to a single, fully translation-invariant, exact geometric object in loop space: the Hodge-dual minimal surface.
Note that the mere Hodge duality condition leaves the freedom of an arbitrary topological charge and arbitrary moduli of the multi-instanton solution. Therefore, our Hodge-dual minimal surface, by satisfying this requirement, represents a general statistical mixture of multi-instanton solutions, implicitly summed over topological charges and integrated over moduli with some unknown measure. This implicit integration is precisely what provides the exact translation invariance and spatial isotropy of the minimal area solution. The profound advantage of the loop-space approach is that we do not need to know the explicit details of this distribution of topological charge and multi-instanton moduli: the loop equations can be solved directly without ever specifying this measure.
Therefore, the rigid background upon which our worldsheet fermions propagate is not merely a mathematical ansatz. It is the exact holographic representation of the translation-invariant, nonperturbative QCD vacuum, naturally incorporating the sum over topological sectors that is analytically prohibitive in the standard gauge-field path integral.
3 Hodge-dual minimal surface
Now, we are going to define the solvable functional $S[C]$ satisfying both the duality and additivity. We are following the definitions and derivations from [1], adapting them for the present purpose of a base for the fermions living on that minimal surface.
3.1 The external and internal geometry
We define the surface coordinates $X^{i}_{\mu}(\xi)$ (for $\xi=(\xi_{1},\xi_{2}),i=1,2,3$.  [^1] This surface was first proposed in [2] as a zero-mode of the loop operator in the context of the Yang–Mills gradient flow, but unfortunately, the wrong number of components (four instead of three) was chosen for the $X^{A}_{\mu},A=1,2,3,4$. The rest of the theory in that paper was validated by our recent work [1], including the loop calculus and the Hodge-duality of the minimal surface. The field $X^{i}_{\mu}(\xi)$ maps the disk $\Sigma$ (Fig 2) to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$.
Figure 2: The unit disk $\Sigma$ is mapped to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$ and then projected to $\mathbb{C}^{4}$ by a holomorphic map, minimizing the area functional.
Our surface area is defined as
$S_{\chi}[C]=\min_{X}\int_{\Sigma}d^{2}\xi\sqrt{\Sigma_{\mu\nu}^{2}/2};$ (3.1)
The area element is defined as
$\Sigma_{\mu\nu}=\epsilon_{ab}\partial_{a}X^{i}_{\mu}\partial_{b}X^{i}_{\nu};$ (3.2)
The Hodge chirality $\chi=\pm 1$ enters through the boundary conditions
$X^{i}_{\mu}(\partial\Sigma)=\eta^{\chi,i}_{\mu\nu}C_{\nu};$ (3.3)
Here $\eta^{\chi,i}_{\mu\nu}$ are ’t Hooft’s matrices corresponding to Hodge duality $\chi=\pm 1$
$\eta^{\chi,i}_{\mu\nu}=\delta_{i\mu}\delta_{\nu 4}-\delta_{i\nu}\delta_{\mu 4}+\chi e_{i\mu\nu 4};$ (3.4)
The coordinate functions $X_{\mu}^{a}(\xi)$ satisfy the general Euler-Lagrange equations for the area density $\mathcal{L}=\sqrt{\Sigma^{2}/2}$:
$\left[\text{E-L Equations}\right]_{\mu}^{a}=\epsilon_{lm}\partial_{l}\left(\frac{\Sigma_{\mu\nu}\partial_{m}X^{a}_{\nu}}{\sqrt{\Sigma^{2}}}\right)=0.$ (3.5)
Assuming the bulk variation vanishes by the Euler-Lagrange equations, we use Stokes theorem and get the boundary term
$\delta S=2\int d\theta\frac{\delta X^{i}_{\mu}\Sigma_{\mu\nu}\partial_{\theta}X^{i}_{\nu}}{\sqrt{\Sigma^{2}}}$ (3.6)
Substituting the boundary conditions (3.3) for $\delta X,\partial_{\theta}X$ and summing over $a=1,2,3$ we find
$\delta S=-2\int d\theta\frac{\delta C_{\mu}\Sigma_{\mu\nu}\partial_{\theta}C_{\nu}}{\sqrt{\Sigma^{2}}};$ (3.7)
$\frac{\delta S}{\delta\sigma_{\mu\nu}}=-2T_{\mu\nu};$ (3.8)
$T_{\mu\nu}=\frac{\Sigma_{\mu\nu}}{\sqrt{\Sigma^{2}}}$ (3.9)
Note the change of sign, compared with the ordinary area derivative of the minimal surface in four Euclidean dimensions.
Regardless of this sign change, the Hodge duality of the area element $\Sigma$ with these boundary conditions for a generic field $X^{a}_{\mu}(z,\bar{z})$ leads to the same Hodge duality of the area derivative.
This self-duality of the area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations [7, 5], as we argue in the next section.
Note that this area functional is invariant under the reparametrizations of the boundary loop $\delta_{\text{param}}C_{\mu}(\theta)=\epsilon(\theta)\partial_{\theta}C_{\mu}(\theta)$
$\delta_{\text{param}}S=-2\int d\theta\epsilon(\theta)\frac{\partial_{\theta}C_{\mu}\Sigma_{\mu\nu}\partial_{\theta}C_{\nu}}{\sqrt{\Sigma^{2}}}=0$ (3.10)
by the skew symmetry of $\Sigma_{\mu\nu}$.
The Hodge-dual area is defined as a minimum of this functional under an extra duality constraint
$\ast\Sigma_{\mu\nu}\equiv\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta}=\chi\Sigma_{\mu\nu}$ (3.11)
In the previous papers [2, 1], we found a complete solution for the Hodge-dual surface in terms of holomorphic map.
3.2 The Holomorphic Map
We resolve the duality constraint using the following holomorphic Ansatz:
$X^{i}_{\mu}=\eta^{\chi,i}_{\mu\nu}Y_{\nu}$ (3.12)
$Y_{\mu}=f_{\mu}(z)+\bar{f}_{\mu}(\bar{z});$ (3.13)
$z=\xi_{1}+\mathrm{i}\xi_{2};$ (3.14)
Geometrically, this is a projection from $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$ to $\mathbb{C}^{4}$. The boundary conditions for the Hodge-dual surface become a conventional boundary condition for the Dirichlet problem
[…]
So, we have a line in four dimensional complex space $f\in\mathbb{C}^{4}$. Our surface area (3.1) reduces to the induced metric in this space
$\Sigma_{\mu\nu}=2(F_{\mu\nu}+\chi\ast F_{\mu\nu});$ (3.16)
$F_{\mu\nu}=\mathrm{i}(f^{\prime}_{\mu}\bar{f}^{\prime}_{\nu}-\bar{f}^{\prime}_{\mu}f^{\prime}_{\nu});$ (3.17)
$\sqrt{\Sigma_{\mu\nu}^{2}/2}=2\sqrt{2}\sqrt{-\det g};$ (3.18)
$g_{ab}=\partial_{z_{a}}Y_{\mu}\partial_{z_{b}}Y_{\mu};\qquad z_{a},z_{b}=(z,\bar{z})$ (3.19)
This area element is Hodge-dual
$\ast\Sigma_{\mu\nu}=\chi\Sigma_{\mu\nu}$ (3.20)
for arbitrary functions $f_{\mu}(z)$.
This duality condition, as argued above, together with the above boundary conditions, provides the self-duality of the area derivative of the minimal area. The self-duality of the area derivative follows from the general formula for variation of the surface $X^{a}_{\mu}(\xi)$ with area (3.1), projected on the holomorphic Ansatz.
3.3 The Virasoro Constraint and Uniformization
We further impose a Virasoro constraint on the holomorphic maps derivatives
$f^{\prime}_{\mu}(z)^{2}=0$ (3.21)
The imposition of the null constraint $(f^{\prime})^{2}=0$ is not an approximation but a rigorous gauge fixing procedure rooted in the geometric symmetries of the problem.
The physical starting point is the geometric Area functional, $S=\int\sqrt{-\det g}\,d^{2}z$, which possesses full reparametrization invariance under general diffeomorphisms $z\to w(z,\bar{z})$. This infinite symmetry group allows us to choose a specific coordinate system on the world-sheet without altering physical results.
We choose to work in isothermal (conformal) coordinates, where the induced metric is diagonal:
$ds^{2}=\rho(z,\bar{z})\,dz\,d\bar{z}\implies g_{zz}=\partial_{z}Y\cdot\partial_{z}Y=0$ (3.22)
For our specific ansatz (3.12), the metric component $g_{zz}$ is explicitly proportional to the square of the holomorphic derivative:
$g_{zz}=(f^{\prime}_{\mu})^{2}=0$ (3.23)
Thus, the geometric condition of conformal gauge fixing $g_{zz}=0$ is algebraically equivalent to the null constraint $(f^{\prime})^{2}=0$.
The existence of such a coordinate system for any surface topology is guaranteed by the Uniformization Theorem [9]. This theorem ensures that we are always free to set $(f^{\prime})^{2}=0$ by a suitable diffeomorphism. In this gauge, the non-linear geometric area density simplifies exactly to the quadratic Lagrangian:
$\mathcal{L}$ $=\sqrt{-\det g}=\sqrt{(g_{z\bar{z}})^{2}-|g_{zz}|^{2}}$
$\xrightarrow{(f^{\prime})^{2}=0}g_{z\bar{z}}\propto|f^{\prime}|^{2}$ (3.24)
This reduction allows us to describe the Minimal Surface using the linear Laplace equation for $f_{\mu}$, subject to the Virasoro constraint. To summarize, we have found that
$S_{\chi}[C]=2\sqrt{2}\int_{D}|f^{\prime}(z)|^{2}d^{2}z;$ (3.25)
$\frac{\delta S_{\chi}[C]}{\delta\sigma_{\mu\nu}}=-2\frac{\Sigma_{\mu\nu}}{\sqrt{\Sigma_{\mu\nu}^{2}}}$ (3.26)
The solution of the linear boundary problem (3.15) for a holomorphic vector function is given by the Hilbert transform
$f_{\mu}=\frac{1}{2}(1+\mathrm{i}\mathcal{H})C_{\mu};$ (3.27)
$H[u](\theta)\;=\;\frac{1}{2\pi}\,\mathrm{P.V.}\!\int_{0}^{2\pi}d\theta^{\prime}\;\cot\!\Big(\frac{\theta-\theta^{\prime}}{2}\Big)\,u(\theta^{\prime})$ (3.28)
$\int_{D}|f^{\prime}(z)|^{2}d^{2}z\;=\;\int_{0}^{2\pi}d\theta\;C^{\prime}_{\mu}(\theta)\,H[C_{\mu}](\theta);$ (3.29)
In terms of the Taylor coefficients
$f_{\mu}(z)=\sum_{n>0}z^{n}C_{\mu,n};$ (3.30)
$C_{\mu}(\theta)=\sum_{n=-\infty}^{\infty}e^{\mathrm{i}n\theta}C_{\mu,n}$ (3.31)
In the presence of the Virasoro constraint (3.21), the problem becomes a challenging one and has been investigated over the last two centuries by great mathematicians, starting with Riemann and Hilbert. We give a brief summary of modern state of this theory in Appendix B.
In Part I [1], we compared the ordinary minimal surface in four dimensions with the Hodge-dual one. This comparison reveals some hidden variables present in the area derivatives of the Hodge dual surface, but not in the extremal value of this area. This minimal value up to a normalization factor $2\sqrt{2}$ equals that of the so called Goldschmidt solution [10, 12, 13] for the Plateau problem. This additive solution represents a local minimum of the Dirichlet functional, with the simplest topology (collection of disks one for each closed part of $C$).
The area derivatives for the Goldschmidt minimal area and Hodge dual minimal area are different (the second one being SD or ASD component of the first). The apparent paradox is resolved by the obvious fact that the area derivative of a functional equals the skew-symmetric part of its second functional derivative, but only the extremal value and its first derivative (=0=0) coincide for the functional and its minimal value.
Also, the Hodge-dual surface is not parity-symmetric, but its minimal value (minimal area) is. The chirality $\chi$ manifests itself in the external geometry of this minimal surface $X^{i}_{\mu}\in\mathbb{R}^{3}\otimes\mathbb{R}^{4}$, including its area derivative, but its internal geometry is determined by the induced metric (3.22),(3.23), which is the only local function related to surface embedding. Moreover, this metric in the holomorphic map that minimizes the area depends only on the bounding loop $C$, through the holomorphic vector function $f_{\mu}(z)$, which solves the Riemann-Hilbert problem (9.4), (3.30). This dependence is identical for the Hodge-dual surface and the Goldschmidt solution for the Plateau problem in 4D, bounded by $C$.
The last property (the identity of the internal geometry in the Hodge-dual surface in relation to that of the 4D Euclidean problem) will play an important role in proving the MM equation for the fermionic determinant. The only remnants of the Hodge-duality are the overall negative sign and the $\chi$-term in the area derivative (3.26) projecting it on the Hodge-dual sector of skew-symmetric tensors in 4D.
4 The Elfin theory on a flat surface
In this section, we deal with the Elfin theory by itself, without any reference to the gauge theory. The equivalence of these theories will be established in the next sections.
Let us consider the flat surface bounded by some contour $C$. Later, we generalize this to the Hodge-dual minimal surface. All the functionals on the surface will depend on the conformal factor in the mass term
$e=\partial_{+}X^{i}_{\mu}\partial_{-}X^{i}_{\mu}.$ (4.1)
On the minimal surface, we have a conformal map (3.12), reducing the conformal metric to $e=2\sqrt{2}|f^{\prime}_{\mu}|^{2}$. This reduction is not needed for the general theory of fermions on the minimal surface, but it will be used in the forthcoming sections when we establish a relation between the fermionic determinants and the MM loop equation.
Remark 4.1 (Holes, poles, and global signs).
Multiply connected worldsheets (holes/handles) may require multivalued primitives $f_{\mu}(z)=\int^{z}f^{\prime}_{\mu}(\zeta)\,d\zeta$ (logarithms/periods), even when the tangents $f^{\prime}_{\mu}$ are single-valued meromorphic. In the applications to the QCD spectrum, we allow such meromorphic $f^{\prime}_{\mu}$ with poles in the unit disk; the induced conformal factor $e(z,\bar{z})\propto|f^{\prime}_{\mu}(z)|^{2}$ is single-valued but may diverge at these poles.
This does not invalidate the analysis of the Elfin determinant in the flat gauge: closed fermion paths are weighted by $\exp\left(\!\right)\big(-m\int_{\Gamma}\sqrt{e}\,|dz|\big)=\exp\left((\right)-m\,\ell_{\Gamma})$, hence any path approaching a pole (where $\sqrt{e}\sim|z-w|^{-1}$ for a simple pole of $f^{\prime}$) has $\ell_{\Gamma}\to\infty$ and is exponentially suppressed. Since $\ell_{\Gamma}$ is a local additive functional, the length weights of the locally paired configurations (touching vs. intersecting) coincide, while their mod-$2$ intersection signs are opposite, yielding the same cancelation mechanism.
Finally, while handles do not affect the local turning-angle contribution to the loop sign, they can affect global signs of the fermion propagator/determinant (spin-structure/holonomy) and complicate the topological analysis of the corresponding minimal-surface embeddings.
Let us now return to the flat surface and introduce the Elfin field as a bispinor
$\psi=\psi^{\alpha}_{\lambda}(\xi),\qquad\alpha=1,2,\qquad\lambda=\pm 1.$ (4.2)
There are various equivalent forms of the action. The simplest form reads
[…]
where $\sigma_{1},\sigma_{2}$ are the Pauli matrices. We shall also use the complex components
$\sigma_{\pm}$ $=\frac{1}{2}(\sigma_{1}\pm\mathrm{i}\sigma_{2});$ (4.4)
$\partial_{\pm}$ $=\partial_{1}\pm\mathrm{i}\partial_{2};$ (4.5)
$\psi^{\pm}_{\lambda}$ $=\frac{1}{2}(1\pm\sigma_{3})\psi_{\lambda}.$ (4.6)
$\bar{\psi}^{\pm}_{\lambda}$ $=\psi^{\dagger}_{\lambda}\frac{1}{2}(1\pm\sigma_{3}).$ (4.7)
This action is covariant with respect to conformal transformations
$e(\xi_{+},\xi_{-})$ $\to f^{\prime}_{+}(\xi_{+})f^{\prime}_{-}(\xi_{-})e(f_{+},f_{-}),$ (4.8)
$\psi^{\pm}_{\lambda}(\xi)$ $\to\sqrt{f^{\prime}_{\pm}}\psi^{\pm}_{\lambda}(f),$ (4.9)
$\bar{\psi}^{\pm}_{\lambda}(\xi)$ $\to\sqrt{f^{\prime}_{\mp}}\bar{\psi}_{\lambda}(f);$ (4.10)
At the boundary $\dot{\Sigma}$, we require that the spin $\sigma_{3}$ takes definite values, namely:
$\sigma_{3}\psi_{\lambda}=\lambda\psi_{\lambda},\qquad\text{at }\partial\Sigma,$ (4.11)
$\bar{\psi}_{\lambda}\sigma_{3}=\lambda\bar{\psi}_{\lambda},\qquad\text{at }\partial\Sigma.$ (4.12)
In other words, at the boundary
$\psi^{-}_{+}=\psi^{+}_{-}=\bar{\psi}^{-}_{+}=\bar{\psi}^{+}_{-}=0.$ (4.13)
Such "chiral bag" boundary conditions were studied in the mathematical literature [14] in the context of the Atiyah-Singer Index Theorem. For our purposes, it is sufficient to note that these boundary conditions are consistent with the topology of the Dirac path integral below.
The remaining half of the components is not restricted at the boundary. In the Schrödinger picture in Minkowski space this implies that the states with $\sigma_{3}=\pm 1$ at both ends of the string are occupied. Such peculiar boundary conditions are necessary to obtain the closed loop equations (see below). Note that these boundary conditions are conformally invariant and do not depend on the form of the boundary.
At the "time" reflection
$\xi_{2}\to-\xi_{2}\qquad\text{or}\qquad\xi_{+}\leftrightarrow\xi_{-}$ (4.14)
the spin $\sigma_{3}$ changes the sign. According to our boundary conditions, this should be accompanied by a change of sign of $\lambda$:
$\psi_{\lambda}\to\psi_{-\lambda}\sigma_{3}.$ (4.15)
One may introduce the oriented states
$\psi^{R,L}_{\lambda}$ $=\psi^{\pm\lambda}_{\lambda},$ (4.16)
$\bar{\psi}^{R,L}_{\lambda}$ $=\bar{\psi}^{\mp\lambda}_{\lambda}.$ (4.17)
The left-hand states are occupied at the boundary. As we shall see, the orientation $\lambda\sigma_{3}$ is related to the orientation of the loop $C$. The right-hand and left-hand states interchange when the loop is reoriented.
The action can be rewritten in the Left-Right (chiral) form. Based on the transformation properties in equations (4.16) and (4.17), and using the identity $\sigma^{k}\partial_{k}=\sigma_{+}\partial_{+}+\sigma_{-}\partial_{-}$, the action becomes:
[…]
$\left.m\sqrt{e}\left(\bar{\psi}^{R}\psi^{R}+\bar{\psi}^{L}\psi^{L}\right)\right].$ (4.18)
In this form, the conformal properties are manifest. Under the transformations $\xi_{+}\to f_{+}(\xi_{+})$ and $\xi_{-}\to f_{-}(\xi_{-})$ the kinetic term transforms as:
$\left(\sqrt{f^{\prime}_{-}}\bar{\psi}^{R}\right)\left(f^{\prime}_{+}\partial_{+}\right)\left(\sqrt{f^{\prime}_{-}}\psi^{L}\right)$
$=f^{\prime}_{-}f^{\prime}_{+}\left(\bar{\psi}^{R}\partial_{+}\psi^{L}\right);$ (4.19)
$\left(\sqrt{f^{\prime}_{+}}\bar{\psi}^{L}\right)\left(f^{\prime}_{-}\partial_{-}\right)\left(\sqrt{f^{\prime}_{+}}\psi^{R}\right)$
$=f^{\prime}_{-}f^{\prime}_{+}\left(\bar{\psi}^{L}\partial_{-}\psi^{R}\right).$ (4.20)
The mass term transforms as:
$\left(\sqrt{f^{\prime}_{+}f^{\prime}_{-}}\sqrt{e}\right)\left(\sqrt{f^{\prime}_{-}}\bar{\psi}^{R}\right)\left(\sqrt{f^{\prime}_{+}}\psi^{R}\right)$
$=(f^{\prime}_{+}f^{\prime}_{-})\sqrt{e}\bar{\psi}^{R}\psi^{R};$
$\left(\sqrt{f^{\prime}_{+}f^{\prime}_{-}}\sqrt{e}\right)\left(\sqrt{f^{\prime}_{+}}\bar{\psi}^{L}\right)\left(\sqrt{f^{\prime}_{-}}\psi^{L}\right)$
$=(f^{\prime}_{+}f^{\prime}_{-})\sqrt{e}\bar{\psi}^{L}\psi^{L}.$ (4.21)
[…]
Note that the Lagrangian density is invariant with respect to internal $U(1)\times SU(2)$ transformations:
$\psi_{\lambda}$ $\to S_{\lambda\lambda^{\prime}}\psi_{\lambda^{\prime}}e^{\mathrm{i}\alpha},$ (4.22)
$\bar{\psi}_{\lambda}$ $\to\bar{\psi}_{\lambda^{\prime}}(S^{-1})_{\lambda^{\prime}\lambda}e^{-\mathrm{i}\alpha}.$ (4.23)
However, the boundary conditions lower this symmetry down to $U_{+}(1)\times U_{-}(1)$:
$\psi^{\pm}$ $\to\psi^{\pm}e^{\mathrm{i}\alpha_{\pm}},$ (4.24)
$\bar{\psi}^{\pm}$ $\to\bar{\psi}^{\pm}e^{-\mathrm{i}\alpha_{\pm}}.$ (4.25)
Both $U(1)$ currents vanish at the boundary:
$j_{k}=\bar{\psi}^{\pm}\sigma_{k}\psi^{\pm}=0\qquad\text{at }\dot{\Sigma};\qquad k=1,2.$ (4.26)
So, there is no flow outside. The derivative $\partial_{k}$ in (2.11) can be replaced by $D_{k}$, since the boundary terms vanish.
At vanishing mass $m$ the theory possesses chiral $\sigma_{3}$ invariance, which is not violated by any anomaly.
Let us now discuss the symmetry of the physical $x$-space. As usual in string theory, the transformations of the internal coordinates $\xi$ have nothing to do with the physical rotations in $x$-space. From the viewpoint of two-dimensional field theory in $\xi$-space, the rotations in $x$-space represent internal $O(4)$ symmetry transformations of the $C_{\mu}$ field,
$\delta C_{\mu}=\omega_{\mu\nu}C_{\nu}.$ (4.27)
The Elf does not transform:
$\delta\psi_{\lambda}=\delta\bar{\psi}_{\lambda}=0.$ (4.28)
This follows from the boundary conditions. The full target surface field $X^{i}_{\mu}$ transforms as one of the $SO(3)_{\pm}$ subgroups of $SO(4)$
$X^{i}_{\mu}\Rightarrow\Omega^{\pm}_{ij}X^{j}_{\nu}W_{\nu\mu};\qquad\Omega^{\pm}\in SO(3),\qquad W\in SO(4)$ (4.29)
Consider now local variations of the loop $C$. The area derivative of $e$ can be found from the known variation of the scalar area
[…]
$=2\int_{S}t_{\mu\nu}(\xi_{2})d^{2}\xi_{2}\delta(\xi_{1}-\xi_{2})$ (4.31)
Using the exact area derivatives of the Hodge-dual minimal area established in the previous section, we find for the area derivative of
$\frac{\delta e(\xi)}{\delta\sigma_{\mu\nu}(\xi_{0})}$ $=-2T_{\mu\nu}\delta^{2}(\xi-\xi_{0});$ (4.32)
$T$ $=\frac{F+\chi\ast F}{\sqrt{2F^{2}}}$ (4.33)
$\frac{\delta\left\langle\exp\left(A_{0}\right)\right\rangle}{\delta\sigma_{\mu\nu}}$ $=-2m\left\langle\frac{T_{\mu\nu}\bar{\psi}\psi}{\sqrt{e}}\right\rangle_{\partial\Sigma}$ (4.34)
where $F$ is given by (3.26) . According to our boundary conditions, only the right-hand components are present in (4.34):
$\bar{\psi}\psi=\bar{\psi}^{R}\psi^{R}\qquad\text{at }\dot{\Sigma}.$ (4.35)
5 Conformal metric as a local gauge parameter
We begin by defining the effective action for the elf (Majorana fermion) field on a two-dimensional surface $S$ with metric
$g_{ab}=\exp\left(2\rho\right)\delta_{ab},$
including the mass term and the dependence on the conformal factor $\rho$. We are reproducing here the construction in Eqs. (2.57)–(2.63) of [8].
The action of the general theory is given by
$A_{\rho}=\int d^{2}\xi$
$\left(e^{2\rho}\bar{\psi_{\lambda}}\sigma^{k}e^{-\rho}\nabla_{k}\psi_{\lambda}+m\sqrt{e}e^{\rho}\bar{\psi_{\lambda}}\psi_{\lambda}-\frac{\partial_{+}\rho\partial_{-}\rho}{3\pi}\right);$ (5.1)
with the spinor connection $\omega_{k}$:
$\nabla_{k}=\partial_{k}+\frac{\mathrm{i}}{2}\omega_{k}\sigma^{3},$ (5.2)
vierbein
$E_{k}^{a}=\delta^{a}_{k}e^{\rho},$ (5.3)
and curvature
$\hat{R}_{kl}$ $=[\nabla_{k},\nabla_{l}]=\mathrm{i}\sigma_{3}e_{kl}R\det E;$ (5.4)
$R$ $=-8e^{-2\rho}\partial_{+}\partial_{-}\rho.$ (5.5)
At $\rho=0$ this is the old action $A_{0}$, whereas at $\rho=\ln\sqrt{e}$ this is the action of the Dirac particle with constant mass at the surface $S$ with the induced metric $g_{ab}$.
Theorem 5.1 (Conformal Invariance of the Elf Determinant).
The functional integral
$Z[S|e,\rho]=\int\mathcal{D}\bar{\psi}\mathcal{D}\psi\exp\left(A_{\rho}\right)$ (5.6)
is invariant under local variations of the conformal metric field $\rho$. That is, the theory depends only on the conformal class (moduli) of the surface, and the induced metric $e(\xi)$ but not on the local gauge parameter $\rho(\xi)$.
Proof.
The main point here is the covariant regularization of the Dirac theory at the surface. We have to define the divergent trace in the vacuum amplitude
$\int\mathcal{D}\bar{\psi}\mathcal{D}\psi\exp\left(A_{\rho}\right)=\exp\left(-\int d^{2}\xi\frac{(\partial\rho)^{2}}{3\pi}\right$.
[…]
Here
$\hat{D_{\rho}}=\sigma^{k}e^{-\rho}(-\mathrm{i}\partial_{k}\omega_{k}),$ (5.8)
is the covariant Dirac operator, and the subscript ++ conditions $\sigma^{3}=+1$.
We employ the Pauli-Villars regularization in the following form:
[…]
=tr[$=\tr\left[\ln(\mathrm{i}\hat{D_{\rho}}+m\sqrt{e}e^{-\rho})-\frac{1}{2}\sum_{i}c_{i}\ln(\hat{D_{\rho}}^{2}+M_{i}^{2})\right],$
$\sum_{i}c_{i}=1.$ (5.9)
The shift of the $\rho$ field
$\rho\to\rho+\delta\rho$ (5.10)
results in the multiplicative transformation of the Dirac operator
$\hat{D_{\rho}}\Rightarrow e^{-\tfrac{3}{2}\delta\rho}\hat{D_{\rho}}e^{-\tfrac{1}{2}\delta\rho}.$ (5.11)
Therefore, the variation of the regularized determinant reads
tr[$\tr\left[-\delta\rho+2\sum_{i}c_{i}\left(\delta\rho\frac{M_{i}^{2}}{M_{i}^{2}+\hat{D_{\rho}}^{2}}\right)\right]$
[…]
The WKB calculations of [8] (reproduced here in the Appendix) yield
$\int d^{2}\xi e^{2\rho}\delta\rho\frac{R}{24\pi}=\delta\int d^{2}\xi\frac{\partial_{+}\rho\partial_{-}\rho}{6\pi}.$ (5.13)
Therefore, the variations of terms in the exponential cancel each other, as claimed. ∎
Corollary 5.1.
As a consequence of this theorem, the theory of massive Majorana fermions on a minimal surface with induced metric $g_{ab}=\delta_{ab}e(\xi)$ , and extra conformal term $\frac{(\partial\rho)^{2}}{3\pi}$ in Lagrangian is equivalent to the theory of Majorana fermions on a flat surface with variable mass $m\sqrt{e(\xi)}$. The first theory corresponds to taking $\rho=\log\sqrt{e}$ in $A_{\rho}$, and the second theory corresponds to taking $\rho=0$. By the proven gauge invariance these theories are equivalent.
Remark 5.1 (Moduli and Topology).
The theorem establishes local independence from $\rho(\xi)$. For a disk, all metrics are conformally equivalent to the flat metric, so the determinant is a a unique functional of external loop via the conformal metric $e=|f^{\prime}_{\mu}|^{2}$. For surfaces with handles or holes, the metric can be written as
$g_{ab}=\exp\left(2\rho\right)\,\hat{g}_{ab}(\tau),$
where $\tau$ are the Teichmüller parameters (moduli). The determinant then depends on $\tau$, i.e., $Z_{\mathrm{elf}}[C]=Z[C|\tau]$. In the present geometric theory, the minimal surface is rigid: the boundary loop $C$ uniquely determines the solution to the Plateau problem and thus fixes the moduli $\tau=\tau[C]$. Therefore, the Dirac determinant is a well-defined functional of $C$, encoding the "gluon cloud" effects for the specific minimal surface spanned by $C$. In the degeneration limit (e.g., infinitesimal bridges), the moduli approach a boundary value, but the locality of the loop equation is preserved.
6 The fermion determinants and the planar topology
Let us consider the vacuum functional $Z$ of the Elfin theory with the boundary conditions discussed in the previous sections. By virtue of conformal invariance, we can choose the local gauge $\rho(\xi)=\log\sqrt{e(\xi)}$ or $rho(\xi)=0$ The following analysis of the vacuum loops in this theory simplifies in the gauge $\rho=0$, where it is similar to the topological solution of the 2D Ising model.
As a functional of the loop $C$, the vacuum functional satisfies a certain non-linear path integral equation. This equation will be derived in this section and investigated in the following sections.
6.1 Mathematical justification of the signed sum over loops
The path integral representation of the Dirac determinant on a finite surface with boundary conditions, as employed in this work, is rigorously justified by several mathematical results and earlier physical constructions. In particular, Ichinose [15] and Gaveau [16] provide theorems establishing the equivalence between the determinant of the Dirac operator and path integrals over spinor-valued paths with appropriate boundary conditions on finite domains. These constructions carefully incorporate the spinor connection and holonomies of spin rotations along geodesic segments, ensuring the correct treatment of spinor phases and curvature effects.
Our representation of the determinant as a sum over oriented closed loops weighted by factors of the form $\prod\exp\left(-m\,dl+\frac{\mathrm{i}\sigma_{3}\delta\theta}{2}\right)$, where $dl$ is an infinitesimal geodesic step and $\delta\theta$ the corresponding trajectory rotation angle, follows directly from these rigorous path integral formulations. This approach generalizes the combinatorial and topological methods pioneered by Vdovichenko [17] in the exact solution of the 2D Ising model, where the fermionic nature and spinor holonomies enforce cancelations of non-planar intersections, leaving only planar loop configurations that contribute to the determinant.
Furthermore, the Gauss-Bonnet theorem relates the total tangent rotation angle around a loop to its self-intersection number, providing a topological invariant that governs the sign factors in the loop expansion. This ensures that the fermionic statistics and boundary conditions are correctly encoded in the loop amplitudes.
Thus, the path integral representation employed here is not only physically intuitive but also mathematically sound, grounded in well-established results in spin geometry, spectral theory of Dirac operators, and rigorous path integral constructions on manifolds with boundary. This justifies its use as the foundation for deriving loop equations and analyzing the planar topology of the fermion determinants in the Elfin theory.
6.2 Random loops at the surface
[…]
This produces two determinants of the Dirac operators with the boundary conditions $\lambda\sigma_{3}=\pm 1$:
$I[S]$ $=\int\mathcal{D}\psi e^{A_{0}}=\Delta_{+}\Delta_{-};$ (6.1)
$\Delta_{\pm}$ $=\det(D+m\sqrt{e})_{\pm}$ (6.2)
On an infinite flat surface, such a determinant would be calculable using the trace of the logarithm of a resolvent in Fourier representation. On a finite surface $S$ bounded by an arbitrary curve $C$, there are nontrivial boundary conditions for a Dirac resolvent, complicating the analytic computation.
Instead, we can represent the Dirac determinant as a sum over vacuum loops at the surface. The similar representation in the context of the 2D Ising model was discussed in my earlier work [18], inspired by studies of the Ising model by Vdovichenko [17]. That pioneering paper laid a framework for a topological solution of the Ising model in terms of self-avoiding paths. The next step of our paper [18] was to generalize this method to a continuum Dirac operator on a Riemann surface using the Gauss-Bonnet theorem to relate the rotation angle to the intersection index. There are also some theorems in the mathematical literature [15, 16] establishing the Dirac and Heat kernel path integral representations in a finite domain. In this paper, we follow the technology described in [8], with proper adjustments from a random surface to a Hodge-dual minimal surface.
We base our study on the superposition principle. The vacuum amplitude for the Dirac particle represents the sum over all configurations of the loop $\Gamma\in S$ of the corresponding amplitude for the given loop. The amplitude of the given loop represents the product of the amplitudes of individual events. The individual events can be regarded as a sequence of classical propagations along little geodesic steps alternating with instant rotations.
Technically, we use the logarithm of a Dirac determinant and represent it as a trace of a resolvent integrated over proper time
$\Delta_{\pm}=$
[…]
$\kappa\propto\log\epsilon;$ (6.4)
The logarithmically divergent term $\log\epsilon$ in the exponential is proportional to the area of the surface by virtue of the locality of this divergent contribution to the trace. The terms in the exponential, proportional to the area of the additive Hodge-dual surface, pass through the loop equation, as we already established. The role of these terms is to regularize the effective string tension (or vacuum energy density) of the Fermi model on the surface.
The path integral representation emerges when this exponential is replaced by a limit of a product (see [19])
$\exp\left(-T(D+m\sqrt{e})\right)$ $\to\prod\exp\left(-dT(D+m\sqrt{e}\right);$ (6.5)
$dT$ $=dl/\sqrt{e}$ (6.6)
The mathematical literature replaces this product with a stochastic process that is equivalent to Feynman’s sum over paths, with the path amplitude being a product of WKB amplitudes for these propagations/rotations [16].
The amplitude for the rotation is given by the rotation matrix
$A(\Delta\theta)=\exp\left(\mathrm{i}\sigma_{3}\Delta\theta/2\right).$ (6.7)
The WKB amplitude for the geodesic step reads
[…]
6.3 The Gauss-Bonnet theorem and self-intersections
The total product of the factors (3.5) and (3.6) along the loop can be calculated by means of the Gauss-Bonnet formula which in our notation reads
$\sum\Delta\theta=2\pi(1-\nu)\mod 2\pi.$ (6.9)
Here $\nu$ is the algebraic number of self-intersections of the loop $\Gamma$ at the surface. The Gauss-Bonnet theorem ensures that the relation is topologically invariant.
Remark 6.1.
Mathematical theorems. For an oriented regular simple closed curve $\gamma\subset\mathbb{R}^{2}$, the rotation (turning) index satisfies $\operatorname{rot}(\gamma)=\frac{1}{2\pi}\int_{\gamma}\kappa\,ds\in\mathbb{Z}$ and equals +1+1 for the positively oriented boundary of a disk [20]. In the planar case, this is exactly the Gauss–Bonnet theorem for the enclosed region (since $K\equiv 0$), yielding $\int_{\gamma}\kappa\,ds=2\pi\,\chi=2\pi$ (hence equality modulo $2\pi$ is automatic) [21]. Moreover, since $\operatorname{rot}(\gamma)$ is the degree of the unit-tangent map $T:S^{1}\to S^{1}$, it depends only on the immersed curve itself and therefore is unchanged if the ambient plane is replaced by a disk with finitely many holes (and the same mod-$2$ intersection sign $(-1)^{\nu}$ follows from Whitney’s regular-homotopy classification of immersed plane curves) [20].
The amplitude for the loop reads:
$A[\Gamma]=(-1)^{\nu}\exp\left(-ml_{\Gamma}\right);\qquad l_{\Gamma}=\int_{\Gamma}\sqrt{e}|dz|$ (6.10)
The extra negative sign comes from Fermi statistics. This amplitude is a scalar, independent of the quantum numbers, so we should multiply it by the number of allowed states. Each two-component Dirac particle has only one allowed state which is projected by the operator
$P=\frac{1}{2}(1+\sigma_{a}v_{a})$ (6.11)
where $v_{a}$ is the normalized 2-velocity of the rest frame. In the rest frame, there should be only one component (in 4 dimensions, there were two components).
So we are left with the factor of 2 from the sum over the orientations $\sigma_{3}=\pm 1$. Yet it will be convenient to distinguish the terms with two orientations, i.e. to sum over oriented loops $\Gamma$. We mark the loop by an (anti)clockwise arrow for (left) right orientation $\sigma_{3}=-1,+1$.
The point is that the loops that touch the external boundary at least once are oriented in the right direction according to our boundary conditions.
The reflection amplitude
$R=\frac{1}{2}(1+\lambda\sigma_{3})$ (6.12)
corresponds to the repulsion of the left-hand side components from the wall (see the previous section). As discussed above, one may occupy the wall with the left-hand side states so that this repulsion would come about automatically due to the Pauli principle.
So, the sum over loops involves the sum over orientations only for the inner loops.
6.4 From exponential of free elfin loops to sum over multiple crossing paths
[…]
The factor $N^{-1}=1/l$ in (6.13) accounts for $N$ equivalent choices for the origin of the parametrization of the loop. These factors will cancel the combinatorial factors that arise at the intersections (see below).
We write the integral over loops with the above prescription at the reflections as follows:
$\Delta$ $=\Delta_{+}\Delta_{-}$
$\propto\exp\left(\sum\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.4965pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.23955pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.45624pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.6875pt}}\!\int\mathcal{D}\Gamma(-1)^{\nu}\exp\left(-ml_{\Gamma}\right)\right).$ (6.14)
We ignore the (singular) normalization factor in the fermionic determinant. This factor on a finite surface amounts to an exponential of the area $\exp\left(-\textit{ const }{}|S|\right)$, which we include in the confining factor that is already present in our solution for the Wilson loop, according to the previous section.
[…]
In what follows, it will be convenient to consider the product $\Delta$ rather than the separate factors $\Delta_{\pm}$.
Let us expand (6.4) in a series of multiple integrals over loops:
$\Delta$ $=\sum\frac{1}{n!}$
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.4965pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.23955pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.45624pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.6875pt}}\!\int D\Gamma_{1}...\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.4965pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.23955pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.45624pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.6875pt}}\!\int D\Gamma_{n}(-1)^{\sum\nu_{i}}\exp\left(-m\sum l_{i}\right).$ (6.15)
These loops are independent by construction, but we are going to rearrange the sum in such a way that they will start to repel. This miraculous transformation, first conjectured in [18] is possible due to the following topological and combinatorial observations.
6.5 Cancellation of intersections: step towards planar graphs
This section reproduces (with some simplifications and improvements) the most important results of our ‘81 paper: the topological and geometrical analysis of the path integrals of the Dirac particle on a disk-like surface.
The arguments of this section were inspired by a geometric solution of the 2D Ising model by Natasha Vdovichenko [17]. The crucial observation made in that pioneering work was that the sum of planar loops splitting the plane into domains of up and down spin in the Ising model is equivalent to the sum over freely intersecting independent loops, with phase factors reflecting the rotation of a spin 1/2 particle. The signs arising for intersecting paths lead to cancelation between such paths, leaving only planar loops. The computation in that paper relied on the geometry of the paths on a square lattice and other specifics of the Ising model. Below, the same topological observations will be stripped of the lattice artifacts and applied to a two-fermion system with oriented loops on a Hodge-dual minimal surface.
Consider the given configuration of loops at the surface as a graph consisting of several disconnected parts. Connected subgraphs contain intersecting and touching lines. The total length in exponential factors $\exp\left(-ml\right)$ can be redistributed among the lines of the graph:
$l_{tot}=\sum l_{loops}=\sum l_{lines}.$ (6.16)
This statement is not as trivial as it seems. The length of the line is not just a geometric length of this line drawn on a planar disk; rather, it is an integral
$l_{line}=\int|dz(t)|\sqrt{e\left(z(t),\bar{z}(t)\right)}$ (6.17)
However, the metric $e(z,\bar{z})$ in this integral being local, this length is additive across the pieces made by intersections of these lines on a planar domain. The factorization also applies the sign factor since the number of mutual intersections is always even in the absence of handles at the surface.
$\nu_{tot}=\sum\nu_{loops}=\sum\nu_{graph}\pmod{2}$. (6.18)
The phase volume of the connected subgraph $G$ can also be rearranged in such a way that the old loops lose their individuality:
[…]
$\differential W(step) =d$$=\differential lDx(l)\delta(\dot{x}^{2}-1).$ (6.20)
The factors $1/l$ in front of the old loops are compensated by the same factors that arise from the sums over the numbers of geodesic steps between the intersections of lines.
It is important that there are no other factors in front of the loop. Were there, say, $k$ internal degrees of freedom, the corresponding factor
$\prod_{loops}k$ (6.21)
could not possibly be distributed among the lines of the graph. In the case of $k=1$ one could return to the non-oriented lines, then the phase volume would be distributed again. This is the case of the Ising model, which was considered in the preliminary version [18] of the Elfin theory. This possibility is ruled out by the boundary conditions. As we shall see soon, the peculiar boundary conditions play a central role in the construction of the loop equations.
Let us proceed with the sum over graphs. Each internal line of our graph can be oriented in both directions, but orientation is conserved in the vertices. We neglect multiple collisions as a violation of the Pauli Principle (Grassmann algebra $(\bar{\psi}^{R}\psi^{R})^{n}=0,(\bar{\psi}^{L}\psi^{L})^{n}=0$ for $n>1$, so we deal with the two pairs of vertices depicted at Fig. 3:
Figure 3: Types of line collision: two planar (upper drawing) and one planar, another non-planar (lower drawing)
Naturally, one could treat these two vertices as a single vertex depending on the angles between the lines; however, in this case, there will be step functions of the angles. When the lines cross each other in coordinate space, the term in the vertex changes sign.
The non-planar vertex (the second one on the lower drawing in Fig. 3 ) cancels the first one, so this configuration of lines drops from the sum. This is the heart of the matter. The lines may touch each other with opposite orientations according to Fig. 3, but the terms with intersecting lines are always canceled by those with parallel touching lines in the graph with the same configuration of lines.
Note that in the general case, this is a cancelation of terms with a different number of loops, so there is no room for extra degrees of freedom. This is the Pauli principle reformulated in terms of the path integrals.
We may now return to our old loops but disregard all the non-planar configurations. We obtain a hierarchy of trapped loops that move in the free space left by the others. With one-component fermions, we would obtain the sum over all decompositions of the whole surface into two different phases, which is equivalent to an Ising model.
The above representation of the Majorana determinant would amount to the Onsager solution of the Ising model in terms of free Majorana fermions. In our case, we have twice the number of components, providing for the oriented planar loops instead of just boundaries of the spin domains.
The time slice ($\xi_{2}=\textit{ const }{}$) of this picture describes a string with particles moving between neighbors. Pairs of $\psi$ and $\bar{\psi}$ particles can be created or annihilated. Only $\psi\bar{\psi}$ neighbors can touch each other. The touching of $\psi\psi$ or $\bar{\psi}\bar{\psi}$ neighbors is forbidden by the Pauli principle. One $\psi,\bar{\psi}$ pair is placed at the two ends of the string.  [^2] This interpretation of the time slice dynamics can be used for a rigorous proof of the equivalence of the free Dirac vacuum loops and the planar loops made of non-intersecting closed paths on a surface. Such a proof would use the Hamiltonian representation of a Fermionic system with conventional creation/annihilation operators and chiral boundary conditions.
This resembles the ’t Hooft string model for planar graphs [22]. Our string has the same topology but takes into account the gauge invariance of the planar graph expansion. This will be demonstrated in the next sections.
6.6 Loop equation for fermion determinants
Let us now find the variation of $\Delta$ with respect to the boundary value of $e$.
The basic relation reads
[…]
The invariant $\delta$-function
$e^{-1}\delta^{2}(\xi-\xi_{0})=\delta_{inv}(x,x_{0}),\qquad x,x_{0}\in S,$ (6.23)
[…]
The variation selects the graph that touches the boundary at $x_{0}$. The touching line is oriented to the right according to our boundary conditions.
The amplitude of touching equals m-m; the remaining factors contribute to the phase volume of the graph.
Let us follow the loop that touches the boundary. This is the correct loop; therefore, each loop that touches it from the outside is also correct. We observe that the outside domain $S_{out}$ between this loop $\Gamma$ and the external boundary $C$ is completely equivalent to the original surface $S$. The loops trapped in this domain reflect from $\Gamma$ in the same way as they do from the external boundary (this is why we need the special boundary conditions).
The sum over configurations and orientations of the loops in $S_{out}$ produces precisely the same Dirac determinant.
$\Delta_{out}=\Delta_{+}^{out}\Delta_{-}^{out}$ (6.24)
The loops that are trapped inside $\Gamma$ reflect from inside, so they are oriented to the left at the boundary $\Gamma^{-1}$. The sum over configurations and orientations of these loops produces the Dirac determinant for the internal domain,
$\Delta_{in}=\Delta_{+}^{in}\Delta_{-}^{in},$ (6.25)
with the opposite boundary conditions,
$\psi^{R}=\bar{\psi}^{R}=0,\qquad\text{at }\Gamma^{-1}.$ (6.26)
The numerical value of this determinant is the same as that for the old boundary conditions.
We see, however, that the orientation $\lambda\sigma_{3}$ is related to the orientation of the boundary in space. This will be important in higher order $1/N$ expansion of QCD where the multiloop propagators depend upon the relative orientation of loops.
Note that the loop $\Gamma$ may touch itself as well as the external loop $C$ (at the surface). In this case, one or both of the domains $S_{in},S_{out}$ reduce to a set of windows touching at the corners. The corresponding functional should be understood as
$\Delta=\prod\Delta(windows).$ (6.27)
This is clear from the original representation of $\Delta$ as an exponential of the sum over configurations of independent loops. This planar hierarchical set of loops is shown in Fig. 4 .
Figure 4: The hierarchical set of loops, some touching, but never intersecting each other nor self-intersecting
7 The Minimal Surface and the Loop Equation
In the 1981 theory, the next step was to integrate over all embeddings $X_{\mu}(\xi)$. This led to the Liouville instability. In the present theory, we replace this integration with the Rigid Hodge-Dual Surface $S_{min}[C]$ defined in [1] and described in the section 3 of this paper.
This surface is the unique additive solution to the minimal area problem with the constraint that the area derivative is self-dual. It is fixed by the boundary loop $C$. Therefore, we do not integrate over embeddings. The functional $Z[C]$ is simply the fermion determinant on this specific manifold:
$Z[C]\equiv Z[S_{\chi}[C]].$ (7.1)
For the time being, we fix the Hodge duality $\chi$ of this surface; later, we will discuss the restoration of parity symmetry in the loop equation.
7.1 Loop equations
We now demonstrate that this $Z[C]$ satisfies the loop equation of the same structure as (2.10).
Using the above planar loop equations for Dirac determinants and area derivatives of the HD minimal surface, we may write the following loop equation for the functional $Z[S]$:
$\delta Z[S]/\delta\sigma_{\mu\nu}(x_{1})=-m$
$\int\mathcal{D}\Gamma_{11}\delta_{I,0}T_{\mu\nu}(1)Z[S_{in}]Z[S_{out}]e^{-ml_{11}}.$ (7.2)
This equation is illustrated in Fig. 5
Figure 5: The first area derivative of Dirac determinant. There is one internal path $\Gamma$ which cut the surface $S$ into two pieces $S_{in}=S_{1},S_{out}=S_{2}$
Here, the tensor $T$ was defined in (10.8). The loop $\Gamma_{11}$ touches $C_{11}$ at $x_{1}$. The integer index factor
$I=$
$\int_{\Gamma}dl_{1}\int_{\Gamma}dl_{2}\theta(l_{1}-l_{2})e_{ab}\dot{\xi}_{a}(l_{1})\dot{\xi}_{b}(l_{2})\delta^{2}(\xi(l_{1})-\xi(l_{2}))$ (7.3)
is the absolute number of self-intersections of the loop $\Gamma_{11}$ at the surface. The Kronecker delta $\delta_{I,0}$ in the integral in (7.1) enforces the self-avoiding loop $\Gamma_{11}$. In the following, we denote path integrals over self-avoiding loops as $\int_{\asymp}\mathcal{D}\Gamma$
The measure of the open path is defined as follows:
[…]
Eq. (7.1) was postulated in the first version of the Fermi String theory [18] from heuristic arguments. At that time, it seemed that the equation applied to the Ising model at the surface, i.e., for the 2-component spinors. However, as we see now, one should introduce two spinors with special boundary conditions in order to obtain this equation for the propagator of the Fermi string.
Note that our theory is constructed in such a way that it is equivalent to the flat theory with a variable fermion mass. In the next section, we shall use the second area derivative of $Z[C]$, which reads
$\frac{\delta^{2}Z}{\delta\sigma_{\mu\nu}(l)\delta\sigma_{\alpha\beta}(r)}=A_{\mu\nu\alpha\beta}+B_{\mu\nu\alpha\beta};$ (7.5)
$A_{\mu\nu\alpha\beta}=(m/2)^{2}T_{\mu\nu}(l)T_{\alpha\beta}(r)\iint_{\asymp}\mathcal{D}\Gamma_{l}\mathcal{D}\Gamma_{r}$
$Z[S_{l}]Z[S_{mid}]Z[S_{r}]\exp\left(-m(l_{l}+l_{r})\right);$ (7.6)
$B_{\mu\nu\alpha\beta}=(m/2)^{2}T_{\mu\nu}(l)T_{\alpha\beta}(r)\iint_{\asymp}\mathcal{D}\Gamma_{up}\mathcal{D}\Gamma_{dn}$
$\times Z[S_{up}]Z[S_{mid}]Z[S_{dn}]\exp\left(-m(l_{up}+l_{dn})\right).$ (7.7)
This equation is illustrated in Fig. 6,7.
Figure 6: The first term in the second area derivative of Dirac determinant. There are two closed loops $\Gamma_{l},\Gamma_{r}$ which cut the surface $S$ into three pieces $S_{1}=S_{l},S_{2}=S_{mid},S_{3}=S_{r}$.
Figure 7: The second term in the second area derivative of Dirac determinant. There is one closed loop $\Gamma=\Gamma_{up}\Gamma_{dn}$, touching the loop $C$ at the left and the right points and cutting the surface $S$ into three pieces $S_{1}=S_{mid},S_{2}=S_{up},S_{3}=S_{dn}$.
The first term arose due to the variation of the last factor in (7.1). The second term arose due to the variation of the factor $\exp\left(-ml\right)$ as has been discussed above. In this case, there are two paths $\Gamma_{up},\Gamma_{dn}$ which cut the surface $S$ into three pieces $S_{up},S_{mid},S_{down}$ with the boundaries
$\partial S_{up}$ $=\Gamma_{up}C_{rl},$ (7.8)
$\partial S_{mid}$ $=\Gamma_{dn}^{-1}\Gamma_{up}^{-1},$ (7.9)
$\partial S_{dn}$ $=\Gamma_{dn}C_{lr}.$ (7.10)
The variation of the factor $Z[S_{in}]$ in (7.1) would require one of the internal loops inside $S_{in}$ to touch the external boundary $C$. This implies that $\Gamma$ touches $C$ at the same point, so this is the triple collision that we rule out as a violation of the Pauli Principle (and Grassmann algebra) $(\bar{\psi}^{R}\psi^{R})^{3}=(\bar{\psi}^{L}\psi^{L})^{3}=0$. As for the remaining factors in (7.1) they do not depend upon $e$.
The operator $\mathcal{L}_{\nu}$ involves the gradient of the first area derivative. Using the translation invariance identity [5]
[…]
we can rewrite $\mathcal{L}_{\nu}(Z[C])$ as a similar double integral with the second area derivative
[…]
Substituting the loop equation (7.5) for $Z[C]$ here, we find two terms corresponding to the two drawings in Fig. 6 and 7.
7.2 Large fermion mass and induced QCD
We assume that the fermion mass $m$ is much larger than the physical scale $\Lambda_{QCD}$, along the lines of the induced QCD scenario [23]. As we shall see, in this limit, the mass of the fermion $m$ serves as a UV cutoff, and the $A$-term reduces to the same loop operator applied to the area of the Hodge-dual minimal surface. This term vanishes by the Bianchi identity, as we have established in our previous work and summarized in the beginning of this paper. The second term reduces, in this limit, to the right side of the MM equation multiplied by a coefficient determined by the local properties of the Dirac determinant on a flat minimal surface in the vicinity of the self-intersection of the loop.
These conclusions will be based on a well-known property of geodesic paths on Riemann surfaces, which we formulate as a Lemma.
Lemma 7.1 (Geodesic Inequality).
For any minimal surface embedded in Euclidean space, the geodesic distance $L_{geo}(x,y)$ is strictly bounded by the Euclidean distance:
$L_{geo}(x,y)\geq|x-y|_{\mathbb{R}^{4}}.$ (7.13)
Proof.
The geodesic distance between two points $x,y$ along the surface embedded in Euclidean space is a constrained minimum of the length of paths connecting these points in Euclidean space. The constraint is that every point of this path belongs to the surface. The set of such paths is a subset of all Euclidean paths, and the line element $dl=|d\xi|\sqrt{e}$ in the induced metric $e=|f^{\prime}|^{2}$ coincides with the Euclidean line element $ds=\sqrt{|df|^{2}}$; therefore, the minimal length of the subset path is greater or equal to the target space distance $\int ds=|x-y|_{\mathbb{R}^{4}}$. ∎
Remark 7.1.
Note that we used the induced metric $e=|f^{\prime}|^{2}$, the same as for the Euclidean minimal surface (Goldschmidt solution to the Plateau problem), rather than the full induced metric on a general surface in $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$. The metrics coincide on a holomorphic minimizer, but for the area derivatives, we need the full external geometry in the 12 dimensional space $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$. Once the area derivatives are taken (as they are in our loop equations, resulting in Hodge-dual tensors $T_{\mu\nu}$) we could use a holomorphic minimizer for the induced metric in path integrals.
We are going to use this theorem for the terms arising in the loop equation for the Dirac determinants.
7.3 The $A$-term
The first term (see Fig. 6) in the limit of large fermion mass (much larger than the external curvature $Q_{ab}=\operatorname{diag}\left\{-\rho,\rho\right\}$ of the minimal surface at its boundary $C$) effectively involves infinitesimal flat loops $\Gamma_{l},\Gamma_{r}$. The factor $Z[S_{mid}]$ reduces to the constant $Z[S[C]]$, which we take out of the path integrals $\iint_{\asymp}\mathcal{D}\Gamma_{l}\mathcal{D}\Gamma_{r}$. These integrals can be computed on an infinite flat semiplane, and they reduce to constants independent of the points $x_{l},x_{r}$. We find
$A_{\mu\nu\alpha\beta}\to T_{\mu\nu}(l)T_{\alpha\beta}(r)Z[C]a^{2};$ (7.14)
$a=m/2\int_{\asymp}\mathcal{D}\Gamma Z[\Gamma]\exp\left(-ml[\Gamma]\right);$ (7.15)
Now, we recall that the area derivative of the minimal surface is also proportional to the tangent tensor $T_{\mu\nu}$, and let us factor out of $Z[C]$ the vacuum energy factor (with $S[C]$ being the Hodge-dual minimal area)
$Z[C]$ $=G[C]Z_{1}[C];$ (7.16)
$G[C]$ $=\exp\left(-\kappa S[C]\right)$ (7.17)
We can rewrite the factor $TT$ as a second area derivative
$A_{\mu\nu\alpha\beta}\to\frac{-a^{2}Z_{1}[C]}{4\kappa^{2}}\frac{\delta^{2}G[C]}{\delta\sigma_{\mu\nu}(l)\delta\sigma_{\alpha\beta}(r)};$ (7.18)
which is valid for arbitrary $\kappa$ and $x(l)\neq x(r)$. Now, we use the translational identity (7.11) backwards
[…]
$\frac{a^{2}Z_{1}[C]}{4\kappa^{2}}\partial_{\mu}\frac{\delta G[C]}{\delta\sigma_{\mu\nu}(l)}$ (7.19)
The last integral, as we already established, vanishes for the Hodge dual minimal surface by the Bianchi identity. Thus, the (unwanted) AA- term in the loop equation vanishes in the local limit.
7.4 The $B$- term
The $B$- term in the Dirac determinant loop equation has a completely different structure, and it is singular in the local limit.
$B_{\mu\nu\alpha\beta}=(m/2)^{2}T_{\mu\nu}(l)T_{\alpha\beta}(r)\iint_{\asymp}\mathcal{D}\Gamma_{up}\mathcal{D}\Gamma_{dn}$
$\times Z[S_{up}]Z[S_{mid}]Z[S_{dn}]\exp\left(-m(l_{up}+l_{dn})\right).$ (7.20)
The $B$-term in the loop equation
[…]
involves the matrix product of two $T$ tensors
$TT_{\alpha\nu}=T_{\alpha\mu}(r)T_{\mu\nu}(l)$ (7.22)
which can be reduced further using the algebra of the $\eta$ tensors (valid for both Hodge-chiralities $\chi=\pm 1$)
$\eta^{i,\chi}_{\alpha\mu}\eta^{j,\chi}_{\mu\nu}=-\delta_{ij}\delta_{\alpha\nu}-e_{ijk}\eta^{k,\chi}_{\alpha\nu}$ (7.23)
Both tensors $T$, being Hodge-dual, can be represented as
$T_{\mu\nu}(x)=\frac{1}{2}\vec{\tau}(x)\cdot\vec{\eta}^{\chi}_{\mu\nu}$ (7.24)
with some unit vector $\vec{\tau}(x)^{2}=1$, depending on the point $x\in C$. This tangent tensor was defined as a limit of the normalized area element $\Sigma_{\mu\nu}$ on the surface when a point $x\in S[C]$ approaches the boundary $C$.
Afterward, the matrix product becomes
$TT_{\alpha\nu}=\frac{-1}{4}\left((\vec{\tau}(r)\cdot\vec{\tau}(l))\delta_{\alpha\nu}+(\vec{\tau}(r)\times\vec{\tau}(l))\cdot\vec{\eta}_{\alpha\nu}\right)$ (7.25)
Now, we explore the consequences of the geodesic inequality. In the local limit, the points $l,r$ belong to the self-intersection of the loop $C$. The Goldschmidt minimal surface (additive in the limit when $C=C_{lr}\cdot C_{rl}$) breaks into two surfaces covering closed loops connected at the self-intersection point, as shown in Fig 8.
Figure 8: Self-intersecting loop.
The zoom into the vicinity of the self-intersection is shown in three dimensions in Fig. 9.
Figure 9: The additive minimal surface in the vicinity of the self-intersection. There is a narrow twisted strip (White’s bridge) connecting the minimal surfaces bounded by two parts of the self-intersecting loop.
The minimal surface is, in general, curved, though the arithmetic mean curvature vanishes everywhere. In the vicinity of the self-intersection, for our Goldschmidt solution, there is an infinitesimal narrow bridge connecting two parts bounded by loops $C_{lr},C_{rl}$. This is proven in White’s bridge theorem [24]. This property places the tensors $T(l),T(r)$ in the infinitesimal vicinity of each other on this bridge. In the local limit, we have $\vec{\tau}(l)\to\vec{\tau}(r)$, so that the matrix product
$T_{\alpha\mu}(r)T_{\mu\nu}(l)\to\frac{-1}{4}\delta_{\alpha\nu}$ (7.26)
As for the rest of the factors, the sum over the short paths $\Gamma_{up},\Gamma_{dn}$ is dominated by the geodesic distance, which is bounded by the Euclidean distance. This precise alignment relies on the isotropic scaling of White’s bridge in the varifold limit $\epsilon\to 0$. In the strict limit of the topology change, the continuity of the area Hessian demands that the tangent planes perfectly align, yielding the required Fierz-like identity for the MM loop equation without generating anomalous contact terms. A rigorous geometric measure-theory proof of the Hessian’s continuity across this topology change is left for future mathematical study, but is physically mandated by the $O(4)$ symmetry of the local loop diffusion. We find:
$B_{\mu\nu\alpha\mu}\to-b\delta_{\alpha\nu}Z[C_{lr}]Z[C_{rl}]\delta^{4}(C(l)-C(r));$ (7.27)
$b\delta^{4}_{m}(C(l)-C(r))=$
$\frac{m^{2}}{16}\iint_{\asymp}\mathcal{D}\Gamma_{l}\mathcal{D}\Gamma_{r}Z[S_{mid}]\exp\left(-m(l_{l}+l_{r})\right).$ (7.28)
Figure 10: Two arks $\Gamma_{l},\Gamma_{r}$ connecting self-intersection points, bounding the crescent-shaped area $S_{mid}$.
This path integral over small surface paths $\Gamma_{l},\Gamma_{r}$ is illustrated in fig.10, where these paths are two arcs bounding the crescent-shaped $S_{mid}$. We use an approximation of the delta-function
$\delta^{4}_{m}(x-y)=4m^{4}\exp\left(-2m|x-y|\right)$ (7.29)
Putting the pieces together, we find the loop equation (2.10) with
$W[C]=\frac{Z[C]}{Z[\mathds{1}]};$ (7.30)
$\lambda\delta^{4}_{m}(C(l)-C(r))=\frac{m^{2}Z[\mathds{1}]^{2}}{16}$
$\iint_{\asymp}\mathcal{D}\Gamma_{up}\mathcal{D}\Gamma_{dn}\exp\left(-m(l_{up}+l_{dn})\right)W[C_{up}\cdot C_{dn}].$ (7.31)
The chirality of the Hodge dual minimal surface dropped from this equation, as it dropped from the whole internal geometry, including the minimal area and the Dirac determinant. Therefore, there is no necessity to symmetrize our solution over Hodge chirality to restore the parity of QCD: the parity is preserved by the Dirac determinant on a Hodge-dual minimal surface.
The sign is positive, as it should be for QCD with a positive ’t Hooft coupling constant $\lambda$. As for the value of this coupling constant, it is determined by normalization $Z[\mathds{1}]$. The string tension normalization factor $G[C]=\exp\left(-\kappa S[C]\right)$ cancels on both sides of the loop equation, so it does not influence $\lambda$, but the normalization factor $Z[\mathds{1}]$ at vanishing loop multiplies this constant twice.
The asymptotically free QCD we are looking for corresponds to the limit when this factor goes to zero. In that limit, the standard RG computations based on planar graphs would produce the QCD mass scale
$\Lambda_{QCD}=m\lambda^{-\tfrac{51}{121}}\exp\left(-\frac{24\pi^{2}}{11\lambda}\right);$ (7.32)
Therefore, all we need is the limit of the Elfin theory such that this factor vanishes
$\lambda\propto Z[\mathds{1}]^{2}\to 0;$ (7.33)
The rest of perturbative QCD will follow from the loop equation by iterations in $\lambda$ starting with $W[\mathds{1}]=1$. Let us now outline this perturbative QCD derivation.
8 Asymptotically free planar QCD from the loop equation
The Elfin theory induces QCD in the limit of large fermion mass, serving as a short-distance cutoff. There are no singularities at finite mass, which eliminates the messy discussion about cusp and perimeter singularities of the Wilson loop. The Wilson loop is defined in the Elfin theory as a regularized solution of the loop equation, with the delta function smeared by a finite fermion mass. This definition of the loop equation makes mathematical sense, in our opinion: regularize this equation by smearing the delta function on the right side while preserving the gauge invariance by keeping all loops closed. The Elfin theory falls into the category of regularized solutions of the loop equation.
[…]
This model will be discussed in detail in the next section; in the rest of this section, we shall derive the planar graph from the loop equation in the context of Elfin theory.
8.1 Bootstrap equation and planar graphs with glassed windows
The planar graphs in Feynman gauge, including the ghost loops, were reproduced from the loop equation in the original MM paper [7] using the so-called Bootstrap equation (section 6 in that paper). This equation was based on the inversion of the point derivative operator $\partial_{\mu}$ in the equations of motion, using the Bianchi identity. The area derivative was split into linear $B_{\alpha}$ and quadratic $B_{\mu\nu}$ terms
$\frac{\delta W[C]}{\delta\sigma_{\mu\nu}(x)}=\left(\partial^{x}_{\nu}\delta_{\mu\alpha}-\partial^{x}_{\mu}\delta_{\nu\alpha}\right)B_{\alpha}[C_{xx}]+B_{\mu\nu}[C_{xx}];$ (8.1)
The linear terms were related to the quadratic terms and the source term $J_{\nu}[C_{xx}]$ on the right side of the MM equation
$B_{\nu}=\left(-\partial^{2}\delta+[\partial,\partial]\right)^{-1}\left(J_{\lambda}-\partial_{\mu}B_{\mu\lambda}\right);$ (8.2)
$J_{\nu}[C_{xx}]=\lambda\int dy_{\nu}\delta^{4}(x-y)W[C_{xy}]W[C_{yx}]$ (8.3)
The next step was to expand the inversion operator in powers of the commutator
$\left(-\partial^{2}\mathds{1}+[\partial,\partial]\right)^{-1}=-\partial^{-2}\mathds{1}$
$-\partial^{-2}[\partial,\partial]\partial^{-2}-\partial^{-2}[\partial,\partial]\partial^{-2}[\partial,\partial]\partial^{-2}+\dots$ (8.4)
and use the identity
$[\partial_{\alpha},\partial_{\beta}]F[C_{xx}]=\left.\frac{\delta F[C_{xx}]}{\delta\sigma_{\alpha_{\beta}}(y)}\right|^{y=x-0}_{y=x+0}$ (8.5)
The last step was to represent the inversion of the point derivative operator $\partial^{-2}$ as a sum over the Brownian path $\Gamma$ in 4D space of the functional $J[C_{xx}]$ transported along this path by adding "wires" to the loop
$-\partial^{-2}J_{\nu}[C_{xx}]=\lambda\int d^{4}z\int\mathcal{D}\Gamma_{xz}$
$\oint\displaylimits_{\Gamma_{zx}\cdot C_{xy}\cdot C_{yx}\cdot\Gamma{xz}}dy_{\nu}\delta^{4}(z-y)W[\Gamma_{zx}\cdot C_{xy}]W[C_{yx}\cdot\Gamma_{xz}]$ (8.6)
The important property of this bootstrap equation is that it involves the path integral over Brownian paths $\Gamma_{xz}$ in Euclidean 4D space, which is the basis for the further reconstruction of the gluon propagators . This integral representation is depicted on Fig.11.
Figure 11: The path integral in (8.1) with delta function represented by a straight double line, and Brownian paths $\Gamma_{xz},\Gamma_{zx}$ by a crescent.
8.2 Gluon graphs by iterations of Bootstrap equation
In the old MM papers [7, 25], the bootstrap equation was iterated in the coupling constant $\lambda$, starting with $W[\mathds{1}]=1$.
The derivation was rather tedious, but it faithfully reproduced conventional planar graphs, including Faddeev-Popov ghost loops. The first order planar graph follows immediately from the (8.1), as shown on Fig. 11, by replacing $W$ by $1$ in the path integral, after which this path integral yields the gluon propagator $1/(x-y)^{2}$. The second order graphs are already nontrivial (see [7], Appendix B). This is where the Faddeev-Popov ghost loops emerge directly from iterations of the bootstrap loop equation; this is also where the beta function appears, leading to asymptotic freedom with its running coupling constant. Perturbative QCD is reproduced by the bootstrap equation to all orders, as it was further investigated in [25].
Let us summarize this section as follows.
  • 1.
    The planar MM equation (2.10) can be transformed into an equivalent path integral equation by inverting the non-commutative operator $\partial{}$. The new equation expresses the area variation as a non-linear functional of $W[C]$, which can be represented as a series of frame diagrams, like the one in Fig. 11.
  • 2.
    Frame diagrams can be constructed by means of a certain operator technique. The frame diagram of an arbitrary order looks like a planar tree with windows glassed with $W$ functionals. Integration over internal paths is implied.
  • 3.
    Perturbatively, when $W[C_{ab}]\Rightarrow 1$ in a framed graph, the sums over Brownian paths produce the gluon propagators, and the commutators $\partial{}$$\partial{}$] generate proper tensor numerators in the gluon graphs.
  • 4.
[…]
  • 5.
    The bootstrap equation is equivalent to the system including the Bianchi identity and the planar equation accompanied by euclidean boundary conditions. For this reason, iterative solution of the bootstrap equation recovers uniquely Faddeev-Popov perturbation theory (see appendices of [7] and [25]).
  • 6.
    Various parts of the loop integral correspond to various graphs of ordinary perturbation theory including ghost loops. Being gauge variant separately, these graphs combine into a manifestly gauge-invariant diagram in the loop space.
9 Effective action, its geometry and its singularities
The formal proof that the Elfin determinant $Z[C]$ in the limit of large Elf mass and vanishing normalization $Z[\mathds{1}]$ satisfies the MM loop equation with an asymptotically free coupling constant $\lambda\to 0$ raises many questions. The loop equation is local, due to the delta function on the right side. However, as we all know, the perturbative expansion of $W[C]$ in powers of the ’t Hooft coupling constant $\lambda$ produces the planar graphs inside the loop $C$, which are nonlocal functionals of the loop because of the massless gluon propagator $\delta_{\mu\nu}/(x-y)^{2}$ in coordinate space. We get a sequence of logarithmically divergent integrals like
$\oint_{C}dx_{\mu}\oint_{C_{xx}}dy_{\mu}/(x-y)^{2},$
leading to logarithmic singularities at the cusps of the loop $C$.
Where is this non-locality hiding in the Elfin theory, and how will these logarithms emerge? In this section, we dig deeper into the formal solution we have found, searching for the perturbative singularities of planar QCD.
9.1 Local expansion of the heavy fermion determinant
[…]
$W_{\text{eff}}[C]\approx\int_{\Sigma}d^{2}z\sigma e^{2\rho}+\frac{|\partial_{z}\rho|^{2}}{3\pi}+O(m^{-2});$ (9.1)
$\rho(z,\bar{z})=\frac{\log(\bar{\lambda}(\bar{z})\lambda(z))+\log(\bar{\mu}(\bar{z})\mu(z))}{2};$ (9.2)
The leading term is the Area Law, where the tension $\sigma\sim m^{2}$ provides the confining potential. The subleading term is the conformal anomaly (Liouville action). There is always another zero mode confining factor $G[C]=\exp\left(-\delta\sigma S[C]\right)$, renormalizing the string tension $\sigma$, as we found in the previous work [1]. Equivalently, the area term generated in the heavy-mass expansion of the determinant provides an additive renormalization of the vacuum-energy/string tension sector, already parametrized by the zero-mode dressing $\exp\left((\right)-\kappa S[C])$; only the sum defines the single physical string tension $\sigma_{phys}=2\sqrt{2}\kappa_{phys}$ in the continuum theory, while $M$ acts strictly as a regulator scale removed in the local limit.
In our framework, we do not integrate over fluctuating metrics (which would lead to the Polyakov-Liouville anomaly). Instead, the surface $\Sigma$ is the rigid solution to the Plateau problem bounded by the loop $C$. Consequently, the anomaly term becomes a functional of the boundary loop itself:
$S_{\text{shape}}[C]=\frac{1}{12\pi}\int_{D}\left|\partial_{z}\left(\log\bar{\lambda}\lambda+\log\bar{\mu}\mu\right)\right|^{2}d^{2}z$ (9.3)
This functional is non-local with respect to the loop coordinates $C_{\mu}(\theta)$, as it requires solving the bulk Laplace equation to determine $\lambda(z),\mu(z)$. It represents the "Loewner Energy" [26] of the loop, acting as a geometric stiffness that penalizes deviations of the flux tube cross-section from circularity.
9.2 Cusp singularities and the Hilbert transform
The analytic structure of the minimal surface provides a natural regularization of the ultraviolet divergences associated with the Wilson loop, while correctly reproducing the universal singularities of Gauge Theory. The holomorphic tangent vector is reconstructed from the boundary loop velocity $\dot{C}_{\mu}$ via the Hilbert transform $H$:
$2\mathrm{i}e^{\mathrm{i}\theta}f^{\prime}_{\mu}(e^{\mathrm{i}\theta})=\dot{C}_{\mu}(\theta)+\mathrm{i}H[\dot{C}_{\mu}](\theta),$ (9.4)
$H[u](\theta)=\frac{1}{2\pi}\text{P.V.}\int d\theta^{\prime}\cot\left(\frac{\theta-\theta^{\prime}}{2}\right)u(\theta^{\prime})$ (9.5)
For smooth loops (such as the circle or helicoid), $H[\dot{C}]$ is smooth, and the area density is finite everywhere. However, for polygonal loops used in scattering amplitudes, the tangent vector $\dot{C}_{\mu}$ has discontinuities (cusps).
At a cusp, the Hilbert transform of the step function generates a logarithmic singularity:
$H[\dot{C}_{\mu}](\theta)\sim\log|\theta-\theta_{cusp}|$ (9.6)
[…]
9.3 New view at the cusp singularity
The logarithmic singularity at the cusps, derived in perturbative QCD by Korchemsky and Radyushkin [27], finds a natural interpretation in our framework. In standard perturbation theory, the Wilson loop with a cusp of angle $\gamma$ (in Euclidean space) exhibits a divergence:
$\log W$ $\sim-\Gamma_{cusp}(\gamma)\log(\Lambda_{UV}L)$
$\sim-\frac{\lambda}{8\pi^{2}}(\gamma\cot\gamma-1)\log(\Lambda_{UV}L)$ (9.7)
In the standard renormalization group approach, this divergence is absorbed into a multiplicative renormalization constant. However, in our geometric framework, the logarithmic divergence arises physically from the energy density of the minimal surface near the corner, generated by the pole in the Hilbert transform.
Rather than removing this term, we observe that the theory possesses an intrinsic ultraviolet cutoff $\Lambda$ determined by the mass scale of the heavy Elf particles, which is equivalent to the regularization of the fermion determinant. In the continuum limit $\Lambda\to\infty$, the bare ’t Hooft coupling $\lambda(\Lambda)$ must vanish according to the law of asymptotic freedom:
$\lambda(\Lambda)\approx\frac{24\pi^{2}}{11\log(\Lambda/\Lambda_{QCD})}$ (9.8)
Consequently, the "divergent" cusp energy becomes the product of a vanishing coupling and a diverging geometric factor. This product tends to a finite, universal limit:
$E_{cusp}$ $\sim\lim_{\Lambda\to\infty}\lambda(\Lambda)\log(\Lambda L)\times(\text{Geometric Factor})$
$=\text{finite}$ (9.9)
Specifically, using the one-loop beta function coefficient, the cusp contribution to the effective action becomes a scale-invariant geometric potential:
$S_{cusp}=\frac{3}{11}(\gamma\cot\gamma-1)$ (9.10)
This mechanism is strictly analogous to the continuum limit in Lattice Gauge Theory, where the bare coupling $g_{0}\to 0$ as the lattice spacing $a\to 0$ to keep physical masses finite. In our induced QCD, the cutoff has a physical and mathematical origin in the regularization of the minimal surface, so the limit $m\to\infty$ (where $\lambda\to 1/(\beta_{0}\log m)$) has a precise mathematical meaning, replacing the formal subtraction of infinities with a finite limiting process.
9.4 Spinor factorization and the geometry of Gauss maps
The twistor factorization [28] of the null tangent vector $f^{\prime}_{a\dot{b}}=\lambda_{a}\mu_{\dot{b}}$ leads to a factorization of the induced metric on the worldsheet, as summarized in our B. Using the isomorphism $SO(4)\cong SL(2)_{L}\times SL(2)_{R}$, the metric density factorizes into the product of invariant norms of the left and right spinors:
$e(z,\bar{z})=2\sqrt{2}f^{\prime}_{\mu}\bar{f}^{\prime}_{\mu}=\sqrt{2}\bar{\lambda}\lambda\bar{\mu}\mu$ (9.11)
The area derivative (3.8) in the spinor representation simplifies as follows
$\frac{\delta S_{\pm}}{\delta\sigma_{\alpha\beta}}=-n^{\pm}_{i}\eta^{i\pm}_{\alpha\beta};$ (9.12)
$n^{\pm}_{i}=\left\{\frac{\bar{\lambda}\sigma_{i}\lambda}{\bar{\lambda}\lambda},\frac{\bar{\mu}\sigma_{i}\mu}{\bar{\mu}\mu}\right\}$ (9.13)
The Liouville field $\rho=\frac{1}{2}\log(f^{\prime}_{\mu}\bar{f}^{\prime}_{\mu})$ splits into a sum of potentials:
$\rho(z,\bar{z})=\Phi_{L}(z,\bar{z})+\Phi_{R}(z,\bar{z}),$
$\text{where }\Phi_{L,R}=\left\{\frac{1}{2}\log\bar{\lambda}\lambda,\frac{1}{2}\log\bar{\mu}\mu\right\};$ (9.14)
Geometrically, $\Phi_{L}$ and $\Phi_{R}$ are the Kähler potentials for the maps from the worldsheet to the projective spinor spaces $\mathbb{CP}^{1}_{L}$ and $\mathbb{CP}^{1}_{R}$. These maps $n_{i}^{\pm}$ are identified with the Left and Right Gauss maps of the minimal surface [29]. In our particular case of the Hodge-dual minimal surface $S_{\chi}$, the area element reduces to a single Gauss map for a given Hodge chirality $\chi=\pm 1$.
$\Sigma_{\mu\nu}\propto n^{\pm}_{i}\eta^{i\pm}_{\alpha\beta}H^{\mp}(z,\bar{z});$ (9.15)
$H^{\pm}(z,\bar{z})=\left\{\bar{\lambda}\lambda,\bar{\mu}\mu\right\};$ (9.16)
The area element still depends on both left and right spinors, but it transforms under $SO(4)$ as a single Gauss map, corresponding to its chirality. The anomalous part of the effective action, $I\propto\int(\nabla\rho)^{2}$, therefore expands into:
$I=I_{L}[\lambda]+I_{R}[\mu]+2\int_{D}\nabla\Phi_{L}\cdot\nabla\Phi_{R}\,d^{2}z$ (9.17)
The first two terms are the actions for the Sigma models on $\mathbb{CP}^{1}$, which are topological invariants (proportional to the winding numbers of the Gauss map). The third term represents a non-trivial interaction between the left and right sectors. Unlike the metric, the action does not decouple; this interaction term measures the integrated correlation between the curvatures of the left and right spinor bundles. It represents the geometric energy cost of the relative orientation of the left and right Gauss maps required to satisfy the boundary conditions.
10 Momentum loops and continuum limit
10.1 Regularization, renormalizability, and the microscopic definition of Planar QCD
Since the inception of QCD half a century ago, the understanding of its regularization and renormalizability has evolved significantly. Currently, Lattice QCD serves as the standard microscopic definition of the theory. In this framework, renormalizability reduces to the mathematical statement that, in the limit of a vanishing bare coupling constant, the physical mass scale $\Lambda_{QCD}$ (measured in units of inverse lattice spacing) exponentially approaches zero. Consequently, the lattice spacing vanishes when measured in observable physical units such as the proton radius.
However, the lattice definition comes at a cost: it explicitly breaks the continuous symmetries of four-dimensional Euclidean space, from the rotation group $O(4)$ to subtle topological structures like the Hodge duality of the area tensor. For instance, exact instanton solutions or Hodge-dual minimal surfaces do not exist on a discrete grid; they only emerge in the continuum limit.
These disadvantages are compensated by the universality of the Renormalization Group (RG) flow. It is well-established (though unproven) that asymptotically free theories reach a universal fixed point at large scales, regardless of the specific microscopic regularization, provided the symmetries are restored in the infrared. This universality allows for various microscopic definitions of QCD, provided they land in the correct phase (the basin of attraction of the asymptotically free fixed point).
History offers cautionary tales. The large-$N$ reduced Eguchi-Kawai model, for example, encountered unphysical fixed points. Similarly, the original Induced QCD proposal [23]—which attempted to induce gauge dynamics via heavy free fermions with vanishing color currents on a lattice—failed because the theory collapsed into the wrong phase rather than asymptotically free QCD. Our Elfin theory is a variation on this theme: heavy free fermions inducing QCD. However, the critical distinction is that our fermions live on a rigid minimal surface rather than a four-dimensional lattice. This geometric constraint keeps the theory non-trivial in the local limit, making it a viable regulator for Planar QCD.
A central question arises regarding the renormalization of the Wilson loop itself. Standard perturbation theory dictates that $W[C]$ requires multiplicative renormalization to remove divergences associated with cusps and self-intersections. Are we obliged to redefine the loop equation to make the coordinate-space Wilson loop finite in the local limit?
We argue that this is a misguided requirement that obscures the underlying dynamics. Physical renormalizability demands only the existence of a universal mass scale $\Lambda_{QCD}$ and the finiteness of observable scattering amplitudes. The Wilson loop $W[C]$ is an intermediate, off-shell quantity. Physical observables involve path integrals over quark trajectories, which are governed by a Brownian measure. A Brownian trajectory is nowhere differentiable; it consists of an infinite density of cusps. Attempting to renormalize the Wilson loop for a smooth contour $C$ is physically artificial because the dominant contributions to the path integral come from fractal paths where the concept of a "cusp angle" is ill-defined.
Therefore, instead of forcing the non-renormalizable coordinate Wilson loop to be finite, we adopt a radically different strategy:
  1. 1.
    We keep the theory regularized at the level of the Wilson loop $W[C]$ by the finite Elf mass $m$ (which acts as the UV cutoff).
  2. 2.
    We perform the summation over quark paths to construct the momentum loop amplitude $W[P]$.
  3. 3.
    Only after transitioning to Momentum Loop Space do we take the local limit ($m\to\infty$).
As demonstrated in this paper, the transition to momentum space integrates out the geometric singularities. The momentum amplitude $W[P]$ remains finite and satisfies an algebraic-differential equation. By taking the bare coupling to zero as $m\to\infty$ in accordance with the RG beta function, we recover the finite, universal observables of Planar QCD without ever needing to define a "renormalized" Wilson loop in coordinate space.
10.2 Momentum loop path integral
As we discussed in the previous section, the ordinary Wilson loop is not directly observable, and its notorious singularities are merely intermediate results that will eventually disappear (cancel or integrate out) in observables.
Unlike the standard Feynman-Schwinger worldline representation in coordinate space, which is inherently plagued by the UV cusp singularities of $W[C]$, we utilize the phase-space path integral representation introduced by the author in [30]. In this formulation, the exact planar quark amplitude factorizes strictly into a universal kinematic Dirac phase-space trace $\mathcal{K}[P]$ and the dynamical gluonic momentum-loop functional $W[P+Q]$.
The amplitude for the propagation of a free Dirac particle in the phase-space loop (the Dirac loop) is:
[…]
The full amplitude for the quark loop in the Planar QCD vacuum, with some external momenta $q_{1},\dots q_{n}$ injected into these amplitudes,is a product of the Dirac loop and Wilson loop times the vertex factor for the injection of external momenta:
[…]
$Q_{\mu}(\tau)=\sum_{k}q_{k}\Theta(\tau-\tau_{k});\qquad\sum_{k}q_{k}=0$ (10.2)
$W[P]=\lambda\int\mathcal{D}CW[C]\exp\left(\mathrm{i}\int d\tau P_{\mu}(\tau)\dot{C}_{\mu}(\tau)\right)$ (10.3)
[…]
The momentum loop amplitude $W[P]$ was introduced in our papers [31, 30, 32] and was recently used in the solution of the Navier-Stokes turbulence problem [33, 3]. The advantage of the momentum loop space is that the loop equation for $W[P]$ is purely algebraic/differential. There are no delta functions; just polynomial functionals of the local momentum $P(\tau)$ and some non-singular functional derivatives of $W[P]$. The coordinate delta function, which leads to all perturbative singularities such as logarithmic divergences at the cusps, disappears in momentum space in the same way that it disappears from the Klein-Gordon equation
$(-\partial^{2}+M^{2})G=\delta(x-x_{0})\Rightarrow(p^{2}+M^{2})\tilde{G}=1$
Figure 12: The momentum loop amplitude, $A(q_{1},\dots q_{n})$ with the inside part of the wheel corresponding to Momentum loop $W[P]$, and the outer rim corresponding to Dirac path amplitude $\mathcal{K}[P]$. The paths $P(t)$ are random, interacting with inner geometry of the string surface represented by amplitude $W[P]$.
10.3 The Momentum loop equation
The MM equation (2.10), when transformed into the momentum loop equation (MLE) by functional Fourier transform, also has a simple algebraic structure ( expanded in so-called Magnus invariant forms [34])
$W[P]=\sum_{n}\mathcal{W}^{(n)}_{\alpha_{1},\dots\alpha_{n}}\Omega^{n}_{\alpha_{1}\dots\alpha_{n}}[P];$ (10.4)
[…]
where $\hat{T}\int$ stands for the ordering of the integration angles on the circle. The MLE becomes an algebraic differential equation in terms of Magnus forms (a.k.a iterated path integrals):
[…]
$Q_{\nu}[P]=T^{\alpha\beta\gamma}{}_{\nu}\Omega^{3}_{\alpha\beta\gamma}[P];$ (10.7)
$T^{\alpha\beta\gamma}{}_{\nu}=\delta_{\alpha\beta}\delta_{\gamma\nu}+\delta_{\gamma\beta}\delta_{\alpha\nu}-2\delta_{\alpha\gamma}\delta_{\beta\nu};$ (10.8)
The algebraic expression on the left side emerged as a Fourier transformation $\frac{\delta}{\delta\dot{C}}\Rightarrow\mathrm{i}P$ of the operator
$\hat{L}_{\nu}(0)$ $=\partial_{\mu}\frac{\delta}{\delta\sigma_{\mu\nu}(0)}$
$=T^{\alpha\beta\gamma}{}_{\nu}\frac{\delta^{3}}{\delta\dot{C}_{\alpha}(\theta-0)\delta\dot{C}_{\beta}(\theta)\delta\dot{C}_{\gamma}(\theta+0)}$ (10.9)
The notorious delta function on the right (which caused all the perturbative singularities in $W[C]$) disappeared in the Fourier integral (see original papers [31, 30, 32] for details).
Here is how this happens in our new framework, with the closed loop described by the position field $C_{\nu}(\theta)$ subject to the closure constraint $C(2\pi)=C(0)$. The linear functional measure in loop space can be written in one of two equivalent forms ( with $v(\theta)=C^{\prime}(\theta)$)
$\mathcal{D}C$ $\propto\delta^{4}\left(C(2\pi)-C(0)\right)\prod_{\theta=0}^{2\pi}d^{4}C(\theta)$
$\propto\delta^{4}\left(\oint v(\theta)d\theta\right)\prod_{\theta=0}^{2\pi}d^{4}v(\theta)$ (10.10)
It is factorizable at self-intersections, which leads to the factorization in the momentum loop equation, with the delta function absorbed by a measure
$\delta^{4}\left(C(2\pi)-C(0)\right)\delta^{4}\left(C(\theta_{1})-C(\theta_{2})\right)\prod_{\theta=0}^{2\pi}d^{4}C(\theta)$
$=\delta^{4}\left(C(\theta_{1})-C(\theta_{2})\right)\prod_{\theta^{\prime}=\theta_{1}}^{\theta_{2}}d^{4}C(\theta^{\prime})$
$\delta^{4}\left(C(\theta_{2})-C(\theta_{1}+2\pi)\right)\prod_{\theta^{\prime\prime}=\theta_{2}}^{\theta_{1}+2\pi}d^{4}C(\theta^{\prime\prime})$
$\implies\mathcal{D}C\delta^{4}\left(C(\theta_{2})-C(\theta_{1})\right)=\mathcal{D}C_{12}\mathcal{D}C_{21}$
$\implies\int\mathcal{D}C\hat{L}_{\nu}(\theta_{1})W[C]\exp\left(\mathrm{i}\oint P\cdot C^{\prime}\right)=$
$\int\mathcal{D}C_{12}\int d\theta_{2}\stackrel{{\scriptstyle\longleftrightarrow}}{{C^{\prime}_{\nu}(\theta_{2})}}\exp\left(\mathrm{i}\int_{\theta_{1}}^{\theta_{2}}P\cdot C^{\prime}\right)W[C_{12}]$
$\int\mathcal{D}C_{21}\exp\left(\mathrm{i}\int_{\theta_{2}}^{\theta_{1}+2\pi}P\cdot C^{\prime}\right)W[C_{21}]$
$\implies L_{\nu}W[P_{11}]=\int d\theta_{2}\stackrel{{\scriptstyle\longleftrightarrow}}{{\frac{\delta}{\delta P_{\nu}(\theta_{2})}}}W[P_{12}]W[P_{21}]$ (10.11)
Here
$\stackrel{{\scriptstyle\longleftrightarrow}}{{f(x)}}=\frac{f(x+0)+f(x-0)}{2}$
After conformal transformation mapping the unit disk to the upper semiplane, we map the circle origin $\theta=\pm\pi$ by $x=\tan(\theta/2)$ to $x=\pm\infty$.
For our final local solution, however, we shall use a different gauge condition for diffeomorphisms, which is more adequate for string theory. The gauge ambiguity of the parametrization of the linear measure (10.3) leads to the overcounting of every loop by an infinite volume of the gauge orbit. This volume must be factored out of the measure to avoid overcounting. This refactoring seems incompatible with the above factorization of the loop measure over the self-intersecting loops, but upon closer inspection, the paradox disappears. The measure must be factored by the total volume of its isometries, which, in the case of a simple loop, reduces to the volume of diffeomorphisms of a unit circle $\text{Vol}(\text{Diff}(S^{1}))$ ; however, in the case of a self-intersecting loop, there are two periodic parts; therefore, such a loop maps twice the unit circle to $\mathbb{R}^{4}$. The volume of these isometries is $\left(\text{Vol}(\text{Diff}(S^{1}))\right)^{2}$, which matches the product of linear measures on the right side of the factorization identity. So, the physical measure, with single counting of each loop, factorizes in the MM equation in momentum loop space.
10.4 Equivalence of loop diffusion operator to the third Magnus form
In momentum loop space, the loop diffusion operator $L_{\nu}$ from (10.9) is recast as a trilinear product of the momentum field $P(\theta)$:
$T^{\alpha\beta\gamma}{}_{\nu}\mathrm{i}^{3}P_{\alpha}(\theta_{1})P_{\beta}(\theta_{2})P_{\gamma}(\theta_{3});$ (10.12)
$\theta_{1}=2\pi-\epsilon,\theta_{2}=0,\theta_{3}=+\epsilon;\qquad\epsilon\to+0;$ (10.13)
By utilizing the integral representation
$P(\theta)=\oint d\tau P^{\prime}(\tau)\Theta(\theta-\tau)$ (10.14)
where $\Theta(\theta-\tau)$ denotes the periodic theta function on the circle, the trilinear product is transformed into a general triple integral of the momentum derivative $P^{\prime}$. The rigorous mathematical foundation for expanding loop functionals in such iterated integrals is provided by Chen’s theorem [35]. Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a strictly complete functional basis. This guaranties that our formal expansion in terms of Magnus forms is exhaustive.
Through direct tensor contraction, the kinematic tensor $T^{\alpha\beta\gamma}{}_{\nu}$ implements the exact Dynkin projector [36] onto the nested double commutator $\hat{D}_{\alpha},[\hat{D}_{\alpha},\hat{D}_{\nu}]]$. Here, $\hat{D}=\partial+A$ represents the covariant derivative operator within the gauge field representation of the Wilson loop. This mapping identifies the associative word structure of the integrated momentum field with its unique Lie algebra representation.
Consequently, by Ree’s Theorem [37] for Free Lie Algebras [38], this operator identically annihilates the symmetric shuffle ideal associated with the triple integral. As the point-split coordinates $\theta_{i}$ collapse to the origin, the periodic step-functions define a strictly cyclically ordered simplex. This unique parametric-invariant Lie projection ensures that the functional converges exactly to the third-order Magnus form $\Omega^{(3)}$. Direct evaluation of this triple integral, mediated by the $T$ projector, confirms this algebraic correspondence and yields the canonical normalization coefficient $1/6$.
10.5 Algebraic recurrence for the Momentum Loop Equation
The Momentum Loop Equation (MLE) exhibits a profound simplification when the momentum functional is represented as a Magnus-ordered exponential:
[…]
where $\hat{X}_{\mu}$ represents the operator position of the string endpoint in spacetime. Expanding this trace in Magnus forms $\Omega^{n}$, we obtain universal tensor coefficients $\mathcal{W}_{\alpha_{1}\dots\alpha_{n}}^{(n)}=\mathrm{tr}\{\hat{X}_{\alpha_{1}}\dots\hat{X}_{\alpha_{n}}\}$. The functional derivative acts algebraically on this structure, bringing down exact commutator insertions:
$\frac{\delta}{\delta P_{\mu}(\theta)}\mathrm{tr}\left\{\hat{\mathcal{T}}\exp\left(\int d\theta^{\prime}\hat{X}_{\alpha}P_{\alpha}^{\prime}\right)\right\}=$
$P_{\nu}^{\prime}(\theta)\mathrm{tr}\left\{[\hat{X}_{\mu},\hat{X}_{\nu}]\hat{\mathcal{T}}\exp\left(\int_{\theta}^{\theta+2\pi}\hat{X}_{\alpha}P_{\alpha}^{\prime}\right)\right\}$ (10.16)
By applying this exact derivative formula, the Makeenko–Migdal loop equation sheds its coordinate-space singularities and collapses into a finite, recursive combinatorial algebra. On the left-hand side, the kinematic operator $Q_{\nu}$ acts via a shuffle product between the 3-index $T^{\alpha\beta\gamma}_{\nu}$ tensor and the lower-order $\mathcal{W}^{(n-3)}$ coefficients. On the right-hand side, the integration over $\theta$ seamlessly sews the split traces back together, yielding an algebraic sum over commutator insertions without requiring cyclic sums or residual functional calculus. These relations are depicted in Fig.13
Figure 13: The recurrent equation for the coefficient tensor parameters $\mathcal{W}^{n}$ in the Magnus expansion (10.4). The external circle represents the unit circle where the angular parameter belong. The inner round blobs with wavy legs landing on a unit circle represent the $\mathcal{W}^{n}$ tensors, the rhombus with four wavy legs represents the 4-tensor $T$ in (10.8). Two arrows on the right side correspond to functional derivatives $\frac{\delta}{\delta P_{\nu}}$ bringing commutators by momentum area derivatives (10.5). These equations relate the higher $\mathcal{W}^{n}$ tensors to the lower ones, which allows the solution depending on some gauge parameters, undetermined by the recurrent relations.
Because the physical momentum loop is closed ($\oint P^{\prime}d\theta=0$), the first-order form identically vanishes: $\Omega^{1}_{\alpha}=0$. This topological constraint generates a formal “shuffle ideal” that relates higher-order Magnus forms. We must therefore equate the left and right sides of the MLE modulo this ideal. The null space of this ideal natively gives rise to symmetric gauge ambiguities that dynamically decouple from physical observables.
10.6 Exact solution up to $\mathcal{O}(P^{\prime 7})$ and non-analyticity of the Taylor expansion
To construct the exact solution to the Momentum Loop Equation (MLE) in the perturbative regime of small loop momenta, we employ the functional Taylor–Magnus expansion of $\mathcal{W}[P]$. Because the MLE evaluates the functional derivative on closed loops, the geometric ansatz must natively reflect this exact topological closure.
When the functional area derivative acts, it splits the integration domain into two subloops on the right-hand side of the equation. To analytically enforce the closure of these subloops within the continuous path integral, we introduce a modified momentum derivative containing a boundary Dirac delta function:
$P^{\prime}_{\text{closed}}(\theta^{\prime})=P^{\prime}(\theta^{\prime})+\Delta P\,\delta(\theta^{\prime}-\theta_{0})$ (10.17)
where $\Delta P=\int P^{\prime}d\theta$ is the boundary gap of the subloop, and $\theta_{0}$ is the position of the gap ($\theta_{0}=\theta$ and $2\pi$ for the respective right-hand side subloops).
When substituted into the iterated integrals of the Magnus expansion, the $\delta$-function explicitly inserts the macroscopic gap vector $\Delta P_{\mu}\equiv\Omega^{(1)}_{\mu}$ precisely at the boundaries of the integration domain. Because $\Delta P$ corresponds to the first-order Magnus form, this topological closure modifies the $n$-th order form by subtracting the product $\Omega^{(1)}\Omega^{(n-1)}$. By the integration-by-parts identity of the Magnus expansion (the shuffle relations), this product translates exactly to the sum of all interleaved operator positions. The loop-closing transformation thus yields the modified Magnus form:
$\widetilde{\Omega}^{(n)}=\Omega^{(n)}-\sum_{k=1}^{n-1}\Omega^{(n)}(\text{interleaved})$ (10.18)
This subtraction mathematically enforces the closed-loop Shuffle Ideal, identically restoring the cyclic symmetry of the invariant tensors $\mathcal{W}^{(n)}$. Operating in the pure closed-loop vacuum, all odd-parity traces identically vanish ($\mathcal{W}^{(2k-1)}=0$). The functional $\mathcal{W}[P]$ is therefore expanded in a basis of cyclic symmetric, invariant tensor polynomials constructed entirely from the Kronecker metric $\delta_{\mu\nu}$.
Substituting this closed-loop basis into the MLE, we dynamically equate the left-hand side kinematic shuffle products against the right-hand side discrete commutator series. Starting from the perturbative vacuum $\mathcal{W}^{(0)}=1$, the algebraic system successfully yields finite, exact solutions for the lowest-order tensors. The fundamental string tension enters at $\mathcal{W}^{(2)}$, and the kinematic stress generated by the loop derivative exactly constrains the cyclic symmetric pairings of $\mathcal{W}^{(4)}$ and $\mathcal{W}^{(6)}$. (The explicit analytic formulas for these invariant tensors are provided in C for small $n$. Crucially, C also contains the exact finite-dimensional linear algebra diagnostics (Table 1) and the explicit Fredholm inconsistency proof demonstrating the absolute lack of solutions at $\mathcal{W}^{(8)}$. The general algorithmic implementation is provided in [39].)
10.7 The Reasons for the Failure of the Magnus Expansion for $W[P]$
The general reason for the breakdown of the expansion at this order is that the linear space of $\mathcal{W}^{(8)}$ invariant tensors—which is strictly constrained by the Shuffle Ideal and cyclic symmetry—possesses a significantly smaller matrix rank than the highly specific multilinear stress generated by the crossing Kronecker deltas of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$.
This is not a heuristic observation, but a rigorously quantified algebraic contradiction. As detailed in C (and summarized in Table 1), extracting the exact linear system at the 8th order yields 379 geometric constraints but only 212 available parameters (which exhaustively includes the full $\mathcal{W}^{(8)}$ basis, all residual lower-order free parameters, and all gauge/ideal freedoms). The resulting system explicitly violates the Fredholm alternative condition, yielding a strictly positive rank deficiency ($\Delta\text{rank}=1$) and a non-zero projection of the inhomogeneous stress onto the left-null space of the kinematic matrix.
Consequently, it is mathematically impossible for any choice of linear coefficients, or the remaining lower-order terms, to absorb and cancel this rigid cross-term stress. This yields an irreducible algebraic contradiction in the tensor matching.
The physical meaning of this failing Taylor–Magnus expansion is of paramount importance. It demonstrates that the MLE is finite, free of UV divergences, and expandable in a functional Taylor series up to $\mathcal{O}(P^{\prime 6})$. However, the geometric contradiction at the 8th order rigorously proves that the exact solution $W[P]$ is not an analytic functional of the bounding loop $P(\theta)$. It cannot be indefinitely expanded in a continuous Taylor–Magnus series.
Remark 10.1 (The Vector vs. Scalar Dimensionality Mismatch).
The fundamental root of this non-analyticity is that the loop equation is inherently a vector equation,
$\hat{L}_{\nu}W=\int\frac{\delta}{\delta P_{\nu}}(W\times W),$ (10.19)
governing a scalar functional $W[P]$. The number of independent tensor structures available for a scalar functional grows strictly slower than those required for a free-index vector equation (growing approximately four times slower in $d=4$). This is why, at some critical level of the Magnus expansion, there will inevitably be more vector equations than available scalar parameters.
This is a foundational feature of the loop equation: it descends directly from the vector Yang–Mills equations of motion, which dictate that a local vector loop equation (evaluated at a fixed point on the contour) must be satisfied. Resolving this overdetermined system demands a deep mathematical “conspiracy” among the internal degrees of freedom within $W[P]$. The generic Taylor–Magnus expansion simply lacks the internal geometric structure to orchestrate this conspiracy.
(Interestingly, this exact same vector-versus-scalar mismatch is fatal to all naive bosonic string theories of QCD: the loop equations are strictly vector equations, but the Nambu–Goto string Laplace operator is a scalar, meaning a scalar area-derivative constraint fundamentally cannot capture the full vector loop equation without an anomaly.)
However, our theory natively provides this required conspiracy! In the previous sections, we derived exact vector equations for the Elfin determinant on the Hodge-dual minimal surface, and these vector equations are satisfied across all vector components. This mathematical miracle is made possible by Hodge duality and the planar factorization of the Majorana determinant on this Hodge dual surface.
The profound physical reasons for this non-analyticity will become fully apparent in later sections, when we construct the continuum solution and reduce it to the exact twistor string path integral. In that formulation, the formal Taylor expansion in $P(\theta)$ corresponds to the perturbative expansion of the exact exponential phase factor:
[…]
For such a functional Taylor series to converge, the boundary twistors $\lambda(z)$ and $\mu(z)$ would need to be strictly bounded. The algebraic contradiction at the 8th order rigorously proves this is not the case; the path integral is not convergent in a perturbative Taylor expansion.
As we will discuss in detail when deriving the twistor string path integral, the integration over the momentum loops is not dominated by smooth, small $P$ variations at all. When integrated alongside the Dirac target-space product $\mathrm{tr}\,\hat{\mathcal{T}}\exp\left(-\mathrm{i}\int d\theta\,\gamma_{\mu}P_{\mu}(\theta)+m_{0}\right)$, we encounter strictly ultra-local integrals. At every angle $\theta$, we integrate out the local value $P(\theta)$. Because there are no worldline kinetic derivatives $\partial_{\theta}P$ in the action, the path integral completely factorizes point-by-point. As a result, the dominant trajectories are neither continuous nor dominated by small $P$ variations. We therefore conclude that the MLE is completely finite and anomaly-free, but its true continuum solution is inherently non-perturbative and cannot be Taylor-expanded beyond the 6th order.
10.8 The General Master Field $\hat{X}$ and the Word Space
To fully grasp the physical and mathematical implications of the geometric breakdown at $\mathcal{W}^{(8)}$, it is instructive to re-examine the loop expansion through the lens of the Master Field construction.
A natural mathematical objection arises when considering the breakdown of the Taylor-Magnus expansion: Does the representation of the momentum loop functional as the trace of a path-ordered exponential,
$W[P]=\mathrm{tr}\left\{\hat{\mathcal{T}}\exp\left(i\int dP_{\mu}\hat{X}_{\mu}\right)\right\}$ (10.21)
restrict the generality of the solution? If $\hat{X}_{\mu}$ were assumed to be finite $N\times N$ matrices, the Cayley-Hamilton theorem and finite polynomial trace identities would artificially constrain the higher-order Magnus tensors $\mathcal{W}^{(n)}$, destroying the generality of the expansion.
However, in the planar limit ($N_{c}\rightarrow\infty$), the Master Field operates in a strictly infinite-dimensional Hilbert space. As established by the Master Field Theorem proven by the author in 1994 [30, 32], any generic parametric-invariant loop functional expandable in iterated path integrals (guaranteed to be topologically complete by Chen’s theorem [35]) can be mapped exactly onto the non-commutative moment space of Free Probability and represented exactly as a Magnus-ordered exponential path integral over a non-commutative word space.
The true non-perturbative Fock space of the string endpoint is the discrete space of “words” generated by $d=4$ Cuntz algebra operators:
$a_{\mu}a_{\nu}^{\dagger}=\delta_{\mu\nu},\qquad a_{\mu}|0\rangle=0$ (10.22)
In this space, the position operator $\hat{X}_{\mu}$ is uniquely constructed as an infinite, cascading sum of creation operators (the Voiculescu expansion [40]):
$\hat{X}_{\mu}=a_{\mu}+\sum_{k=1}^{\infty}Q_{\mu,\mu_{1}\dots\mu_{k}}a_{\mu_{1}}^{\dagger}\dots a_{\mu_{k}}^{\dagger}$ (10.23)
By the Free Moment-Cumulant Theorem [41], this single, order-independent operator contains an infinite tower of algebraically unconstrained parameters (the free planar cumulants $Q$). Because the free Cuntz algebra contains no finite trace identities, our Master Field Theorem guaranties that the trace of the path-ordered exponential is mathematically unrestricted. It exhaustively spans the continuous functional space of any closed loop without truncation.
These $Q$ tensors are related to the invariant $\mathcal{W}$ tensors by a free cumulant expansion, uniquely solvable term by term for $Q(\mathcal{W})$, assuming the $\mathcal{W}$ tensors are finite. Therefore, the breakdown at the eighth order is definitively not an artifact of a restrictive algebraic ansatz. Rather, this lack of a solution for $\mathcal{W}^{(8)}$ rigorously implies that the assumption of an analytic functional $W[P]$ is violated beyond the sixth order of the Magnus expansion. The physical resolution is that the true solution must be a singular functional possessing a non-local dependence on the momentum loop $P$.
The theory of minimal surfaces provides a precise geometric precedent for this behavior: the Douglas functional for the area of a minimal surface bounded by a loop $C$ [12]. As established by Douglas, the area becomes a singular functional of the loop only after it is minimized over all possible boundary parametrizations.
Before this minimization, the area is simply a quadratic Dirichlet-Hilbert functional 3.28 governed by the singular principal-value kernel $\mathcal{H}$. However, the strict minimization over parametrizations transforms it into a fundamentally singular functional because the optimal (isothermal) parametrization becomes implicitly and non-locally tied to the geometric shape of the loop. Any local change to the shape of the loop (for instance, tracking the normalized momentum $\tilde{P}^{\prime}=P^{\prime}/\int|P^{\prime}|d\theta$) forces a global readjustment of this parametrization. This continuous global geometric shift modifies the analytic structure (though not the position) of the pole at $\theta=\theta^{\prime}$ of the Hilbert kernel on the unit circle. Such shape-dependent, non-local singularities are fundamentally incompatible with a continuous, analytic Taylor-Magnus expansion in terms of iterated path integrals.
Our exact continuum solution—the fermion determinant evaluated on the rigid Hodge-dual minimal surface—belongs precisely to this class of singular geometric functionals. The required reparametrization invariance is dynamically enforced by the null-twistor (Virasoro) constraint, which acts as the exact algebraic mechanism implementing the Douglas minimization over parametrizations. The algebraic catastrophe at $\mathcal{W}^{(8)}$ is thus the mathematical signature of the string geometry asserting its non-local rigidity and justifying the transition from a 1D algebraic Master Field to the 4D continuous geometry of the rigid twistor string.
11 The twistor parametrization of the loop space measure
The construction of the momentum loop path integral (10.3) in planar QCD begins with a linear measure (10.3) over the space of closed loops. However, after gauge fixing the reparametrization symmetry ($\text{Diff}(S^{1})$) via the Virasoro constraint, the correct measure on the reduced phase space becomes nonlinear. In this section, we derive the Jacobian $J=\sqrt{\det Q}$ arising from the twistor parametrization of the metric in loop space. In addition to this determinant, there is a Faddeev-Popov Jacobian arising from the gauge condition that fixes the residual local rescaling invariance of the twistor parametrization.
11.1 Gauge Symmetry and the Need for Gauge Fixing
The linear integration measure $\mathcal{D}C$ (10.3) is redundant due to the reparametrization invariance of the Wilson loop $W[C]$. To define the path integral rigorously, we must fix this gauge freedom and factor the measure by the volume of diffeomorphisms of the unit circle.
For the computation of the momentum loop $W[P]$, we choose the conformal gauge, where the loop parametrization is identified with the boundary value of a holomorphic null curve $f_{\mu}(z)$ in the bulk. The tangent vector of this complex curve factorizes into spinors (referred to as twistors in this context):
$f^{\prime}_{\mu}(z)\sigma_{\mu}=\lambda(z)\otimes\mu(z)$ (11.1)
The physical loop $C_{\mu}(\theta)$ is recovered as the real part of the boundary value of $2f_{\mu}(e^{\mathrm{i}\theta})$. The Virasoro constraint $(f^{\prime})^{2}=0$ imposed by this ansatz does not restrict the shape of the loop in $\mathbb{R}^{4}$; rather, it fixes the parametrization to be isothermal (where $|\dot{C}|=|H[\dot{C}]|$).
The original measure $\mathcal{D}C$ is invariant under the infinite-dimensional diffeomorphism group $\text{Diff}(S^{1})$. Physically, this is a redundancy: different parametrizations describe the same geometric loop. To obtain a well-defined path integral, we must factor out this gauge volume:
$\mathcal{D}C=\mathcal{D}C_{\text{phys}}\,\text{Vol}(\text{Diff}(S^{1}))$ (11.2)
where $\mathcal{D}C_{\text{phys}}$ is the measure on the space of unparametrized loops. Imposing the Virasoro constraint is equivalent to choosing a specific gauge slice. In the twistor language, the physical degrees of freedom are encoded in the fields $(\lambda(z),\mu(z))$, modulo residual gauge symmetries.
Remark 11.1.
The scalar products between tilde spinors and spinors are defined as
$\tilde{\lambda}\lambda=\tilde{\lambda}_{1}\lambda_{1}+\tilde{\lambda}_{2}\lambda_{2};$ (11.3)
$\tilde{\mu}\mu=\tilde{\mu}_{1}\mu_{1}+\tilde{\mu}_{2}\mu_{2};$ (11.4)
These products are $O(4)$ -invariant, unlike products
$v_{\alpha}=\lambda\sigma_{\alpha}\mu=\sum_{a,\dot{b}}\lambda_{a}\sigma_{\alpha}^{a\dot{b}}\mu_{\dot{b}}$ (11.5)
which are normal matrix-spinor products and transform as four- vectors. By design, these are null-vectors
$(v_{\alpha})^{2}=0$ (11.6)
11.2 The residual $U(1)$ gauge symmetry and normalized twistors
A residual symmetry remains in the general twistor parametrization of the solution to the Virasoro constraint. The parametrization (11.1) is invariant under the gauge transformation with a complex function $u(z)$:
$\delta\lambda(z)=u(z)\lambda(z);\qquad\delta\mu(z)=-u(z)\mu(z)$ (11.7)
This transformation leaves the product $\lambda\otimes\mu$ (and thus the velocity $v$) invariant. To fix this redundancy, we impose the normalization constraint at the boundary:
$\left(\bar{\lambda}\lambda-\bar{\mu}\mu\right)_{|z|=1}=0$ (11.8)
This constraint fixes the real part of the local rescaling (where $u(z)$ is real). Considering a variation $\delta r(z)$ corresponding to this real scaling:
$\delta\lambda(z)=\delta r(z)\lambda(z);\qquad\delta\mu(z)=-\delta r(z)\mu(z);$ (11.9)
$\delta\left(\bar{\lambda}\lambda-\bar{\mu}\mu\right)=2\delta r(z)\left(\bar{\lambda}\lambda+\bar{\mu}\mu\right)$ (11.10)
This gauge constraint requires a Faddeev-Popov Jacobian. Since the transformation is local, the Jacobian is diagonal in the position basis, leading to an extra local factor in the measure:
[…]
$=\prod_{|z|=1}d\lambda d\mu\,\delta\left(\bar{\lambda}\lambda-\bar{\mu}\mu\right)\left(\bar{\lambda}\lambda+\bar{\mu}\mu\right)$ (11.11)
Remark 11.2.
[…]
11.3 Twistor-induced metric in loop space
We treat the linear measure in loop space (10.3) as a measure in the space of velocities $v_{\mu}=C^{\prime}_{\mu}$, subject to the periodicity constraint:
$\mathcal{D}C=\delta^{4}\left(\oint d\theta v(\theta)\right)\prod_{\theta=0}^{2\pi}d^{4}v(\theta)$ (11.12)
The boundary condition provides a local relation between the loop velocity and twistor boundary values
[…]
This projection induces a metric on the tangent space of the twistors. The norm of a variation in loop space is:
$\lVert\delta v\rVert^{2}=\oint d\theta(\delta v_{\alpha})^{2}\propto$
$\oint d\theta\left|z\delta\lambda\sigma_{\alpha}\mu+z\lambda\sigma_{\alpha}\delta\mu-\text{c.c.}\right|^{2}$ (11.14)
Expanding the square and using Fierz identities, this norm can be written in a quadratic form acting on the spinor variations $\delta\Lambda=(\delta\lambda,\delta\mu)^{T}$:
$\lVert\delta v\rVert^{2}\propto\oint d\theta\,\delta\Lambda^{\dagger}\cdot\hat{Q}\cdot\delta\Lambda$ (11.15)
The kernel is the $4\times 4$ block matrix (in spinor indices):
$\hat{Q}=\begin{pmatrix}(\bar{\mu}\mu)\mathds{1}_{2}&\,\lambda\bar{\mu}\\ \mu\bar{\lambda}&\,(\bar{\lambda}\lambda)\mathds{1}_{2}\\ \end{pmatrix}$ (11.16)
where $\lambda\bar{\mu}$ denotes the dyadic product (rank-1 matrix). To determine the measure, we analyze the eigenvalues of $\hat{Q}$. Let $\Lambda=\bar{\lambda}\lambda=\bar{\mu}\mu$ on the constraint surface. The eigenvalues are:
  1. 1.
    $\omega_{1}=0$: The zero mode, corresponding to the gauge symmetry $\delta\lambda=\lambda,\delta\mu=-\mu$.
  2. 2.
    $\omega_{2}=2\Lambda$: The dilatation mode $\delta\lambda=\lambda,\delta\mu=\mu$.
  3. 3.
    $\omega_{3,4}=\Lambda$: Rotation modes orthogonal to the spinors.
The determinant of $\hat{Q}$ vanishes due to the zero mode, which reflects the redundancy fixed by the constraint (11.8). Following the Faddeev-Popov procedure, the correct Jacobian is the square root of the determinant restricted to the physical subspace (orthogonal to the gauge orbit). This is given by the product of the non-zero eigenvalues:
[…]
Evaluating the determinant on the physical subspace yields the factor:
[…]
11.4 Normalization of the Zero Mode and the Full Measure
[…]
The zero mode corresponds to the generator of the real rescaling symmetry. The variation along this gauge orbit is given by:
$\delta\lambda=\delta r\lambda,\qquad\delta\mu=-\delta r\mu$ (11.19)
where $\delta r$ is the infinitesimal parameter of the transformation. The norm of this tangent vector in the flat spinor space is:
$\lVert\delta_{\text{gauge}}\rVert^{2}$ $=|\delta\lambda|^{2}+|\delta\mu|^{2}$
$=(\delta r)^{2}\bar{\lambda}\lambda+(\delta r)^{2}\bar{\mu}\mu$ (11.20)
Imposing the constraint $\bar{\lambda}\lambda=\bar{\mu}\mu=\Lambda$, this simplifies to:
$\lVert\delta_{\text{gauge}}\rVert^{2}=2\Lambda(\delta r)^{2}$ (11.21)
Thus, the metric factor (or "length") associated with the gauge direction is:
$\sqrt{\det G_{\parallel}}=\sqrt{2\Lambda}$ (11.22)
The total measure on the spinor space is the product of the measure on the physical subspace and the measure along the gauge orbit. Combining the physical Jacobian (11.18) with the gauge normalization (11.22), we recover the full scaling factor:
[…]
$=\left(\sqrt{2}\Lambda^{3/2}\right)\times\left(\sqrt{2\Lambda}\right)=2\Lambda^{2}$ (11.23)
This confirms that the effective measure element on the twistor space, prior to the Faddeev-Popov division of the gauge group volume, scales as $2\Lambda^{2}$. This factor will later cancel with another local scale factor coming from the momentum loop measure.
11.5 Summary and Physical Interpretation
To summarize:
  • 1.
    The original linear measure $\mathcal{D}C$ is overparameterized due to reparameterization invariance.
  • 2.
    Gauge fixing via the Virasoro constraint and twistor parametrization reduces the integration to the physical moduli space.
  • 3.
    The physically correct path integral measure is defined by the Faddeev-Popov Jacobian for the null twistor constraint (11.8). This ensures a unique counting of twistor states.
  • 4.
    Additionally, the change of variables from velocity $C^{\prime}(x)$ to twistors introduces the Jacobian (11.4), representing the volume of the physical tangent space.
  • 5.
    While this measure covers the moduli space uniformly, there remains a local $U(1)$ gauge invariance: $(\lambda,\mu)\Rightarrow(e^{\mathrm{i}\phi}\lambda,e^{-\mathrm{i}\phi}\mu)$. The corresponding volume $|U(1)|=2\pi$ will be simply factored out of the measure for the normalized spinors $\xi,\eta,\bar{\xi}\xi=\bar{\eta}\eta=1$. These spinors vary on $S^{3}\times S^{3}$, therefore, the space becomes $S^{3}\times S^{3}/S^{1}$ after this final factorization.
11.6 Local limit of $W[P]$ using effective action of Elfin theory
With the non-singular functional $W[P]$, we can utilize the local limit of our continuum solution. Parametrizing the loop $C$ by twistors $\lambda,\mu$, we compute the Fourier integral for $W[\hat{P}$] (with $\hat{P}=P_{\alpha}\sigma^{\dagger}_{\alpha}$):
$W[P]=\int\mathcal{D}\lambda\mathcal{D}\mu\,2\Lambda^{3}\,\delta\left(\bar{\lambda}\lambda-\bar{\mu}\mu\right)\exp\left(-\int d^{2}zL(z,\bar{z})\right)$
[…]
where the combined measure factor $2\Lambda^{2}\cdot\Lambda\sim\Lambda^{3}$ (depending on normalization conventions) is absorbed into the effective Lagrangian:
$L(z,\bar{z})=\frac{\sigma}{2}\Lambda^{2}+\frac{1}{12\pi}\left|\partial_{z}\left(\log\Lambda\right)\right|^{2}$ (11.25)
The extra integration over the "zero mode" $k_{\mu}$ reproduces the delta function $\delta(\oint vd\theta)$ for the periodicity of the loop ($C(\infty)=C(-\infty)$ in the upper plane map). This results in a nonlinear theory involving spinor fields $\lambda(z),\mu(z)$ holomorphic inside the unit circle.
12 From null twistor to Liouville coupled to $S^{1}\mapsto(S^{3}\otimes S^{3})$ sigma model
12.1 Polar Coordinate Parametrization of the Null Twistor Measure
Consider the twistor variables $(\lambda_{\alpha},\mu^{\dot{\alpha}})$ on a unit circle $|z|=1$, subject to the null constraint:
$\bar{\lambda}^{\alpha}\lambda_{\alpha}=\bar{\mu}_{\dot{\alpha}}\mu^{\dot{\alpha}}$ (12.1)
We parametrize the spinors in polar coordinates as:
$\lambda_{\alpha}=u\,\xi_{\alpha},$
$\mu^{\dot{\alpha}}=v\,\eta^{\dot{\alpha}}$ (12.2)
where $u,v\in\mathbb{R}_{+}$ are positive real scales, and $\xi,\eta$ are normalized spinors:
$\bar{\xi}^{\alpha}\xi_{\alpha}=\bar{\eta}_{\dot{\alpha}}\eta^{\dot{\alpha}}=1$ (12.3)
The constraint implies $u=v$. The measure on the null twistor space in these coordinates transforms as:
$\mathcal{D}\lambda\,\mathcal{D}\mu\,\delta(\bar{\lambda}\lambda-\bar{\mu}\mu)\propto u^{5}\,du\,d\Omega_{\xi}\,d\Omega_{\eta}$ (12.4)
where $d\Omega_{\xi}$ and $d\Omega_{\eta}$ are the Haar measures on $S^{3}$.
There is one subtle issue left to settle. The local gauge $U(1)$ transformation (11.7) leaves the physical variables invariant. This $U(1)$ subgroup acts diagonally on the spinors and should be factored out, defining the angular manifold as the coset space:
$\xi,\eta\in\frac{S^{3}\times S^{3}}{U(1)};$ (12.5)
$d\Omega_{\xi\eta}=\frac{d\Omega_{\xi}d\Omega_{\eta}}{2\pi}$ (12.6)
12.2 Reduction to $u$, $\xi$, $\eta$ Variables
The path integral over null twistors reduces to:
$\int\mathcal{D}\lambda\,\mathcal{D}\mu\,\delta(\bar{\lambda}\lambda-\bar{\mu}\mu)\,F(\lambda,\mu)$
$=\int_{0}^{\infty}u^{5}du\int_{(S^{3}\times S^{3})/U(1)}d\Omega_{\xi\eta}\,F(u\xi,u\eta)$ (12.7)
This makes explicit the separation between the radial scale $u$ and the angular geometry.
12.3 Transformation to Liouville Field
We define the Liouville field $\rho$ by the logarithmic map:
$\rho=2\log u\hskip 18.49988pt\Rightarrow\hskip 18.49988ptu=e^{\rho/2}$ (12.8)
The radial measure transforms as:
$u^{5}du=e^{5\rho/2}(e^{\rho/2}d\rho)=e^{3\rho}d\rho$ (12.9)
Consequently, the path integral measure becomes:
$\int\mathcal{D}\lambda\,\mathcal{D}\mu\,\dots=\int_{-\infty}^{\infty}d\rho\,e^{3\rho}\int d\Omega_{\xi\eta}\,\dots$ (12.10)
12.4 Phase space path integral for quark loop amplitudes
The physical amplitudes in (10.2) require one final functional integration: the path integral over the momentum field $P_{\mu}(\theta)$ (defined on the boundary).
The parametric invariance $\theta\to\phi\theta)$ applies to the complete phase space integral. This invariance is broken only by the trace of the ordered path exponential of the Dirac operator. To restore it fully in the effective theory, we introduce an einbein $e(\theta)>0$ multiplying the Dirac operator:
[…]
While the conventional gauge choice is $e(\theta)=\text{const}$, we have already fixed the reparametrization gauge via the Virasoro constraint on the holomorphic vector $f_{\mu}(z)$. Since the reparametrization acts simultaneously on $P$ and $C$, and we have factored out the volume of $\text{Diff}(S^{1})$, there are no more gauge fixing to be done. We are left with integration over the arbitrary positive local einbein field $e(\theta)$.
This integration yields the inverse matrix structure.
$\int_{0}^{\infty}de\exp\left(-ed\theta(\mathrm{i}\gamma_{\mu}P_{\mu}+m_{q})\right)\propto\frac{1}{(\mathrm{i}\gamma_{\mu}P_{\mu}(\theta)+m_{q})}$ (12.12)
Taking the limit of zero quark mass and focusing on the local factor dependent on $P$, we compute the integral over the local momentum variable $q_{\mu}(\theta)=P_{\mu}(\theta)d\theta$:
$I(\tau)=\int d^{4}q\,\frac{\exp\left(\mathrm{i}q_{\alpha}\tau_{\alpha}\right)}{\mathrm{i}\gamma\cdot q}$ (12.13)
where the source term is given by the twistor bilinear:
[…]
This Fourier integral is calculable. Up to normalization, we obtain:
$I(\tau)\propto\frac{\mathrm{i}\gamma_{\mu}\tau_{\mu}}{(\tau^{2})^{2}}$ (12.15)
Using the normalization of the spinors $\xi,\eta$, and the Fiertz identity, the square of the source vector simplifies:
[…]
$=\frac{-1}{4}\left(z\xi\sigma\eta-\bar{z}\overline{\xi\sigma\eta}\right)^{2}e^{2\rho}=\frac{1}{4}(4)e^{2\rho}=e^{2\rho}$ (12.16)
Substituting this back into the result:
[…]
Remarkably, the factor $e^{-3\rho}$ from the momentum integration exactly cancels the measure factor $e^{3\rho}$ derived from the twistor transformation, rendering the effective local volume element scale-invariant.
Remark 12.1 (Scale Invariance of the Phase Space Measure).
The exact cancelation of the Liouville factors, $e^{3\rho}\cdot e^{-3\rho}=1$, is a necessary consequence of the symplectic geometry of the loop phase space. The canonical path integral measure $\mathcal{D}P\mathcal{D}C$ is invariant under canonical transformations, including the scaling dilatation $C_{\mu}\to sC_{\mu}$ and $P_{\mu}\to s^{-1}P_{\mu}$. In our construction, the twistor transformation introduces a specific scaling of the coordinate velocity $v\sim e^{\rho}$. The Jacobian of this transformation ($e^{3\rho}$) represents the compression of the coordinate measure. However, the integration over the conjugate momentum $P_{\mu}$ acts as an inverse scaling operator (a Fourier transform), generating the counter-factor $e^{-3\rho}$. The resulting scale invariance of the effective volume element ensures that the theory remains conformal at the classical level, consistent with the dimensionless nature of the phase space volume element $dP\wedge dC\sim\hbar$.
12.5 The Holographic Liouville Theory
The effective action now reduces to a Liouville-type term for the field $\rho$:
$S_{\text{Liouville}}=\int d^{2}z\,\left(\frac{1}{3\pi}|\partial_{z}\rho|^{2}+\sigma e^{2\rho}\right);$ (12.18)
$\rho(z,\bar{z})=\frac{\log(\bar{\lambda}(\bar{z})\lambda(z))+\log(\bar{\mu}(\bar{z})\mu(z))}{2};$ (12.19)
and the amplitude is proportional to
$\mathcal{D}\Omega_{FP}[\lambda,\mu]\exp\left(-S_{\text{Liouville}}\right)$
[…]
$Q_{k}(\theta)=\left(\Theta(\theta-\alpha_{k})-\Theta(\theta-\alpha_{k+1})\right)$ (12.21)
The closure of the loop: $\oint d\theta v_{\mu}=0$ is satisfied identically in our parametrization
[…]
This contour integral over a unit circle of a holomorphic function vanishes by the Cauchy theorem (no singularities inside the circle).
It is important to distinguish this construction from the conventional Liouville field theory encountered in 2D quantum gravity. Here, the Liouville field $\rho(z,\bar{z})$ in the bulk is not an independent fluctuating quantum field; rather, it is rigidly determined by the boundary data via the analytic continuation of the underlying holomorphic twistors $\lambda(z),\mu(z)$. Thus, this is not the standard dynamical Liouville theory. Instead, it represents a hidden one-dimensional theory of twistors on the boundary, disguised as a 2D Liouville theory—a simpler, albeit less familiar, mathematical object.
In our framework, the path integral measure over loop space is rigorously constructed through a systematic procedure: we begin with a linear measure, impose the Virasoro constraint via twistor parametrization, and fix the residual local scaling symmetry. The resulting measure is manifestly real and positive. Crucially, this measure does not contain an exponential of the KKS symplectic form—a hallmark of conventional string quantization and Conformal Field Theory (CFT). We emphasize that our approach does not involve quantizing a fundamental string in the Nambu-Goto sense, but rather constitutes a faithful implementation of the solution to the Makeenko–Migdal loop equations in momentum loop space.
12.6 Block Matrix Structure and Diagonalization via Staggered Spinors
[…]
$\gamma_{\alpha}=\begin{pmatrix}0&\sigma_{\alpha}\\ \bar{\sigma}_{\alpha}&0\end{pmatrix}$ (12.23)
where $\sigma_{\alpha}=(1,\vec{\sigma})$ and $\bar{\sigma}_{\alpha}=(1,-\vec{\sigma})$ in Euclidean signature (or $\sigma_{\alpha}$ are the standard Pauli matrices). The velocity vector is given explicitly by the projection of the loop momentum onto the spinor dyadics:
$v_{\alpha}=iz(\xi\sigma_{\alpha}\eta)-i\bar{z}(\eta^{\dagger}\sigma_{\alpha}\xi)$ (12.24)
Substituting this into the link matrix $\mathcal{M}$ and using the Fierz completeness relation $(\psi^{\dagger}\sigma_{\alpha}\chi)\sigma^{\alpha}=2\chi\psi^{\dagger}$, the operator takes the specific block form:
$\mathcal{M}=i\begin{pmatrix}0&\mathcal{P}\\ \mathcal{P}&0\end{pmatrix}$ (12.25)
where $\mathcal{P}$ is the Hermitian operator on the 2-component spinor space, explicitly dependent on the loop coordinate $z=e^{i\theta}$:
$\mathcal{P}(\theta)=2\mathrm{i}\left[z\eta\xi^{\dagger}-\bar{z}\xi\eta^{\dagger}\right]_{z=e^{\mathrm{i}\theta}}$ (12.26)
This off-diagonal structure implies a “chirality hopping” mechanism: the operator maps a left-handed spinor at point $\theta$ to a right-handed spinor at $\theta+d\theta$, and vice versa. To compute the trace of the ordered product efficiently, we employ a staggered spinor transformation (analogous to the Kawamoto-Smit transformation in lattice gauge theory).
We discretize the loop into $2N$ points and define a transformed spinor basis $\Psi^{\prime}_{n}$ related to the original basis $\Psi_{n}$ by a site-dependent rotation $\Omega_{n}$:
$\Psi_{n}=\Omega_{n}\Psi^{\prime}_{n},\qquad\Omega_{n}=\begin{cases}\mathds{1}_{4}&\text{for even }n\\ \gamma_{0}&\text{for odd }n\end{cases}$ (12.27)
where $\gamma_{0}=\begin{pmatrix}0&\mathds{1}_{2}\\ \mathds{1}_{2}&0\end{pmatrix}$ is the chirality-flipping matrix. The effective link matrix between sites $n$ and $n+1$ transforms as $\tilde{\mathcal{M}}_{n}=\Omega_{n+1}^{\dagger}\mathcal{M}_{n}\Omega_{n}$.
Due to the bipartite nature of the 1D loop, there are two alternating cases:
  1. 1.
    Even to Odd ($n\to n+1$):
[…]
  2. 2.
    Odd to Even ($n\to n+1$):
[…]
The transformation completely diagonalizes the operator in the Dirac space, decoupling the 4-component spinor into two identical 2-component sectors.
The trace of the path-ordered product along the closed loop splits into two equal $2\times 2$ traces:
[…]
This reduction from $4\times 4$ to $2\times 2$ matrices significantly reduces the computational complexity for numerical evaluation. The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates. Therefore, the Complex Langevin method [42] required to compute this path integral.
13 The Linear Regge trajectories in the WKB limit
13.1 The Helicoid Saddle Point
In the semiclassical limit (large momentum or large tension $\sigma$), the path integral is dominated by the saddle point of the effective action. We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
Consider a momentum distribution rotating in the $(1,2)$ plane with frequency $\omega$:
$P_{1}(\tau)+\mathrm{i}P_{2}(\tau)=P_{0}e^{\mathrm{i}\omega\tau},\qquad P_{3}=P_{4}=0.$ (13.1)
This source term in the action (11.6) drives the spinor fields. We search for a solution that respects this rotational symmetry. The natural ansatz corresponds to the helicoid minimal surface described in Appendix B.4.
In the spinor formalism, with rescaling $w(z)=z^{-1/2}$ the Helicoid is generated by:
$\lambda(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}1\\ -z\end{pmatrix}$ (13.2)
Substituting this into the null vector definition $f^{\prime}_{\mu}\sigma_{\mu}=\lambda\otimes\mu$, we find:
$f^{\prime}_{1}+\mathrm{i}f^{\prime}_{2}\sim\frac{1}{z},\qquad f^{\prime}_{3}=f^{\prime}_{4}=0.$ (13.3)
Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid. The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
13.2 The Minkowski Helicoid
The Minkowski Helicoid is parameterized by time $\tau\in[0,T]$ and radius $r\in[-R,R]$, see [43]. The time $T$ tends to infinity in the end.
$X(\tau,r)=(\ r\cos\omega\tau,\ r\sin\omega\tau,0,\tau)$ (13.4)
The induced Minkowski metric is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
In Minkowski space we are not seeking to minimize the area in the geometric sense, as the area functional is not positive definite and the minimization problem is distinct. Instead, our criterion is algebraic: we require a solution to the Loop Equations for the Wilson loop. This demands the self-duality of the area derivative and the additivity of the area functional. We utilize the fact that the Euclidean Helicoid is a minimal surface for arbitrary angular velocity $\omega$. We obtain the Minkowski Helicoid via a double Wick rotation:
$x_{4}\to\mathrm{i}t,\qquad\omega_{E}\to-\mathrm{i}\omega.$ (13.5)
Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space. Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity). By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space. Thus, the Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
Figure 14: The helicoid spanned by rotating $\bar{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, was discovered by Meusnier in 1785.
Another important point is that the Helicoid boundary loop (double helix) is infinitely smooth, lacking any cusps or self-intersections. This means that no UV divergent renormalization factors could potentially affect our computation of the meson Regge trajectory.
13.3 Semiclassical Motivation: Twisted Boundary Conditions
Before solving the spectral equation, we justify the choice of the Helicoid geometry from the perspective of the semiclassical quark path integral.
Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$. In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor333the perimeter renormalization terms in that factor are included in the physical quark mass. This is relevant to our WKB approximation in the Wilson loop at large loops:
$\mathcal{Z}\sim\int\mathcal{D}C\,\exp\left(-\mathrm{i}m_{q}\oint_{C}ds-\mathrm{i}\sigma S_{\mathrm{Mink}}[\Sigma_{C}]\right)$ (13.6)
To extract the spectrum of states with angular momentum $J$, we impose twisted boundary conditions. We require the quark positions at time $t=T$ to be a rotation of their positions at $t=0$ by an angle $\Theta=\omega T$ in the rotation plane $xy$:
$(x+\mathrm{i}y)(T)=e^{\mathrm{i}\omega T}(x+\mathrm{i}y)(0)$ (13.7)
The angular momentum $J$ arises as the thermodynamic conjugate to the twist frequency $\omega$.
The ground state of this system is determined by the saddle point of the effective action. This requires the simultaneous minimization of the boundary length (quark term) and the surface area (gluon term).
  • 1.
    The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
  • 2.
    The minimal surface spanning two twisted helical boundaries is the helicoid.
Thus, the Helicoid is not merely an ansatz; it is the unique geometric solution mandated by the kinematics of angular momentum in the semiclassical limit. The balance between the surface tension $\sigma$ (pulling the quarks inward) and the centrifugal barrier (pushing the twisted trajectories outward) determines the physical radius of the meson.
13.4 Effective Lagrangean
We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks. The confining factor introduces an effective Minkowski area term. We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m\int ds$. The total Lagrangian is:
$L=L_{\text{helicoid}}+L_{\text{quarks}}$ (13.8)
The Helicoid contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is:
$L_{\text{helicoid}}$ $=-\sigma\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
$=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (13.9)
Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional. Adding the particle term $2m\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
$L(v,\omega)=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m_{q}\sqrt{1-v^{2}}$ (13.10)
It is worth noting that this mass term is proportional to the perimeter of the loop in Minkowski space. Therefore, the potential perturbative terms proportional to the loop perimeter are included in this physical quark mass $m_{q}$. We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m_{q}$, which , as the experiment tells us, is much smaller than the QCD mass scale, so that $m_{q}^{2}\ll\kappa$.
We proceed to compute the conserved quantities via the Legendre transform. The angular momentum $J$ is the conjugate momentum to the rotation frequency:
$J=\frac{\partial L}{\partial\omega}$ (13.11)
Taking into account that $v$ depends on $\omega$ via the fixed radial extent ($v=\omega R$), the differentiation yields:
$J=\frac{\sigma}{\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2mv^{2}}{\omega\sqrt{1-v^{2}}}$ (13.12)
The Energy $E$ is defined by the Hamiltonian $E=\omega J-L$. Substituting the expressions for $J$ and $L$ and simplifying the algebra, we obtain:
$E=\frac{2\sigma}{\omega}\arcsin v+\frac{2m_{q}}{\sqrt{1-v^{2}}}$ (13.13)
13.5 Balance of quark forces and boundary condition
The relationship between the endpoint velocity $v$ and the string length $R$ is not arbitrary; it is determined by the balance of forces at the string endpoint, which depends on the quark mass $m_{q}$. We derive this boundary condition by minimizing the action with respect to the geometric radius $R$ of the loop, keeping the angular velocity $\omega$ fixed.
Recall that the total Lagrangian $L$ is composed of the bulk integral for the helicoid and the boundary terms for the quarks:
$L=-2\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m_{q}\sqrt{1-\omega^{2}R^{2}}$ (13.14)
The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$. Differentiating the integral using the Leibniz rule yields the tension force, while differentiating the boundary term yields the centrifugal force:
$\frac{\partial L}{\partial R}=-2\sigma\sqrt{1-\omega^{2}R^{2}}-2m_{q}\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (13.15)
Substituting $v=\omega R$ and computing the derivative of the mass term:
$-2\sigma\sqrt{1-v^{2}}-2m_{q}\left(\frac{-\omega^{2}R}{\sqrt{1-\omega^{2}R^{2}}}\right)=0$ (13.16)
Simplifying the expression, we obtain the exact force balance equation for finite mass:
$\sigma\sqrt{1-v^{2}}=\frac{m_{q}v^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+m_{q}}$ (13.17)
This equation balances the confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
We now consider the limit $m_{q}\to 0$, and we find the boundary condition
$v^{2}\to 1$ (13.18)
Thus, as the quark mass tends to zero, the endpoint velocity $v$ must approach the speed of light ($v\to 1$).
These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$. To recover the Regge trajectories, we consider the limit of massless quarks ($m_{q}\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$). In this limit, the mass terms vanish, and the system is governed purely by the confining geometry:
$E\approx\frac{2\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{\omega^{2}}\frac{\pi}{2}$ (13.19)
Eliminating $\omega$ between these two relations:
$J=\frac{\sigma(\pi/2)}{(\pi\sigma/E)^{2}}\cdot\frac{\pi}{2}\implies J=\frac{E^{2}}{2\pi\sigma}$ (13.20)
We thus recover the linear Regge trajectory:
$E^{2}=2\pi\sigma J$ (13.21)
This result confirms that the linear confinement arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space, without invoking independent string vibrations or ad-hoc constants.
14 Twistor Holography and the Rigid Twistor String
14.1 Twistor Holography (Gauge Holography)
In our framework, twistor holography refers to the geometric duality in which the solution to the Makeenko–Migdal loop equation is realized as a rigid, Hodge-dual minimal surface embedded in flat $\mathbb{R}^{3}\times\mathbb{R}^{4}$ (or, more generally, a flat 12-dimensional space). This surface is uniquely determined by the boundary Wilson loop and does not fluctuate quantum mechanically. The holographic principle here projects the internal geometry of the 12D target space onto the observable boundary, with the physical Wilson loop acting as the gauge-invariant observable. Unlike AdS/CFT gravity holography, this construction is entirely geometric and does not involve a gravitational bulk or fluctuating worldsheet metrics.
14.2 Analytic Twistor String vs. Topological Twistor String
The use of spinor variables and null twistors in this framework is closely related to the Penrose transform, which encodes solutions to spacetime field equations in terms of cohomological data on twistor space. By imposing the specific gauge condition on boundary values of twistor map: $\bar{\lambda}\lambda=\bar{\mu}\mu$, we reduce the theory to a Liouville field interacting with a one-dimensional sigma model on $S^{3}\times S^{3}/S^{1}$ on the boundary. The bulk values of the Liouville field depend on the twistor map. This parametrization makes the structure transparent: the minimal surface in the bulk is entirely encoded by the holomorphic data (the twistors), while the physical observables are determined by the boundary curve.
A crucial distinction must be made regarding the nature of the resulting bulk action. In the context of Witten’s celebrated Twistor String theory (and related formulations such as the Berkovits model), the extension of boundary fields into the bulk typically generates a Wess-Zumino-Witten (WZW) term. The WZW functional is topological: it measures a winding number and is invariant under smooth deformations of the extension inside the disk.
In contrast, our bulk functional is not topological. While it is determined by the boundary data, its value depends explicitly on the unique analytic continuation of the holomorphic twistors into the disk. In this sense, our framework is closer in spirit to Penrose’s Non-Linear Graviton construction than to topological string theory. In the Non-Linear Graviton, the curved geometry of the bulk spacetime is encoded by the global complex structure of twistor space, which arises as the unique holomorphic extension compatible with boundary (patching) data. Similarly, our action represents the geometric energy (akin to a Kähler potential) of the specific holomorphic disk filling the boundary loop. This non-topological dependence is essential: it encodes the geometric rigidity of the minimal surface, leading to the emergence of the area law and confinement in the rigid twistor string.
The general issues of the discretization of our theory and its direct numerical simulation (DNS) are discussed in D, where we present formulas and algorithms for the analytic extension of holomorphic twistor fields from a discrete set of boundary values. We attempted the direct numerical simulation of this path integral with oscillating weight using the Complex Langevin equation (CLE) on E. The results are encouraging but not convincing; perhaps some additional numerical algorithms are needed to overcome the instabilities of the CLE and obtain reliable results.
15 Numerical Simulation of Confining String spectrum
The analytic twistor-string formulation of the previous sections reduces the meson spectroscopy problem to the analytic structure of the quark two-point amplitude $\Pi(E)$ (after analytic continuation to Minkowski energy $E$). Since the corresponding boundary-twistor path integral is not localized, the evaluation of the spectrum becomes primarily computational. In this section, we summarize the numerical strategy and present preliminary proof-of-concept results. Full implementation details, discretization formulas, drift terms, and numerical diagnostics are given in E.
Spectral observables and mass extraction
The physical quark two-point amplitude $\Pi(E)$ is a meromorphic function of the energy, characterized by physical poles at the bound-state meson masses. However, directly searching for vertical asymptotes (poles) in a stochastic Monte Carlo simulation is a numerical trap: statistical variance near a singularity diverges uncontrollably, producing massive fat-tail excursions and destroying the Complex Langevin evolution.
To completely bypass this fatal instability, we apply an analytic transformation that converts the dangerous meromorphic observable into a perfectly bounded, smooth entire function. Rather than evaluating $\Pi(E)$ directly, we consider the logarithmic derivative
$R(E)=\frac{\Pi^{\prime}(E)}{\Pi(E)},$ (15.1)
and reconstruct the spectrum from the zero-crossings of its exact inverse, $1/R(E)$.
This inversion maps vertical asymptotes into well-behaved roots. Near a physical mass pole, $\Pi(E)\sim(E-m_{n})^{-1}$, the inverse resolvent becomes $1/R(E)\sim m_{n}-E$. The singularity is mathematically flattened into a smooth zero-crossing with a strictly negative slope. Conversely, near an artifact zero of the vacuum polarization ($\Pi(E)\sim E-E_{0}$), one obtains $1/R(E)\sim E-E_{0}$, yielding a zero-crossing with a strictly positive slope.
Because $1/R(E)$ is an entire function with no singularities anywhere in the complex plane (as detailed in Appendix D.7), the Complex Langevin process can sample it safely. Therefore, the $\bar{q}q$ masses are unambiguously identified as the zero-crossings of $\text{Re}[1/R(E)]$ possessing a negative slope. Furthermore, because the true physical signal is bounded, we know a priori that any massive spikes in our simulation data are purely statistical fat-tail excursions, rigorously justifying aggressive outlier filtering (such as median pooling) without fear of clipping physical resonances.
Reweighted Complex Langevin strategy (summary)
The effective twistor-string functional integral has an oscillatory weight; at present, Complex Langevin evolution (CLE) provides a practical route to evaluating it. A key technical point is that pole-finding is numerically unstable in a stochastic computation. We convert the problem into a robust zero-finding problem by a reweighted formulation, summarized here and derived in Appendix D.
Let $S_{\rm eff}$ denote the effective action for the discretized boundary-twistor variables at fixed $E$, and define the energy-derivative observable
$O\;\equiv\;\frac{\partial S_{\rm eff}}{\partial E}\,.$ (15.2)
Then one has the exact identity
$\frac{1}{R(E)}\;=\;\frac{\Pi(E)}{\Pi^{\prime}(E)}\;=\;\frac{\int\exp\left(S_{\rm eff}\right)}{\int O\,\exp\left(S_{\rm eff}\right)}\,.$ (15.3)
Introducing the modified (reweighted) action
$\widetilde{S}\;=\;S_{\rm eff}+\log O\,$, (15.4)
this becomes
$\frac{1}{R(E)}\;=\;\left\langle\frac{1}{O}\right\rangle_{\widetilde{S}}\,$. (15.5)
[…]
Topological stabilization and non-compact excursions
There is a deep geometric reason for this stabilization, naturally understood through the lens of Picard-Lefschetz theory, introduced in QFT by Witten [44], and successfully used in recent QCD simulations [45, 46]. In the complexified phase space, the exact path integral is formally defined on a middle-dimensional integration contour composed of steepest descent manifolds (Lefschetz thimbles). While the deterministic drift of the Complex Langevin Equation naturally targets these thimbles, stochastic noise inevitably kicks the trajectory off these stable manifolds and into adjacent “Stokes wedges,” where the action diverges and the Markov chain suffers runaway excursions.
To counteract this, modern lattice simulations typically rely on computationally expensive algorithmic interventions. As detailed in the review [46], standard approaches involve explicitly deforming the integration contour via holomorphic gradient flow (which incurs a prohibitive $\mathcal{O}(V^{3})$ Jacobian cost), utilizing machine-learning-optimized manifolds, or applying ad-hoc dynamic stabilization and gauge cooling to actively force the system to remain near the thimbles.
Our logarithmic reweighting, however, achieves this stabilization natively for the observable. By shifting the effective action ($\tilde{S}=S_{eff}+\log\mathcal{O}$), we fundamentally alter the topology of the complexified phase space. The associated CLE drift force acquires essential meromorphic poles exactly at the zero-locus of the observable ($\mathcal{O}=0$):
$-\nabla\tilde{S}=-\nabla S_{eff}-\frac{\nabla\mathcal{O}}{\mathcal{O}}$ (15.6)
Crucially, in the Picard-Lefschetz framework, steepest descent flows cannot cross meromorphic poles. By explicitly injecting these poles into the drift dynamics, we puncture the complex manifold. The singular repulsive force acts as an absolute topological obstacle that dynamically corrals the stochastic Markov chain. It safely traps the trajectory within the basin of the physical thimbles and mathematically forbids it from escaping into the runaway regions of the complex plane associated with the zeros of $\mathcal{O}$. This organically enforces dynamic stabilization without the need for computationally devastating Jacobian evaluations.
The intermittent excursions that do occur in our simulation are not numerical instabilities, but rather the exact analytic consequence of the complexified noise exploring the vicinity of this meromorphic pole. As shown in the smooth histogram of the micro-states (see Figure 18), the probability density perfectly matches the analytically predicted $1/|u|^{3}$ power-law fat tails (derived in detail in Appendix D). This exact geometric matching confirms that the finite-sample Cauchy Principal Value is perfectly stable, and that the CLE is genuinely exploring the correct complexified geometry dictated by the topological barrier.
[…]
To prevent the system from exploring these unphysical infinite-volume regimes, we currently impose explicit numerical clipping (hard bounding walls) on the twistor variables. While this pragmatic regularization successfully bounds the metric, there is no a priori mathematical guaranty that the CLE trajectory will not systematically drift outward and continuously collide with these artificial walls, which would artificially truncate the sampled distribution.
Remark 15.1.
It should be noted that the use of explicit numerical clipping (hard bounding walls) to tame the non-compact Liouville directions technically violates the strict holomorphy required for the formal proof of correctness of the Complex Langevin Equation, as it introduces non-analytic boundaries into the Fokker–Planck evolution. A theoretically cleaner approach would involve introducing a globally holomorphic confining potential to the effective action that generates a smooth restoring force. For example, one could augment the action to produce a holomorphic drift force such as $K_{\text{wall}}\propto-\tanh(c(\rho-\rho_{\max}))$. Because the hyperbolic tangent is a meromorphic function of the complex variable $\rho$, this force smoothly pushes the trajectory back toward the physical region $\operatorname{Re}(\rho)<\rho_{\max}$ at large distances without breaking the analyticity of the Langevin evolution. Implementing such holomorphic barriers represents a natural analytical refinement for future high-precision simulations, safely taming the infinite-volume runaway channels while preserving the exact geometric justification of the CLE.
To explicitly defend our simulation against accidental discrete jumps over the topological barriers, we replaced the naive Euler integration with a second-order Stochastic Heun (predictor-corrector) SDE solver. Furthermore, we implemented an adaptive time-step algorithm designed to dynamically brake the CLE evolution upon approaching singular regions. Together, these algorithmic safeguards yield highly stable trajectories that remain predominantly confined to the physical integration cycles (Lefschetz thimbles), consistently producing finite expectation values with rigorously controlled absolute errors.
[…]
Discretization and parallel ensemble
We discretize the boundary circle by a spectral grid (angular nodes) and evaluate the bulk (holographic) Liouville contribution by Gauss–Legendre radial quadrature, leading to a finite-dimensional holomorphic system for the complexified boundary variables evolved by CLE. To reduce sensitivity to non-Gaussian (fat-tailed) fluctuations and to monitor reproducibility, we run an ensemble of independent replicas on the Typhon cluster. In the current production setting, we use 64 nodes (independent seeds) per energy value and pool the results across nodes using robust estimators (median central value and absolute-deviation-based error estimates). The precise implementation and diagnostics are documented in Appendix D.
16 Topological Decomposition of the Meson Spectrum via Information Clustering
16.1 Lefschetz Thimbles and Multiple Spectral Branches
In the planar limit, the exact solution of the Makeenko–Migdal loop equation involves evaluating a path integral over a highly complex configuration space. The effective action landscape of this theory is not dominated by a single, featureless ground state, but rather by a multitude of complex saddle points, or Lefschetz thimbles. Physically, these distinct topological sectors correspond to the multiple branches of the meson spectrum. In our numerical approach, we explore this landscape using stochastic Langevin dynamics. Because each independent simulation is initialized from a completely random configuration, the system naturally falls into the basin of attraction of one of these specific thimbles as it relaxes.
16.2 Topological Trapping and the Failure of Naive Pooling
Due to the presence of immense topological (or simply numerical) action barriers in the complexified space, a given simulation history generally remains confined to its respective thimble for the duration of the run. This dynamical trapping has profound implications for data analysis. If one were to follow standard numerical practice and naively pool all independent stochastic histories at a given energy $E$ to compute a global ensemble average, the procedure would physically mix distinct topological sectors. Consequently, this blind averaging does not yield a single, smooth resolvent curve for $1/\mathcal{R}(E)$; rather, it produces a chaotic, smeared superposition of states. The raw numerical data are actually a collection of distinct histories confined to various thimbles, each tracing out its own unbroken $1/\mathcal{R}(E)$ curve separated by topological barriers.
16.3 Phase-Information as a Topological Fingerprint
To extract the true physical spectrum, we must mathematically untangle these mixed histories and group them by the specific saddle point they occupy. As established in our analysis of the Complex Langevin dynamics in E, the Cartesian observables $\mathrm{Re}[1/\mathcal{O}]$ and $\mathrm{Im}[1/\mathcal{O}]$ suffer from severe infinite-variance Cauchy tails ($1/|u|^{3}$), which generate massive radial excursions. Consequently, attempting to compare the Cartesian coordinates forces one to arbitrarily trim and shift the central moments to prevent the variance from dominating the signal.
We can completely bypass this pathological radial variance by projecting the data onto the complex phase angle $\alpha=\arg(\mathcal{O})$. This maps the entire complexified phase space onto a finite, compact support $\pi,\pi]$, safely absorbing infinite radial fluctuations without discarding a single data point. The probability density function of this phase angle is intrinsically scale-invariant and serves as an exact, bounded topological fingerprint of the underlying Lefschetz thimble. To compare these fingerprints, we compute the symmetric Kullback–Leibler (KL) divergence (the Jensen–Shannon distance) between their finite-support phase distributions. This strict information-theoretic metric allows us to rigorously identify pairs of histories with identical quantum geometries, entirely independent of their macroscopic radial drift.
16.4 Isolating True Mass Poles
This phase-matching provides a precise, unsupervised mechanism for identifying true meson masses. A physical mass corresponds to a strictly downward zero-crossing of the resolvent, where the sequence transitions from $\mathrm{Re}[1/\mathcal{R}(E_{i})]>0$ to $\mathrm{Re}[1/\mathcal{R}(E_{i+1})]<0$. By examining pairs of runs at neighboring energy points that bracket the zero axis and demanding that they possess the absolute minimum symmetric KL distance between their phase distributions, we effectively filter out random topological jumps and numerical noise. This strict requirement of phase-geometry invariance unequivocally identifies the true crossing of the $1/\mathcal{R}(E)$ curve for a particular topological branch.
16.5 Reconstructing the Spectrum via Minimax Clustering
Having isolated a pool of high-confidence zero-crossings, we reconstruct the global spectrum of the different branches. To achieve this, we cluster the crossing events based on a strict minimax distance criterion. The topological distance between any two distinct crossing events—say, a pair $(A,B)$ and a pair (C,D)(C,D)—is defined as the maximum of all four cross-comparisons:
$\Delta=\max\left(d_{JS}(A,C),d_{JS}(A,D),d_{JS}(B,C),d_{JS}(B,D)\right)$ (16.1)
where $d_{JS}$ is the symmetric Kullback–Leibler distance (Jensen–Shannon divergence) between the phase distributions. This mathematically merciless constraint guaranties that a pole is only clustered into a specific sequence if the quantum phase geometry, both before and after the crossing, is practically identical to the rest of the cluster.
16.6 First results of numerical solution of CLE
The CLE was coded into a Mathematica ® package, tested on the laptop and desktop computers, and then run on the Typhon IAS cluster for 48 hours on 64 nodes in parallel, with 24 kernels working in parallel on each node. The code and the results are published on Wolfram cloud, and results are stored on the authors’ Google Drive with open access (see references in the Data section in the end of this paper). Using these Mathematica ® files, one can verify and reproduce our results at a larger scale with higher precision.
Here is the summary of the results (all technical details are described in E).
This unsupervised information-clustering natively reconstructs the spectrum into distinct branches, mapping out different Regge trajectories. Crucially, these trajectories are strictly labeled by an internal quantum number inherent to the Twistor String path integral complex saddle points, serving as a replacement for the “daughter trajectories” of string models.
Remark 16.1.
Note that all the usual quantum numbers are identical for these two branches: this is a scalar $\bar{q}q$ channel, no spin, no flavor quantum numbers specified. In the future, more detailed simulations, spin, parity, and flavor $SU(3)$ representation will be specified, but this internal quantum number will exist on top of other quantum numbers, in the same way as internal excitations of the string (vibrational modes) lead to daughter trajectories after quantization.
Applying this algorithm to our unvarnished stochastic phase data successfully isolated two continuous principal Regge trajectories spanning $14$ and $15$ consecutive negative crossings. The negative crossing of zero between two consecutive points of energy on a grid was identified with accuracy $\Delta E=\pm 1/480\approx\pm 0.002$. Both branches exhibit non-linearity at low energies—reflecting the complex boundary dynamics and mass gaps of the true physical ground state—before transitioning into linear Regge trajectories at higher mass squared. By relying purely on the phase-information geometry, this topological extraction is designed to reveal the true spectrum of Planar QCD.
Figure 15 shows preliminary cluster-scale results for the extracted trajectory $n$ vs. $m_{n}^{2}$. These plots are intended as a proof of concept: the crossings of zero were selected by information clustering of the phase distributions we described above. The precision of higher excitations and the control of systematic effects (step size, resolution, and CLE correctness diagnostics) will be improved in larger-scale runs; the present computation establishes feasibility and provides a concrete computational pipeline for continuum planar-QCD spectroscopy in this framework.
Figure 15: Plot of $n$ vs $m_{n}^{2}$ (in units of the string tension $\sigma$) as a result of Complex Langevin Simulation of quark two point function in the Geometric QCD Theory. These two Regge trajectories were identified by information clustering of phase distributions of $1/O$ for $240$ different energy points between $2.$ and $3.$ on $64$ independent CLE simulations with different initial values. The optimal clustering identified a pair of Regge trajectories in this interval of $m^{2}$
17 The Catastrophe Theory and the Geometric Mass Spectrum
17.1 Geometric Quantization via Degenerate Lefschetz Thimbles
While the Complex Langevin Equation provides a robust numerical method to evaluate the twistor-string path integral, the holomorphic nature of the effective action reveals a profound mathematical pathway to compute the exact meson spectrum analytically, entirely bypassing stochastic simulation. This approach relies on the asymptotic evaluation of multiple complex integrals via Picard-Lefschetz theory and the singularity theory of differentiable maps.
Consider the original, un-reweighted path integral for the quark two-point amplitude $\Pi(E)$ as a function of the Minkowski energy parameter $E$:
$\Pi(E)=\int_{\Gamma}\mathcal{D}\Phi_{|z|=1}\,\exp\left(E\,\mathcal{O}(\Phi)-S_{eff}(\Phi)\right)$ (17.1)
where $\Phi_{|z|=1}=\left\{z_{2},\left\{\rho,\xi,\eta,\bar{\xi},\bar{\eta},\gamma,\bar{\gamma}\right\}_{k}\right\}$ encompasses the boundary twistor variables extended analytically into the complexified phase space $\Phi=(\lambda(z),\mu(z))$, and $\Gamma$ is the middle-dimensional integration cycle (the principal Lefschetz thimble) bounding the physical vacuum.
In the semiclassical (WKB) limit, this integral is dominated by the saddle points $\Phi_{\star}(E)$ satisfying the classical boundary equations of motion:
$\nabla S_{eff}(\Phi_{\star})-E\nabla\mathcal{O}(\Phi_{\star})=0$ (17.2)
(which, as shown in Section 13, yields the rigid Helicoid geometry in the special case of twisted boundary conditions with large angular momentum $J$). The Gaussian fluctuations around this saddle point yield an amplitude proportional to the inverse square root of the complex Hessian determinant:
$\Pi(E)\propto\frac{\exp\left(E\,\mathcal{O}(\Phi_{\star})-S_{eff}(\Phi_{\star})\right)}{\sqrt{\det\mathcal{H}(E)}}$ (17.3)
where the fluctuation operator is $\mathcal{H}(E)=\nabla^{2}S_{eff}(\Phi_{\star})-E\nabla^{2}\mathcal{O}(\Phi_{\star})$.
As the continuous energy parameter $E$ is varied, the saddle points drift through the complexified phase space. Physical bound states (mesons) correspond to macroscopic divergences in this amplitude. Mathematically, these occur at critical energies $E=E_{n}$ where the saddle point becomes strictly shallow or degenerate (a non-Morse critical point). At these bifurcation points, the classical stability matrix develops a null space, and the determinant of the Hessian exactly vanishes: $\det\mathcal{H}(E_{n})=0$. At this point, the WKB approximation becomes exact, as it dominates the divergent term in the path integral. According to the classification of singular integrals in Catastrophe Theory [47], the nature of the resulting divergence depends strictly on the corank of the singularity (the number of simultaneously vanishing eigenvalues). If only a single real eigenvalue vanishes (corank 1), the local action along the flat direction scales as $\epsilon x^{2}$ (with $\epsilon\propto E_{n}-E$). The integration over this 1D valley yields:
$\int dx\,e^{-\epsilon x^{2}}\propto\frac{1}{\sqrt{\epsilon}}\propto\frac{1}{\sqrt{E_{n}-E}}$ (17.4)
Physically, a square-root singularity corresponds to a branch cut, generating a multi-particle continuum threshold, not a bound state.
However, our twistor-string geometry orchestrates a remarkable mathematical conspiracy. Because the integration variables $\Phi$ are coupled complex holomorphic twistors, $\mathcal{H}(E)$ is a complex operator. The vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
This alters the local topology of the steepest descent path. The Gaussian integral over this 2D flat valley yields:
$\iint dxdy\exp\left(-\epsilon(x^{2}+y^{2})\right)\propto\frac{1}{\epsilon}\propto\frac{1}{E_{n}-E}$ (17.5)
[…]
Remark 17.1.
Corank-2 from a single complex zero mode, and the pinch condition.
[…]
so that the vanishing of one complex eigenvalue generically flattens two real directions on the thimble (corank =2=2) and produces the pole-like local scaling (17.5). In this sense, a “single complex zero mode” is already a corank-2 degeneracy of the real thimble geometry.
Thus, the complex holomorphic geometry natively upgrades the standard square-root branch cut into a strictly simple pole. This rigorously explains how the degenerate saddle points of the Twistor String path integral explicitly generate the discrete mass poles of the meson spectrum. (For this pole to manifest macroscopically without being capped by higher-order terms, the steepest descent contour must become infinitely long. In the framework of Landau pinch singularities [48], this infinite integration volume is provided dynamically by the Liouville zero-mode stretching the string worldsheet into an infinite cylinder).
Consequently, the extraction of the exact Planar QCD meson spectrum is formally reduced from a stochastic path integral to a deterministic Generalized Eigenvalue Problem (akin to the Gutzwiller trace formula [49] and the Dashen-Hasslacher-Neveu soliton quantization [50]). The physical masses $E_{n}$ are precisely the discrete values for which the classical stability matrix develops a complex null space:
$\det_{\mathbb{C}}\Big[\nabla^{2}S_{eff}(\Phi_{\star})-E_{n}\nabla^{2}\mathcal{O}(\Phi_{\star})\Big]=0$ (17.7)
Solving this boundary-twistor Fredholm equation in general case provides a direct analytical roadmap for the exact, noise-free computation of the geometric spectrum in future work.
17.2 Catastrophes, flat valleys, and topological barriers
In the general theory of functional integration, hadronic resonances emerge when complex saddle points of the effective action become degenerate. A vanishing Hessian matrix indicates a local flat valley in the functional landscape, which is the hallmark of Catastrophe Theory. However, a local flat valley alone is fundamentally insufficient to produce a true $S$-matrix bound state pole. A generic Taylor expansion of the action around a degenerate saddle point will eventually be capped by higher-order terms (e.g., cubic or quartic corrections). These terms force the steepest descent contour to curve upwards, trapping the integral and yielding finite fractional powers or branch cuts rather than macroscopic divergences.
To dynamically generate an exact simple mass pole $\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat zero-mode that extends infinitely far in configuration space without ever curving upwards. All higher-order corrections along this specific trajectory must identically vanish, which is a known localization phenomenon in the WKB approximation for path integrals [11].
In our twistor string framework, this exact infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields. The zero-mode trajectory in functional space corresponds to the topological winding of the boundary parameter as it wraps around the enclosed twistor poles. Because each winding around the complex pole maps holographically to a uniform translation in target-space Minkowski time, the effective action accumulates strictly additively. The total action is exactly linear in the winding number $N$, which serves as the non-compact integration variable along the steepest descent path. Consequently, all higher-order nonlinear corrections for this particular field are strictly absent. The contour along this monodromy direction remains perfectly flat as $N\to\infty$, allowing the integral to diverge into the exact geometric series that analytically generates the simple meson pole.
Furthermore, this geometric mechanism reveals that the number of twistor poles enclosed within the unit circle ($|z_{k}|<1$) serves as a strict topological quantum number that classifies the resonance states.
The integrity of this integer quantum number is rigorously protected by an infinite topological action barrier. One might intuitively ask whether continuous fluctuations could cause twistor poles to smoothly drift across the unit circle boundary ($|z|=1$), thereby changing the topological sector. However, the holomorphic boundary geometry explicitly forbids this. If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside. Exactly at the moment of crossing, these two singularities inevitably collide directly on the physical integration contour $|z|=1$.
This collision creates a classic pinch singularity. The integration loop becomes trapped between the colliding poles $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them. Consequently, the action strictly diverges at the boundary, creating an impenetrable topological barrier. The twistor path integral is therefore dynamically locked into a fixed topological sector, mathematically guaranteeing that the number of enclosed poles is a rigorously conserved topological quantum number.
17.3 Poles inside the disk and the monodromies
We are looking for a saddle point with a single pole located at a complex coordinate $z_{0}$ strictly inside the unit disk ($|z_{0}|<1$). Why inside? Because in this case, there is a non-trivial winding number. As the target space surface $X$ winds up in the time direction, it means the quark goes around the loop infinitely many times, creating a bound state of a rotating string. This infinite motion is precisely the saddle point we are looking for—it will create the resonance pole through a geometric series over the multiple windings.
The total effective action evaluated over these multiple windings consists of two parts. First, the Liouville bulk action is simply additive because its Lagrangian density is a single-valued, explicitly periodic function of the angle on the boundary circle. Therefore, for $k$ windings, its contribution simply duplicates: $S^{(k)}_{bulk}=kS^{(1)}_{bulk}$. Second, the boundary energy term in the effective action involves a logarithmic primitive. Because the pole $z_{0}$ is enclosed by the contour, analytic continuation around the loop causes this term to pick up a strict topological phase shift at every extra cycle. This is the nontrivial monodromy of the effective action.
Together, these terms lead to an infinite sum over the winding number $k$ of the form $\sum\exp\left(\mathrm{i}k\Delta S_{\text{eff}}\right)$. This infinite geometric series is a direct manifestation of the flat valley near the degenerate fixed point. The valley is flat near the energy pole because the monodromy of the effective action $\Delta S_{\text{eff}}$ approaches $2\pi n$. This makes $\exp\left(\mathrm{i}\Delta S_{\text{eff}}\right)=1$, which leaves the amplitude invariant under winding, so that the sum over windings macroscopically diverges as $\sim 1/(1-\exp\left(\mathrm{i}\Delta S_{\text{eff}}\right))\sim 1/(E-E_{n})$.
17.4 One pole and its symmetries
To construct the exact saddle point in the one-pole sector, we leverage the unbroken symmetries of the system. First, we fix the external kinematics by moving to the rest frame of the meson, orienting the injected momentum as $q=(0,0,0,\mathrm{i}E)$. While this specific frame choice explicitly breaks the full Lorentz group, the $O(3)$ spatial rotational symmetry remains intact.
We may utilize this unbroken space rotation to align the twistor dynamics. Specifically, we can always choose a spatial frame such that the symmetric bilinear combination vanishes:
$\lambda^{T}\sigma_{1}\mu=0$ (17.8)
Because the spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, this strict algebraic constraint forces one component of the tangent vector to be zero everywhere on the worldsheet. This dimensionally reduces the physical minimal surface from a generic 4-dimensional embedding into a 3-dimensional hyperplane, exactly matching the geometric prerequisite for a canonical Helicoid.
Furthermore, we possess a global $U(1)$ symmetry on the worldsheet integration domain (the unit disk). The extremum equation for the effective action with respect to the pole position $z_{0}$ must reflect this. Because of this $U(1)$ symmetry, we may claim that there is always a symmetric solution where the pole is centered at the origin:
$z_{0}=0$ (17.9)
The string dynamically centers itself to restore the $U(1)$ symmetry. By locking the pole to the origin, the minimal surface reduces exactly to the canonical, maximally symmetric Helicoid. The remaining computations then proceed in the exact same way as for the canonical Helicoid in Minkowski space. As we derive in a new section in the Appendix, the worldsheet metric following from these canonical twistors $\lambda,\mu$ transforms precisely into the canonical Nambu-Goto rotating string metric. In the text below, we simply use this metric.
17.5 The exact spectrum in the one pole sector
With the twistor saddle point locked to $z_{0}=0$ and the surface constrained to the 3D Helicoid, we repeat the action computation. In this one-pole sector, there are no other terms; the entire worldsheet dynamics are governed by exactly two terms within the Liouville action: the classical area term ($\sigma$) and the kinetic anomaly ($|\nabla\rho|^{2}$).
Because the conformal density is strictly proportional to $1/|z|^{2}$, the kinetic term acts as a rigid, local shift to the effective Helicoid string tension. Evaluating the 1-cycle Minkowski bulk action, we obtain the effective Helicoid energy, complete with the anomalous Lüscher term:
$S^{(1)}_{bulk}=-2\pi^{2}\left(\sigma+\frac{1}{12\pi R^{2}}\right)R^{2}=-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$ (17.10)
Simultaneously, the boundary term coupling the string to the external energy $E$ integrates to yield the exact phase monodromy per winding:
$\Delta S_{E}=4\pi ER$ (17.11)
Summing these to form the total phase of one physical cycle, we enforce the geometric series resonance condition ($\Delta S_{\text{eff}}=2\pi n$):
$4\pi ER-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$
$=2\pi n\implies 2ER-\pi\sigma R^{2}-\frac{1}{12}=n$ (17.12)
To locate the physical vacuum, we minimize this effective base action with respect to the spatial radius $R$ to balance the string tension against the injected energy:
$\frac{\partial}{\partial R}\left(2ER-\pi\sigma R^{2}\right)=2E-2\pi\sigma R=0\implies R=\frac{E}{\pi\sigma}$ (17.13)
Substituting this exact saddle-point radius back into the quantization condition, we obtain the exact algebraic spectral equation:
$2E\left(\frac{E}{\pi\sigma}\right)-\pi\sigma\left(\frac{E}{\pi\sigma}\right)^{2}-\frac{1}{12}=n$ (17.14)
Simplifying the squares yields:
$\frac{E^{2}}{\pi\sigma}=n+\frac{1}{12}\implies E^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (17.15)
Since we are in the rest frame, the invariant mass squared is $m^{2}=E^{2}$, resulting in the final string spectrum:
$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (17.16)
We got the famous Lüsher term, without any string models or assumptions. We obtained this term from the conformal anomaly left from the Elfin determinant, which we used to induce MM loop equations. This remarkable coincidence tells us that our solution for planar QCD matches accepted phenomenology as well as the results of the real and numerical experiments. QED! (or rather QCD!).
17.6 QCD as a Geometric Theory
The derivation of the exact spectral equation in the previous section represents a significant shift in our understanding of the large-$N_{c}$ meson spectrum. We have moved from a stochastic description (random surfaces) to a deterministic geometric principle. Let us summarize the key physical insights that have emerged from this analysis.
Our approach avoids the usual “graviton problem” of gravity holography: since our bulk is not dynamical, there are no propagating bulk graviton modes (or other unwanted massless bulk fields) that would have no counterpart in confining QCD.
More precisely, the propagating degrees of freedom in the present formulation are confined to the boundary: (i) the quark loop degrees of freedom, represented by the matrix trace of the Dirac propagator around the boundary loop, and (ii) the holomorphic moduli (twistors) that parametrize the rigid minimal surface. The twistor moduli are specified by their boundary data and extended uniquely into the disk by analytic continuation (Plemelj/Hilbert reconstruction in the spirit of the Penrose transform). Accordingly, bulk quantities such as the Liouville field are not integrated over as independent quantum fields; they are determined functionally by the analytically continued twistor data.
It is also important not to conflate this construction with Witten/Berkovits twistor strings or other topological twistor-string models. In those theories the extension of boundary data into the bulk typically contributes a topological functional (e.g. a WZW term) whose value depends only on the homotopy class of the extension. In contrast, our bulk functional is not topological: its value depends on the specific holomorphic disk selected by the boundary loop, and this non-topological dependence is essential for confinement and for the existence of a mass scale. In this sense, our “rigid twistor string” is best viewed as gauge holography (bulk as a rigid encoding of boundary dynamics), rather than gravity holography or a topological string.
Finally, the observation relating the QCD mass spectrum to the degenerate saddle points of the complexified twistor-string action introduces the concept of localization of the string path integral, but on an entirely different level. Our spectral equation is exactly localized, much like the computation of path integrals in geometric quantization [51], but this is not a familiar fixed point in the moduli space. Instead, the path integral localizes exclusively on degenerate saddle points governed by catastrophe theory. We thus have uncovered a novel, purely geometric mechanism of localization that deterministically dictates the QCD mass spectrum.
17.7 Future Directions: Glueballs, Baryons and Scattering
The formulation of Planar QCD as an Analytic Twistor String opens several avenues for future research:
  • 1.
    Glueballs: the whole construction, starting with the loop equations to the fermion determinant at the minimal surface, to the rigid twistor string and the localization of the twistor path integral for quark amplitudes at the boundary of the moduli space, naturally generalizes to the higher $1/N_{c}$ corrections. There are some important mathematical issues to be resolved for that program to materialize. This is a natural next step for this Geometric QCD theory.
  • 2.
    Baryons: In the large-$N_{c}$ limit, baryons appear as heavy states composed of $N_{c}$ quarks. Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a "Y-junction" or a surface with multiple boundaries) [5]. Extending the twistor ansatz to surfaces with higher connectivity is a natural next step.
  • 3.
    Scattering Amplitudes: We have focused here on the two-point function (spectrum). The four-point function (meson-meson scattering) involves a minimal surface bounded by four segments. The crossing symmetry and Regge behavior of these amplitudes should emerge from the analytic continuation of the twistor data, potentially connecting this framework to the modern "Amplituhedron" methods.
18 Conclusion
In this paper, we developed the dynamical part of the Geometric QCD program. Building on [1], where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ arises from the rigid Hodge-dual additive minimal surface, we formulated a concrete large-$N_{c}$ framework in which the remaining degrees of freedom can be treated without summing over fluctuating worldsheet metrics.
The first main outcome is conceptual and structural: by working in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum interpretation of the loop equations. In this representation, the planar loop equation becomes an algebraic–differential equation. Furthermore, by mapping the Master Field to the infinite-dimensional Cuntz-algebra “word space” (Section 10.8), we rigorously establish that a purely one-dimensional analytic Taylor–Magnus (finite-moment) solution of the exact vector momentum-loop equation cannot exist: the expansion is consistent through $\mathcal{O}(P^{\prime 6})$ but becomes algebraically inconsistent at $\mathcal{W}^{(8)}$ (C, Table 1). This violent breakdown is a theorem-level obstruction showing that a 1D closed-loop scalar Taylor–Magnus framework lacks sufficient internal degrees of freedom to absorb the macroscopic spatial stress generated by planar QCD. This breakdown prompts the introduction of additional continuous degrees of freedom; in our construction these are the 4D boundary-twistor variables that parametrize the rigid Hodge-dual minimal surface and yield the confining twistor-string representation.
The second main outcome is dynamical: we showed that the fermionic (Majorana) worldsheet degrees of freedom of the Fermi string provide the correct algebraic mechanism for planar factorization at intersections, while the rigidity of the Hodge-dual minimal surface removes the Liouville instability of the original random-surface formulation. The resulting theory may be viewed as an analytic (confining) twistor-string description in which the reduced loop-space measure is most naturally written in twistor variables, leading to an effective holographic Liouville functional determined by holomorphic extension of boundary data.
The third main outcome concerns spectroscopy and the profound realization that the quantum path integral geometrically localizes. We initially demonstrated the computability of the boundary-twistor functional integral by evaluating it on the IAS Typhon supercluster via Complex Langevin (CLE) evolution with logarithmic reweighting. To overcome the severe topological trapping and infinite-variance $1/|u|^{3}$ Cauchy tails inherent to the complexified phase space, we introduced an unsupervised phase-information clustering algorithm. Applying this to our massive stochastic histories successfully isolated two distinct, approximately linear Regge trajectories. This numerical exploration served a crucial purpose: it proved that the Confining Twistor String natively possesses internal quantum numbers (analogous to string “daughter trajectories”) emerging directly from the multiple complex saddle points (Lefschetz thimbles) of the path integral.
However, the ultimate realization of this work goes far beyond numerical stochasticity. By analyzing the analytic structure of the Twistor String, we discovered that the physical mass spectrum is deterministically governed by Catastrophe Theory. The meson poles correspond exactly to degenerate Lefschetz thimbles—bifurcation points where the complex stability matrix of the classical twistor trajectory develops a null space. Because the integration variables are coupled holomorphic twistors, this degeneracy inherently possesses a corank of 2. As dictated by the singularity theory of multiple integrals, this complex degeneracy mathematically upgrades standard semiclassical square-root branch cuts into the exact simple poles of the meson states. Summary
To summarize, what started as a rigorous continuum formulation of planar QCD via the Makeenko–Migdal loop equations culminated in an exact analytical solution for the fundamental meson spectrum. Although our numerical investigations using the Complex Langevin approach successfully provided proof-of-concept evidence for linear Regge trajectories, the true predictive power of this framework lies in its analytical semiclassical limit, which provides exact equations for the mass spectrum beyond the WKB approximation at the non-perturbative level. It is a known phenomenon that some WKB spectral equations are valid beyond the WKB approximation, because they reflect deeper hidden symmetries leading to infinite flat valleys for the steepest descent path, making the WKB estimate an exact localization of the path integral [11]. We have shown that the $S$-matrix mass poles of the theory are naturally captured by Catastrophe Theory, corresponding to non-compact flat valleys in the complexified effective action. These flat valleys, where the steepest descent integral diverges without higher-order corrections, are driven entirely by the topological monodromies of twistor poles enclosed within the unit disk.
The number of such poles serves as a rigorously conserved topological quantum number, protected by an absolute pinch-singularity barrier at the conformal boundary. The crossing of the unit circle by a twistor pole leads to a divergent action, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop. By analytically continuing the Euclidean area action to the boundary and summing over the exact purely imaginary monodromy phases, we obtain a Bohr–Sommerfeld quantization condition. For the fundamental one-pole trajectory, this yields the exact simple mass poles and the linear Regge spectrum:
$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (18.1)
This analytical result not only confirms the perfectly linear Regge slope but uniquely derives the exact open-string Lüscher intercept $(+1/12)$ directly from the 2D anomaly of the Elfin determinant on rigid minimal surface. Thus, by mapping the momentum loop-space dynamics to boundary twistors, the theory does not merely approximate the confining dynamics of planar QCD; it provides its exact algebraic geometric solution. The exact analytical classification of higher topological sectors (multi-pole configurations) and their corresponding daughter trajectories is now a fully accessible problem. Consequently, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry. The extraction of the mass spectrum is reduced from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem for the degenerate saddle points of the effective action.
Acknowledgements
I am grateful to Nima Arkani-Hamed for the invitation to present this work as a series of lectures at the IAS Particle Physics (Pizza) Seminar, which provided a stimulating environment for discussing these results. I also thank Edward Witten for valuable discussions regarding the distinction between this confining geometry and the holographic dualities of conformal field theories.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author used Google Gemini to fix typographical errors, identify relevant mathematical terminology, clarify the notations, and improve the overall language and style. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Data Availability.
The spectral data reported in this paper are generated by numerical simulation of the Complex Langevin equation (CLE) for the discretized Confining Twistor String spectral observable, including the reweighted estimator for $1/R(E)=\Pi(E)/\Pi^{\prime}(E)$. To enable reproduction and extension of these calculations, I have made the full Mathematica ® workflow publicly available on the Wolfram Cloud, including (i) the CLE simulator and effective-action implementation, (ii) SLURM submission scripts for running independent replicas on an HPC cluster, and (iii) the analysis tools used to pool independent runs and extract the spectrum:
  • 1.
    TwistorString.wl, [52] (main CLE simulator):
  • 2.
    run_node.m (SLURM/cluster driver) [53]
  • 3.
    VerifyTwistorHistory.nb [54] (fat tail analysis of distributions for a CLE history):
  • 4.
    TwistorInformationClustering.nb [55] (spectrum extraction and plots)
  • 5.
    GeometricQCD CLE results [56] (the full set of histories of the evolution of $R(E)$ for $240$ values of energy $E\in(2.,3.)$ ran on 64 nodes of Typhon cluster, with random initial states)
The raw per-replica output consists of node-indexed binary arrays (e.g. TwistorHistory*.mx) storing, for each sampled energy value, the whole stochastic trajectory of observable resolvent $R(E)$. These files are verified by VerifyTwistorHistory.nb, testing fat tails of distribution against theoretical prediction. Then TwistorInformationClustering.nb performs robust pooling across independent replicas, analyzes the history of the phase of the resolvent, builds the statistical distribution, clusters the data by information distance between these phase distributions and produces the final spectral curves and extracted mass tables. This public release is intended to provide a practical basis for future large-scale runs (larger ensembles, finer discretization, and improved CLE diagnostics and stabilization methods).
Appendix A Conformal anomaly in determinants
A.1 Laplace operator
Let us consider the logarithm of the determinant of the scalar Laplace operator:
[…]
We apply the standard Pauli-Villars regularization
$(\ln\hat{L})_{reg}=\ln\hat{L}-\sum c_{i}\ln(\hat{L}+M_{i}^{2}).$ (A.2)
Now, at finite regulator masses $M_{i}$ the trace converges provided the constants $c_{i}$ satisfy the sum rules
$\sum c_{i}M_{i}^{2k}=\delta_{k0},\qquad k=0,1,$... (A.3)
Consider now the conformal variation
$\delta\hat{L}$ $=-\delta ee^{-1}\hat{L}$, (A.4)
$\delta\ln\hat{L}$ $=\delta\hat{L}\hat{L}^{-1}=-\delta ee^{-1};$ (A.5)
$\delta\ln(\hat{L}+M^{2})$ $=-\delta ee^{-1}(1-M^{2}(\hat{L}+M^{2})^{-1}).$ (A.6)
The variation of the regularized trace (A.2) reads-
[…]
Now only the regulator terms contribute, so that we may apply the WKB expansion
$\hat{L}=-\frac{1}{e}(\partial_{\xi})^{2}\to-\frac{1}{e_{0}}\left(1+\frac{e_{0}R_{0}}{4}(\xi-\xi_{0})^{2}\right)(\partial_{\xi})^{2}$ (A.8)
where $R_{0}$ is the local curvature. Here $\xi_{0}$ is the running point in the integral
[…]
Using the Fourier expansion at the tangent plane,
$\left\langle\xi_{0}|(\hat{L}+M^{2})^{-1}|\xi_{0}\right\rangle$
[…]
and calculating the Fourier integral, we find [by virtue of (A.3)]
[…]
This equation can be easily integrated backwards:
[…]
with
$a_{1}=\frac{1}{4\pi}\sum c_{i}M_{i}^{2}\ln M_{i}^{2},$ (A.13)
and arbitrary constant $a_{0}$.
A.2 Dirac operator
In case of the square of the Dirac operator $D^{2}$ the WKB expansion starts as follows:
$D^{2}\approx\mathds{1}(\hat{L}+R/4)+\frac{\mathrm{i}\sigma_{3}}{4}R\epsilon_{ab}\xi_{a}\partial_{\xi_{b}}.$ (A.14)
The definition of the Dirac operator was given in the text. The same line of argument as with the Laplace operator yields
$\sum_{\alpha=1,2}\left\langle\alpha\xi_{0}|(D^{2}+M^{2})^{-1}|\alpha\xi_{0}\right\rangle$
[…]
$\left(1+1/4R_{0}(-1+(\partial_{p})^{2}p^{2})(M^{2}+p^{2})^{-1})\right),$ (A.15)
[…]
The additional factor of 2 in these relations comes from two spin states $a=1,2$. The term with the $\sigma_{3}$ matrix in (A.14) did not contribute.
Note the remarkable relation
$\det D^{2}\det\hat{L}=const\cdot\exp\left(\textit{ const }{}Area\right)$ (A.18)
which follows from the above formulas.
Appendix B Spinor Factorization and Analytical Structure of Minimal surface
The mathematical problem of minimization of the Dirichlet functional with Virasoro constraint in four dimensions is a well known problem in modern mathematics. Let us briefly summarize the existing theory.
B.1 Twistor Framework and the Klein Quadric
The fundamental constraint of the theory is the null condition on the holomorphic tangent vector of the minimal surface, $(f^{\prime})^{2}=\sum_{\mu=1}^{4}(f^{\prime}_{\mu})^{2}=0$. Geometrically, this defines the Klein Quadric in the complexified tangent space $\mathbb{CP}^{3}$. Using the isomorphism $SO(4,\mathbb{C})\cong SL(2,\mathbb{C})_{L}\times SL(2,\mathbb{C})_{R}$, we map the vector $f^{\prime}_{\mu}$ to a $2\times 2$ matrix $f^{\prime}_{a\dot{b}}=f^{\prime}_{\mu}(\sigma_{\mu})_{a\dot{b}}$. The null condition corresponds to $\det(f^{\prime})=0$, ensuring the matrix has rank 1. Thus, the tangent vector factorizes globally into a product of two spinors:
$f^{\prime}_{a\dot{b}}(z)=\lambda_{a}(z)\mu_{\dot{b}}(z)$ (B.1)
where $\lambda_{a}$ and $\mu_{\dot{b}}$ are meromorphic sections of spinor bundles over the worldsheet. This factorization linearizes the quadratic Virasoro constraint, replacing it with the problem of determining the analytic properties of the spinors.
B.2 Gauge Invariance and Birkhoff Factorization
The physical vector $f^{\prime}$ is invariant under the local complex gauge transformation:
$\lambda_{a}(z)\to w(z)\lambda_{a}(z),\qquad\mu_{\dot{b}}(z)\to w^{-1}(z)\mu_{\dot{b}}(z)$ (B.2)
where $w(z)$ is a meromorphic function. This redundancy is governed by the Birkhoff-Grothendieck theorem on the splitting of vector bundles over $\mathbb{P}^{1}$. The gauge function $w(z)$ allows us to distribute the zeroes and poles between $\lambda$ and $\mu$, classifying solutions by integer topological indices (partial indices) related to the winding number of the gauge field required to regularize the spinors at infinity.
B.3 Analytic Structure via Plemelj Projection
The connection to the physical loop $C(\sigma)$ (where $\sigma=e^{\mathrm{i}\theta}$) is achieved via the Plemelj-Sokhotski decomposition. We expand the loop coordinates into positive and negative frequency modes:
$C(\sigma)=C_{+}(\sigma)+C_{-}(\sigma)=\sum_{n\geq 0}c_{n}\sigma^{n}+\sum_{n<0}c_{n}\sigma^{n}$ (B.3)
If the loop is algebraic, $C_{-}(\sigma)$ extends to a rational function $C_{-}(z)$ in the complex plane $|z|>1$. The poles of the spinor solution $f^{\prime}(z)$ are dictated by the singularities of $C_{-}(z)$. Analytically, the problem is to find the "closest" null vector $f^{\prime}$ to the derivative $\partial_{z}C_{+}$. This is equivalent to finding a subspace $W$ in the Sato Grassmannian that satisfies the isotropic condition and matches the boundary data.
B.4 Explicit Algebraic Solutions
The general theory of the spinor factorization predicts that rational choices for the sections $\lambda_{a}(z)$ and $\mu_{\dot{b}}(z)$ generate algebraic minimal surfaces. These correspond to "Finite Gap" solutions where the algebraic curve has a finite genus. The distribution of poles between the spinors determines the topology and boundary behavior of the surface.
We illustrate this with four classical examples, reinterpreted in our spinor language: 1. The Enneper Surface (Polynomial Solution). This is the simplest solution with trivial topology (genus zero, one boundary at infinity), corresponding to polynomial spinors with a single pole at $z=\infty$.
$\lambda(z)=\begin{pmatrix}1\\ z\end{pmatrix},\qquad\mu(z)=\begin{pmatrix}1\\ -z\end{pmatrix}$ (B.4)
The resulting null vector $f^{\prime}(z)\sim(1-z^{2},\mathrm{i}(1+z^{2}),2z,0)$ integrates to a cubic polynomial map. This surface, discovered by Enneper in 1864 [57], represents the fundamental "ground state" of the algebraic solutions. 2. The Catenoid (Rational Solution with Poles). To generate a surface with the topology of an annulus (the only minimal surface of revolution), one must introduce simple poles into the spinors.
$\lambda(z)=\begin{pmatrix}z^{-1/2}\\ z^{1/2}\end{pmatrix},\qquad\mu(z)=\begin{pmatrix}z^{-1/2}\\ -z^{1/2}\end{pmatrix}$ (B.5)
Using the gauge freedom $\lambda\to w\lambda,\mu\to w^{-1}\mu$ with $w=z^{-1/2}$, we can redistribute the poles to find the standard rational form. Euler [58] originally proved the minimality of this surface, which corresponds to the physical shape of a soap film stretched between two rings. 3. The Henneberg Surface (Non-Orientable). A more complex rational ansatz generates the Henneberg surface [59], which is a non-orientable minimal surface (a realization of the projective plane in 3D space). It arises from a rational map of degree 4:
$\lambda(z)=\begin{pmatrix}1-z^{-2}\\ z-z^{-1}\end{pmatrix},\qquad\mu(z)=\begin{pmatrix}1\\ -1\end{pmatrix}$ (B.6)
4. The Helicoid (Transcendental Solution). The Helicoid is the locally isometric conjugate surface to the Catenoid. In the spinor formalism, it arises from the same meromorphic data but with a phase shift that exposes the period of the complex logarithm.
$\lambda(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}z^{-1/2}\\ z^{1/2}\end{pmatrix},\qquad\mu(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}z^{-1/2}\\ -z^{1/2}\end{pmatrix}$ (B.7)
[…]
Appendix C Explicit invariant tensor solutions for the MLE
In this appendix, we provide the explicit polynomial solutions for the finite functional Taylor–Magnus expansion of the Momentum Loop Equation, valid up to the non-analytic boundary at $\mathcal{W}^{(8)}$.
C.1 Solving for the lowest terms $W^{(n<8)}$
The vacuum is normalized to $\mathcal{W}^{(0)}=1$. With the topological closure of the boundary gaps enforced via the subtraction of interleaved permutations, all odd traces identically vanish, $\mathcal{W}^{(2k-1)}=0$.
The fully symmetric even-parity invariant tensors are constructed from all possible pair-wise contractions of the Kronecker delta $\delta_{\mu\nu}$. The lowest-order non-trivial tensors, parameterized by the unconstrained fundamental mass scales, evaluate to:
$\mathcal{W}^{(2)}_{\mu_{1}\mu_{2}}$ $=\delta_{\mu_{1}\mu_{2}}$
$\mathcal{W}^{(4)}_{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}$ $=\frac{1}{6}\delta_{\mu_{1}\mu_{3}}\delta_{\mu_{2}\mu_{4}}$
$+c_{4,1}\left(\delta_{\mu_{1}\mu_{2}}\delta_{\mu_{3}\mu_{4}}+\delta_{\mu_{1}\mu_{4}}\delta_{\mu_{2}\mu_{3}}\right)$ (C.1)
The tensor $\mathcal{W}^{(6)}_{\mu_{1}\dots\mu_{6}}$ is completely determined by the geometric stress cascading from the left-hand side loop derivative acting on the lower-order terms, successfully absorbing the kinematic constraints without contradiction. Its explicit expansion in terms of the 15 independent 6-point Kronecker pairings is uniquely fixed by the lower-order parameters $c_{2,1}\equiv 1$, $c_{4,1}$, and $c_{4,2}$.
At $\mathcal{W}^{(8)}$, this linear absorption fails. The expansion truncates due to the irreducible non-linear $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$ stress, which exceeds the tensor rank of the pure cyclic Shuffle Ideal basis, rigorously proving the non-analyticity of the functional $\mathcal{W}[P]$.
We compute these terms in [39], where we prove lack of solution for $\mathcal{W}^{(8)}$.
C.2 Counting of unknowns/equations and Fredholm inconsistency at $\mathcal{W}^{(8)}$
Order $n$ $N_{eq}(n)$ $N_{unk}(n)$ $\text{rank}(A_{n})$ $\dim\text{LeftNull}(A_{n})$ $\Delta\text{rank}$ Solvable?
4 4 6 4 0 0 yes
6 31 30 26 5 0 yes
8 379 212 82 297 1 no
Table 1: Linear algebra diagnostics for the exact Vector Momentum Loop Equation. Here $\Delta\text{rank}:=\text{rank}([A_{n}|b_{n}])-\text{rank}(A_{n})$. A strict positivity, $\Delta\text{rank}>0$, rigorously proves the system is mathematically inconsistent.
To rigorously demonstrate the breakdown of the Taylor-Magnus expansion at the 8th order ($\mathcal{W}^{(8)}$), we cast the Momentum Loop Equation at each order $n$ into a finite linear system. After imposing the closed-loop shuffle-ideal reduction (loop-closing prescription), cyclic symmetry, and parity constraints (all odd orders vanish), the exact functional matching reduces to:
$A_{n}x_{n}=b_{n}$ (C.2)
where the vector of unknowns $x_{n}$ is explicitly decomposed as:
$x_{n}=(c^{(n)};\;\alpha^{(<n)};\;g^{(n)})$ (C.3)
Here, $c^{(n)}$ are the coefficients of the invariant-tensor basis of $\mathcal{W}^{(n)}$, $\alpha^{(<n)}$ are the residual free parameters surviving from lower-order solutions (e.g., those left unconstrained in $\mathcal{W}^{(4)}$ and $\mathcal{W}^{(6)}$), and $g^{(n)}$ are the gauge parameters associated with solving modulo the closed-loop shuffle ideal.
The total number of unknown variables solved for at order $n$ is therefore $N_{unk}(n)=N_{\mathcal{W}}(n)+N_{free}(<n)+N_{gauge}(n)$. We analyzed the exact linear algebra of these systems using a specialized symbolic script. The exact matrix ranks and subsystem counts are summarized in Table 1.
The finite-dimensional system $A_{n}x_{n}=b_{n}$ is solvable if and only if it satisfies the Fredholm alternative consistency condition:
$u^{T}b_{n}=0\qquad\forall u\in\text{LeftNull}(A_{n})$ (C.4)
At $n=8$, the matrix $A_{8}$ possesses 297 left-null vectors. Projecting the inhomogeneous right-hand side $b_{8}$ (the geometric stress generated by the discrete non-linear cross-terms of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$) onto these null directions yields explicit, non-zero rational obstructions. For example, selecting the first canonical left-null basis vector $u_{1}$ yields the exact rational projection:
$u_{1}^{T}b_{8}=\frac{1}{36}\neq 0$ (C.5)
This establishes a decisive mathematical contradiction. Two critical robustness statements follow directly from this matrix representation:
  • 1.
    Because the vector of unknowns $x_{8}$ explicitly includes all residual lower-order parameters $\alpha^{(<8)}$, this inconsistency proves that no possible choice of lower-order free parameters can restore solvability.
  • 2.
    Setting the physical scale parameter to zero does not alleviate the structural rank deficiency.
The full basis construction, shuffle-ideal reduction, and exact linear-algebra verification—including the generation of the matrices $A_{n}$ and the explicit null-vector projections—are provided in the executable Mathematica notebook [39] (see Data Availability).
Appendix D Algorithm for Fast Spectral-Gaussian Integration
To evaluate the Liouville action in the bulk, we employ a spectral-product quadrature rule that combines the spectral accuracy of the boundary discretization with high-precision Gaussian integration in the radial direction. This method avoids coordinate singularities at the origin and preserves the holomorphicity of the extension.
The integral of the Liouville Lagrangian density over the unit disk is computed in polar coordinates:
$S_{\text{Liouville}}=\int_{0}^{1}rdr\int_{0}^{2\pi}d\theta\,L\left(\rho(r,\theta),|\partial_{z}\rho(r,\theta)|^{2}\right)$ (D.1)
We discretize this domain using a product grid defined by the $N$ roots of unity in the angular direction and $M$ Gauss-Legendre nodes in the radial direction.
1. The Radial Grid (Gauss-Legendre): We map the standard Legendre nodes $x_{m}\in[-1,1]$ (roots of the Legendre polynomial $P_{M}(x)$) to the radial interval $r\in(0,1]$:
$r_{m}=\frac{1+x_{m}}{2},\qquad w_{m}^{\prime}=\frac{w_{m}}{2}$ (D.2)
where $w_{m}$ are the standard weights. This quadrature is exact for polynomials of degree 2M12M-1. Since the origin $r=0$ is not a node, the coordinate singularity is naturally avoided.
2. The Angular Grid (Roots of Unity): For the angular integration, we use the trapezoidal rule on the $N$ vertices of the polygon, $\theta_{k}=2\pi k/N,\omega_{k}=e^{\mathrm{i}\theta_{k}}$. For periodic functions (such as our fields on circles of constant radius), this rule exhibits exponential convergence (spectral accuracy).
3. Bulk Field Evaluation: The values of the (anti)holomorphic null-vector fields $W(z)=\left\{V(z),\tilde{V}(\bar{z})\right\}$ at a grid point $z_{mk}=r_{m}\omega_{k}$ are computed using the analytic extension formula (E.3). Crucially, the term $(z_{mk}/\omega_{j})^{N}$ reduces to $r_{m}^{N}$, which is independent of the angular indices. This simplifies the kernel significantly:
$V(r_{m}\omega_{k})=\frac{1-r_{m}^{N}}{N}\sum_{j=0}^{N-1}\frac{V(\omega_{j})}{1-r_{m}\omega_{k}/\omega_{j}};$ (D.3)
$\tilde{V}(r_{m}\bar{\omega}_{k})=\frac{1-r_{m}^{N}}{N}\sum_{j=0}^{N-1}\frac{\tilde{V}(\omega_{j})}{1-r_{m}\bar{\omega}_{k}/\omega_{j}}$ (D.4)
This summation takes the form of a discrete convolution. By choosing $N$ to be a power of 2 (e.g., $N=128$), this convolution can be accelerated via Fast Fourier Transform (FFT) to $O(N\log N)$ operations per radial ring, providing a massive speedup over the naive $O(N^{2})$ summation.
4. Hybrid Algorithm for Analytic Derivative: The kinetic term in the Liouville action requires the holomorphic derivative $\partial_{z}\Lambda$. To maintain precision, we differentiate the interpolating polynomial analytically. Differentiating (E.3) with respect to $z$ yields:
$\partial_{z}V(z)\Big|_{z_{mk}}=\frac{1}{N}\sum_{j=0}^{N-1}V(\omega_{j})/\omega_{j}K^{\prime}\left(z_{mk}/\omega_{j}\right)$ (D.5)
where $K^{\prime}(w)$ is the derivative of the geometric series kernel. To optimize computational complexity without sacrificing stability near the boundary, we employ a hybrid algorithm:
a) The Fractional Form (Bulk): For the majority of grid points in the bulk ($1-r_{m}>\epsilon\sim 0.1$), we use the closed fractional form:
$K^{\prime}(w)=\frac{(1-w^{N})+Nw^{N-1}(1-w)}{(1-w)^{2}}$ (D.6)
This form is algebraically equivalent to the standard derivative but is numerically rearranged to avoid cancellations. Its evaluation is $O(1)$, ensuring the total derivative calculation remains efficient.
b) The Geometric Doubling Algorithm (Boundary Layer): For the outermost radial nodes where $r_{m}>1-\epsilon$, the fractional form suffers from numerical instability due to the denominator $(1-w)^{2}$. In this "skin" layer, we switch to the exact polynomial sum $K^{\prime}(w)=\sum_{n=1}^{N-1}nw^{n-1}$. Crucially, for $N=2^{k}$, this sum can be evaluated in $O(\log N)$ operations using the recursive doubling identity for geometric series:
$S_{2N}(w)=(1+P_{N}(w))S_{N}(w);$ (D.7)
$P_{2N}(w)=P^{2}_{N}(w);$ (D.8)
$S_{1}(w)=1;\qquad P_{1}(w)=w;$ (D.9)
and its derivative by $w$. Here is the iterative algorithm:
Algorithm 1 Fast Geometric Doubling Algorithm ($O(\log N)$)
0: Complex argument $w$, Iterations $k$ (where $N=2^{k}$)
0: Kernel values $K=\sum_{n=0}^{N-1}w^{n}$ and $K^{\prime}=\sum_{n=1}^{N-1}nw^{n-1}$
1:Initialize:
[…]
6:for $i=1$ to $k$ do
7:  $K^{\prime}\leftarrow(1+P)K^{\prime}+P^{\prime}K$ {Update Derivative of Sum (Product Rule)}
8:  $K\leftarrow(1+P)K$ {Update Sum ($S_{2M}=(1+w^{M})S_{M}$)}
9:  $P^{\prime}\leftarrow 2\cdot P\cdot P^{\prime}$ {Update Derivative of Power}
10:  $P\leftarrow P\cdot P$ {Update Power ($w^{2M}=(w^{M})^{2}$)}
11:end for
12:return $K,K^{\prime}$
This iterative algorithm computes both the kernel and its derivative with logarithmic complexity $O(\log N)$, avoiding the $O(N)$ cost of the direct sum while using $O(1)$ memory and maintaining perfect numerical stability near the boundary. The numerical stability of this algorithm is verified in our Mathematica ® code [52].
5. Summation: The total action is obtained by summing the weighted contributions:
$S\approx\frac{2\pi}{N}\sum_{m=1}^{M}w_{m}^{\prime}r_{m}\sum_{k=0}^{N-1}L_{\text{Liouville}}(r_{m},\theta_{k})$ (D.10)
This algorithm ensures that the bulk integration error is on the same order as the boundary discretization error, providing a stable and consistent numerical definition of the rigid twistor string. In our implementation (using Mathematica), the kernel data is precomputed into parallel tables, further accelerating the simulation.
Appendix E Numerical Implementation of CLE
To resolve the severe sign problem inherent in the Twistor String path integral, we employ Stochastic Quantization via the Complex Langevin Equation (CLE) [42, 60].
E.1 Complexification of the parameter space
To ensure the strict analyticity of the stochastic evolution required by the Parisi condition [42], we must eliminate all non-holomorphic operations from the Langevin equations. We achieve this by complexifying the target manifold and analytically continuing the action integral itself. This ensures that the drift forces are partial derivatives of a globally holomorphic function.
We replace the original real and Hermitian variables with a set of independent holomorphic variables. The angular coordinates $\alpha_{k}$ are mapped to the complex plane via $z_{1}=e^{i\alpha_{1}},,z_{2}=e^{i\alpha_{2}}$. The spinor fields and their conjugates are treated as independent complex vectors. The full set of dynamical variables for the discretized loop is:
$\mathcal{V}=\{\{z_{1},z_{2}\},\{\rho_{k},\xi_{k},\eta_{k},\tilde{\xi}_{k},\tilde{\eta}_{k},\gamma_{1,k},\gamma_{2,k}\}_{k=1}^{N}\}.$ (E.1)
Here, $\tilde{\xi}$ and $\tilde{\eta}$ replace $\bar{\xi}$ and $\bar{\eta}$ as independent variables. The geometric constraints are enforced by the complex Lagrange multipliers $\gamma_{1,k}$ and $\gamma_{2,k}$, leading to the holomorphic constraint terms in the effective action:
$S_{\text{constr}}=i\sum_{k}\left[\gamma_{1,k}(\tilde{\xi}_{k}\xi_{k}-1)+\gamma_{2,k}(\tilde{\eta}_{k}\eta_{k}-1)\right].$ (E.2)
E.2 Spectral Interpolation of Fields
To evaluate the action integrals analytically, we construct global holomorphic polynomials that interpolate the values at the vertices $z_{k}$.
E.2.1 Spinor Fields
Let $\Lambda^{(k)}=e^{\rho_{k}/2}(\xi_{k},\eta_{k})$ denote the spinor data $\Lambda(z)=\{\lambda(z),\mu(z)\}$ at vertex $k$. The Lagrange interpolation polynomial on the spectral unit circle is:
$\Lambda(z)=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}K\left(ze^{-\mathrm{i}\theta_{k}}\right);$ (E.3)
$K(z)=\frac{1-z^{N}}{1-z};$ (E.4)
Using the Fast Fourier Transform (FFT), we extract the spectral coefficients $L_{n},M_{n}$ (and $\tilde{L}_{n},\tilde{M}_{n}$ from the independent $\tilde{\xi},\tilde{\eta}$ variables) such that:
$\lambda(z)=\sum_{n=0}^{N-1}L_{n}z^{n},\qquad\tilde{\lambda}(z)=\sum_{n=0}^{N-1}\tilde{L}_{n}z^{n}.$ (E.5)
$\{L_{n},M_{n}\}=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}e^{-\mathrm{i}\theta_{k}}$ (E.6)
E.2.2 Liouville Field
The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an exact formula
$\rho(z,\bar{z})=\frac{\log\tilde{\lambda}(\bar{z})\lambda(z)+\log\tilde{\mu}(\bar{z})\mu(z)}{2}$ (E.7)
where $\lambda(z),\mu(z)$} is extended by interpolation (E.3). The derivatives of the above kernel $K^{\prime}(z)$ involved in the Liouville Lagrangian with this $\rho$ were computed in the previous sections.
These expressions provide the values of the fields and derivatives at any complex $z$, not just at the vertices. These expressions also depend upon complex parameters in $\mathcal{V}$ through the Fourier coefficients $\Lambda(k),\tilde{\Lambda}(k)$. This dependence is exponential for $\rho_{k}$ and polynomial for the rest of variables in $\mathcal{V}$, which allows for required analytic continuation of the CLE.
E.3 The Boundary Terms in the action
[…]
$G_{p}(z_{1},z_{2})\equiv\int_{z_{1}}^{z_{2}}z^{p}\frac{dz}{iz}=\frac{z_{2}^{p}-z_{1}^{p}}{ip}.$ (E.8)
With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is strictly positive, ensuring there are no singularities in the complex plane.
The contribution to the action of external momentum $q$ injected into the quark loop from the arc between $z_{1}$ and $z_{2}$ is:
$S_{q}=-\mathrm{i}q_{\alpha}$
$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}\sigma_{\alpha}M_{m})-(\tilde{L}_{n}\sigma^{\dagger}_{\alpha}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (E.9)
Then, there are terms related to the constraints on the $\xi,\eta$ variables. These terms are just discrete sums
$S_{\gamma}=\mathrm{i}\sum_{k=0}^{N-1}\gamma_{1,k}\left(\tilde{\xi}_{k}\xi_{k}-1\right)+\gamma_{2,k}\left(\tilde{\eta}_{k}\eta_{k}-1\right)$ (E.10)
In addition, there are terms related to the trace (12.30). Our complexification reduces this trace to
[…]
$\mathcal{P}(\theta)=2\mathrm{i}\left[z\eta\tilde{\xi}^{T}-\bar{z}\xi\eta^{T}\right]_{z=e^{\mathrm{i}\theta}}$ (E.13)
Finally, our reweighting adds one more complex term to the action,
$S_{O}=\log\mathcal{O};$ (E.14)
$\mathcal{O}=\frac{\partial S_{q}}{\partial q_{4}}=$
$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}M_{m})-(\tilde{L}_{n}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (E.15)
All the above expressions are purely holomorphic functions of all the variables in $\mathcal{V}$.
E.4 The Bulk Action (Gauss-Legendre Integration)
The bulk action $S_{\rho}$ involves the kinetic and potential terms for the Liouville field of the form: $\int d^{2}z[\partial_{z}\rho\partial_{\bar{z}}\rho+V(\rho)]$. Unlike the boundary term, this cannot be integrated in closed form due to the nonlinear dependence of both variables $z,\bar{z}$.
We evaluate this integral using $M$-point Gauss-Legendre quadrature coupled with spectral (FFT) summation over the roots of unity for the angular integration. The integration domain is divided into $N$ rays $R:(0,z_{k});\,z_{k}=e^{\mathrm{i}\theta_{k}}$. Then, $M$ quadrature nodes $z_{q}\in(0,1)$ are placed on each ray: $z_{k,q}=z_{q}e^{\mathrm{i}\theta_{k}}$.
The bulk action is approximated as:
$S_{\rho}\to\frac{1}{N}\sum_{k=0}^{N-1}\sum_{q=1}^{M}w^{\prime}_{q}\mathcal{L}_{\text{bulk}}\left(\rho(z_{k,q}),\rho^{\prime}(z_{k,q})\right).$ (E.16)
Here, the values of the field $\rho(z,\bar{z})$ and its derivative $\rho^{\prime}(z,\bar{z})$ at the quadrature points $z_{k,q}$ are computed exactly using the spectral Lagrange interpolation formula (Eq. E.3) derived in the previous section. This combination ensures exact integration of the polynomials in $z,\bar{z}$ of degree up to $2M$ and exponential accuracy of angular integrations. For practical purposes, this means that the integration errors will be negligible or $N,M=64$. This, of course, does not imply the exponential conversion of the CLE to the path integral at large $N,M$. The CLE introduces errors of its own, and statistical errors $O(1/\sqrt{T})$ related to the cancelation of oscillating terms are the dominant.
E.5 Holomorphic Langevin Equations
The time evolution is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{\text{total}}=S_{\alpha}+S_{\gamma}+S_{T}+S_{\rho}+S_{O}$.
  1. 1.
    Coordinate Evolution ($z_{k}$): The force on $z_{k}$ arises from the boundary terms of the $G$ functions in $S_{\text{int}}$ and the integration limits in $S_{\text{bulk}}$.
    $\partial_{\tau}z_{k}=-z_{k}^{2}\frac{\partial S_{\text{total}}}{\partial z_{k}}+iz_{k}\nu_{z_{k}}.$ (E.17)
    (The factor $z_{k}^{2}$ arises from the metric of the map $\alpha=-i\log z$).
  2. 2.
    Spinor Evolution ($\xi_{k},\tilde{\xi}_{k}$, etc.): The forces are computed via the chain rule through the spectral coefficients $L_{n}$. For example:
    $\partial_{\tau}\xi_{k}=-\frac{\partial S_{\text{total}}}{\partial\xi_{k}}-i\gamma_{1,k}\tilde{\xi}_{k}+\nu_{\xi_{k}}.$ (E.18)
    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables.
  3. 3.
    Liouville Evolution ($\rho_{k}$): The force on $\rho_{k}$ is nonlocal, mediated by the interpolation kernel $K_{k}(z)$ appearing in both the bulk quadrature points and the spectral coefficients of the boundary term.
    $\partial_{\tau}\rho_{k}=-\frac{\partial S_{\text{total}}}{\partial\rho_{k}}+\nu_{\rho_{k}}.$ (E.19)
  4. 4.
    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the constraints:
    $\partial_{\tau}\gamma_{1,k}=-i(\tilde{\xi}_{k}\xi_{k}-1)+\nu_{\gamma}.$ (E.20)
By utilizing this spectral action, we ensure that the drift terms are analytic functions of all variables $\mathcal{V}$ throughout the complex plane. This prevents the "excursion" problem common in standard CLE, where non-analytic projections onto the real axis lead to incorrect convergence.
E.6 Reweighted Spectral Search for Meson Masses
The poles of $\Pi(E)$ correspond to the bound state masses (the meson spectrum), and these poles also become poles of its log derivative $R(E)$.
As discussed in Section 15, directly searching for these poles in a stochastic simulation is numerically disastrous, as the infinite variance near the singularity destroys the Markov chain. Instead, we employ a reweighting method to compute the inverse ratio $1/R(E)$ and search for its zeros. This inversion mathematically transforms the meromorphic amplitude into an entire function, effectively neutralizing the singularities. We define the observable as the derivative of the action with respect to energy:
$\mathcal{O}=\frac{\partial S_{\text{eff}}}{\partial E}$
$=-\mathrm{i}\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}M_{m})-(\tilde{L}_{n}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2})$ (E.21)
Then, we have an identity
$\Pi^{\prime}(E)/\Pi(E)=\frac{\int\mathcal{O}e^{S_{\text{eff}}}}{\int e^{S_{\text{eff}}}}$ (E.22)
We then introduce a Modified Action $\tilde{S}$ by shifting the probability measure:
$\tilde{S}=S_{\text{eff}}+\ln\mathcal{O}$ (E.23)
Using this modified action, the inverse ratio can be expressed as a statistical average:
$\frac{1}{R(E)}=\frac{\Pi(E)}{\Pi^{\prime}(E)}=\frac{\int e^{S_{\text{eff}}}}{\int\mathcal{O}e^{S_{\text{eff}}}}=\frac{\int\frac{1}{\mathcal{O}}e^{\tilde{S}}}{\int e^{\tilde{S}}}=\left\langle\frac{1}{\mathcal{O}}\right\rangle_{\tilde{S}}$ (E.24)
Thus, the algorithm proceeds as follows:
  1. 1.
    We simulate the Complex Langevin evolution governed by the modified action $\tilde{S}$. This introduces an additional "logarithmic force" to the drift: $\nabla\tilde{S}=-\nabla S_{\text{eff}}-\frac{1}{\mathcal{O}}\nabla\mathcal{O}$.
  2. 2.
    At each time step, we measure the quantity $1/\mathcal{O}$.
  3. 3.
    We compute the time average $\langle 1/\mathcal{O}\rangle_{\tilde{S}}$.
  4. 4.
    We scan the energy $E$ and identify the meson masses as the points where this average vanishes (crosses zero).
This method transforms the search for poles into a search for zeros, which is numerically stable and precise.
E.7 Spectral Extraction and the Sign of the Slope
A crucial subtlety arises in interpreting the zeros of the measured observable. The simulation computes the expectation value of the inverse resolvent operator, $\langle\mathcal{O}^{-1}\rangle\sim\text{Re}[1/R(E)]$. To distinguish physical mass states from artifacts, we must analyze the pole structure of the resolvent.
The resolvent is related to the logarithmic derivative of the vacuum polarization $\Pi(E)$:
$R(E)\sim\frac{\Pi^{\prime}(E)}{\Pi(E)}=\frac{d}{dE}\ln\Pi(E)$ (E.25)
This function has singularities of two types:
  1. 1.
    Physical Poles: At a particle mass $E\approx m$, the polarization behaves as $\Pi(E)\sim\frac{A}{E-m}$. Consequently, the resolvent behaves as a simple pole with residue 1-1:
    $R(E)\approx\frac{-1}{E-m}\implies\frac{1}{R(E)}\approx m-E$ (E.26)
  2. 2.
    Polarization Zeros: Between two physical poles, $\Pi(E)$ must cross zero. Near such a zero $E\approx E_{0}$, $\Pi(E)\sim C(E-E_{0})$. The resolvent behaves as a simple pole with residue +1+1:
    $R(E)\approx\frac{1}{E-E_{0}}\implies\frac{1}{R(E)}\approx E-E_{0}$ (E.27)
Therefore, when plotting the real part of the inverse resolvent $\text{Re}[1/R(E)]$ against energy, both physical masses and polarization zeros appear as zero-crossings. However, they are distinguished by the sign of the slope:
  • 1.
    Negative Slope ($\frac{d}{dE}<0$): Corresponds to a physical particle mass (Pole of $\Pi$).
  • 2.
    Positive Slope ($\frac{d}{dE}>0$): Corresponds to a zero of the polarization operator (Artifact).
Consequently, the meson mass spectrum is identified exclusively by the zero-crossings of $1/R(E)$ possessing a negative gradient.
Let us note in passing that this function $\Pi(E)$ in the confined planar QCD we are solving is a meromorphic function of $q^{2}=-E^{2}$ with poles on the negative axis with positive residues:
$\Pi(E)=\sum_{n}\frac{Z_{n}}{m_{n}^{2}-E^{2}}$ (E.28)
We also know that asymptotically, the residues tend to $Z_{n}\propto\partial_{n}(m_{n}^{2})$ which leads to logarithmic asymptotic behavior in the deep Euclidean limit (Asymptotic freedom):
$\Pi(E)\to\int dn\frac{Z_{n}}{m_{n}^{2}+q^{2}}$
$\propto\int_{\Lambda^{2}_{QCD}}^{\infty}d\sigma\frac{1}{q^{2}+\sigma}\propto\textit{ const }{}+\log\frac{q^{2}}{\Lambda^{2}_{QCD}}$ (E.29)
For our resolvent, these analytic properties mean that $1/R(E)$ is an odd entire function, with zeros along the real axis and no singularities in the complex plane.
E.8 Stabilizing CLE by Reweighting
The application of the Complex Langevin Equation (CLE) to theories with severe sign problems is often described as more art than science. Unlike the real Langevin equation, where the drift is derived from a real, growing at infinity action and convergence to the Boltzmann distribution is guaranteed by rigorous theorems, the CLE operates in a complexified phase space. There is no a priori guarantee that the stochastic process will converge to a finite limit, nor that it will sample the correct distribution. A common failure mode is the “excursion problem,” where the holomorphic variables wander into non-physical regions of the complex plane where the action is not holomorphic or simply diverges, causing the simulation to explode.
We observed precisely this instability when attempting a direct simulation of the standard effective action $S_{\text{eff}}$. As shown in our control tests, without modification, the magnitude of the observable $\mathcal{O}$ diverges exponentially (reaching magnitudes of $10^{40}$), rendering the extraction of physical data impossible. The drift forces derived solely from $S_{\text{eff}}$ are insufficient to confine the system to the relevant integration domain.
The reweighting method, introduced primarily to convert the search for poles into a search for zeros, provides a crucial secondary benefit: dynamical stabilization. By simulating the modified action $\tilde{S}=S_{\text{eff}}+\ln\mathcal{O}$, the drift force receives an additional contribution:
$K_{\text{drift}}=-\nabla S_{\text{eff}}-\frac{1}{\mathcal{O}}\nabla\mathcal{O}.$ (E.30)
This “logarithmic force” acts as a dynamical potential barrier. When the system approaches a region where $\mathcal{O}$ vanishes (a zero of the inverse ratio), the force becomes singular and repulsive, preventing the trajectory from crossing the zero but allowing it to explore the vicinity. Conversely, the modification of the measure appears to suppress the runaway excursions seen in the unmodified theory.
E.9 Numerical Integration: Second-Order SDE and Adaptive Braking
While the logarithmic reweighting ($\tilde{S}=S_{eff}+\log\mathcal{O}$) provides a continuous topological barrier against runaway trajectories, the discrete nature of numerical integration introduces a severe algorithmic vulnerability. In continuous Langevin time $\tau$, steepest descent flows cannot cross the meromorphic poles at $\mathcal{O}=0$. However, in a discrete simulation, the holomorphic drift force $K(\mathcal{V})=-\nabla\tilde{S}$ strictly diverges near the zero-locus. A naive first-order Euler-Maruyama discretization with a fixed time step $\Delta\tau$ would generate a massive coordinate displacement $\Delta\mathcal{V}=K(\mathcal{V})\Delta\tau$ in the vicinity of the pole. This would cause the discrete trajectory to violently “jump over” the topological barrier, catapulting the system directly into the runaway Stokes wedges or triggering arithmetic overflow.
To rigorously preserve the integrity of the complexified geometry and enforce the topological barriers during discrete evolution, our numerical implementation relies on two synergistic stabilization algorithms: a second-order Stochastic Differential Equation (SDE) integrator and an adaptive drift clipping mechanism.
First, we upgrade the integration scheme from the standard Euler method to the second-order Stochastic Heun (predictor-corrector) algorithm. For a generic complex state vector $\mathcal{V}$, the discrete update proceeds in two stages:
$\mathcal{V}_{pred}=\mathcal{V}(\tau)+\Delta\mathcal{V}_{1}+\eta(\tau)\sqrt{\Delta\tau}\qquad\text{(Predictor)}$ (E.31)
$\mathcal{V}(\tau+\Delta\tau)=$
$\mathcal{V}(\tau)+\frac{1}{2}\Big(\Delta\mathcal{V}_{1}+\Delta\mathcal{V}_{2}\Big)+\eta(\tau)\sqrt{\Delta\tau}\qquad\text{(Corrector)}$ (E.32)
where $\Delta\mathcal{V}_{1}=K(\mathcal{V}(\tau))\Delta\tau$ and $\Delta\mathcal{V}_{2}=K(\mathcal{V}_{pred})\Delta\tau$. The complexified Gaussian noise $\eta(\tau)$ is identical in both steps. By sampling the drift at both the current state and the projected future state, the predictor-corrector averaging intrinsically accounts for the severe local curvature of the effective action, providing vastly superior stability when navigating the highly non-linear valleys of the Lefschetz thimbles.
Second, to absolutely forbid discrete overshoots near the meromorphic poles and the bounding walls of the non-compact Liouville directions, we implement dynamic “braking” via adaptive drift clipping. At both the predictor and corrector stages, we compute the maximum proposed deterministic displacement, $\delta_{max}=\max|\Delta\mathcal{V}_{i}|$, across all degrees of freedom. If this proposed step exceeds a strict geometric safety threshold $\epsilon$ (in our production runs, $\epsilon=0.05$), we isotropically rescale the entire drift vector:
$\Delta\mathcal{V}_{i}\to\Delta\mathcal{V}_{i}\left(\frac{\epsilon}{\delta_{max}}\right)\qquad\text{for }\delta_{max}>\epsilon$ (E.33)
This operation is mathematically equivalent to an adaptive, localized reduction of the Langevin time step ($\Delta\tau\to\Delta\tau_{eff}$) during violent excursions.
Physically, this acts as an automatic braking system. In the flat, safe regions of the bulk phase space, the simulation explores efficiently with the full time step $\Delta\tau$. But the moment the complexified noise kicks the trajectory toward a singularity, the repulsive drift spikes, and the adaptive clipping safely throttles the velocity. The trajectory slows to a crawl, taking infinitesimally fine integration steps that smoothly trace the repulsive contours of the meromorphic pole. This ensures that the discrete trajectory gently reflects off the $1/\mathcal{O}$ topological barriers rather than numerically tearing through them, preserving the structural integrity of the complexified path integral over billions of consecutive updates.
E.10 Analyzing intermittent histories of the Complex Langevin Equation
The numerical evaluation of the twistor-string path integral via the Complex Langevin Equation (CLE) introduces profound statistical challenges. The physical observable required for the spectrum extraction, $1/\mathcal{O}$, is meromorphic. Consequently, during the stochastic evolution, the complexified trajectory occasionally wanders near the singularity at $\mathcal{O}=0$. These close encounters generate violent, intermittent numerical spikes—or excursions—in the micro-state history. Extracting the pristine physical signal from this complexified noise requires a rigorous pipeline of burn-in truncation, dynamic symmetry restoration, and Cauchy Principal Value error bounding.
Thermalization and the Measurement Criterion
Because the CLE simulation is initialized from an arbitrary complex configuration, the early micro-state history is dominated by a massive coordinate drift as the system relaxes toward the true vacuum. Including this non-equilibrium initial phase would severely pollute the physical averages.
To rigorously determine the exact step $N_{\text{cut}}$ where this “burn-in” phase ends, we utilize the raw imaginary trajectory, Im[$\imaginary[1/\mathcal{O}]$, as an exact equilibrium thermometer. The exact, uncomplexified path integral respects a global $Z\leftrightarrow Z^{*}$ parity. Consequently, the steady-state Fokker-Planck distribution of the CLE must eventually become symmetric with respect to the real axis. The strict criterion for thermalization is therefore the cessation of macroscopic secular drift in the imaginary component. As illustrated in the full raw history (Fig. 16, left), the measurement phase is initiated only after this trajectory stabilizes into a horizontal stochastic oscillation. We discard all data prior to this threshold (e.g., $N_{\text{cut}}=100000$).
Figure 16: Left: The raw intermittent history of the CLE trajectory. The vertical dashed line marks the strict measurement criterion ($N_{\text{cut}}$) where macroscopic secular drift ceases. All prior data is discarded. Right: The naive cumulative arithmetic mean. Post-thermalization excursions permanently offset the mean, demonstrating the fragility of unweighted averaging.
Dynamic Symmetry Restoration
At finite Langevin times, the stochastic trajectory can become temporarily trapped in one complex half-plane, leading to a delayed restoration of the $Z\leftrightarrow Z^{*}$ parity. We bypass this topological tunneling lag through Dynamic Symmetry Restoration. Because the theoretical ensemble is invariant under complex conjugation, we explicitly symmetrize the post-burn-in history. By adding the complex conjugate of every generated micro-state to the dataset, we artificially and instantly enforce the parity symmetry, projecting the stochastic fluctuations entirely onto the real physical slice (Fig. 17).
Figure 17: Kernel Density Estimation of Im[$\imaginary[1/\mathcal{O}]$ during the valid measurement phase. The bounded peak near zero confirms statistical equilibrium. Any residual skewness is instantly annihilated by the $Z\leftrightarrow Z^{*}$ double-history symmetrization.
The $1/|u|^{3}$ Singularity and the Principal Value
The physical core of the real distribution, Re[$\real[1/\mathcal{O}]$, is highly multimodal (Fig. 18, left), reflecting instanton-like tunneling between discrete topological sectors (Lefschetz thimbles). However, the excursions generate a severe statistical pathology.
[…]
$P(u)\approx\int p_{0}\,\delta\left(u-\frac{\cos\theta}{r}\right)r\,dr\,d\theta=\frac{\pi p_{0}}{2|u|^{3}}$ (E.34)
This exact $1/|u|^{3}$ power-law behavior dictates that while the theoretical mean is finite, the variance diverges logarithmically ($\int u^{2}/|u|^{3}du\to\infty$). Consequently, the standard empirical arithmetic mean is formally ill-defined in finite samples; it is continuously hijacked by infinite-variance sample noise (Fig. 16, right).
Because the leading asymptotic tail is exactly symmetric (P(u)=P(u)P(u)=P(-u)), the divergent variance contributions from the extreme left and right excursions mathematically cancel. Therefore, the physical expectation value of the amplitude must be rigorously defined via the Cauchy Principal Value (PV). To extract this regularized observable, we utilize a Symmetric Trimmed Mean (clipping the extreme $q=5\%$ of positive and negative outliers). By symmetrically shearing off the infinite-variance tails, it acts as the exact numerical implementation of the Principal Value, isolating the finite thermodynamic center of mass over the central $90\%$ of the physical bulk.
Figure 18: Upper: The multimodal physical core of Re[$\real[1/\mathcal{O}]$. The robust trimmed mean (red dashed line) isolates the true Principal Value. Lower: Log-scaled probability density demonstrating the severe asymptotic $1/|u|^{3}$ power-law tails (black dashed lines), which cause infinite variance and necessitate Cauchy Principal Value regularization.
Exact Analytic Error Bounding and Regression Weights
To rigorously bound the systematic statistical error introduced by discarding these numerical excursions, we exploit the analytic form of the asymptotic tails $P(u)=C/|u|^{3}$. For the right tail (values exceeding the 1q1-q quantile threshold $u_{R}$), the normalization constant is fixed by the trim fraction $q$:
$\int_{u_{R}}^{\infty}\frac{C_{R}}{u^{3}}du=\frac{C_{R}}{2u_{R}^{2}}=q\implies C_{R}=2qu_{R}^{2}$ (E.35)
The exact thermodynamic contribution of this infinite right tail evaluates to $\int_{u_{R}}^{\infty}u\left(2qu_{R}^{2}/u^{3}\right)du=2qu_{R}$. By identical logic, the left tail contribution evaluates to $2qu_{L}$. Thus, the exact reconstructed Principal Value mean is a parameter-free sum of the trimmed core ($\mu_{\text{core}}$) and the analytic tail boundaries:
Re[$\langle\real[1/\mathcal{O}]\rangle_{\text{total}}=(1-2q)\mu_{\text{core}}+2q(u_{R}+u_{L})$ (E.36)
The difference between the symmetric trimmed mean and this full analytic reconstruction, $\Delta_{\text{tail}}=2q|u_{R}+u_{L}-\mu_{\text{core}}|$, provides a strict mathematical measure of the finite-sample Principal Value asymmetry (Table 2).
Beyond bounding the mean, the exact $1/|u|^{3}$ geometry strictly determines the statistical variance of the Markov chain. This variance is required for the optimal inverse-variance pooling of independent cluster nodes and for weighting the polynomial regression used to extract the mass spectrum. Because the probability density decays as $1/|u|^{3}$, the formal variance diverges logarithmically with the maximum Langevin excursion cutoff $L$. Evaluating the second moment for the right tail yields:
$\sigma_{R}^{2}=\int_{u_{R}}^{L}u^{2}\left(\frac{2qu_{R}^{2}}{u^{3}}\right)du=$
$2qu_{R}^{2}\log\left(\frac{L}{u_{R}}\right)\to 2qu_{R}^{2}\log L$ (E.37)
Combining the identically regulated contribution from the left tail, the total empirical variance of the numerical history diverges logarithmically as:
$\sigma^{2}\simeq 2q(u_{R}^{2}+u_{L}^{2})\log L$ (E.38)
Crucially, because the maximum Langevin excursion extent $L$ acts as a universal scale factor across the parallel Monte Carlo ensemble, the formally divergent $\log L$ completely cancels out when constructing the normalized inverse-variance weights ($w_{i}\propto 1/\sigma_{i}^{2}$) for the polynomial regression. Therefore, the optimal statistical weight for each independent energy point is governed entirely by the parameter-free relative variance of its physical core:
$\sigma_{\text{rel}}^{2}=q(u_{R}^{2}+u_{L}^{2})$ (E.39)
By replacing heuristic error estimation with this exact geometric variance, the fitting procedure becomes statistically bulletproof. Any node that suffers a violent, parity-violating topological excursion will exhibit drastically widened boundaries ($u_{R}^{2}+u_{L}^{2}\to\infty$), naturally driving its statistical weight to exactly zero. This rigorously guarantees that the final extracted Regge trajectory is determined strictly by the thermalized, parity-preserving vacuum configurations of the Complex Langevin dynamics.
Parameter Sym. Value
Trimmed Core $\mu_{\text{core}}$ 0.210137-0.210137
Left Cut ($5\%$) $u_{L}$ 2.103169-2.103169
Right Cut ($95\%$) $u_{R}$ $1.230750$
Analytic PV Mean $\mu_{\text{tot}}$ 0.276365-0.276365
Asym. PV Error $\Delta_{\text{tail}}$ $\mathbf{\pm 0.066228}$
Table 2: Analytic bounds on finite-sample Cauchy Principal Value (PV) asymmetry, derived from exact $1/|u|^{3}$ Langevin excursions.
E.11 First computations of the meson spectrum
Empirically, we find that this reweighted CLE is numerically stable over extremely long integration times. In production runs extending to $10^{6}$ time steps per node per energy value (corresponding to a Langevin time of $\tau\sim 10^{4}$ ), the observable $1/\mathcal{O}$ remains finite and oscillates within a bounded range (typically $\in[-20,20]$), with no sign of the divergences characterizing the unweighted dynamics. While there is no general theorem proving that logarithmic reweighting stabilizes all complex actions, in the specific context of the Twistor String, it effectively regularizes the stochastic evolution, allowing for high-precision spectroscopy. Our cumulative statistics on all nodes for all $480$ energy points per energy value is about one billion data points. The lack of instabilities in such a large dataset speaks for some form of statistical equilibrium.
The code implementing the above algorithm was written in Mathematica ® [52] and simulated on the IAS supercluster Typhon using a high-precision grid: $64$ angular nodes and $32$ radial nodes in the Gauss-Legendre computation of the Liouville energy integral. We employed a Langevin time step $dt=0.0005$ and averaged the inverse resolvent $1/R(m)$ over $48$ physical hours, following a thermalization period of $2.4$ hours. The number of time steps was approximately $T_{tot}=17M$, and the total Langevin time spent was $\tau=T_{tot}*dt\approx 8500$.
The computation was parallelized across $64$ nodes of the Typhon cluster. Each node, equipped with $24$ cores, ran an independent Langevin evolution initialized with a unique random seed. The energy spectrum was sampled over $N_{E}=10\times N_{\text{cores}}=240$ values equally spaced between $2.$ and $3.0$ in units of string tension $\sigma$. The calculation was computationally intensive, requiring two days of wall-clock time on each of the 64 nodes to achieve the necessary convergence.
The results of this simulation are presented in Figures 16, 18 and 15. The first plot displays the full history of real part of the inverse resolvent, $\text{Re}[1/R(m)]$, for a particular energy parameter $E=2.27197$. The second plot is the distribution of this real part over the measurement period (after first $100K$ steps). The third plot displays the roots of $1/R(E)$ with a negative slope in a form of the Regge trajectory $n$ vs $m_{n}^{2}$. As derived in Section E.7, the physical meson mass spectrum is identified by the zero-crossings of this function that possess a negative slope. There are two different branches of Regge trajectory, both close to linear (note that $n$ represents the internal excitation number, not the angular momentum). The label distinguishing these two branches corresponds to the internal quantum number of the $\bar{q}q$ states in our Confining Twistor theory. We identified these two branches by the information clustering method described in Section 16.
The simulation demonstrates certain stability:
  • 1.
    Robustness: The process remained bounded without "runaways" or numerical explosions, confirming the efficacy of the holographic bulk forces in stabilizing the non-compact gauge variables.
  • 2.
    Precision: The histograms are smooth and have symmetric power tails, matching theoretical prediction $1/|u|^{3}$
  • 3.
    Spectrum: Information clustering of negative crossings by information produced two Regge trajectories, approximately linear with the same slope.
While the current simulation parameters are sufficient to establish a proof of concept, we acknowledge that a detailed spectroscopic analysis—particularly to resolve the fine structure of high-mass resonances—requires higher resolution and significantly larger statistics. However, the existing code [52] is fully capable of addressing this challenge without modification. The implementation is highly efficient and parallelized, maintaining numerical stability even at the high grid densities required for Gauss-Legendre integration, angular FFTs, and Lagrange interpolation.
Appendix F Canonical Twistors and the Minkowski Helicoid Metric
In this appendix, we explicitly derive the macroscopic worldsheet metric of the rotating string directly from the microscopic twistor fields. We demonstrate how the canonical twistor parameterization enforces a dimensional reduction to a 3D hyperplane, and how the double Wick rotation analytically continues the Euclidean twistor metric into the physical Minkowski Nambu-Goto metric. We then pedantically evaluate the resulting bulk action integral.
F.1 Canonical Twistors and Dimensional Reduction
The canonical Helicoid is generated by twistors endowed with fractional powers of the worldsheet coordinate $z$. This fractional scaling is a spin-structure gauge choice that explicitly enforces the Neveu-Schwarz anti-periodic boundary conditions for the worldsheet fermions as they complete a full rotation around the string center. We define the left and right twistors as:
$\lambda(z)=e^{\mathrm{i}\pi/4}z^{-1/2}\begin{pmatrix}1/z\\ 1\end{pmatrix}$,
$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ -z\end{pmatrix}$ (F.1)
Because $\lambda\propto z^{-1/2}$ and $\mu\propto z^{1/2}$, their product has no overall fractional branching. The resulting bosonic physical observables are strictly rational, making this configuration geometrically equivalent to a rational single-pole state.
The defining geometric characteristic of this canonical configuration is its behavior under the symmetric Pauli matrices. The spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$. Evaluating the symmetric bilinear for the first spatial component:
$f^{\prime}_{1}\propto\lambda^{T}\sigma_{1}\mu=\lambda_{1}\mu_{2}+\lambda_{2}\mu_{1}$
$=\mathrm{i}\left[\left(\frac{1}{z}\right)(-z)+(1)(1)\right]=\mathrm{i}(-1+1)=0$ (F.2)
Because the tangent vector component $f^{\prime}_{1}$ vanishes identically everywhere on the worldsheet, the target space coordinate $X_{1}$ is strictly constant. This exact algebraic constraint dynamically crushes the generic 4-dimensional minimal surface into a 3-dimensional hyperplane, forming the exact topological skeleton of the macroscopic 3D Helicoid.
F.2 The Euclidean Metric and Double Wick Rotation
To extract the physical string metric, we compute the induced conformal factor $\exp\left(2\rho\right)=|\lambda|^{2}|\mu|^{2}$. Restoring the dimensionful scale parameter $R$, the squared norms are:
$|\lambda|^{2}=\frac{R}{|z|}\left(\frac{1}{|z|^{2}}+1\right)=R\frac{1+|z|^{2}}{|z|^{3}},$
$|\mu|^{2}=R|z|\left(1+|z|^{2}\right)$ (F.3)
When multiplied, the fractional $|z|$ weights cancel perfectly:
$\exp\left(2\rho\right)=R^{2}\frac{(1+|z|^{2})^{2}}{|z|^{2}}=R^{2}\left(|z|+\frac{1}{|z|}\right)^{2}$ (F.4)
To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$. The flat integration measure transforms as $dzd\bar{z}/|z|^{2}=d\sigma_{E}^{2}+d\tau_{E}^{2}$. Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
$\exp\left(2\rho\right)=R^{2}\left(e^{\sigma_{E}}+e^{-\sigma_{E}}\right)^{2}=4R^{2}\cosh^{2}(\sigma_{E})$ (F.5)
Thus, the induced Euclidean worldsheet metric is exactly:
$ds^{2}_{E}=\exp\left(2\rho\right)\frac{dzd\bar{z}}{|z|^{2}}=4R^{2}\cosh^{2}(\sigma_{E})\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (F.6)
To evaluate the physical states, we analytically continue this metric into Minkowski spacetime via a double Wick rotation. The angular parameter $\tau_{E}$ serves as the target space time. We perform a Wick rotation on the spatial worldsheet coordinate into the physical light-cone regime:
$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (F.7)
Under this transformation, the hyperbolic cosine strictly rotates into the trigonometric cosine: $\cosh(\mathrm{i}\sigma_{M})=\cos(\sigma_{M})$. Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact Minkowski metric is:
$ds^{2}_{M}=4R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (F.8)
To explicitly verify that this describes the canonical rotating string, we change from the conformal spatial coordinate $\sigma_{M}$ to the physical target space radial coordinate $r$. For a rigidly rotating string, the fractional radius is $r=\sin(\sigma_{M})$. The differentials transform as:
$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (F.9)
Simultaneously, the time component of the metric incorporates the trigonometric identity $\cos^{2}(\sigma_{M})=1-\sin^{2}(\sigma_{M})=1-r^{2}$. Substituting these directly into the Minkowski metric, we obtain:
$ds^{2}_{M}=4R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (F.10)
This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})dt^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light.
F.3 Action Evaluation and the Lüscher Shift
Finally, we pedantically evaluate the total effective action over the fundamental domain of this Minkowski metric. The determinant of the conformal Minkowski metric $ds^{2}_{M}=4R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ is $g=16R^{4}\cos^{4}(\sigma_{M})$, yielding the exact proper area measure:
$\sqrt{-g}=4R^{2}\cos^{2}(\sigma_{M})$ (F.11)
The center of the rotating string lies at $r=0$ ($\sigma_{M}=0$), and the physical speed-of-light boundary is at $r=1$ ($\sigma_{M}=\pi/2$). For one full spatial period of the meson, the time coordinate wraps exactly one cycle: $\tau_{E}\in[0,2\pi]$. The exact classical Minkowski area integral is therefore:
Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}4R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
$=(2\pi)(4R^{2})\left(\frac{\pi}{4}\right)=2\pi^{2}R^{2}$ (F.12)
Simultaneously, the Liouville action dictates that the worldsheet carries the kinetic anomaly $\frac{1}{12\pi}|\nabla\rho|^{2}$. Because the kinetic anomaly density shares the exact same spatial singularity as the conformal area density at the twistor pole, it mathematically merges into the classical area term. As established in the main text, it acts strictly as a rigid, local shift to the effective string tension:
$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (F.13)
Multiplying the classical Minkowski area by this effective tension, we obtain the complete 1-cycle bulk action:
$S^{(1)}_{bulk}$ $=-\sigma_{\text{eff}}\times\text{Area}$
$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(2\pi^{2}R^{2}\right)$
$=-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$ (F.14)
This rigorous derivation explicitly demonstrates that the universally mandated $\pi/6$ Lüscher phase is identically generated by the exact geometric mapping of the twistor string to the Minkowski cylinder.

### v2

@@
 Abstract
-We develop a continuum framework for confining planar QCD ($N_{c}\to\infty$) by quantizing the Fermi-string (1981) degrees of freedom on the rigid Hodge-dual minimal surface constructed in Geometric QCD I.
-The internal Majorana fermions on the worldsheet provide the algebraic mechanism required by the Makeenko–Migdal loop equations: by the Pauli principle, they enforce cancellations of non-planar self-intersection contributions, leaving the planar factorization structure.
-The Liouville instability of the original random-surface formulation is avoided because the bulk geometry is rigid and fixed holographically by the boundary loop, with no summation over fluctuating worldsheet metrics.
-A key outcome is a formulation in momentum loop space, where coordinate-space cusp/contact singularities are integrated out, and the relevant quark-loop amplitudes admit a finite local limit.
-A central technical step is an explicit change of variables in the quark-loop phase-space path integral: after gauge fixing the reparametrization symmetry (Virasoro constraint), the reduced loop-space measure is parametrized by boundary twistors and an induced Jacobian, yielding a concrete twistor-string functional integral derived directly from the planar QCD amplitude rather than postulated as an abstract string model.
-In the local limit, the theory becomes a confining (analytic) twistor string: a boundary sigma model $S^{1}\to(S^{3}\times S^{3})/S^{1}$ coupled to a holographically determined Liouville field $\rho(z,\bar{z})=\log\!\bigl(|\lambda(z)||\mu(z)|\bigr)$ obtained by analytic continuation of the boundary twistor data (Hilbert/Plemelj reconstruction).
-The meson spectrum is encoded in a one-dimensional functional integral over the boundary twistor trajectory $\lambda(z),\mu(z),|z|=1$.
-While we evaluate this integral numerically using a reweighted Complex Langevin approach—successfully extracting approximately linear Regge trajectories as a proof of concept—our ultimate finding is analytical.
-By analyzing the monodromies of complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles inside the unit circle.
-The simplest sector with just one pole has the spectrum $m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$.
+We present the exact analytic solution of the Makeenko-Migdal loop equations, thereby solving planar QCD ($N_{c}\rightarrow\infty$) in the continuum limit.
+The solution is constructed by quantizing internal Majorana fermions (elves) on a rigid Hodge-dual minimal surface.
+We demonstrate that these worldsheet degrees of freedom provide the exact algebraic mechanism required to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing the planar factorization.
+In the local limit, the theory reduces to a confining analytic twistor string.
+By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles.
+The simplest sector with just one pole yields the exact linear Regge spectrum $m^{2}=\frac{\pi\sigma}{4}(n+1/12)$, matching the experimental Regge trajectories for light mesons ($\pi,K$ trajectories with $n=4J$ and $\rho$ trajectory with n=4J2n=4J-2) within 95% confidence intervals.
+The empirically correct intercepts $\alpha_{\pi}(0)=-1/48$ and $\alpha_{\rho}(0)=+23/48$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations.
+Ultimately, this exact geometric quantization explicitly realizes Witten’s Master Field as a critical classical trajectory in twistor space.
 1 Introduction
 The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the loop equations.
-At leading order in the $1/N_{c}$ expansion, the Makeenko–Migdal hierarchy exhibits the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
+At leading order in the $1/N_{c}$ expansion, the Makeenko-Migdal equations  [^1] Referred to hereafter as the MM equations.
+While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible.
+exhibit the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
 Historically, the main obstruction is the ultraviolet singular nature of the coordinate-space Wilson loop functional $W[C]$.
@@
 This leads to an effective “analytic twistor string” description: a boundary sigma-model (a coset built from two $S^{3}$ spinors modulo a local $U(1)$) coupled to a holographic Liouville field whose bulk values are determined by the unique holomorphic extension of the boundary twistors.
-Remark 1.1 (Equation solving vs. model building).
-The starting point of this work is not a proposed “string model of QCD,” but the planar Makeenko–Migdal loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD.
+Remark 1.1.
+Equation solving vs.
+model building The starting point of this work is not a proposed “string model of QCD,” but the planar MM loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD.
 The aim is constructive: to exhibit explicit variables and an explicit functional measure in which the loop equations admit a finite continuum solution and yield physical observables.
@@
 The culminating result of this paper is the exact reduction of the meson spectroscopy problem to a one-dimensional functional integral over boundary twistor data $(\lambda(z),\mu(z),|z|=1)$.
-While we initially evaluate this integral numerically via a reweighted Complex Langevin (CLE) approach—successfully extracting approximately linear Regge trajectories as a computational proof of concept—our ultimate finding is analytical.
-By evaluating the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory.
+By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory.
 The path integral exactly localizes at critical resonance energies where the Lefschetz thimbles become degenerate.
@@
 Planar QCD thus ceases to be a theory of stochastic quantum fluctuations; the extraction of the exact spectrum reduces to a deterministic, classical generalized eigenvalue problem.
-While we initially evaluate the one-dimensional functional integral over the boundary twistor trajectory numerically using a reweighted Complex Langevin approach—successfully extracting approximately linear Regge trajectories as a proof of concept—our ultimate finding in this paper is analytical.
-By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory.
 The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral.
@@
 For the simplest sector with just one pole, this yields the spectrum:
-$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (1.1)
-Remarkably, the universally correct Lüscher open-string intercept (+1/12+1/12) emerges directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations.
+$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (1.1)
+where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory.
+Remarkably, the intercepts 1/48,+23/48-1/48,+23/48 emerge directly from the 2D Liouville kinetic anomaly.
+In a profound sense, the culminating result of this construction is the explicit geometric realization of Witten’s original "Master Field" hypothesis for large-$N_{c}$ QCD.
+While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, we discover its exact representation in loop space.
+The true classical limit of Planar QCD is identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory and accessible via modern geometric quantization.
 Organization of the paper and navigation points.
-To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with four specific navigation points that address anticipated physical and computational questions.
+To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with five specific navigation points that address anticipated physical and computational questions.
 Section 2 summarizes the loop-equation zero modes and recalls why the Hodge-dual minimal surface provides the appropriate rigid geometric background.
@@
 Addressing (iii) the parametrization of the quark loop in phase space by boundary values of holomorphic twistors, Sec.
-11 details how the linear phase-space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
+11 details how the linear phase space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
+Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
 Section 15 discusses the semiclassical (WKB) limit and the emergence of linear Regge behavior.
-Section 14.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
-Then, we address (iv) the discrete approximation and numerical implementation of the path integral.
-Section 15 summarizes the CLE evaluation, while E provides exhaustive details on the CLE implementation and topological analysis.
-Section 16 details the unsupervised phase-information clustering used to isolate the spectral branches.
-Finally, we highlight (v) the exact analytic extraction of the mass spectrum.
-Section 17.1 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory.
-This last observation reveals the true geometric nature of Planar QCD: the main theoretical result of this work.
-Section 17.6 discusses the geometric origin of QCD, and Section 18 concludes.
-The remaining appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C.
+Finally, we highlight (iv) the exact analytic extraction of the mass spectrum.
+Section 16 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory.
+Section 17 then reframes this localization as an infinite-dimensional realization of Picard-Lefschetz theory, Resurgence, and Exact WKB analysis.
+We establish a strict structural correspondence between the twistor monodromies of Geometric QCD and the spectral networks of Seiberg-Witten theory, identifying the exact resonance condition as the twistor-string analogue of an Argyres-Douglas point.
+Section 18 concludes by synthesizing these results into the final geometric Master Field.
+The appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C and computations of the Minkowski action for our chiral helicoid minimal surface.
+Remark 1.2.
+Conceptual Demarcation: Bypassing 2D Quantum Gravity
+Because the mathematical vocabulary of this paper (strings, twistors, minimal surfaces) intersects with decades of literature on string theory, a reader may instinctively attempt to map this construction onto the conventional quantization of a fluctuating 2D worldsheet.
+To prevent fundamental conceptual misunderstandings, it is vital to explicitly state that we are not proposing a phenomenological string model, nor are we quantizing a fundamental string in the Polyakov sense.Consequently, the reader will not find integrations over fluctuating 2D bulk metrics, conformal ghost sectors, or integrations over ambiguous Teichmüller moduli spaces.
+Attempting to evaluate this work by the standard criteria of abstract 2D string modeling is a category error.Our absolute starting point is the rigorously defined, 1-dimensional phase space of Planar QCD: the momentum-loop representation of the MM equations, governed by the exact linear measure in phase space $\int\mathcal{D}C^{\prime}\mathcal{D}P$.
+The transition to "twistor space" is strictly a local, 1-dimensional algebraic change of variables on the boundary circle.
+The twistors do not dynamically fluctuate across a 2D bulk; they are localized strictly at the 1D boundary to parametrize the loop velocity, and are analytically continued inward solely as a tool of complex analysis.
+We are transforming the intractable functional geometry of a 1D gauge-theory phase space into a rigorously solvable problem of 1D complex analysis and linear algebra.
 2 Zero modes of Loop equation
@@
 Before we formulate the dynamics of the Fermi string on the minimal surface, we must clarify the physical origin of this rigid geometric background.
-The Makeenko-Migdal (MM) loop equations serve as the exact quantum equations of motion for Planar QCD.
+The MM (MM) loop equations serve as the exact quantum equations of motion for Planar QCD.
 Like any equations of motion, they possess an inherent ambiguity corresponding to the choice of the physical vacuum.
@@
 In the presence of the Virasoro constraint (3.21), the problem becomes a challenging one and has been investigated over the last two centuries by great mathematicians, starting with Riemann and Hilbert.
-We give a brief summary of modern state of this theory in Appendix B.
+We give a brief summary of modern state of this theory in B.
 In Part I [38], we compared the ordinary minimal surface in four dimensions with the Hodge-dual one.
@@
 $\Delta=\prod\Delta(windows).$ (6.27)
-This is clear from the original representation of $\Delta$ as an exponential of the sum over configurations of independent loops. This planar hierarchical set of loops is shown in Fig. 4 .
-Figure 4: The hierarchical set of loops, some touching, but never intersecting each other nor self-intersecting
+This is clear from the original representation of $\Delta$ as an exponential of the sum over configurations of independent loops.
+This planar hierarchical set of loops is shown in Fig.
+4, with the red line illustrating the time slice: the number of Elf pairs flying between quarks.
+Figure 4: The hierarchical set of loops, some touching, but never intersecting each other nor self-intersecting. The time slice (red line) represents some number of Elf pairs.
 7 The Minimal Surface and the Loop Equation
@@
 […]
+Remark 10.1.
+Phase Space Normalization and the Departure from the Brownian Gauge.
+To appreciate the mathematical control within this phase-space formulation, it is instructive to examine its exact canonical normalization.
+The functional measure is strictly normalized by the standard symplectic volume element $\mathcal{D}P\mathcal{D}C^{\prime}/(2\pi)$ per continuous degree of freedom.
+Consequently, in the absence of the interacting Wilson loop (the free-particle limit, $W[C]=1$), integrating over the loop coordinates $C$ natively yields a functional Dirac delta function, strictly enforcing the conservation of local momentum:
+$\int d^{4}k\int\frac{\mathcal{D}C^{\prime}\mathcal{D}P}{2\pi}\exp\left(\mathrm{i}\oint(P_{\mu}(\tau)-k_{\mu})\dot{C}_{\mu}d\tau\right)F[P]$
+$=\int d^{4}k\int\mathcal{D}P\,\delta[P(\cdot)-k]F[P]=\int d^{4}k\,F[k]$ (10.4)
+This demonstrates that the phase-space integral is rigorously well-defined from standard quantum mechanics, unique strictly modulo the infinite 1D diffeomorphism volume $\text{Vol}(\text{Diff}(S^{1}))$ associated with the simultaneous reparametrization of $P(\tau)$ and $C(\tau)$.Conventionally, this $\text{Diff}(S^{1})$ volume is factored out by imposing a proper-time gauge-fixing via an $\int(C^{\prime})^{2}d\tau$ kinetic term.
+This specific gauge choice inevitably generates the Feynman-Schwinger representation with its standard Brownian (Wiener) measure in loop space—a measure inherently dominated by everywhere-discontinuous fractal paths that artificially generate the notorious UV cusp singularities.
+In our framework, we recognize that the Brownian measure is just one possible choice of gauge fixing for the 1D diffeomorphisms.
+We will deliberately abandon it.
+Instead, as detailed in Section 11, we fix $\text{Diff}(S^{1})$ by imposing the Virasoro constraint $(f^{\prime})^{2}=0$ on the holomorphic twistors parametrizing the loop velocity.
+This specific gauge choice entirely avoids the fractal artifacts of the Brownian measure and perfectly locks the 1D phase-space kinematics of the quark loop to the 2D internal confining geometry of the minimal surface.
 The momentum loop amplitude $W[P]$ was introduced in our papers [32, 33, 34] and was recently used in the solution of the Navier-Stokes turbulence problem [35, 39].
@@
 Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a strictly complete functional basis.
-This guaranties that our formal expansion in terms of Magnus forms is exhaustive.
+This guarantees that our formal expansion in terms of Magnus forms is exhaustive.
 Through direct tensor contraction, the kinematic tensor $T^{\alpha\beta\gamma}{}_{\nu}$ implements the exact Dynkin projector [11] onto the nested double commutator $\hat{D}_{\alpha},[\hat{D}_{\alpha},\hat{D}_{\nu}]]$.
@@
 $P_{\nu}^{\prime}(\theta)\mathrm{tr}\left\{[\hat{X}_{\mu},\hat{X}_{\nu}]\hat{\mathcal{T}}\exp\left(\int_{\theta}^{\theta+2\pi}\hat{X}_{\alpha}P_{\alpha}^{\prime}\right)\right\}$ (10.17)
-By applying this exact derivative formula, the Makeenko–Migdal loop equation sheds its coordinate-space singularities and collapses into a finite, recursive combinatorial algebra.
+By applying this exact derivative formula, the MM loop equation sheds its coordinate-space singularities and collapses into a finite, recursive combinatorial algebra.
 On the left-hand side, the kinematic operator $Q_{\nu}$ acts via a shuffle product between the 3-index $T^{\alpha\beta\gamma}_{\nu}$ tensor and the lower-order $\mathcal{W}^{(n-3)}$ coefficients.
@@
 By the Free Moment-Cumulant Theorem [48], this single, order-independent operator contains an infinite tower of algebraically unconstrained parameters (the free planar cumulants $Q$).
-Because the free Cuntz algebra contains no finite trace identities, our Master Field Theorem guaranties that the trace of the path-ordered exponential is mathematically unrestricted.
+Because the free Cuntz algebra contains no finite trace identities, our Master Field Theorem guarantees that the trace of the path-ordered exponential is mathematically unrestricted.
 It exhaustively spans the continuous functional space of any closed loop without truncation.
@@
 In addition to this determinant, there is a Faddeev-Popov Jacobian arising from the gauge condition that fixes the residual local rescaling invariance of the twistor parametrization.
+Remark 11.1.
+Measure Transformation via Algebraic Jacobians.
+As emphasized in Remark 1.2, we are not evaluating an abstract functional measure over a fluctuating 2D surface.
+Therefore, the measure we derive here does not encounter the ambiguities of Teichmüller space.
+Our starting point is the rigorously defined, linear measure of the 1D quark-loop phase space: $\int\mathcal{D}C^{\prime}\mathcal{D}P$.
+The transition to twistor space is strictly a local (in $\theta$) change of variables, mapping the bounding curve velocity $C^{\prime}(\theta)\propto\text{Im}(z\lambda\sigma\mu)|_{z=e^{i\theta}}$ directly to the boundary twistors $\lambda,\mu$.
+Because the twistors are localized at the boundary, this transformation involves no functional anomalies.
+Once the zero modes are fixed, we derive the corresponding twistor measure via standard, explicitly calculable algebraic matrix Jacobians.
 11.1 Gauge Symmetry and the Need for Gauge Fixing
@@
 […]
+11.3 Euclidean Self-Intersection vs. Minkowski Helicity
+A common objection to minimal surfaces in the context of QCD strings arises from Euclidean intuition, where a surface of finite area bounded by a loop must satisfy strict embedding theorems (e.g., the Meeks-Yau theorem).
+In the Euclidean "rubber band" analogy, a surface cannot pass through itself without creating a physical self-intersection, which would typically be singular or unstable.
+Euclidean Space: The surface is constrained to a finite sheet. Any attempt to introduce a logarithmic branching point would force the sheet to intersect itself.
+Minkowski/Twistor Space: The surface behaves like a helicoid.
+As the trajectory winds around the twistor pole $\lambda(z)\propto 1/(z-z_{0})$ or $\mu(z)\propto 1/(z-z_{0})$, it moves to a "different sheet" in the complexified coordinate space.
+This is why the surface can have an infinite area in the Minkowski sense—it is effectively a Riemann surface with an infinite number of sheets: the coordinate map function $f_{\mu}(z)\propto\log(z-z_{0})$.
+This multiple-valued coordinate is what allows the poles to exist without the "self-intersection" catastrophe.
+The surface doesn’t intersect itself for the same reason a spiral staircase doesn’t intersect the floor below it: it has moved in an orthogonal dimension.
+The "infinite area" is simply the total area of all these sheets, and the monodromy is the "height" gained between each floor of the staircase, which we quantize as:
+$\Delta S_{eff}=\int_{sheet_{n}}^{sheet_{n+1}}\mathcal{L},d\tau=2\pi n$ (11.12)
+In this complexified geometry, the ’infinite area’ is not a divergence of the physical action, but the sum over the topological winding numbers that localizes the energy poles $E_{n}$ at the discrete values where the monodromy satisfies the phase coherence condition (see the Section 15).
 11.4 Twistor-induced metric in loop space
@@
 This integration yields the inverse matrix structure.
-$\int_{0}^{\infty}de\exp\left(-ed\theta(\mathrm{i}\gamma_{\mu}P_{\mu}+m_{q})\right)\propto\frac{1}{(\mathrm{i}\gamma_{\mu}P_{\mu}(\theta)+m_{q})}$ (12.12)
+[…]
 Taking the limit of zero quark mass and focusing on the local factor dependent on $P$, we compute the integral over the local momentum variable $q_{\mu}(\theta)=P_{\mu}(\theta)d\theta$:
@@
 The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates.
-Therefore, the Complex Langevin method [42] required to compute this path integral.
-13 The Linear Regge trajectories in the WKB limit
-13.1 The Helicoid Saddle Point
-In the semiclassical limit (large momentum or large tension $\sigma$), the path integral is dominated by the saddle point of the effective action.
-We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
-Consider a momentum distribution rotating in the $(1,2)$ plane with frequency $\omega$:
-$P_{1}(\tau)+\mathrm{i}P_{2}(\tau)=P_{0}e^{\mathrm{i}\omega\tau},\qquad P_{3}=P_{4}=0.$ (13.1)
-This source term in the action (11.6) drives the spinor fields.
-We search for a solution that respects this rotational symmetry.
-The natural ansatz corresponds to the helicoid minimal surface described in Appendix B.4.
-In the spinor formalism, with rescaling $w(z)=z^{-1/2}$ the Helicoid is generated by:
-$\lambda(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=e^{\mathrm{i}\pi/4}\begin{pmatrix}1\\ -z\end{pmatrix}$ (13.2)
-Substituting this into the null vector definition $f^{\prime}_{\mu}\sigma_{\mu}=\lambda\otimes\mu$, we find:
-$f^{\prime}_{1}+\mathrm{i}f^{\prime}_{2}\sim\frac{1}{z},\qquad f^{\prime}_{3}=f^{\prime}_{4}=0.$ (13.3)
-Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid.
-The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
-13.2 The Minkowski Helicoid
-The Minkowski Helicoid is parameterized by time $\tau\in[0,T]$ and radius $r\in[-R,R]$, see [43]. The time $T$ tends to infinity in the end.
-$X(\tau,r)=(\ r\cos\omega\tau,\ r\sin\omega\tau,0,\tau)$ (13.4)
-The induced Minkowski metric is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
-In Minkowski space we are not seeking to minimize the area in the geometric sense, as the area functional is not positive definite and the minimization problem is distinct.
-Instead, our criterion is algebraic: we require a solution to the Loop Equations for the Wilson loop.
-This demands the self-duality of the area derivative and the additivity of the area functional.
-We utilize the fact that the Euclidean Helicoid is a minimal surface for arbitrary angular velocity $\omega$.
-We obtain the Minkowski Helicoid via a double Wick rotation:
-$x_{4}\to\mathrm{i}t,\qquad\omega_{E}\to-\mathrm{i}\omega.$ (13.5)
-Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space.
-Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity).
-By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space.
-Thus, the Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
-Figure 14: The helicoid spanned by rotating $\bar{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, was discovered by Meusnier in 1785.
-Another important point is that the Helicoid boundary loop (double helix) is infinitely smooth, lacking any cusps or self-intersections.
-This means that no UV divergent renormalization factors could potentially affect our computation of the meson Regge trajectory.
-13.3 Semiclassical Motivation: Twisted Boundary Conditions
-Before solving the spectral equation, we justify the choice of the Helicoid geometry from the perspective of the semiclassical quark path integral.
-Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$.
-In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor333the perimeter renormalization terms in that factor are included in the physical quark mass.
-This is relevant to our WKB approximation in the Wilson loop at large loops:
-$\mathcal{Z}\sim\int\mathcal{D}C\,\exp\left(-\mathrm{i}m_{q}\oint_{C}ds-\mathrm{i}\sigma S_{\mathrm{Mink}}[\Sigma_{C}]\right)$ (13.6)
-To extract the spectrum of states with angular momentum $J$, we impose twisted boundary conditions.
-We require the quark positions at time $t=T$ to be a rotation of their positions at $t=0$ by an angle $\Theta=\omega T$ in the rotation plane $xy$:
-$(x+\mathrm{i}y)(T)=e^{\mathrm{i}\omega T}(x+\mathrm{i}y)(0)$ (13.7)
-The angular momentum $J$ arises as the thermodynamic conjugate to the twist frequency $\omega$.
-The ground state of this system is determined by the saddle point of the effective action.
-This requires the simultaneous minimization of the boundary length (quark term) and the surface area (gluon term).
-  • 1.
-    The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
-  • 2.
-    The minimal surface spanning two twisted helical boundaries is the helicoid.
-Thus, the Helicoid is not merely an ansatz; it is the unique geometric solution mandated by the kinematics of angular momentum in the semiclassical limit.
-The balance between the surface tension $\sigma$ (pulling the quarks inward) and the centrifugal barrier (pushing the twisted trajectories outward) determines the physical radius of the meson.
-13.4 Effective Lagrangean
-We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks.
-The confining factor introduces an effective Minkowski area term.
-We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m\int ds$.
-The total Lagrangian is:
-$L=L_{\text{helicoid}}+L_{\text{quarks}}$ (13.8)
-The Helicoid contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is:
-$L_{\text{helicoid}}$ $=-\sigma\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
-$=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (13.9)
-Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional.
-Adding the particle term $2m\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
-$L(v,\omega)=-\frac{\sigma}{\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m_{q}\sqrt{1-v^{2}}$ (13.10)
-It is worth noting that this mass term is proportional to the perimeter of the loop in Minkowski space.
-Therefore, the potential perturbative terms proportional to the loop perimeter are included in this physical quark mass $m_{q}$.
-We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m_{q}$, which , as the experiment tells us, is much smaller than the QCD mass scale, so that $m_{q}^{2}\ll\kappa$.
-We proceed to compute the conserved quantities via the Legendre transform. The angular momentum $J$ is the conjugate momentum to the rotation frequency:
-$J=\frac{\partial L}{\partial\omega}$ (13.11)
-Taking into account that $v$ depends on $\omega$ via the fixed radial extent ($v=\omega R$), the differentiation yields:
-$J=\frac{\sigma}{\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2mv^{2}}{\omega\sqrt{1-v^{2}}}$ (13.12)
-The Energy $E$ is defined by the Hamiltonian $E=\omega J-L$. Substituting the expressions for $J$ and $L$ and simplifying the algebra, we obtain:
-$E=\frac{2\sigma}{\omega}\arcsin v+\frac{2m_{q}}{\sqrt{1-v^{2}}}$ (13.13)
-13.5 Balance of quark forces and boundary condition
-The relationship between the endpoint velocity $v$ and the string length $R$ is not arbitrary; it is determined by the balance of forces at the string endpoint, which depends on the quark mass $m_{q}$.
-We derive this boundary condition by minimizing the action with respect to the geometric radius $R$ of the loop, keeping the angular velocity $\omega$ fixed.
-Recall that the total Lagrangian $L$ is composed of the bulk integral for the helicoid and the boundary terms for the quarks:
-$L=-2\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m_{q}\sqrt{1-\omega^{2}R^{2}}$ (13.14)
-The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$.
-Differentiating the integral using the Leibniz rule yields the tension force, while differentiating the boundary term yields the centrifugal force:
-$\frac{\partial L}{\partial R}=-2\sigma\sqrt{1-\omega^{2}R^{2}}-2m_{q}\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (13.15)
-Substituting $v=\omega R$ and computing the derivative of the mass term:
-$-2\sigma\sqrt{1-v^{2}}-2m_{q}\left(\frac{-\omega^{2}R}{\sqrt{1-\omega^{2}R^{2}}}\right)=0$ (13.16)
-Simplifying the expression, we obtain the exact force balance equation for finite mass:
-$\sigma\sqrt{1-v^{2}}=\frac{m_{q}v^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+m_{q}}$ (13.17)
-This equation balances the confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
-We now consider the limit $m_{q}\to 0$, and we find the boundary condition
-$v^{2}\to 1$ (13.18)
-Thus, as the quark mass tends to zero, the endpoint velocity $v$ must approach the speed of light ($v\to 1$).
-These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$.
-To recover the Regge trajectories, we consider the limit of massless quarks ($m_{q}\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$).
-In this limit, the mass terms vanish, and the system is governed purely by the confining geometry:
-$E\approx\frac{2\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{\omega^{2}}\frac{\pi}{2}$ (13.19)
-Eliminating $\omega$ between these two relations:
-$J=\frac{\sigma(\pi/2)}{(\pi\sigma/E)^{2}}\cdot\frac{\pi}{2}\implies J=\frac{E^{2}}{2\pi\sigma}$ (13.20)
-We thus recover the linear Regge trajectory:
-$E^{2}=2\pi\sigma J$ (13.21)
-This result confirms that the linear confinement arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space, without invoking independent string vibrations or ad-hoc constants.
+It is precisely this highly oscillatory nature of the path integral that mandates its exact analytical evaluation via Picard-Lefschetz theory and Catastrophe Theory.
 13 Twistor Holography and the Rigid Twistor String
@@
 This non-topological dependence is essential: it encodes the geometric rigidity of the minimal surface, leading to the emergence of the area law and confinement in the rigid twistor string.
-The general issues of the discretization of our theory and its direct numerical simulation (DNS) are discussed in D, where we present formulas and algorithms for the analytic extension of holomorphic twistor fields from a discrete set of boundary values.
-We attempted the direct numerical simulation of this path integral with oscillating weight using the Complex Langevin equation (CLE) on E.
-The results are encouraging but not convincing; perhaps some additional numerical algorithms are needed to overcome the instabilities of the CLE and obtain reliable results.
-15 Numerical Simulation of Confining String spectrum
-The analytic twistor-string formulation of the previous sections reduces the meson spectroscopy problem to the analytic structure of the quark two-point amplitude $\Pi(E)$ (after analytic continuation to Minkowski energy $E$).
-Since the corresponding boundary-twistor path integral is not localized, the evaluation of the spectrum becomes primarily computational.
-In this section, we summarize the numerical strategy and present preliminary proof-of-concept results.
-Full implementation details, discretization formulas, drift terms, and numerical diagnostics are given in E.
-Spectral observables and mass extraction
-The physical quark two-point amplitude $\Pi(E)$ is a meromorphic function of the energy, characterized by physical poles at the bound-state meson masses.
-However, directly searching for vertical asymptotes (poles) in a stochastic Monte Carlo simulation is a numerical trap: statistical variance near a singularity diverges uncontrollably, producing massive fat-tail excursions and destroying the Complex Langevin evolution.
-To completely bypass this fatal instability, we apply an analytic transformation that converts the dangerous meromorphic observable into a perfectly bounded, smooth entire function.
-Rather than evaluating $\Pi(E)$ directly, we consider the logarithmic derivative
-$R(E)=\frac{\Pi^{\prime}(E)}{\Pi(E)},$ (15.1)
-and reconstruct the spectrum from the zero-crossings of its exact inverse, $1/R(E)$.
-This inversion maps vertical asymptotes into well-behaved roots.
-Near a physical mass pole, $\Pi(E)\sim(E-m_{n})^{-1}$, the inverse resolvent becomes $1/R(E)\sim m_{n}-E$.
-The singularity is mathematically flattened into a smooth zero-crossing with a strictly negative slope.
-Conversely, near an artifact zero of the vacuum polarization ($\Pi(E)\sim E-E_{0}$), one obtains $1/R(E)\sim E-E_{0}$, yielding a zero-crossing with a strictly positive slope.
-Because $1/R(E)$ is an entire function with no singularities anywhere in the complex plane (as detailed in Appendix D.7), the Complex Langevin process can sample it safely.
-Therefore, the $\bar{q}q$ masses are unambiguously identified as the zero-crossings of $\text{Re}[1/R(E)]$ possessing a negative slope.
-Furthermore, because the true physical signal is bounded, we know a priori that any massive spikes in our simulation data are purely statistical fat-tail excursions, rigorously justifying aggressive outlier filtering (such as median pooling) without fear of clipping physical resonances.
-Reweighted Complex Langevin strategy (summary)
-The effective twistor-string functional integral has an oscillatory weight; at present, Complex Langevin evolution (CLE) provides a practical route to evaluating it.
-A key technical point is that pole-finding is numerically unstable in a stochastic computation.
-We convert the problem into a robust zero-finding problem by a reweighted formulation, summarized here and derived in Appendix D.
-Let $S_{\rm eff}$ denote the effective action for the discretized boundary-twistor variables at fixed $E$, and define the energy-derivative observable
-$O\;\equiv\;\frac{\partial S_{\rm eff}}{\partial E}\,.$ (15.2)
-Then one has the exact identity
-$\frac{1}{R(E)}\;=\;\frac{\Pi(E)}{\Pi^{\prime}(E)}\;=\;\frac{\int\exp\left(S_{\rm eff}\right)}{\int O\,\exp\left(S_{\rm eff}\right)}\,.$ (15.3)
-Introducing the modified (reweighted) action
-$\widetilde{S}\;=\;S_{\rm eff}+\log O\,$, (15.4)
-this becomes
-$\frac{1}{R(E)}\;=\;\left\langle\frac{1}{O}\right\rangle_{\widetilde{S}}\,$. (15.5)
+Because the path integral over the twistor target space is highly oscillatory, one might assume that extracting the physical spectrum requires massive stochastic lattice simulations.
+However, by analyzing the holomorphic structure of the effective action, we find that the discrete mass spectrum can be extracted exactly and analytically.
+Planar QCD ceases to be a theory of stochastic quantum fluctuations; it is a theory of pure classical complex geometry.
+14 Analytic continuation of Twistor String Path integral
+The Euclidean path integral of Twistor String theory, as formulated so far, is similar to a partition function of a statistical system with positive definite energy , given by the Liouville bulk integral.
+There are, however, some pre-exponential factors that violate this statistical interpretation.
+First, there is a "vertex operator" factor
 […]
-Topological stabilization and non-compact excursions
-There is a deep geometric reason for this stabilization, naturally understood through the lens of Picard-Lefschetz theory, introduced in QFT by Witten [44], and successfully used in recent QCD simulations [45, 46].
-In the complexified phase space, the exact path integral is formally defined on a middle-dimensional integration contour composed of steepest descent manifolds (Lefschetz thimbles).
-While the deterministic drift of the Complex Langevin Equation naturally targets these thimbles, stochastic noise inevitably kicks the trajectory off these stable manifolds and into adjacent “Stokes wedges,” where the action diverges and the Markov chain suffers runaway excursions.
-To counteract this, modern lattice simulations typically rely on computationally expensive algorithmic interventions.
-As detailed in the review [46], standard approaches involve explicitly deforming the integration contour via holomorphic gradient flow (which incurs a prohibitive $\mathcal{O}(V^{3})$ Jacobian cost), utilizing machine-learning-optimized manifolds, or applying ad-hoc dynamic stabilization and gauge cooling to actively force the system to remain near the thimbles.
-Our logarithmic reweighting, however, achieves this stabilization natively for the observable.
-By shifting the effective action ($\tilde{S}=S_{eff}+\log\mathcal{O}$), we fundamentally alter the topology of the complexified phase space.
-The associated CLE drift force acquires essential meromorphic poles exactly at the zero-locus of the observable ($\mathcal{O}=0$):
-$-\nabla\tilde{S}=-\nabla S_{eff}-\frac{\nabla\mathcal{O}}{\mathcal{O}}$ (15.6)
-Crucially, in the Picard-Lefschetz framework, steepest descent flows cannot cross meromorphic poles.
-By explicitly injecting these poles into the drift dynamics, we puncture the complex manifold.
-The singular repulsive force acts as an absolute topological obstacle that dynamically corrals the stochastic Markov chain.
-It safely traps the trajectory within the basin of the physical thimbles and mathematically forbids it from escaping into the runaway regions of the complex plane associated with the zeros of $\mathcal{O}$.
-This organically enforces dynamic stabilization without the need for computationally devastating Jacobian evaluations.
-The intermittent excursions that do occur in our simulation are not numerical instabilities, but rather the exact analytic consequence of the complexified noise exploring the vicinity of this meromorphic pole.
-As shown in the smooth histogram of the micro-states (see Figure 18), the probability density perfectly matches the analytically predicted $1/|u|^{3}$ power-law fat tails (derived in detail in Appendix D).
-This exact geometric matching confirms that the finite-sample Cauchy Principal Value is perfectly stable, and that the CLE is genuinely exploring the correct complexified geometry dictated by the topological barrier.
+Second, there is a trace factor (12.30).
+Both of these factors are not positive, and they oscillate as functionals of our dynamical variables: i.e., the complex curve $\lambda(z),\mu(z)$ in the Twistor space $\mathcal{C}P^{3}$.
+These oscillations become even more obvious when the path integral is analytically continued in Minkowski space, where the effective action is imaginary, and the exponential decay of the statistical weight becomes a quantum oscillation of the Feynman-Dirac weight $\exp\left(\mathrm{i}S/\bar{h}\right)$.
+In ordinary quantum mechanics, this phenomenon is well known: the path integral, when formulated in terms of physical, real trajectories in phase space, is exponentially small as a result of the interference of the classical trajectories in the Feynman path integral.
+However, when reformulated as a complex path integral by analytic continuation, it is dominated by complex saddle points, where there is no interference along the steepest descent path.
+The spectrum of certain quantum mechanical systems can be exactly computed by finding the dominant complex trajectory winding around the complex branch point of the analytically continued effective Action.
+This is a mechanism for the conversion of the Bohr-Sommerfeld quantization condition from the WKB approximation to an exact equation for the spectrum [6, 19, 9].
+We are looking for a similar phenomenon in our Twistor String theory, and the first step towards such a solution is the analytic continuation of the path integral into the complex plane in parameter space.
+14.1 Complexification of the parameter space
+To ensure the strict analyticity of the path integral, we must remove complex conjugation from the effective action and its parameter space.
+We achieve this by complexifying the target manifold and analytically continuing the action integral itself.
+This ensures that the drift forces are partial derivatives of a globally holomorphic function.
+Then the path integral can be regarded as a complex integral in multi-dimensional space of holomorphic functions in Twistor space.
+We replace the original real and Hermitian variables with a set of independent holomorphic variables.
+The angular coordinates $\alpha_{k}$ are mapped to the complex plane via $z_{1}=e^{i\alpha_{1}},\;z_{2}=e^{i\alpha_{2}}$.
+The spinor fields and their conjugates are treated as independent complex vectors.
+The full set of dynamical variables for the discretized loop is:
+$\mathcal{V}=\{\{z_{1},z_{2}\},\{\rho_{k},\xi_{k},\eta_{k},\tilde{\xi}_{k},\tilde{\eta}_{k},\gamma_{1,k},\gamma_{2,k}\}_{k=1}^{N}\}.$ (14.1)
+Here, $\tilde{\xi}$ and $\tilde{\eta}$ replace $\bar{\xi}$ and $\bar{\eta}$ as independent variables.
+The geometric constraints are enforced by the complex Lagrange multipliers $\gamma_{1,k}$ and $\gamma_{2,k}$, leading to the holomorphic constraint terms in the effective action:
+$S_{\text{constr}}=i\sum_{k}\left[\gamma_{1,k}(\tilde{\xi}_{k}\xi_{k}-1)+\gamma_{2,k}(\tilde{\eta}_{k}\eta_{k}-1)\right].$ (14.2)
+14.2 Spectral Interpolation of Fields
+To evaluate the action integrals analytically, we construct global holomorphic polynomials that interpolate the values at the vertices $z_{k}$.
+14.2.1 Spinor Fields
+Let $\Lambda^{(k)}=e^{\rho_{k}/2}(\xi_{k},\eta_{k})$ denote the spinor data $\Lambda(z)=\{\lambda(z),\mu(z)\}$ at vertex $k$. The Lagrange interpolation polynomial on the spectral unit circle is:
+$\Lambda(z)=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}K\left(ze^{-\mathrm{i}\theta_{k}}\right);$ (14.3)
+$K(z)=\frac{1-z^{N}}{1-z};$ (14.4)
+Using the Fast Fourier Transform (FFT), we extract the spectral coefficients $L_{n},M_{n}$ (and $\tilde{L}_{n},\tilde{M}_{n}$ from the independent $\tilde{\xi},\tilde{\eta}$ variables) such that:
+$\lambda(z)=\sum_{n=0}^{N-1}L_{n}z^{n},\qquad\tilde{\lambda}(z)=\sum_{n=0}^{N-1}\tilde{L}_{n}z^{n}.$ (14.5)
+$\{L_{n},M_{n}\}=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}e^{-\mathrm{i}\theta_{k}}$ (14.6)
+Remark 14.1.
+Reduction to Pole Moduli and the Algebraic Jacobian
+While the path integral is formulated using the discrete boundary values at the roots of unity $\theta_{k}=2\pi k/N$, the physical saddle points and their monodromies are governed by the meromorphic structure of the twistors—specifically, the positions $z_{i}$ and residues $R_{i}$ of their poles inside the unit disk.The transformation of the integration measure from the boundary values to these pole moduli is exact and purely algebraic.
+For a twistor field parameterized by a finite number of poles, its evaluation on the boundary circle $z=e^{i\theta}$ yields a geometric series:
+$\lambda(e^{i\theta})=\sum_{i}\frac{R_{i}}{e^{\mathrm{i}\theta}-z_{i}}=-\sum_{i}\frac{R_{i}}{z_{i}}\sum_{n=0}^{\infty}z_{i}^{-n}e^{\mathrm{i}n\theta}$ (14.7)
+Comparing this to our spectral expansion, the Fourier coefficients $L_{n}$ are exactly determined as algebraic superpositions of the pole parameters:
+$L_{n}=-\sum_{i}\frac{R_{i}}{z_{i}^{n+1}}$ (14.8)
+Consequently, transforming the functional measure from the $N$ discrete boundary values to the physical moduli $(R_{i},z_{i})$ involves exactly two finite, calculable matrix Jacobians:1.
+The constant Jacobian of the Discrete Fourier Transform matrix (the Fourier matrix $e^{2\pi\mathrm{i}kn/N}$), which maps the boundary values to the Fourier coefficients $L_{n}$.2.
+The purely algebraic Jacobian mapping the Fourier coefficients to the pole parameters $(R_{i},z_{i})$.We do not explicitly compute these algebraic Jacobians here, because they contribute strictly to the pre-exponential fluctuation determinant of the path integral.
+For the extraction of the exact geometric mass spectrum, the relevant physics is governed entirely by the stationary phase (the saddle-point equations) and the topological monodromies of the effective action.
+As we will demonstrate, the S-matrix poles emerge from the constructive interference of the geometric series over the topological winding number $k$.
+In this resonant limit, the multi-valued action scales as $k\Delta S_{eff}$.
+Consequently, the transverse quantum fluctuations around the saddle point are dynamically damped by a factor of $1/k$.
+This perfectly crushes the pre-exponential factor, rendering the explicit form of the Jacobian irrelevant to the exact localization of the S-matrix poles.
+However, the explicit existence of this algebraic mapping mathematically guarantees that our twistor measure is strictly finite and constructible, firmly rooting the evaluation of the path integral in standard finite-dimensional complex analysis rather than abstract string-theoretic functional geometry.
+14.2.2 Liouville Field
+The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an exact formula
+$\rho(z,\bar{z})=\frac{\log\tilde{\lambda}(\bar{z})\lambda(z)+\log\tilde{\mu}(\bar{z})\mu(z)}{2}$ (14.9)
+where $\lambda(z),\mu(z)$} is extended by interpolation (14.3).
+The derivatives of the above kernel $K^{\prime}(z)$ involved in the Liouville Lagrangian with this $\rho$ were computed in the previous sections.
+These expressions provide the values of the fields and derivatives at any complex $z$, not just at the vertices.
+These expressions also depend upon complex parameters in $\mathcal{V}$ through the Fourier coefficients $\Lambda(k),\tilde{\Lambda}(k)$.
+This dependence is exponential for $\rho_{k}$ and polynomial for the rest of variables in $\mathcal{V}$, which allows for required analytic continuation of the CLE.
+14.3 The Boundary Terms in the action
 […]
-To prevent the system from exploring these unphysical infinite-volume regimes, we currently impose explicit numerical clipping (hard bounding walls) on the twistor variables.
-While this pragmatic regularization successfully bounds the metric, there is no a priori mathematical guaranty that the CLE trajectory will not systematically drift outward and continuously collide with these artificial walls, which would artificially truncate the sampled distribution.
-Remark 15.1.
-It should be noted that the use of explicit numerical clipping (hard bounding walls) to tame the non-compact Liouville directions technically violates the strict holomorphy required for the formal proof of correctness of the Complex Langevin Equation, as it introduces non-analytic boundaries into the Fokker–Planck evolution.
-A theoretically cleaner approach would involve introducing a globally holomorphic confining potential to the effective action that generates a smooth restoring force.
-For example, one could augment the action to produce a holomorphic drift force such as $K_{\text{wall}}\propto-\tanh(c(\rho-\rho_{\max}))$.
-Because the hyperbolic tangent is a meromorphic function of the complex variable $\rho$, this force smoothly pushes the trajectory back toward the physical region $\operatorname{Re}(\rho)<\rho_{\max}$ at large distances without breaking the analyticity of the Langevin evolution.
-Implementing such holomorphic barriers represents a natural analytical refinement for future high-precision simulations, safely taming the infinite-volume runaway channels while preserving the exact geometric justification of the CLE.
-To explicitly defend our simulation against accidental discrete jumps over the topological barriers, we replaced the naive Euler integration with a second-order Stochastic Heun (predictor-corrector) SDE solver.
-Furthermore, we implemented an adaptive time-step algorithm designed to dynamically brake the CLE evolution upon approaching singular regions.
-Together, these algorithmic safeguards yield highly stable trajectories that remain predominantly confined to the physical integration cycles (Lefschetz thimbles), consistently producing finite expectation values with rigorously controlled absolute errors.
+$G_{p}(z_{1},z_{2})\equiv\int_{z_{1}}^{z_{2}}z^{p}\frac{dz}{iz}=\frac{z_{2}^{p}-z_{1}^{p}}{ip}.$ (14.10)
+With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is strictly positive, ensuring there are no singularities in the complex plane.
+The contribution to the action of external momentum $q$ injected into the quark loop from the arc between $z_{1}$ and $z_{2}$ is:
+$S_{q}=-\mathrm{i}q_{\alpha}$
+$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}\sigma_{\alpha}M_{m})-(\tilde{L}_{n}\sigma^{\dagger}_{\alpha}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (14.11)
+Then, there are terms related to the constraints on the $\xi,\eta$ variables. These terms are just discrete sums
+$S_{\gamma}=\mathrm{i}\sum_{k=0}^{N-1}\gamma_{1,k}\left(\tilde{\xi}_{k}\xi_{k}-1\right)+\gamma_{2,k}\left(\tilde{\eta}_{k}\eta_{k}-1\right)$ (14.12)
+In addition, there are terms related to the trace (12.30). Our complexification reduces this trace to
 […]
-Discretization and parallel ensemble
-We discretize the boundary circle by a spectral grid (angular nodes) and evaluate the bulk (holographic) Liouville contribution by Gauss–Legendre radial quadrature, leading to a finite-dimensional holomorphic system for the complexified boundary variables evolved by CLE.
-To reduce sensitivity to non-Gaussian (fat-tailed) fluctuations and to monitor reproducibility, we run an ensemble of independent replicas on the Typhon cluster.
-In the current production setting, we use 64 nodes (independent seeds) per energy value and pool the results across nodes using robust estimators (median central value and absolute-deviation-based error estimates).
-The precise implementation and diagnostics are documented in Appendix D.
-16 Topological Decomposition of the Meson Spectrum via Information Clustering
-16.1 Lefschetz Thimbles and Multiple Spectral Branches
-In the planar limit, the exact solution of the Makeenko–Migdal loop equation involves evaluating a path integral over a highly complex configuration space.
-The effective action landscape of this theory is not dominated by a single, featureless ground state, but rather by a multitude of complex saddle points, or Lefschetz thimbles.
-Physically, these distinct topological sectors correspond to the multiple branches of the meson spectrum.
-In our numerical approach, we explore this landscape using stochastic Langevin dynamics.
-Because each independent simulation is initialized from a completely random configuration, the system naturally falls into the basin of attraction of one of these specific thimbles as it relaxes.
-16.2 Topological Trapping and the Failure of Naive Pooling
-Due to the presence of immense topological (or simply numerical) action barriers in the complexified space, a given simulation history generally remains confined to its respective thimble for the duration of the run.
-This dynamical trapping has profound implications for data analysis.
-If one were to follow standard numerical practice and naively pool all independent stochastic histories at a given energy $E$ to compute a global ensemble average, the procedure would physically mix distinct topological sectors.
-Consequently, this blind averaging does not yield a single, smooth resolvent curve for $1/\mathcal{R}(E)$; rather, it produces a chaotic, smeared superposition of states.
-The raw numerical data are actually a collection of distinct histories confined to various thimbles, each tracing out its own unbroken $1/\mathcal{R}(E)$ curve separated by topological barriers.
-16.3 Phase-Information as a Topological Fingerprint
-To extract the true physical spectrum, we must mathematically untangle these mixed histories and group them by the specific saddle point they occupy.
-As established in our analysis of the Complex Langevin dynamics in E, the Cartesian observables $\mathrm{Re}[1/\mathcal{O}]$ and $\mathrm{Im}[1/\mathcal{O}]$ suffer from severe infinite-variance Cauchy tails ($1/|u|^{3}$), which generate massive radial excursions.
-Consequently, attempting to compare the Cartesian coordinates forces one to arbitrarily trim and shift the central moments to prevent the variance from dominating the signal.
-We can completely bypass this pathological radial variance by projecting the data onto the complex phase angle $\alpha=\arg(\mathcal{O})$.
-This maps the entire complexified phase space onto a finite, compact support $\pi,\pi]$, safely absorbing infinite radial fluctuations without discarding a single data point.
-The probability density function of this phase angle is intrinsically scale-invariant and serves as an exact, bounded topological fingerprint of the underlying Lefschetz thimble.
-To compare these fingerprints, we compute the symmetric Kullback–Leibler (KL) divergence (the Jensen–Shannon distance) between their finite-support phase distributions.
-This strict information-theoretic metric allows us to rigorously identify pairs of histories with identical quantum geometries, entirely independent of their macroscopic radial drift.
-16.4 Isolating True Mass Poles
-This phase-matching provides a precise, unsupervised mechanism for identifying true meson masses.
-A physical mass corresponds to a strictly downward zero-crossing of the resolvent, where the sequence transitions from $\mathrm{Re}[1/\mathcal{R}(E_{i})]>0$ to $\mathrm{Re}[1/\mathcal{R}(E_{i+1})]<0$.
-By examining pairs of runs at neighboring energy points that bracket the zero axis and demanding that they possess the absolute minimum symmetric KL distance between their phase distributions, we effectively filter out random topological jumps and numerical noise.
-This strict requirement of phase-geometry invariance unequivocally identifies the true crossing of the $1/\mathcal{R}(E)$ curve for a particular topological branch.
-16.5 Reconstructing the Spectrum via Minimax Clustering
-Having isolated a pool of high-confidence zero-crossings, we reconstruct the global spectrum of the different branches.
-To achieve this, we cluster the crossing events based on a strict minimax distance criterion.
-The topological distance between any two distinct crossing events—say, a pair $(A,B)$ and a pair (C,D)(C,D)—is defined as the maximum of all four cross-comparisons:
-$\Delta=\max\left(d_{JS}(A,C),d_{JS}(A,D),d_{JS}(B,C),d_{JS}(B,D)\right)$ (16.1)
-where $d_{JS}$ is the symmetric Kullback–Leibler distance (Jensen–Shannon divergence) between the phase distributions.
-This mathematically merciless constraint guaranties that a pole is only clustered into a specific sequence if the quantum phase geometry, both before and after the crossing, is practically identical to the rest of the cluster.
-16.6 First results of numerical solution of CLE
-The CLE was coded into a Mathematica ® package, tested on the laptop and desktop computers, and then run on the Typhon IAS cluster for 48 hours on 64 nodes in parallel, with 24 kernels working in parallel on each node.
-The code and the results are published on Wolfram cloud, and results are stored on the authors’ Google Drive with open access (see references in the Data section in the end of this paper).
-Using these Mathematica ® files, one can verify and reproduce our results at a larger scale with higher precision.
-Here is the summary of the results (all technical details are described in E).
-This unsupervised information-clustering natively reconstructs the spectrum into distinct branches, mapping out different Regge trajectories.
-Crucially, these trajectories are strictly labeled by an internal quantum number inherent to the Twistor String path integral complex saddle points, serving as a replacement for the “daughter trajectories” of string models.
-Remark 16.1.
-Note that all the usual quantum numbers are identical for these two branches: this is a scalar $\bar{q}q$ channel, no spin, no flavor quantum numbers specified.
-In the future, more detailed simulations, spin, parity, and flavor $SU(3)$ representation will be specified, but this internal quantum number will exist on top of other quantum numbers, in the same way as internal excitations of the string (vibrational modes) lead to daughter trajectories after quantization.
-Applying this algorithm to our unvarnished stochastic phase data successfully isolated two continuous principal Regge trajectories spanning $14$ and $15$ consecutive negative crossings.
-The negative crossing of zero between two consecutive points of energy on a grid was identified with accuracy $\Delta E=\pm 1/480\approx\pm 0.002$.
-Both branches exhibit non-linearity at low energies—reflecting the complex boundary dynamics and mass gaps of the true physical ground state—before transitioning into linear Regge trajectories at higher mass squared.
-By relying purely on the phase-information geometry, this topological extraction is designed to reveal the true spectrum of Planar QCD.
-Figure 15 shows preliminary cluster-scale results for the extracted trajectory $n$ vs.
-$m_{n}^{2}$.
-These plots are intended as a proof of concept: the crossings of zero were selected by information clustering of the phase distributions we described above.
-The precision of higher excitations and the control of systematic effects (step size, resolution, and CLE correctness diagnostics) will be improved in larger-scale runs; the present computation establishes feasibility and provides a concrete computational pipeline for continuum planar-QCD spectroscopy in this framework.
-Figure 15: Plot of $n$ vs $m_{n}^{2}$ (in units of the string tension $\sigma$) as a result of Complex Langevin Simulation of quark two point function in the Geometric QCD Theory.
-These two Regge trajectories were identified by information clustering of phase distributions of $1/O$ for $240$ different energy points between $2.$ and $3.$ on $64$ independent CLE simulations with different initial values.
-The optimal clustering identified a pair of Regge trajectories in this interval of $m^{2}$
+[…]
+$\mathcal{P}(\theta)=2\mathrm{i}\left[z\eta\tilde{\xi}^{T}-\bar{z}\xi\eta^{T}\right]_{z=e^{\mathrm{i}\theta}}$ (14.15)
+Finally, our reweighting adds one more complex term to the action,
+$S_{O}=\log\mathcal{O};$ (14.16)
+$\mathcal{O}=\frac{\partial S_{q}}{\partial q_{4}}=$
+$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}M_{m})-(\tilde{L}_{n}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (14.17)
+All the above expressions are purely holomorphic functions of all the variables in $\mathcal{V}$.
+14.4 Stochastic Quantization and Complex Langevin Dynamics
+The exact mathematical evaluation of this highly oscillatory complexified path integral is rigorously achieved via Stochastic Quantization.
+The quantum statistical evolution of the twistor string is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{total}=S_{q}+S_{\gamma}+S_{T}+S_{\rho}+S_{\mathcal{O}}$.
+We emphasize that the CLE is utilized here not as a numerical simulation technique, but as an exact continuous mathematical reformulation of the quantum path integral.
+  1. 1.
+    Coordinate Evolution ($z_{k}$): The drift force on the boundary points $z_{k}$ arises from the boundary terms of the $G$ functions in $S_{q}$ and the integration limits in $S_{bulk}$:
+    $\frac{dz_{k}}{d\tau}=-z_{k}^{2}\frac{\partial S_{total}}{\partial z_{k}}+\mathrm{i}z_{k}\nu_{z_{k}}$ (14.18)
+    (The factor $z_{k}^{2}$ arises from the metric of the map $\alpha=-\mathrm{i}\log z$, and $\nu$ represents complex Gaussian white noise).
+  2. 2.
+    Spinor Evolution ($\xi_{k},\tilde{\xi}_{k}$, etc.): The drift forces are computed via the chain rule through the spectral coefficients $L_{n}$. For example:
+    $\frac{d\xi_{k}}{d\tau}=-\frac{\partial S_{total}}{\partial\xi_{k}}-\mathrm{i}\gamma_{1,k}\tilde{\xi}_{k}+\nu_{\xi_{k}}$ (14.19)
+    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables representing the exact quantum fluctuations of the twistor fields.
+  3. 3.
+    Liouville Evolution ($\rho_{k}$): The force on $\rho_{k}$ is nonlocal, mediated by the interpolation kernel $K(z)$ appearing in both the bulk quadrature points and the spectral coefficients of the boundary term:
+    $\frac{d\rho_{k}}{d\tau}=-\frac{\partial S_{total}}{\partial\rho_{k}}+\nu_{\rho_{k}}$ (14.20)
+  4. 4.
+    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the exact algebraic constraints:
+    $\frac{d\gamma_{1,k}}{d\tau}=-\mathrm{i}(\tilde{\xi}_{k}\xi_{k}-1)+\nu_{\gamma_{1,k}}$ (14.21)
+In this mathematical framework, the deterministic drift terms ($\nabla S$) represent the exact holomorphic gradient flow in complexified phase space.
+In Picard-Lefschetz theory, the union of all these deterministic flow lines originating from a complex saddle point constructs a middle-dimensional real manifold known as a Lefschetz thimble.
+If the noise $\nu$ were zero, the system would trivially collapse onto a 1D classical trajectory (the WKB approximation).
+However, the multidimensional stochastic noise $\nu$ dynamically forces the probability distribution to completely explore the transverse width of these $N$-dimensional thimbles, strictly incorporating all exact quantum fluctuations (the “vicinity” of the saddle points).
+The phenomenon of exact S-matrix quantization we are looking for corresponds to a topological bifurcation of this stochastic system.
+At certain critical resonance energies, the complex Hessian develops a null space and the classical restoring drift strictly vanishes along a continuous topological valley.
+Unimpeded by any restoring gradient force, the stochastic noise $\nu$ drives the complex trajectory to infinity along this flat zero mode.
+This infinite flat-valley integration geometrically generates the exact macroscopic S-matrix pole of the path integral, governed by Catastrophe Theory.
+Before studying this phenomenon in general form, we consider the simple solvable limit: the WKB quantization corresponding to the large rotational momentum of our rigid string.
+15 The Linear Regge trajectories in the WKB limit
+15.1 The Chiral Helicoid Saddle Point
+In the semiclassical limit (large effective action), the path integral is dominated by the saddle point of the effective action.
+We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
+Consider a momentum distribution rotating in the $(1,2)$ plane with frequency $\omega$:
+$P_{1}(\tau)+\mathrm{i}P_{2}(\tau)=P_{0}e^{\mathrm{i}\omega\tau},\qquad P_{3}=P_{4}=0.$ (15.1)
+This source term in the action (11.7) drives the spinor fields.
+We search for a solution that respects this rotational symmetry.
+The natural ansatz corresponds to the helicoid minimal surface described in Appendix B.4.
+In the generic spinor formalism, with rescaling $w(z)=z^{-1/2}$, the conventional Helicoid is generated by $\mu(z)\propto(1,-z)^{T}$.
+However, as established in Appendix F, our geometric prescription requires enforcing the exact chiral (Hodge-dual) configuration.
+We freeze the right-handed twistor to a constant basis spinor:
+$\lambda(z)=\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=\mathrm{i}\begin{pmatrix}1\\ 0\end{pmatrix}$ (15.2)
+Substituting this into the null vector definition $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, we find that the spatial tangent vectors no longer identically vanish in the extra dimensions, like they did with conventional helicoid.
+Our Hodge dual minimal surface, with only one of two chiral Gauss maps (9.12), being active, and other being frozen, permanently breaks the 3D embedding constraint and allows the minimal surface to span the 4-dimensional space.
+Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid.
+The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
+15.2 The Symmetrized Minkowski Helicoid
+The Minkowski Helicoid is parameterized by time $\tau\in[0,T]$ and radius $r\in[-R,R]$, see [7]. The time $T$ tends to infinity in the end.
+$X(\tau,r)=(\ r\cos\omega\tau,\ r\sin\omega\tau,0,\tau)$ (15.3)
+The induced Minkowski metric for a generic string is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
+Crucially, as derived in Appendix F, the conformal factor for the generic surface evaluates to $4\cos^{2}(\sigma_{M})$.
+In stark contrast, our chiral twistor evaluates to an induced Euclidean conformal factor $\exp\left(2\rho\right)\propto(1+e^{-2\sigma_{E}})$.
+Applying our physical prescription—taking the parity-even arithmetic mean of the Left-chiral (SD) and Right-chiral (ASD) configurations—upon double Wick rotation ($\sigma_{E}\to\mathrm{i}\sigma_{M}$), precisely isolates the real part:
+$\text{Re}\left(1+e^{-2\mathrm{i}\sigma_{M}}\right)=1+\cos(2\sigma_{M})=2\cos^{2}(\sigma_{M})$ (15.4)
+Thus, the effective proper area of the symmetrized chiral Minkowski Helicoid is exactly halved relative to the conventional generic string.
+In Minkowski space we are not seeking to minimize the area in the geometric sense, as the area functional is not positive definite and the minimization problem is distinct.
+Instead, our criterion is algebraic: we require a solution to the Loop Equations for the Wilson loop.
+This demands the self-duality of the area derivative and the additivity of the area functional.
+We utilize the fact that the Euclidean Helicoid is a minimal surface for arbitrary angular velocity $\omega$.
+We obtain the Minkowski Helicoid via a double Wick rotation:
+$x_{4}\to\mathrm{i}t,\qquad\omega_{E}\to-\mathrm{i}\omega.$ (15.5)
+Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space.
+Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity).
+By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space.
+Thus, the symmetrized Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
+Figure 14: The helicoid spanned by rotating $\bar{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, was discovered by Meusnier in 1785.
+Another important point is that the Helicoid boundary loop (double helix) is infinitely smooth, lacking any cusps or self-intersections.
+This means that no UV divergent renormalization factors could potentially affect our computation of the meson Regge trajectory.
+15.3 Semiclassical Motivation: Twisted Boundary Conditions
+Before solving the spectral equation, we justify the choice of the Helicoid geometry from the perspective of the semiclassical quark path integral.
+Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$.
+In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor444the perimeter renormalization terms in that factor are included in the physical quark mass..
+This is relevant to our WKB approximation in the Wilson loop at large loops:
+$\mathcal{Z}\sim\int\mathcal{D}C\,\exp\left(-\mathrm{i}m_{q}\oint_{C}ds-\mathrm{i}\sigma S_{\mathrm{Mink}}[\Sigma_{C}]\right)$ (15.6)
+To extract the spectrum of states with angular momentum $J$, we impose twisted boundary conditions.
+We require the quark positions at time $t=T$ to be a rotation of their positions at $t=0$ by an angle $\Theta=\omega T$ in the rotation plane $xy$:
+$(x+\mathrm{i}y)(T)=e^{\mathrm{i}\omega T}(x+\mathrm{i}y)(0)$ (15.7)
+The angular momentum $J$ arises as the thermodynamic conjugate to the twist frequency $\omega$.
+The ground state of this system is determined by the saddle point of the effective action.
+This requires the simultaneous minimization of the boundary length (quark term) and the surface area (gluon term).
+  • 1.
+    The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
+  • 2.
+    The minimal surface spanning two twisted helical boundaries is the helicoid.
+Thus, the Helicoid is not merely an ansatz; it is the unique geometric solution mandated by the kinematics of angular momentum in the semiclassical limit.
+The balance between the surface tension $\sigma$ (pulling the quarks inward) and the centrifugal barrier (pushing the twisted trajectories outward) determines the physical radius of the meson.
+15.4 Effective Lagrangean
+We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks.
+The confining factor introduces an effective Minkowski area term.
+We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m_{q}\int ds$.
+The total Lagrangian is:
+$L=L_{\text{helicoid}}+L_{\text{quarks}}$ (15.8)
+Because the arithmetic mean of the SD and ASD surfaces rigorously halves the classical area, the bulk contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is precisely halved relative to the generic string:
+$L_{\text{helicoid}}$ $=-\frac{\sigma}{2}\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}=-\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
+$=-\frac{\sigma}{2\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (15.9)
+Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional.
+Adding the particle term $2m_{q}\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
+$L(v,\omega)=-\frac{\sigma}{2\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m_{q}\sqrt{1-v^{2}}$ (15.10)
+It is worth noting that this mass term is proportional to the perimeter of the loop in Minkowski space.
+Therefore, the potential perturbative terms proportional to the loop perimeter are included in this physical quark mass $m_{q}$.
+We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m_{q}$, which, as the experiment tells us, is much smaller than the QCD mass scale, so that $m_{q}^{2}\ll\sigma$.
+We proceed to compute the conserved quantities via the Legendre transform. The angular momentum $J$ is the conjugate momentum to the rotation frequency:
+$J=\frac{\partial L}{\partial\omega}$ (15.11)
+Taking into account that $v$ depends on $\omega$ via the fixed radial extent ($v=\omega R$), the differentiation yields:
+$J=\frac{\sigma}{2\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2m_{q}v^{2}}{\omega\sqrt{1-v^{2}}}$ (15.12)
+The Energy $E$ is defined by the Hamiltonian $E=\omega J-L$. Substituting the expressions for $J$ and $L$ and simplifying the algebra, we obtain:
+$E=\frac{\sigma}{\omega}\arcsin v+\frac{2m_{q}}{\sqrt{1-v^{2}}}$ (15.13)
+15.5 Balance of quark forces and boundary condition
+The relationship between the endpoint velocity $v$ and the string length $R$ is not arbitrary; it is determined by the balance of forces at the string endpoint, which depends on the quark mass $m_{q}$.
+We derive this boundary condition by minimizing the action with respect to the geometric radius $R$ of the loop, keeping the angular velocity $\omega$ fixed.
+Recall that the total Lagrangian $L$ is composed of the bulk integral for the halved helicoid and the boundary terms for the quarks:
+$L=-\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m_{q}\sqrt{1-\omega^{2}R^{2}}$ (15.14)
+The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$.
+Differentiating the integral using the Leibniz rule yields the effective tension force, while differentiating the boundary term yields the centrifugal force:
+$\frac{\partial L}{\partial R}=-\sigma\sqrt{1-\omega^{2}R^{2}}-2m_{q}\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (15.15)
+Substituting $v=\omega R$ and computing the exact derivative of the mass term (properly applying the chain rule):
+$-\sigma\sqrt{1-v^{2}}-2m_{q}\left(\frac{-\omega^{2}R}{\sqrt{1-v^{2}}}\right)=0$ (15.16)
+Simplifying the expression, we obtain the exact force balance equation for finite mass:
+$\sigma\sqrt{1-v^{2}}=\frac{2m_{q}v^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+2m_{q}}$ (15.17)
+This equation balances the physically halved confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
+We now consider the limit $m_{q}\to 0$, and we find the boundary condition
+$v^{2}\to 1$ (15.18)
+Thus, as the quark mass tends to zero, the endpoint velocity $v$ must approach the speed of light ($v\to 1$).
+These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$.
+To recover the Regge trajectories, we consider the limit of massless quarks ($m_{q}\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$).
+In this limit, the mass terms vanish, and the system is governed purely by the symmetrized confining geometry:
+$E\approx\frac{\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{2\omega^{2}}\frac{\pi}{2}$ (15.19)
+Eliminating $\omega$ between these two relations:
+$J=\frac{\pi\sigma}{4}\frac{1}{\omega^{2}}=\frac{\pi\sigma}{4}\left(\frac{2E}{\pi\sigma}\right)^{2}\implies J=\frac{E^{2}}{\pi\sigma}$ (15.20)
+We thus flawlessly recover the exact linear Regge trajectory for the symmetrized chiral string:
+$E^{2}=\pi\sigma J$ (15.21)
+This result confirms that the linear confinement natively arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space.
 16 The Catastrophe Theory and the Geometric Mass Spectrum
 16.1 Geometric Quantization via Degenerate Lefschetz Thimbles
-While the Complex Langevin Equation provides a robust numerical method to evaluate the twistor-string path integral, the holomorphic nature of the effective action reveals a profound mathematical pathway to compute the exact meson spectrum analytically, entirely bypassing stochastic simulation.
+The holomorphic nature of the effective action reveals a profound mathematical pathway to compute the exact meson spectrum analytically, entirely bypassing stochastic simulation.
 This approach relies on the asymptotic evaluation of multiple complex integrals via Picard-Lefschetz theory and the singularity theory of differentiable maps.
 Consider the original, un-reweighted path integral for the quark two-point amplitude $\Pi(E)$ as a function of the Minkowski energy parameter $E$:
-$\Pi(E)=\int_{\Gamma}\mathcal{D}\Phi_{|z|=1}\,\exp\left(E\,\mathcal{O}(\Phi)-S_{eff}(\Phi)\right)$ (17.1)
-where $\Phi_{|z|=1}=\left\{z_{2},\left\{\rho,\xi,\eta,\bar{\xi},\bar{\eta},\gamma,\bar{\gamma}\right\}_{k}\right\}$ encompasses the boundary twistor variables extended analytically into the complexified phase space $\Phi=(\lambda(z),\mu(z))$, and $\Gamma$ is the middle-dimensional integration cycle (the principal Lefschetz thimble) bounding the physical vacuum.
-In the semiclassical (WKB) limit, this integral is dominated by the saddle points $\Phi_{\star}(E)$ satisfying the classical boundary equations of motion:
-$\nabla S_{eff}(\Phi_{\star})-E\nabla\mathcal{O}(\Phi_{\star})=0$ (17.2)
+$\Pi(E)=\int_{\Gamma}\mathcal{D}\mathcal{V}_{|z|=1}\,\exp\left(E\,\mathcal{O}(\mathcal{V})-S_{eff}(\mathcal{V})\right)$ (16.1)
+where $\mathcal{V}_{|z|=1}=\left\{z_{2},\left\{\rho,\xi,\eta,\bar{\xi},\bar{\eta},\gamma,\bar{\gamma}\right\}_{k}\right\}$ encompasses the boundary twistor variables extended analytically into the complexified phase space $\mathcal{V}=(\lambda(z),\mu(z))$, and $\Gamma$ is the middle-dimensional integration cycle (the principal Lefschetz thimble) bounding the physical vacuum.
+In the semiclassical (WKB) limit, this integral is dominated by the saddle points $\mathcal{V}_{\star}(E)$ satisfying the classical boundary equations of motion:
+$\nabla S_{eff}(\mathcal{V}_{\star})-E\nabla\mathcal{O}(\mathcal{V}_{\star})=0$ (16.2)
 (which, as shown in Section 13, yields the rigid Helicoid geometry in the special case of twisted boundary conditions with large angular momentum $J$).
 The Gaussian fluctuations around this saddle point yield an amplitude proportional to the inverse square root of the complex Hessian determinant:
-$\Pi(E)\propto\frac{\exp\left(E\,\mathcal{O}(\Phi_{\star})-S_{eff}(\Phi_{\star})\right)}{\sqrt{\det\mathcal{H}(E)}}$ (17.3)
-where the fluctuation operator is $\mathcal{H}(E)=\nabla^{2}S_{eff}(\Phi_{\star})-E\nabla^{2}\mathcal{O}(\Phi_{\star})$.
+$\Pi(E)\propto\frac{\exp\left(E\,\mathcal{O}(\mathcal{V}_{\star})-S_{eff}(\mathcal{V}_{\star})\right)}{\sqrt{\det\mathcal{H}(E)}}$ (16.3)
+where the fluctuation operator is $\mathcal{H}(E)=\nabla^{2}S_{eff}(\mathcal{V}_{\star})-E\nabla^{2}\mathcal{O}(\mathcal{V}_{\star})$.
 As the continuous energy parameter $E$ is varied, the saddle points drift through the complexified phase space.
@@
 However, our twistor-string geometry orchestrates a remarkable mathematical conspiracy.
-Because the integration variables $\Phi$ are coupled complex holomorphic twistors, $\mathcal{H}(E)$ is a complex operator.
+Because the integration variables $\mathcal{V}$ are coupled complex holomorphic twistors, $\mathcal{H}(E)$ is a complex operator.
 The vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
@@
 The physical masses $E_{n}$ are precisely the discrete values for which the classical stability matrix develops a complex null space:
-$\det_{\mathbb{C}}\Big[\nabla^{2}S_{eff}(\Phi_{\star})-E_{n}\nabla^{2}\mathcal{O}(\Phi_{\star})\Big]=0$ (17.7)
+$\det_{\mathbb{C}}\Big[\nabla^{2}S_{eff}(\mathcal{V}_{\star})-E_{n}\nabla^{2}\mathcal{O}(\mathcal{V}_{\star})\Big]=0$ (16.7)
 Solving this boundary-twistor Fredholm equation in general case provides a direct analytical roadmap for the exact, noise-free computation of the geometric spectrum in future work.
@@
 This makes $\exp\left(\mathrm{i}\Delta S_{\text{eff}}\right)=1$, which leaves the amplitude invariant under winding, so that the sum over windings macroscopically diverges as $\sim 1/(1-\exp\left(\mathrm{i}\Delta S_{\text{eff}}\right))\sim 1/(E-E_{n})$.
+Remark 16.2.
+Localization of the spectrum vs. localization of path integral
+It is crucial to distinguish between the exact spectral condition and the exact computation of the path integral.
+In the complexified functional space, there are infinitely many directions away from the critical saddle point (where a complex eigenvalue vanishes).
+In the framework of Stochastic Quantization and the Complex Langevin Equation (CLE), the stochastic trajectory is effectively exploring a Lefschetz Thimble—the stable manifold of the gradient flow [5].
+The trajectory passes through the saddle and proceeds along an infinitely long, "flat" minimal valley, while stochastic noise causes it to fluctuate against the steep potential walls surrounding this valley.These transverse fluctuations contribute to the pre-exponential factor (comprising the functional determinant and higher-order expansion terms).
+However, as established in the topological study of Landau singularities by [41], the global properties of the integral are determined by the monodromy of the integration cycle.
+The trajectory must wind infinitely many times around the logarithmic branching point of the effective action $S_{eff}(z)$.Each circuit around this singularity shifts the Minkowski action by a fixed amount, $\mathrm{i}\Delta S_{eff}$.
+The condition for a pole $E=E_{n}$ in the total amplitude corresponds to the constructive interference of these windings, leading to the quantization condition:
+$\Delta S_{eff}=2\pi n$ (16.8)
+This monodromy is a topological invariant of the flow and is independent of the "shape" of the valley sides or the specific stochastic time, which are non-universal.
+This separation of local geometry from global topology is a manifestation of the spectral flow and the index theorems of [2], where the "integer" part of the quantization (the pole) remains robust against the local deformations of the path caused by Langevin noise.
+As demonstrated in our WKB analysis of the helicoid, this mechanism ensures that the energy spectrum is localized by the classical monodromy, even while the path integral remains delocalized by the CLE.
 16.4 One pole and its symmetries
@@
 While this specific frame choice explicitly breaks the full Lorentz group, the $O(3)$ spatial rotational symmetry remains intact.
-We may utilize this unbroken space rotation to align the twistor dynamics. Specifically, we can always choose a spatial frame such that the symmetric bilinear combination vanishes:
+In a conventional approach modeling a generic “rubber band” string, one might utilize this unbroken space rotation to align the twistor dynamics such that the symmetric bilinear combination vanishes:
 $\lambda^{T}\sigma_{1}\mu=0$ (16.9)
-Because the spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, this strict algebraic constraint forces one component of the tangent vector to be zero everywhere on the worldsheet.
-This dimensionally reduces the physical minimal surface from a generic 4-dimensional embedding into a 3-dimensional hyperplane, exactly matching the geometric prerequisite for a canonical Helicoid.
+Because the spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, this strict algebraic constraint would force one component of the tangent vector to be zero everywhere on the worldsheet.
+This would dimensionally reduce the physical minimal surface from a generic 4-dimensional embedding into a 3-dimensional hyperplane.
+However, here is where the profound difference between the conventional “rubber band” string and our twistor formulation manifests itself.
+Instead of crushing the surface into 3D, we enforce the exact chiral (Hodge-dual) configuration.
+Specifically, we freeze one of the chiral sectors by setting the right-handed twistor to a constant basis spinor, such as $\mu=(1,0)^{T}$.
+This choice explicitly breaks the 3D embedding constraint, causing the minimal surface to genuinely span a 4-dimensional space and flawlessly evade the finite-area self-intersection theorems.
+Because macroscopic QCD is strictly parity-conserving, the physical target-space configuration cannot remain purely left- or right-chiral.
+The true physical minimal surface is given by our exact prescription: the parity-even arithmetic mean of the left-chiral (Self-Dual or SD) and right-chiral (Anti-Self-Dual or ASD) surfaces.
+As derived in Appendix F, this symmetrization rigorously projects out the imaginary geometric artifacts of the 4D embedding, yielding a strictly real Nambu-Goto string geometry whose fundamental classical area is exactly halved relative to the conventional generic string.
 Furthermore, we possess a global $U(1)$ symmetry on the worldsheet integration domain (the unit disk).
@@
 The string dynamically centers itself to restore the $U(1)$ symmetry.
-By locking the pole to the origin, the minimal surface reduces exactly to the canonical, maximally symmetric Helicoid.
-The remaining computations then proceed in the exact same way as for the canonical Helicoid in Minkowski space.
-As we derive in a new section in the Appendix, the worldsheet metric following from these canonical twistors $\lambda,\mu$ transforms precisely into the canonical Nambu-Goto rotating string metric.
-In the text below, we simply use this metric.
-17.5 The exact spectrum in the one pole sector
-With the twistor saddle point locked to $z_{0}=0$ and the surface constrained to the 3D Helicoid, we repeat the action computation.
-In this one-pole sector, there are no other terms; the entire worldsheet dynamics are governed by exactly two terms within the Liouville action: the classical area term ($\sigma$) and the kinetic anomaly ($|\nabla\rho|^{2}$).
-Because the conformal density is strictly proportional to $1/|z|^{2}$, the kinetic term acts as a rigid, local shift to the effective Helicoid string tension.
-Evaluating the 1-cycle Minkowski bulk action, we obtain the effective Helicoid energy, complete with the anomalous Lüscher term:
-$S^{(1)}_{bulk}=-2\pi^{2}\left(\sigma+\frac{1}{12\pi R^{2}}\right)R^{2}=-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$ (17.10)
-Simultaneously, the boundary term coupling the string to the external energy $E$ integrates to yield the exact phase monodromy per winding:
-$\Delta S_{E}=4\pi ER$ (17.11)
-Summing these to form the total phase of one physical cycle, we enforce the geometric series resonance condition ($\Delta S_{\text{eff}}=2\pi n$):
-$4\pi ER-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$
-$=2\pi n\implies 2ER-\pi\sigma R^{2}-\frac{1}{12}=n$ (17.12)
-To locate the physical vacuum, we minimize this effective base action with respect to the spatial radius $R$ to balance the string tension against the injected energy:
-$\frac{\partial}{\partial R}\left(2ER-\pi\sigma R^{2}\right)=2E-2\pi\sigma R=0\implies R=\frac{E}{\pi\sigma}$ (17.13)
-Substituting this exact saddle-point radius back into the quantization condition, we obtain the exact algebraic spectral equation:
-$2E\left(\frac{E}{\pi\sigma}\right)-\pi\sigma\left(\frac{E}{\pi\sigma}\right)^{2}-\frac{1}{12}=n$ (17.14)
-Simplifying the squares yields:
-$\frac{E^{2}}{\pi\sigma}=n+\frac{1}{12}\implies E^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (17.15)
-Since we are in the rest frame, the invariant mass squared is $m^{2}=E^{2}$, resulting in the final string spectrum:
-$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (17.16)
-We got the famous Lüsher term, without any string models or assumptions.
-We obtained this term from the conformal anomaly left from the Elfin determinant, which we used to induce MM loop equations.
-This remarkable coincidence tells us that our solution for planar QCD matches accepted phenomenology as well as the results of the real and numerical experiments.
-QED!
-(or rather QCD!).
-17.6 QCD as a Geometric Theory
-The derivation of the exact spectral equation in the previous section represents a significant shift in our understanding of the large-$N_{c}$ meson spectrum.
-We have moved from a stochastic description (random surfaces) to a deterministic geometric principle.
-Let us summarize the key physical insights that have emerged from this analysis.
-Our approach avoids the usual “graviton problem” of gravity holography: since our bulk is not dynamical, there are no propagating bulk graviton modes (or other unwanted massless bulk fields) that would have no counterpart in confining QCD.
-More precisely, the propagating degrees of freedom in the present formulation are confined to the boundary: (i) the quark loop degrees of freedom, represented by the matrix trace of the Dirac propagator around the boundary loop, and (ii) the holomorphic moduli (twistors) that parametrize the rigid minimal surface.
-The twistor moduli are specified by their boundary data and extended uniquely into the disk by analytic continuation (Plemelj/Hilbert reconstruction in the spirit of the Penrose transform).
-Accordingly, bulk quantities such as the Liouville field are not integrated over as independent quantum fields; they are determined functionally by the analytically continued twistor data.
-It is also important not to conflate this construction with Witten/Berkovits twistor strings or other topological twistor-string models.
-In those theories the extension of boundary data into the bulk typically contributes a topological functional (e.g.
-a WZW term) whose value depends only on the homotopy class of the extension.
-In contrast, our bulk functional is not topological: its value depends on the specific holomorphic disk selected by the boundary loop, and this non-topological dependence is essential for confinement and for the existence of a mass scale.
-In this sense, our “rigid twistor string” is best viewed as gauge holography (bulk as a rigid encoding of boundary dynamics), rather than gravity holography or a topological string.
-Finally, the observation relating the QCD mass spectrum to the degenerate saddle points of the complexified twistor-string action introduces the concept of localization of the string path integral, but on an entirely different level.
-Our spectral equation is exactly localized, much like the computation of path integrals in geometric quantization [51], but this is not a familiar fixed point in the moduli space.
-Instead, the path integral localizes exclusively on degenerate saddle points governed by catastrophe theory.
-We thus have uncovered a novel, purely geometric mechanism of localization that deterministically dictates the QCD mass spectrum.
-17.7 Future Directions: Glueballs, Baryons and Scattering
-The formulation of Planar QCD as an Analytic Twistor String opens several avenues for future research:
+By locking the pole to the origin, the minimal surface reduces exactly to the maximally symmetric canonical Helicoid.
+The remaining computations then proceed utilizing the arithmetic mean of the SD and ASD areas.
+In the text below, we simply use this symmetrized Hodge-dual metric.
+16.5 Quark Spin Monodromy and the Unified Regge Spectrum
+The exact meson spectrum is completed by accounting for the discrete spin monodromy of the boundary quarks.
+As algebraically demonstrated in F, the discretization of the quark loop into $2L$ steps yields a path-ordered Dirac trace $T\propto(-1)^{L}$.
+This phase factor provides an intrinsic topological index that modifies the constructive interference condition in the worldsheet amplitude, selecting specific sectors of the geometric series $\sum\exp\left(iS_{k}\right)$.
+The resulting Bohr-Sommerfeld resonance condition is $\Delta S_{\text{eff}}=\pi n_{BS}$.
+Based on the trace $T$, physical light mesons ($\pi$ and $\rho$) are found to occupy the even-$n_{BS}$ sectors.
+Defining the worldsheet topological quantum number $n=n_{BS}$ and evaluating the 1-cycle Minkowski bulk action for the symmetrized chiral Helicoid (Appendix F), we obtain the unified mass-squared formula:
+$m^{2}=\frac{\pi\sigma_{e,o}}{4}\left(n+\frac{1}{12}\right)$ (16.11)
+For even $L$, the projector supports quantization with $n\equiv 0\pmod{4}$ and $n\equiv 2\pmod{4}$.
+To map these topological sectors to the macroscopic meson spectrum and define the intercepts $\alpha(0)=J|_{m^{2}=0}$, we adopt the semi-classical WKB mapping $n=4J+q$:
   • 1.
-    Glueballs: the whole construction, starting with the loop equations to the fermion determinant at the minimal surface, to the rigid twistor string and the localization of the twistor path integral for quark amplitudes at the boundary of the moduli space, naturally generalizes to the higher $1/N_{c}$ corrections.
-There are some important mathematical issues to be resolved for that program to materialize.
-This is a natural next step for this Geometric QCD theory.
+    Unnatural Parity ($\pi$ and $K$): $q=0\implies n=4J$, placing the sequence in the $n\equiv 0\pmod{4}$ sector and yielding the intercept $\alpha_{\pi}(0)=-1/48$.
+  • 2.
+    Natural Parity ($\rho$): $q=-2\implies n=4J-2$, placing the sequence in the $n\equiv 2\pmod{4}$ sector and yielding the intercept $\alpha_{\rho}(0)=23/48$.
+As demonstrated in Fig.
+15, the experimental data confirms this topological identification.
+Because the $K$ mesons share the unnatural parity of the pion, they belong to the identical topological sector $n=4J$, with their trajectory cleanly displaced by a positive mass-squared shift $\Delta m_{s}^{2}\approx 0.17$ GeV2 reflecting the explicit $SU(3)$ flavor symmetry breaking by the heavier strange quark.
+The observed variation of the slopes ($\pi\sigma_{\pi}\approx 1.38$, $\pi\sigma_{\rho}\approx 1.16$, and $\pi\sigma_{K}\approx 1.58$ GeV2) remains an open dynamical question.
+It may arise from the interaction between a pole in $\lambda$ and a pole in $\mu$, which would locally renormalize the effective string tension.
+We mention this possibility here as a direction for future exploration but retain the distinct empirical fits to maintain contact with the PDG data.
+We must also address a natural phenomenological question regarding our selection of experimental candidates.
+A critical reader comparing our trajectories with the Particle Data Group (PDG) tables will note that we have not included all listed resonances possessing the same quantum numbers.
+This selection was deliberate, but it is not a mere phenomenological convenience; rather, it is strictly dictated by the scope of our analytic solution.
+The linear Regge trajectories plotted in Fig.
+15 correspond exclusively to the absolute simplest topological sector of the theory: the exact one-pole twistor solution.
+Our geometric framework guarantees that the remaining, unmapped hadronic resonances in the PDG are not anomalies, but are dynamically required to populate the daughter trajectories generated by multi-pole twistor configurations.
+Solving the complexified effective action—specifically, solving the saddle-point equations for the moduli $R_{i},z_{i}$ of the multi-pole solutions and classifying their monodromies—constitutes a deterministic, yet computationally formidable, nonlinear optimization problem in complex geometry.
+Systematically extracting these higher-topology solutions and mapping them onto the complete PDG database falls beyond the foundational scope of the present paper.
+Instead, we present this as a rigorous, data-rich research program—an open invitation to apply the sophisticated analytic machinery of string theory directly to the rich empirical data of hadronic spectroscopy.
+Figure 15: Unified Chew-Frautschi plot ($J$ vs $m^{2}$) showing the topological identification of the $\pi$ ($n=4J$), $K$ ($n=4J$) and $\rho$ (n=4J2n=4J-2) trajectories.
+The horizontal error bars represent propagated mass uncertainties, and labels indicate the experimental masses in MeV.
+The $K_{3}(2320)$ is marked with a question mark as it is currently omitted from the PDG Summary Tables pending further experimental confirmation.
+16.6 The microscopic origin of the Lüscher term: elves instead of string vibrations
+Our exact derivation of the strictly bounded $\pi/12$ topological phase unveils a fundamentally different, purely microscopic physical origin for the universally mandated Lüscher term, representing the absolute culmination of our theory.
+In standard Effective String Theory, this famous constant is interpreted strictly via the macroscopic Casimir effect.
+The standard calculation dictates that one must manually quantize the zero-point fluctuations of a macroscopic, highly excited “rubber band” string bending and vibrating across the transverse spacetime dimensions.
+The $\pi/12R$ effective potential is then forcefully extracted from the ultraviolet divergence of these vibrating spatial standing waves via Riemann zeta regularization.
+Our twistor-string formulation completely upends this paradigm.
+We bypassed the macroscopic quantization of target-space wiggles entirely.
+Instead, the quantum mechanics takes place directly on the 2-dimensional worldsheet.
+The Lüscher term originates intrinsically from the quantum anomaly of the microscopic zero-point fluctuations of the heavy fermions—the “elves”—that are confined to our minimal surface.
+When evaluating the fermion determinant (required to reproduce the planar factorization of the MM equation), integrating out these fundamental elfin degrees of freedom generates the quantum anomaly, related to the curvature.
+Because these quantum elves live on a highly curved minimal surface governed by the singular twistor conformal factor $\rho$, regulating their vacuum fluctuations inherently breaks conformal symmetry, generating the Weyl/Polyakov trace anomaly.
+The direct mathematical consequence of this trace anomaly is the rigid addition of the Liouville kinetic term $\frac{1}{3\pi}|\nabla\rho|^{2}$ to the effective action.
+Crucially, the numerical coefficient of this anomaly is strictly proportional to the central charge of the worldsheet—it literally counts the active quantum degrees of freedom of the elves.
+In a conventional generic 3-dimensional Helicoid, both chiral sectors dynamically fluctuate, which overcounts the anomaly (yielding an unphysical $\pi/6$).
+However, by enforcing the chiral (Hodge-dual) configuration and taking the parity-even arithmetic mean of the SD and ASD minimal areas, we dynamically halve the internal fluctuating degrees of freedom compared to the generic minimal surface.
+This exact symmetrization prescription makes the day: it mathematically halves the anomaly integral over the minimal surface, flawlessly dropping the topological phase to precisely $\pi/12$.
+[…]
+17 Planar QCD as a Geometric theory
+The dynamical generation of S-matrix mass poles through the degeneracy of the twistor-string effective action is not an isolated mathematical curiosity; it is an infinite-dimensional realization of Exact WKB analysis, Picard-Lefschetz resurgence, and the geometry of Spectral Networks [15, 10].
+The complexified path integral over the boundary twistors exhibits localization phenomena that strictly mirror the topological mechanisms governing Seiberg-Witten (SW) theory and the analytic continuation of gauge theories [55].
+To establish this physical correspondence, we analyze the three fundamental geometric mechanisms driving the twistor string quantization:
+17.1 Topological Barriers and Infinite-Dimensional Picard-Lefschetz Theory
+By analytically continuing the twistor path integral into the complexified phase space, the convergence of the quantum amplitude is governed by middle-dimensional integration cycles, or Lefschetz thimbles [5].
+As formulated by Witten in the analytic continuation of Chern-Simons theory [55], the topology of these thimbles—and their intersection pairings—changes strictly when the system crosses a Stokes line, resulting in a discontinuous jump in the physical vacuum state.
+In our twistor string framework, non-perturbative Stokes phenomena are rigorously suppressed by the conformal boundary geometry.
+If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour.
+Exactly at the moment of crossing, these two singularities inevitably collide directly on the physical integration contour $|z|=1$.
+As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity.
+The integration loop becomes trapped between the colliding singularities at $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them.
+This generates a severe pinch singularity, creating an infinite action barrier.
+In the language of Picard-Lefschetz theory, this impenetrable boundary dynamically locks the Twistor String path integral into a strictly conserved topological sector defined by the enclosed twistor pole number $n$, forbidding unphysical Stokes jumps.
+17.2 Complex Langevin, Cauchy’s Shield, and Flat Valleys
+The rigorous exploration of these complex integration cycles is executed via the holomorphic gradient drift of the Complex Langevin Equation (CLE).
+In standard semiclassical formulations, evaluating the path integral at a degenerate saddle point requires formally isolating a zero mode and calculating the fluctuation determinant of the Hessian.
+In the CLE framework, this static abstraction is bypassed entirely: the path integral expectation value is simply the long-time average $\int d\tau$ over a single, noisy stochastic trajectory $\mathcal{V}(\tau)$.
+At the critical resonance energies $E=E_{n}$, the twistor effective action develops a non-compact flat valley.
+The single CLE trajectory does not rest at a classical minimum; driven by the stochastic noise, it constantly climbs the steep transverse walls of the complexified functional space and bounces back.
+Simultaneously, experiencing no deterministic restoring force along the zero mode, the trajectory drifts unimpeded along the flat valley.
+In this representation, the macroscopic pre-exponential factor (the transition residue) is simply proportional to the total Langevin time the trajectory spends accumulating windings, completely eliminating the need for formal Hessian factorization.
+A fundamental mathematical question arises: how can this violently stochastic trajectory $\mathcal{V}(\tau)$ generate a pristine, exact Bohr-Sommerfeld quantization phase without acquiring infinite quantum corrections from the nonlinear effects amplifying the noise in transverse directions?
+The resolution lies in the analytic continuation of the twistor fields.
+The CLE directly evolves the boundary data $\mathcal{V}_{bound}(\tau)=\mathcal{V}(z,\tau)\big|_{|z|=1}$.
+However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk.
+If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{k}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
+$\lambda(\theta)=\sum_{k}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (17.1)
+The stochastic CLE evolution simply forces these boundary parameters to evolve with Langevin time $\tau$:
+$\lambda(\theta,\tau)=\sum_{k}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (17.2)
+Upon analytic continuation back into the bulk, we obtain:
+$\lambda(z,\tau)=\sum_{k}\frac{R_{i}(\tau)}{z-z_{i}(\tau)}$ (17.3)
+The crucial observation is that the function remains strictly meromorphic at every instant of Langevin time.
+We obtain moving poles with evolving residues, but the fundamental meromorphic structure is absolutely preserved.
+Therefore, the infinite-dimensional stochastic evolution of the boundary fields rigorously and exactly reduces to the finite-dimensional, continuous $\tau$-dependence of the moduli $\{R_{i}(\tau),z_{i}(\tau)\}$.
+The infinite-dimensional phase accumulation thereby rigidly reduces to a standard one-dimensional contour integral over the physical boundary:
+$\Delta S=\oint_{|z|=1}\omega(z,\tau)\,dz$ (17.4)
+Here, the absolute geometric power of Cauchy’s integral theorem manifests itself.
+At any instant $\tau$, this contour integral is completely determined by the sum of the residues at the enclosed twistor poles.
+To evaluate the full physical amplitude, the path integral must be decomposed into a discrete sum over distinct topological sectors.
+These sectors are labeled by the winding number $k$ of the target-space minimal surface (or equivalently, the number of times the physical boundary parameter traverses the twistor poles).
+The full quantum amplitude is thus a superposition of path integrals, each mathematically explored by its own Complex Langevin evolution.
+Because the twistor effective action is strictly additive over these cycles, the total action for the $k$-th topological sector is precisely $S_{k}=S_{eff}^{(1)}+(k-1)\Delta S_{eff}$.
+In the CLE formulation for the $k$-th sector, the stochastic noise is opposed by the deterministic drift and the transverse Hessian, which now scale linearly with the topological winding number: $\mathcal{H}_{k}\to k\nabla^{2}\Delta S_{eff}$.
+Consequently, the residual stochastic variance in this sector is inversely proportional to the Hessian, scaling exactly as $\langle\delta\mathcal{V}^{2}\rangle\propto 1/k$.
+If the CLE noise were to continuously jitter the moduli $\{R_{i},z_{i}\}$ across the different sectors, the topological action increment $\Delta S_{eff}$ would randomly vary from one winding to the next.
+This would destroy the perfect phase coherence required for the geometric series to diverge, preventing the formation of a resonance.
+However, the macroscopic S-matrix pole is analytically generated exclusively by the asymptotic tail of this superposition, where $k\to\infty$.
+In this asymptotic limit, the transverse restoring force becomes strictly infinite, and the stochastic variance vanishes identically.
+The stochastic noise is completely crushed by the macroscopic weight of the multiple windings.
+The moduli $\{R_{i},z_{i}\}$ rigidly freeze at their exact classical saddle-point values for all large-$k$ sectors (validating the algebraic minimization of the single-pole radius $R$ in our spectral derivation).
+The topological phase is thus completely shielded from quantum contamination, ensuring that $\Delta S_{eff}$ remains strictly constant.
+The superposition over topological sectors safely diverges as a pristine geometric series $\sim(1-\exp\left(i\Delta S_{eff}\right))^{-1}$, analytically generating the exact S-matrix pole.
+18 Conclusion: The Solution of Planar QCD
+In this paper, we have presented the exact analytic solution of the planar MM loop equations, thereby solving large-$N_{c}$ QCD in the continuum limit.
+Building on Part I, where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ was shown to arise from the rigid Hodge-dual additive minimal surface, we have constructed the complete dynamics without summing over fluctuating worldsheet metrics.
+Our first main outcome is structural: by working natively in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum limit of the MM loop equations.
+We rigorously established that a purely one-dimensional analytic Taylor-Magnus solution of the exact vector momentum-loop equation cannot exist.
+The expansion is mathematically consistent through $\mathcal{O}(P^{\prime 6})$ but violently breaks down at $\mathcal{W}^{(8)}$ due to a Fredholm rank deficiency.
+This topological obstruction proves that a 1D closed-loop scalar framework lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD.
+This anomaly natively forces the introduction of 4D continuous boundary-twistor variables to parametrize the rigid minimal surface, yielding the confining twistor-string representation.
+Our second main outcome is dynamical: we demonstrated that the internal Majorana fermions (the “elves”) on the worldsheet provide the precise algebraic mechanism required for planar factorization at intersections.
+Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions.
+The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/48-1/48 and $23/48$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
+In a profound sense, this construction explicitly realizes Edward Witten’s original “Master Field” hypothesis—the existence of a purely classical description for the large-$N_{c}$ limit of QCD.
+While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, we have discovered its exact geometric realization in loop space.
+The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface.
+Consequently, the phase-space path integral over quark loops, once transformed into two-dimensional boundary twistor variables, exhibits exact geometric localization governed by Catastrophe Theory.
+The true Master Field of Planar QCD is thus a critical classical trajectory in twistor space—a geometric structure conceptually parallel to the spectral curves of Seiberg-Witten theory.
+By establishing the Master Field as this rigid twistor geometry, our approach naturally avoids the “graviton problem” typical of AdS/CFT gravity holography.
+Because our bulk geometry is not dynamical, there are no propagating bulk graviton modes that would lack a physical counterpart in confining QCD.
+All propagating degrees of freedom are strictly confined to the boundary, while bulk quantities (such as the Liouville field) are rigidly determined by the analytic continuation of boundary data.
+Therefore, our framework is best understood as gauge holography, where the bulk serves as a rigid encoding of the boundary dynamics rather than gravity holography or a topological string theory.
+Ultimately, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry.
+The extraction of the exact mass spectrum reduces from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem.
+By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition.
+For the fundamental one-pole trajectory, this yields the unified linear Regge spectrum:
+$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (18.1)
+where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory.
+This analytic solution passes flawlessly through experimental data on Figure 15.
+The only outlier ($3\sigma$) is $K_{3}(2320)$ which is marked as "unconfirmed" in the PDG tables.
+The exact mass spectrum of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
+18.1 Future Directions: Glueballs, Baryons, and Scattering
+The representation of Planar QCD as an Analytic Twistor String opens several immediate avenues for future research:
+  • 1.
+    Glueballs: The entire construction, from the loop equations and the fermion determinant on the minimal surface to the localization of the twistor path integral, naturally generalizes to higher $1/N_{c}$ corrections.
+Resolving the mathematical specifics of this expansion is the natural next step to analytically extract the glueball spectrum.
   • 2.
     Baryons: In the large-$N_{c}$ limit, baryons appear as heavy states composed of $N_{c}$ quarks.
-Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a "Y-junction" or a surface with multiple boundaries) [5].
-Extending the twistor ansatz to surfaces with higher connectivity is a natural next step.
+Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a “Y-junction” or a surface with multiple boundaries).
+Extending the twistor ansatz to surfaces with higher connectivity will allow for the exact geometric quantization of baryon masses.
   • 3.
-    Scattering Amplitudes: We have focused here on the two-point function (spectrum).
-The four-point function (meson-meson scattering) involves a minimal surface bounded by four segments.
-The crossing symmetry and Regge behavior of these amplitudes should emerge from the analytic continuation of the twistor data, potentially connecting this framework to the modern "Amplituhedron" methods.
-18 Conclusion
-In this paper, we developed the dynamical part of the Geometric QCD program.
-Building on [1], where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ arises from the rigid Hodge-dual additive minimal surface, we formulated a concrete large-$N_{c}$ framework in which the remaining degrees of freedom can be treated without summing over fluctuating worldsheet metrics.
-The first main outcome is conceptual and structural: by working in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum interpretation of the loop equations.
-In this representation, the planar loop equation becomes an algebraic–differential equation.
-Furthermore, by mapping the Master Field to the infinite-dimensional Cuntz-algebra “word space” (Section 10.8), we rigorously establish that a purely one-dimensional analytic Taylor–Magnus (finite-moment) solution of the exact vector momentum-loop equation cannot exist: the expansion is consistent through $\mathcal{O}(P^{\prime 6})$ but becomes algebraically inconsistent at $\mathcal{W}^{(8)}$ (C, Table 1).
-This violent breakdown is a theorem-level obstruction showing that a 1D closed-loop scalar Taylor–Magnus framework lacks sufficient internal degrees of freedom to absorb the macroscopic spatial stress generated by planar QCD.
-This breakdown prompts the introduction of additional continuous degrees of freedom; in our construction these are the 4D boundary-twistor variables that parametrize the rigid Hodge-dual minimal surface and yield the confining twistor-string representation.
-The second main outcome is dynamical: we showed that the fermionic (Majorana) worldsheet degrees of freedom of the Fermi string provide the correct algebraic mechanism for planar factorization at intersections, while the rigidity of the Hodge-dual minimal surface removes the Liouville instability of the original random-surface formulation.
-The resulting theory may be viewed as an analytic (confining) twistor-string description in which the reduced loop-space measure is most naturally written in twistor variables, leading to an effective holographic Liouville functional determined by holomorphic extension of boundary data.
-The third main outcome concerns spectroscopy and the profound realization that the quantum path integral geometrically localizes.
-We initially demonstrated the computability of the boundary-twistor functional integral by evaluating it on the IAS Typhon supercluster via Complex Langevin (CLE) evolution with logarithmic reweighting.
-To overcome the severe topological trapping and infinite-variance $1/|u|^{3}$ Cauchy tails inherent to the complexified phase space, we introduced an unsupervised phase-information clustering algorithm.
-Applying this to our massive stochastic histories successfully isolated two distinct, approximately linear Regge trajectories.
-This numerical exploration served a crucial purpose: it proved that the Confining Twistor String natively possesses internal quantum numbers (analogous to string “daughter trajectories”) emerging directly from the multiple complex saddle points (Lefschetz thimbles) of the path integral.
-However, the ultimate realization of this work goes far beyond numerical stochasticity.
-By analyzing the analytic structure of the Twistor String, we discovered that the physical mass spectrum is deterministically governed by Catastrophe Theory.
-The meson poles correspond exactly to degenerate Lefschetz thimbles—bifurcation points where the complex stability matrix of the classical twistor trajectory develops a null space.
-Because the integration variables are coupled holomorphic twistors, this degeneracy inherently possesses a corank of 2.
-As dictated by the singularity theory of multiple integrals, this complex degeneracy mathematically upgrades standard semiclassical square-root branch cuts into the exact simple poles of the meson states.
-Summary
-To summarize, what started as a rigorous continuum formulation of planar QCD via the Makeenko–Migdal loop equations culminated in an exact analytical solution for the fundamental meson spectrum.
-Although our numerical investigations using the Complex Langevin approach successfully provided proof-of-concept evidence for linear Regge trajectories, the true predictive power of this framework lies in its analytical semiclassical limit, which provides exact equations for the mass spectrum beyond the WKB approximation at the non-perturbative level.
-It is a known phenomenon that some WKB spectral equations are valid beyond the WKB approximation, because they reflect deeper hidden symmetries leading to infinite flat valleys for the steepest descent path, making the WKB estimate an exact localization of the path integral [11].
-We have shown that the $S$-matrix mass poles of the theory are naturally captured by Catastrophe Theory, corresponding to non-compact flat valleys in the complexified effective action.
-These flat valleys, where the steepest descent integral diverges without higher-order corrections, are driven entirely by the topological monodromies of twistor poles enclosed within the unit disk.
-The number of such poles serves as a rigorously conserved topological quantum number, protected by an absolute pinch-singularity barrier at the conformal boundary.
-The crossing of the unit circle by a twistor pole leads to a divergent action, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop.
-By analytically continuing the Euclidean area action to the boundary and summing over the exact purely imaginary monodromy phases, we obtain a Bohr–Sommerfeld quantization condition.
-For the fundamental one-pole trajectory, this yields the exact simple mass poles and the linear Regge spectrum:
-$m^{2}=\pi\sigma\left(n+\frac{1}{12}\right)$ (18.1)
-This analytical result not only confirms the perfectly linear Regge slope but uniquely derives the exact open-string Lüscher intercept $(+1/12)$ directly from the 2D anomaly of the Elfin determinant on rigid minimal surface.
-Thus, by mapping the momentum loop-space dynamics to boundary twistors, the theory does not merely approximate the confining dynamics of planar QCD; it provides its exact algebraic geometric solution.
-The exact analytical classification of higher topological sectors (multi-pole configurations) and their corresponding daughter trajectories is now a fully accessible problem.
-Consequently, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry.
-The extraction of the mass spectrum is reduced from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem for the degenerate saddle points of the effective action.
+    Scattering Amplitudes: Meson-meson scattering involves a minimal surface bounded by four segments.
+The crossing symmetry and Regge behavior of these amplitudes should emerge naturally from the analytic continuation of the twistor data, potentially connecting this confining geometric framework directly to modern Amplituhedron methods.
 Acknowledgements
-I am grateful to Nima Arkani-Hamed for the invitation to present this work as a series of lectures at the IAS Particle Physics (Pizza) Seminar, which provided a stimulating environment for discussing these results.
-I also thank Edward Witten for valuable discussions regarding the distinction between this confining geometry and the holographic dualities of conformal field theories.
-Declaration of generative AI and AI-assisted technologies in the writing process
-During the preparation of this work, the author used Google Gemini to fix typographical errors, identify relevant mathematical terminology, clarify the notations, and improve the overall language and style.
-After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
+I wish to express my deepest gratitude to Nima Arkani-Hamed for his hospitality and for the invitation to present the complete derivation of this solution in a series of lectures at the IAS Particle Physics (Pizza) Seminar.
+The intellectual rigor and stimulating environment he fosters at the Institute were instrumental in the final crystallization of these results.
+I am especially indebted to Edward Witten for his profound insights and for the invaluable discussions that clarified the correspondence between this twistor-geometric solution and the phenomenological QCD string observed in lattice simulations.
+His generous engagement with the topological aspects of this work has been a source of significant inspiration.
+Declaration of Generative AI in the Writing Process
+During the preparation of this work, the author utilized the generative AI Gemini 3 Flash to assist in the technical refinement of the manuscript. Specifically, the tool was employed to:
+  • 1.
+    Identify and correct typographical errors and grammatical inconsistencies.
+  • 2.
+    Standardize mathematical notation and ensure the rigorous formatting of LaTeX equations across the text and appendices.
+  • 3.
+    Improve the clarity and stylistic flow of the scientific narrative while preserving the original physical derivation.
+  • 4.
+    Cross-verify the consistency of the unified mass-squared formula with experimental data fits and topological sector identifications.
+The author has thoroughly reviewed and edited all AI-generated refinements and takes full responsibility for the physical content, mathematical accuracy, and original conclusions presented in this work.
 Data Availability.
-The spectral data reported in this paper are generated by numerical simulation of the Complex Langevin equation (CLE) for the discretized Confining Twistor String spectral observable, including the reweighted estimator for $1/R(E)=\Pi(E)/\Pi^{\prime}(E)$.
-To enable reproduction and extension of these calculations, I have made the full Mathematica ® workflow publicly available on the Wolfram Cloud, including (i) the CLE simulator and effective-action implementation, (ii) SLURM submission scripts for running independent replicas on an HPC cluster, and (iii) the analysis tools used to pool independent runs and extract the spectrum:
-  • 1.
-    TwistorString.wl, [52] (main CLE simulator):
-  • 2.
-    run_node.m (SLURM/cluster driver) [53]
-  • 3.
-    VerifyTwistorHistory.nb [54] (fat tail analysis of distributions for a CLE history):
-  • 4.
-    TwistorInformationClustering.nb [55] (spectrum extraction and plots)
-  • 5.
-    GeometricQCD CLE results [56] (the full set of histories of the evolution of $R(E)$ for $240$ values of energy $E\in(2.,3.)$ ran on 64 nodes of Typhon cluster, with random initial states)
-The raw per-replica output consists of node-indexed binary arrays (e.g.
-TwistorHistory*.mx) storing, for each sampled energy value, the whole stochastic trajectory of observable resolvent $R(E)$.
-These files are verified by VerifyTwistorHistory.nb, testing fat tails of distribution against theoretical prediction.
-Then TwistorInformationClustering.nb performs robust pooling across independent replicas, analyzes the history of the phase of the resolvent, builds the statistical distribution, clusters the data by information distance between these phase distributions and produces the final spectral curves and extracted mass tables.
-This public release is intended to provide a practical basis for future large-scale runs (larger ensembles, finer discretization, and improved CLE diagnostics and stabilization methods).
+No data was created in this paper. The analytic computations of the Magnus expanstion of the Momentum loop equations were performed in the Mathematica ® notebook published on the Wolfram Cloud[37].
 Appendix A Conformal anomaly in determinants
@@
 The full basis construction, shuffle-ideal reduction, and exact linear-algebra verification—including the generation of the matrices $A_{n}$ and the explicit null-vector projections—are provided in the executable Mathematica notebook [37] (see Data Availability).
-Appendix D Algorithm for Fast Spectral-Gaussian Integration
-To evaluate the Liouville action in the bulk, we employ a spectral-product quadrature rule that combines the spectral accuracy of the boundary discretization with high-precision Gaussian integration in the radial direction.
-This method avoids coordinate singularities at the origin and preserves the holomorphicity of the extension.
-The integral of the Liouville Lagrangian density over the unit disk is computed in polar coordinates:
-$S_{\text{Liouville}}=\int_{0}^{1}rdr\int_{0}^{2\pi}d\theta\,L\left(\rho(r,\theta),|\partial_{z}\rho(r,\theta)|^{2}\right)$ (D.1)
-We discretize this domain using a product grid defined by the $N$ roots of unity in the angular direction and $M$ Gauss-Legendre nodes in the radial direction.
-1. The Radial Grid (Gauss-Legendre): We map the standard Legendre nodes $x_{m}\in[-1,1]$ (roots of the Legendre polynomial $P_{M}(x)$) to the radial interval $r\in(0,1]$:
-$r_{m}=\frac{1+x_{m}}{2},\qquad w_{m}^{\prime}=\frac{w_{m}}{2}$ (D.2)
-where $w_{m}$ are the standard weights. This quadrature is exact for polynomials of degree 2M12M-1. Since the origin $r=0$ is not a node, the coordinate singularity is naturally avoided.
-2.
-The Angular Grid (Roots of Unity): For the angular integration, we use the trapezoidal rule on the $N$ vertices of the polygon, $\theta_{k}=2\pi k/N,\omega_{k}=e^{\mathrm{i}\theta_{k}}$.
-For periodic functions (such as our fields on circles of constant radius), this rule exhibits exponential convergence (spectral accuracy).
-3.
-Bulk Field Evaluation: The values of the (anti)holomorphic null-vector fields $W(z)=\left\{V(z),\tilde{V}(\bar{z})\right\}$ at a grid point $z_{mk}=r_{m}\omega_{k}$ are computed using the analytic extension formula (E.3).
-Crucially, the term $(z_{mk}/\omega_{j})^{N}$ reduces to $r_{m}^{N}$, which is independent of the angular indices.
-This simplifies the kernel significantly:
-$V(r_{m}\omega_{k})=\frac{1-r_{m}^{N}}{N}\sum_{j=0}^{N-1}\frac{V(\omega_{j})}{1-r_{m}\omega_{k}/\omega_{j}};$ (D.3)
-$\tilde{V}(r_{m}\bar{\omega}_{k})=\frac{1-r_{m}^{N}}{N}\sum_{j=0}^{N-1}\frac{\tilde{V}(\omega_{j})}{1-r_{m}\bar{\omega}_{k}/\omega_{j}}$ (D.4)
-This summation takes the form of a discrete convolution.
-By choosing $N$ to be a power of 2 (e.g., $N=128$), this convolution can be accelerated via Fast Fourier Transform (FFT) to $O(N\log N)$ operations per radial ring, providing a massive speedup over the naive $O(N^{2})$ summation.
-4.
-Hybrid Algorithm for Analytic Derivative: The kinetic term in the Liouville action requires the holomorphic derivative $\partial_{z}\Lambda$.
-To maintain precision, we differentiate the interpolating polynomial analytically.
-Differentiating (E.3) with respect to $z$ yields:
-$\partial_{z}V(z)\Big|_{z_{mk}}=\frac{1}{N}\sum_{j=0}^{N-1}V(\omega_{j})/\omega_{j}K^{\prime}\left(z_{mk}/\omega_{j}\right)$ (D.5)
-where $K^{\prime}(w)$ is the derivative of the geometric series kernel. To optimize computational complexity without sacrificing stability near the boundary, we employ a hybrid algorithm:
-a) The Fractional Form (Bulk): For the majority of grid points in the bulk ($1-r_{m}>\epsilon\sim 0.1$), we use the closed fractional form:
-$K^{\prime}(w)=\frac{(1-w^{N})+Nw^{N-1}(1-w)}{(1-w)^{2}}$ (D.6)
-This form is algebraically equivalent to the standard derivative but is numerically rearranged to avoid cancellations.
-Its evaluation is $O(1)$, ensuring the total derivative calculation remains efficient.
-b) The Geometric Doubling Algorithm (Boundary Layer): For the outermost radial nodes where $r_{m}>1-\epsilon$, the fractional form suffers from numerical instability due to the denominator $(1-w)^{2}$.
-In this "skin" layer, we switch to the exact polynomial sum $K^{\prime}(w)=\sum_{n=1}^{N-1}nw^{n-1}$.
-Crucially, for $N=2^{k}$, this sum can be evaluated in $O(\log N)$ operations using the recursive doubling identity for geometric series:
-$S_{2N}(w)=(1+P_{N}(w))S_{N}(w);$ (D.7)
-$P_{2N}(w)=P^{2}_{N}(w);$ (D.8)
-$S_{1}(w)=1;\qquad P_{1}(w)=w;$ (D.9)
-and its derivative by $w$. Here is the iterative algorithm:
-Algorithm 1 Fast Geometric Doubling Algorithm ($O(\log N)$)
-0: Complex argument $w$, Iterations $k$ (where $N=2^{k}$)
-0: Kernel values $K=\sum_{n=0}^{N-1}w^{n}$ and $K^{\prime}=\sum_{n=1}^{N-1}nw^{n-1}$
-1:Initialize:
-[…]
-[…]
-[…]
-[…]
-6:for $i=1$ to $k$ do
-7:  $K^{\prime}\leftarrow(1+P)K^{\prime}+P^{\prime}K$ {Update Derivative of Sum (Product Rule)}
-8:  $K\leftarrow(1+P)K$ {Update Sum ($S_{2M}=(1+w^{M})S_{M}$)}
-9:  $P^{\prime}\leftarrow 2\cdot P\cdot P^{\prime}$ {Update Derivative of Power}
-10:  $P\leftarrow P\cdot P$ {Update Power ($w^{2M}=(w^{M})^{2}$)}
-11:end for
-12:return $K,K^{\prime}$
-This iterative algorithm computes both the kernel and its derivative with logarithmic complexity $O(\log N)$, avoiding the $O(N)$ cost of the direct sum while using $O(1)$ memory and maintaining perfect numerical stability near the boundary.
-The numerical stability of this algorithm is verified in our Mathematica ® code [52].
-5. Summation: The total action is obtained by summing the weighted contributions:
-$S\approx\frac{2\pi}{N}\sum_{m=1}^{M}w_{m}^{\prime}r_{m}\sum_{k=0}^{N-1}L_{\text{Liouville}}(r_{m},\theta_{k})$ (D.10)
-This algorithm ensures that the bulk integration error is on the same order as the boundary discretization error, providing a stable and consistent numerical definition of the rigid twistor string.
-In our implementation (using Mathematica), the kernel data is precomputed into parallel tables, further accelerating the simulation.
-Appendix E Numerical Implementation of CLE
-To resolve the severe sign problem inherent in the Twistor String path integral, we employ Stochastic Quantization via the Complex Langevin Equation (CLE) [42, 60].
-E.1 Complexification of the parameter space
-To ensure the strict analyticity of the stochastic evolution required by the Parisi condition [42], we must eliminate all non-holomorphic operations from the Langevin equations.
-We achieve this by complexifying the target manifold and analytically continuing the action integral itself.
-This ensures that the drift forces are partial derivatives of a globally holomorphic function.
-We replace the original real and Hermitian variables with a set of independent holomorphic variables.
-The angular coordinates $\alpha_{k}$ are mapped to the complex plane via $z_{1}=e^{i\alpha_{1}},,z_{2}=e^{i\alpha_{2}}$.
-The spinor fields and their conjugates are treated as independent complex vectors.
-The full set of dynamical variables for the discretized loop is:
-$\mathcal{V}=\{\{z_{1},z_{2}\},\{\rho_{k},\xi_{k},\eta_{k},\tilde{\xi}_{k},\tilde{\eta}_{k},\gamma_{1,k},\gamma_{2,k}\}_{k=1}^{N}\}.$ (E.1)
-Here, $\tilde{\xi}$ and $\tilde{\eta}$ replace $\bar{\xi}$ and $\bar{\eta}$ as independent variables.
-The geometric constraints are enforced by the complex Lagrange multipliers $\gamma_{1,k}$ and $\gamma_{2,k}$, leading to the holomorphic constraint terms in the effective action:
-$S_{\text{constr}}=i\sum_{k}\left[\gamma_{1,k}(\tilde{\xi}_{k}\xi_{k}-1)+\gamma_{2,k}(\tilde{\eta}_{k}\eta_{k}-1)\right].$ (E.2)
-E.2 Spectral Interpolation of Fields
-To evaluate the action integrals analytically, we construct global holomorphic polynomials that interpolate the values at the vertices $z_{k}$.
-E.2.1 Spinor Fields
-Let $\Lambda^{(k)}=e^{\rho_{k}/2}(\xi_{k},\eta_{k})$ denote the spinor data $\Lambda(z)=\{\lambda(z),\mu(z)\}$ at vertex $k$. The Lagrange interpolation polynomial on the spectral unit circle is:
-$\Lambda(z)=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}K\left(ze^{-\mathrm{i}\theta_{k}}\right);$ (E.3)
-$K(z)=\frac{1-z^{N}}{1-z};$ (E.4)
-Using the Fast Fourier Transform (FFT), we extract the spectral coefficients $L_{n},M_{n}$ (and $\tilde{L}_{n},\tilde{M}_{n}$ from the independent $\tilde{\xi},\tilde{\eta}$ variables) such that:
-$\lambda(z)=\sum_{n=0}^{N-1}L_{n}z^{n},\qquad\tilde{\lambda}(z)=\sum_{n=0}^{N-1}\tilde{L}_{n}z^{n}.$ (E.5)
-$\{L_{n},M_{n}\}=\frac{1}{N}\sum_{k=0}^{N-1}\Lambda^{(k)}e^{-\mathrm{i}\theta_{k}}$ (E.6)
-E.2.2 Liouville Field
-The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an exact formula
-$\rho(z,\bar{z})=\frac{\log\tilde{\lambda}(\bar{z})\lambda(z)+\log\tilde{\mu}(\bar{z})\mu(z)}{2}$ (E.7)
-where $\lambda(z),\mu(z)$} is extended by interpolation (E.3).
-The derivatives of the above kernel $K^{\prime}(z)$ involved in the Liouville Lagrangian with this $\rho$ were computed in the previous sections.
-These expressions provide the values of the fields and derivatives at any complex $z$, not just at the vertices.
-These expressions also depend upon complex parameters in $\mathcal{V}$ through the Fourier coefficients $\Lambda(k),\tilde{\Lambda}(k)$.
-This dependence is exponential for $\rho_{k}$ and polynomial for the rest of variables in $\mathcal{V}$, which allows for required analytic continuation of the CLE.
-E.3 The Boundary Terms in the action
-[…]
-$G_{p}(z_{1},z_{2})\equiv\int_{z_{1}}^{z_{2}}z^{p}\frac{dz}{iz}=\frac{z_{2}^{p}-z_{1}^{p}}{ip}.$ (E.8)
-With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is strictly positive, ensuring there are no singularities in the complex plane.
-The contribution to the action of external momentum $q$ injected into the quark loop from the arc between $z_{1}$ and $z_{2}$ is:
-$S_{q}=-\mathrm{i}q_{\alpha}$
-$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}\sigma_{\alpha}M_{m})-(\tilde{L}_{n}\sigma^{\dagger}_{\alpha}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (E.9)
-Then, there are terms related to the constraints on the $\xi,\eta$ variables. These terms are just discrete sums
-$S_{\gamma}=\mathrm{i}\sum_{k=0}^{N-1}\gamma_{1,k}\left(\tilde{\xi}_{k}\xi_{k}-1\right)+\gamma_{2,k}\left(\tilde{\eta}_{k}\eta_{k}-1\right)$ (E.10)
-In addition, there are terms related to the trace (12.30). Our complexification reduces this trace to
-[…]
-[…]
-$\mathcal{P}(\theta)=2\mathrm{i}\left[z\eta\tilde{\xi}^{T}-\bar{z}\xi\eta^{T}\right]_{z=e^{\mathrm{i}\theta}}$ (E.13)
-Finally, our reweighting adds one more complex term to the action,
-$S_{O}=\log\mathcal{O};$ (E.14)
-$\mathcal{O}=\frac{\partial S_{q}}{\partial q_{4}}=$
-$\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}M_{m})-(\tilde{L}_{n}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2}).$ (E.15)
-All the above expressions are purely holomorphic functions of all the variables in $\mathcal{V}$.
-E.4 The Bulk Action (Gauss-Legendre Integration)
-The bulk action $S_{\rho}$ involves the kinetic and potential terms for the Liouville field of the form: $\int d^{2}z[\partial_{z}\rho\partial_{\bar{z}}\rho+V(\rho)]$.
-Unlike the boundary term, this cannot be integrated in closed form due to the nonlinear dependence of both variables $z,\bar{z}$.
-We evaluate this integral using $M$-point Gauss-Legendre quadrature coupled with spectral (FFT) summation over the roots of unity for the angular integration.
-The integration domain is divided into $N$ rays $R:(0,z_{k});\,z_{k}=e^{\mathrm{i}\theta_{k}}$.
-Then, $M$ quadrature nodes $z_{q}\in(0,1)$ are placed on each ray: $z_{k,q}=z_{q}e^{\mathrm{i}\theta_{k}}$.
-The bulk action is approximated as:
-$S_{\rho}\to\frac{1}{N}\sum_{k=0}^{N-1}\sum_{q=1}^{M}w^{\prime}_{q}\mathcal{L}_{\text{bulk}}\left(\rho(z_{k,q}),\rho^{\prime}(z_{k,q})\right).$ (E.16)
-Here, the values of the field $\rho(z,\bar{z})$ and its derivative $\rho^{\prime}(z,\bar{z})$ at the quadrature points $z_{k,q}$ are computed exactly using the spectral Lagrange interpolation formula (Eq.
-E.3) derived in the previous section.
-This combination ensures exact integration of the polynomials in $z,\bar{z}$ of degree up to $2M$ and exponential accuracy of angular integrations.
-For practical purposes, this means that the integration errors will be negligible or $N,M=64$.
-This, of course, does not imply the exponential conversion of the CLE to the path integral at large $N,M$.
-The CLE introduces errors of its own, and statistical errors $O(1/\sqrt{T})$ related to the cancelation of oscillating terms are the dominant.
-E.5 Holomorphic Langevin Equations
-The time evolution is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{\text{total}}=S_{\alpha}+S_{\gamma}+S_{T}+S_{\rho}+S_{O}$.
-  1. 1.
-    Coordinate Evolution ($z_{k}$): The force on $z_{k}$ arises from the boundary terms of the $G$ functions in $S_{\text{int}}$ and the integration limits in $S_{\text{bulk}}$.
-    $\partial_{\tau}z_{k}=-z_{k}^{2}\frac{\partial S_{\text{total}}}{\partial z_{k}}+iz_{k}\nu_{z_{k}}.$ (E.17)
-    (The factor $z_{k}^{2}$ arises from the metric of the map $\alpha=-i\log z$).
-  2. 2.
-    Spinor Evolution ($\xi_{k},\tilde{\xi}_{k}$, etc.): The forces are computed via the chain rule through the spectral coefficients $L_{n}$. For example:
-    $\partial_{\tau}\xi_{k}=-\frac{\partial S_{\text{total}}}{\partial\xi_{k}}-i\gamma_{1,k}\tilde{\xi}_{k}+\nu_{\xi_{k}}.$ (E.18)
-    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables.
-  3. 3.
-    Liouville Evolution ($\rho_{k}$): The force on $\rho_{k}$ is nonlocal, mediated by the interpolation kernel $K_{k}(z)$ appearing in both the bulk quadrature points and the spectral coefficients of the boundary term.
-    $\partial_{\tau}\rho_{k}=-\frac{\partial S_{\text{total}}}{\partial\rho_{k}}+\nu_{\rho_{k}}.$ (E.19)
-  4. 4.
-    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the constraints:
-    $\partial_{\tau}\gamma_{1,k}=-i(\tilde{\xi}_{k}\xi_{k}-1)+\nu_{\gamma}.$ (E.20)
-By utilizing this spectral action, we ensure that the drift terms are analytic functions of all variables $\mathcal{V}$ throughout the complex plane.
-This prevents the "excursion" problem common in standard CLE, where non-analytic projections onto the real axis lead to incorrect convergence.
-E.6 Reweighted Spectral Search for Meson Masses
-The poles of $\Pi(E)$ correspond to the bound state masses (the meson spectrum), and these poles also become poles of its log derivative $R(E)$.
-As discussed in Section 15, directly searching for these poles in a stochastic simulation is numerically disastrous, as the infinite variance near the singularity destroys the Markov chain.
-Instead, we employ a reweighting method to compute the inverse ratio $1/R(E)$ and search for its zeros.
-This inversion mathematically transforms the meromorphic amplitude into an entire function, effectively neutralizing the singularities.
-We define the observable as the derivative of the action with respect to energy:
-$\mathcal{O}=\frac{\partial S_{\text{eff}}}{\partial E}$
-$=-\mathrm{i}\sum_{n=0}^{N-1}\sum_{m=0}^{N-1}\left[(L_{n}M_{m})-(\tilde{L}_{n}\tilde{M}_{m})\right]G_{n+m+1}(z_{1},z_{2})$ (E.21)
-Then, we have an identity
-$\Pi^{\prime}(E)/\Pi(E)=\frac{\int\mathcal{O}e^{S_{\text{eff}}}}{\int e^{S_{\text{eff}}}}$ (E.22)
-We then introduce a Modified Action $\tilde{S}$ by shifting the probability measure:
-$\tilde{S}=S_{\text{eff}}+\ln\mathcal{O}$ (E.23)
-Using this modified action, the inverse ratio can be expressed as a statistical average:
-$\frac{1}{R(E)}=\frac{\Pi(E)}{\Pi^{\prime}(E)}=\frac{\int e^{S_{\text{eff}}}}{\int\mathcal{O}e^{S_{\text{eff}}}}=\frac{\int\frac{1}{\mathcal{O}}e^{\tilde{S}}}{\int e^{\tilde{S}}}=\left\langle\frac{1}{\mathcal{O}}\right\rangle_{\tilde{S}}$ (E.24)
-Thus, the algorithm proceeds as follows:
-  1. 1.
-    We simulate the Complex Langevin evolution governed by the modified action $\tilde{S}$.
-This introduces an additional "logarithmic force" to the drift: $\nabla\tilde{S}=-\nabla S_{\text{eff}}-\frac{1}{\mathcal{O}}\nabla\mathcal{O}$.
-  2. 2.
-    At each time step, we measure the quantity $1/\mathcal{O}$.
-  3. 3.
-    We compute the time average $\langle 1/\mathcal{O}\rangle_{\tilde{S}}$.
-  4. 4.
-    We scan the energy $E$ and identify the meson masses as the points where this average vanishes (crosses zero).
-This method transforms the search for poles into a search for zeros, which is numerically stable and precise.
-E.7 Spectral Extraction and the Sign of the Slope
-A crucial subtlety arises in interpreting the zeros of the measured observable.
-The simulation computes the expectation value of the inverse resolvent operator, $\langle\mathcal{O}^{-1}\rangle\sim\text{Re}[1/R(E)]$.
-To distinguish physical mass states from artifacts, we must analyze the pole structure of the resolvent.
-The resolvent is related to the logarithmic derivative of the vacuum polarization $\Pi(E)$:
-$R(E)\sim\frac{\Pi^{\prime}(E)}{\Pi(E)}=\frac{d}{dE}\ln\Pi(E)$ (E.25)
-This function has singularities of two types:
-  1. 1.
-    Physical Poles: At a particle mass $E\approx m$, the polarization behaves as $\Pi(E)\sim\frac{A}{E-m}$. Consequently, the resolvent behaves as a simple pole with residue 1-1:
-    $R(E)\approx\frac{-1}{E-m}\implies\frac{1}{R(E)}\approx m-E$ (E.26)
-  2. 2.
-    Polarization Zeros: Between two physical poles, $\Pi(E)$ must cross zero. Near such a zero $E\approx E_{0}$, $\Pi(E)\sim C(E-E_{0})$. The resolvent behaves as a simple pole with residue +1+1:
-    $R(E)\approx\frac{1}{E-E_{0}}\implies\frac{1}{R(E)}\approx E-E_{0}$ (E.27)
-Therefore, when plotting the real part of the inverse resolvent $\text{Re}[1/R(E)]$ against energy, both physical masses and polarization zeros appear as zero-crossings.
-However, they are distinguished by the sign of the slope:
-  • 1.
-    Negative Slope ($\frac{d}{dE}<0$): Corresponds to a physical particle mass (Pole of $\Pi$).
-  • 2.
-    Positive Slope ($\frac{d}{dE}>0$): Corresponds to a zero of the polarization operator (Artifact).
-Consequently, the meson mass spectrum is identified exclusively by the zero-crossings of $1/R(E)$ possessing a negative gradient.
-Let us note in passing that this function $\Pi(E)$ in the confined planar QCD we are solving is a meromorphic function of $q^{2}=-E^{2}$ with poles on the negative axis with positive residues:
-$\Pi(E)=\sum_{n}\frac{Z_{n}}{m_{n}^{2}-E^{2}}$ (E.28)
-We also know that asymptotically, the residues tend to $Z_{n}\propto\partial_{n}(m_{n}^{2})$ which leads to logarithmic asymptotic behavior in the deep Euclidean limit (Asymptotic freedom):
-$\Pi(E)\to\int dn\frac{Z_{n}}{m_{n}^{2}+q^{2}}$
-$\propto\int_{\Lambda^{2}_{QCD}}^{\infty}d\sigma\frac{1}{q^{2}+\sigma}\propto\textit{ const }{}+\log\frac{q^{2}}{\Lambda^{2}_{QCD}}$ (E.29)
-For our resolvent, these analytic properties mean that $1/R(E)$ is an odd entire function, with zeros along the real axis and no singularities in the complex plane.
-E.8 Stabilizing CLE by Reweighting
-The application of the Complex Langevin Equation (CLE) to theories with severe sign problems is often described as more art than science.
-Unlike the real Langevin equation, where the drift is derived from a real, growing at infinity action and convergence to the Boltzmann distribution is guaranteed by rigorous theorems, the CLE operates in a complexified phase space.
-There is no a priori guarantee that the stochastic process will converge to a finite limit, nor that it will sample the correct distribution.
-A common failure mode is the “excursion problem,” where the holomorphic variables wander into non-physical regions of the complex plane where the action is not holomorphic or simply diverges, causing the simulation to explode.
-We observed precisely this instability when attempting a direct simulation of the standard effective action $S_{\text{eff}}$.
-As shown in our control tests, without modification, the magnitude of the observable $\mathcal{O}$ diverges exponentially (reaching magnitudes of $10^{40}$), rendering the extraction of physical data impossible.
-The drift forces derived solely from $S_{\text{eff}}$ are insufficient to confine the system to the relevant integration domain.
-The reweighting method, introduced primarily to convert the search for poles into a search for zeros, provides a crucial secondary benefit: dynamical stabilization.
-By simulating the modified action $\tilde{S}=S_{\text{eff}}+\ln\mathcal{O}$, the drift force receives an additional contribution:
-$K_{\text{drift}}=-\nabla S_{\text{eff}}-\frac{1}{\mathcal{O}}\nabla\mathcal{O}.$ (E.30)
-This “logarithmic force” acts as a dynamical potential barrier.
-When the system approaches a region where $\mathcal{O}$ vanishes (a zero of the inverse ratio), the force becomes singular and repulsive, preventing the trajectory from crossing the zero but allowing it to explore the vicinity.
-Conversely, the modification of the measure appears to suppress the runaway excursions seen in the unmodified theory.
-E.9 Numerical Integration: Second-Order SDE and Adaptive Braking
-While the logarithmic reweighting ($\tilde{S}=S_{eff}+\log\mathcal{O}$) provides a continuous topological barrier against runaway trajectories, the discrete nature of numerical integration introduces a severe algorithmic vulnerability.
-In continuous Langevin time $\tau$, steepest descent flows cannot cross the meromorphic poles at $\mathcal{O}=0$.
-However, in a discrete simulation, the holomorphic drift force $K(\mathcal{V})=-\nabla\tilde{S}$ strictly diverges near the zero-locus.
-A naive first-order Euler-Maruyama discretization with a fixed time step $\Delta\tau$ would generate a massive coordinate displacement $\Delta\mathcal{V}=K(\mathcal{V})\Delta\tau$ in the vicinity of the pole.
-This would cause the discrete trajectory to violently “jump over” the topological barrier, catapulting the system directly into the runaway Stokes wedges or triggering arithmetic overflow.
-To rigorously preserve the integrity of the complexified geometry and enforce the topological barriers during discrete evolution, our numerical implementation relies on two synergistic stabilization algorithms: a second-order Stochastic Differential Equation (SDE) integrator and an adaptive drift clipping mechanism.
-First, we upgrade the integration scheme from the standard Euler method to the second-order Stochastic Heun (predictor-corrector) algorithm.
-For a generic complex state vector $\mathcal{V}$, the discrete update proceeds in two stages:
-$\mathcal{V}_{pred}=\mathcal{V}(\tau)+\Delta\mathcal{V}_{1}+\eta(\tau)\sqrt{\Delta\tau}\qquad\text{(Predictor)}$ (E.31)
-$\mathcal{V}(\tau+\Delta\tau)=$
-$\mathcal{V}(\tau)+\frac{1}{2}\Big(\Delta\mathcal{V}_{1}+\Delta\mathcal{V}_{2}\Big)+\eta(\tau)\sqrt{\Delta\tau}\qquad\text{(Corrector)}$ (E.32)
-where $\Delta\mathcal{V}_{1}=K(\mathcal{V}(\tau))\Delta\tau$ and $\Delta\mathcal{V}_{2}=K(\mathcal{V}_{pred})\Delta\tau$.
-The complexified Gaussian noise $\eta(\tau)$ is identical in both steps.
-By sampling the drift at both the current state and the projected future state, the predictor-corrector averaging intrinsically accounts for the severe local curvature of the effective action, providing vastly superior stability when navigating the highly non-linear valleys of the Lefschetz thimbles.
-Second, to absolutely forbid discrete overshoots near the meromorphic poles and the bounding walls of the non-compact Liouville directions, we implement dynamic “braking” via adaptive drift clipping.
-At both the predictor and corrector stages, we compute the maximum proposed deterministic displacement, $\delta_{max}=\max|\Delta\mathcal{V}_{i}|$, across all degrees of freedom.
-If this proposed step exceeds a strict geometric safety threshold $\epsilon$ (in our production runs, $\epsilon=0.05$), we isotropically rescale the entire drift vector:
-$\Delta\mathcal{V}_{i}\to\Delta\mathcal{V}_{i}\left(\frac{\epsilon}{\delta_{max}}\right)\qquad\text{for }\delta_{max}>\epsilon$ (E.33)
-This operation is mathematically equivalent to an adaptive, localized reduction of the Langevin time step ($\Delta\tau\to\Delta\tau_{eff}$) during violent excursions.
-Physically, this acts as an automatic braking system.
-In the flat, safe regions of the bulk phase space, the simulation explores efficiently with the full time step $\Delta\tau$.
-But the moment the complexified noise kicks the trajectory toward a singularity, the repulsive drift spikes, and the adaptive clipping safely throttles the velocity.
-The trajectory slows to a crawl, taking infinitesimally fine integration steps that smoothly trace the repulsive contours of the meromorphic pole.
-This ensures that the discrete trajectory gently reflects off the $1/\mathcal{O}$ topological barriers rather than numerically tearing through them, preserving the structural integrity of the complexified path integral over billions of consecutive updates.
-E.10 Analyzing intermittent histories of the Complex Langevin Equation
-The numerical evaluation of the twistor-string path integral via the Complex Langevin Equation (CLE) introduces profound statistical challenges.
-The physical observable required for the spectrum extraction, $1/\mathcal{O}$, is meromorphic.
-Consequently, during the stochastic evolution, the complexified trajectory occasionally wanders near the singularity at $\mathcal{O}=0$.
-These close encounters generate violent, intermittent numerical spikes—or excursions—in the micro-state history.
-Extracting the pristine physical signal from this complexified noise requires a rigorous pipeline of burn-in truncation, dynamic symmetry restoration, and Cauchy Principal Value error bounding.
-Thermalization and the Measurement Criterion
-Because the CLE simulation is initialized from an arbitrary complex configuration, the early micro-state history is dominated by a massive coordinate drift as the system relaxes toward the true vacuum.
-Including this non-equilibrium initial phase would severely pollute the physical averages.
-To rigorously determine the exact step $N_{\text{cut}}$ where this “burn-in” phase ends, we utilize the raw imaginary trajectory, Im[$\imaginary[1/\mathcal{O}]$, as an exact equilibrium thermometer.
-The exact, uncomplexified path integral respects a global $Z\leftrightarrow Z^{*}$ parity.
-Consequently, the steady-state Fokker-Planck distribution of the CLE must eventually become symmetric with respect to the real axis.
-The strict criterion for thermalization is therefore the cessation of macroscopic secular drift in the imaginary component.
-As illustrated in the full raw history (Fig.
-16, left), the measurement phase is initiated only after this trajectory stabilizes into a horizontal stochastic oscillation.
-We discard all data prior to this threshold (e.g., $N_{\text{cut}}=100000$).
-Figure 16: Left: The raw intermittent history of the CLE trajectory.
-The vertical dashed line marks the strict measurement criterion ($N_{\text{cut}}$) where macroscopic secular drift ceases.
-All prior data is discarded.
-Right: The naive cumulative arithmetic mean.
-Post-thermalization excursions permanently offset the mean, demonstrating the fragility of unweighted averaging.
-Dynamic Symmetry Restoration
-At finite Langevin times, the stochastic trajectory can become temporarily trapped in one complex half-plane, leading to a delayed restoration of the $Z\leftrightarrow Z^{*}$ parity.
-We bypass this topological tunneling lag through Dynamic Symmetry Restoration.
-Because the theoretical ensemble is invariant under complex conjugation, we explicitly symmetrize the post-burn-in history.
-By adding the complex conjugate of every generated micro-state to the dataset, we artificially and instantly enforce the parity symmetry, projecting the stochastic fluctuations entirely onto the real physical slice (Fig.
-17).
-Figure 17: Kernel Density Estimation of Im[$\imaginary[1/\mathcal{O}]$ during the valid measurement phase.
-The bounded peak near zero confirms statistical equilibrium.
-Any residual skewness is instantly annihilated by the $Z\leftrightarrow Z^{*}$ double-history symmetrization.
-The $1/|u|^{3}$ Singularity and the Principal Value
-The physical core of the real distribution, Re[$\real[1/\mathcal{O}]$, is highly multimodal (Fig.
-18, left), reflecting instanton-like tunneling between discrete topological sectors (Lefschetz thimbles).
-However, the excursions generate a severe statistical pathology.
-[…]
-$P(u)\approx\int p_{0}\,\delta\left(u-\frac{\cos\theta}{r}\right)r\,dr\,d\theta=\frac{\pi p_{0}}{2|u|^{3}}$ (E.34)
-This exact $1/|u|^{3}$ power-law behavior dictates that while the theoretical mean is finite, the variance diverges logarithmically ($\int u^{2}/|u|^{3}du\to\infty$).
-Consequently, the standard empirical arithmetic mean is formally ill-defined in finite samples; it is continuously hijacked by infinite-variance sample noise (Fig.
-16, right).
-Because the leading asymptotic tail is exactly symmetric (P(u)=P(u)P(u)=P(-u)), the divergent variance contributions from the extreme left and right excursions mathematically cancel.
-Therefore, the physical expectation value of the amplitude must be rigorously defined via the Cauchy Principal Value (PV).
-To extract this regularized observable, we utilize a Symmetric Trimmed Mean (clipping the extreme $q=5\%$ of positive and negative outliers).
-By symmetrically shearing off the infinite-variance tails, it acts as the exact numerical implementation of the Principal Value, isolating the finite thermodynamic center of mass over the central $90\%$ of the physical bulk.
-Figure 18: Upper: The multimodal physical core of Re[$\real[1/\mathcal{O}]$.
-The robust trimmed mean (red dashed line) isolates the true Principal Value.
-Lower: Log-scaled probability density demonstrating the severe asymptotic $1/|u|^{3}$ power-law tails (black dashed lines), which cause infinite variance and necessitate Cauchy Principal Value regularization.
-Exact Analytic Error Bounding and Regression Weights
-To rigorously bound the systematic statistical error introduced by discarding these numerical excursions, we exploit the analytic form of the asymptotic tails $P(u)=C/|u|^{3}$.
-For the right tail (values exceeding the 1q1-q quantile threshold $u_{R}$), the normalization constant is fixed by the trim fraction $q$:
-$\int_{u_{R}}^{\infty}\frac{C_{R}}{u^{3}}du=\frac{C_{R}}{2u_{R}^{2}}=q\implies C_{R}=2qu_{R}^{2}$ (E.35)
-The exact thermodynamic contribution of this infinite right tail evaluates to $\int_{u_{R}}^{\infty}u\left(2qu_{R}^{2}/u^{3}\right)du=2qu_{R}$.
-By identical logic, the left tail contribution evaluates to $2qu_{L}$.
-Thus, the exact reconstructed Principal Value mean is a parameter-free sum of the trimmed core ($\mu_{\text{core}}$) and the analytic tail boundaries:
-Re[$\langle\real[1/\mathcal{O}]\rangle_{\text{total}}=(1-2q)\mu_{\text{core}}+2q(u_{R}+u_{L})$ (E.36)
-The difference between the symmetric trimmed mean and this full analytic reconstruction, $\Delta_{\text{tail}}=2q|u_{R}+u_{L}-\mu_{\text{core}}|$, provides a strict mathematical measure of the finite-sample Principal Value asymmetry (Table 2).
-Beyond bounding the mean, the exact $1/|u|^{3}$ geometry strictly determines the statistical variance of the Markov chain.
-This variance is required for the optimal inverse-variance pooling of independent cluster nodes and for weighting the polynomial regression used to extract the mass spectrum.
-Because the probability density decays as $1/|u|^{3}$, the formal variance diverges logarithmically with the maximum Langevin excursion cutoff $L$.
-Evaluating the second moment for the right tail yields:
-$\sigma_{R}^{2}=\int_{u_{R}}^{L}u^{2}\left(\frac{2qu_{R}^{2}}{u^{3}}\right)du=$
-$2qu_{R}^{2}\log\left(\frac{L}{u_{R}}\right)\to 2qu_{R}^{2}\log L$ (E.37)
-Combining the identically regulated contribution from the left tail, the total empirical variance of the numerical history diverges logarithmically as:
-$\sigma^{2}\simeq 2q(u_{R}^{2}+u_{L}^{2})\log L$ (E.38)
-Crucially, because the maximum Langevin excursion extent $L$ acts as a universal scale factor across the parallel Monte Carlo ensemble, the formally divergent $\log L$ completely cancels out when constructing the normalized inverse-variance weights ($w_{i}\propto 1/\sigma_{i}^{2}$) for the polynomial regression.
-Therefore, the optimal statistical weight for each independent energy point is governed entirely by the parameter-free relative variance of its physical core:
-$\sigma_{\text{rel}}^{2}=q(u_{R}^{2}+u_{L}^{2})$ (E.39)
-By replacing heuristic error estimation with this exact geometric variance, the fitting procedure becomes statistically bulletproof.
-Any node that suffers a violent, parity-violating topological excursion will exhibit drastically widened boundaries ($u_{R}^{2}+u_{L}^{2}\to\infty$), naturally driving its statistical weight to exactly zero.
-This rigorously guarantees that the final extracted Regge trajectory is determined strictly by the thermalized, parity-preserving vacuum configurations of the Complex Langevin dynamics.
-Parameter Sym. Value
-Trimmed Core $\mu_{\text{core}}$ 0.210137-0.210137
-Left Cut ($5\%$) $u_{L}$ 2.103169-2.103169
-Right Cut ($95\%$) $u_{R}$ $1.230750$
-Analytic PV Mean $\mu_{\text{tot}}$ 0.276365-0.276365
-Asym. PV Error $\Delta_{\text{tail}}$ $\mathbf{\pm 0.066228}$
-Table 2: Analytic bounds on finite-sample Cauchy Principal Value (PV) asymmetry, derived from exact $1/|u|^{3}$ Langevin excursions.
-E.11 First computations of the meson spectrum
-Empirically, we find that this reweighted CLE is numerically stable over extremely long integration times.
-In production runs extending to $10^{6}$ time steps per node per energy value (corresponding to a Langevin time of $\tau\sim 10^{4}$ ), the observable $1/\mathcal{O}$ remains finite and oscillates within a bounded range (typically $\in[-20,20]$), with no sign of the divergences characterizing the unweighted dynamics.
-While there is no general theorem proving that logarithmic reweighting stabilizes all complex actions, in the specific context of the Twistor String, it effectively regularizes the stochastic evolution, allowing for high-precision spectroscopy.
-Our cumulative statistics on all nodes for all $480$ energy points per energy value is about one billion data points.
-The lack of instabilities in such a large dataset speaks for some form of statistical equilibrium.
-The code implementing the above algorithm was written in Mathematica ® [52] and simulated on the IAS supercluster Typhon using a high-precision grid: $64$ angular nodes and $32$ radial nodes in the Gauss-Legendre computation of the Liouville energy integral.
-We employed a Langevin time step $dt=0.0005$ and averaged the inverse resolvent $1/R(m)$ over $48$ physical hours, following a thermalization period of $2.4$ hours.
-The number of time steps was approximately $T_{tot}=17M$, and the total Langevin time spent was $\tau=T_{tot}*dt\approx 8500$.
-The computation was parallelized across $64$ nodes of the Typhon cluster.
-Each node, equipped with $24$ cores, ran an independent Langevin evolution initialized with a unique random seed.
-The energy spectrum was sampled over $N_{E}=10\times N_{\text{cores}}=240$ values equally spaced between $2.$ and $3.0$ in units of string tension $\sigma$.
-The calculation was computationally intensive, requiring two days of wall-clock time on each of the 64 nodes to achieve the necessary convergence.
-The results of this simulation are presented in Figures 16, 18 and 15.
-The first plot displays the full history of real part of the inverse resolvent, $\text{Re}[1/R(m)]$, for a particular energy parameter $E=2.27197$.
-The second plot is the distribution of this real part over the measurement period (after first $100K$ steps).
-The third plot displays the roots of $1/R(E)$ with a negative slope in a form of the Regge trajectory $n$ vs $m_{n}^{2}$.
-As derived in Section E.7, the physical meson mass spectrum is identified by the zero-crossings of this function that possess a negative slope.
-There are two different branches of Regge trajectory, both close to linear (note that $n$ represents the internal excitation number, not the angular momentum).
-The label distinguishing these two branches corresponds to the internal quantum number of the $\bar{q}q$ states in our Confining Twistor theory.
-We identified these two branches by the information clustering method described in Section 16.
-The simulation demonstrates certain stability:
-  • 1.
-    Robustness: The process remained bounded without "runaways" or numerical explosions, confirming the efficacy of the holographic bulk forces in stabilizing the non-compact gauge variables.
-  • 2.
-    Precision: The histograms are smooth and have symmetric power tails, matching theoretical prediction $1/|u|^{3}$
-  • 3.
-    Spectrum: Information clustering of negative crossings by information produced two Regge trajectories, approximately linear with the same slope.
-While the current simulation parameters are sufficient to establish a proof of concept, we acknowledge that a detailed spectroscopic analysis—particularly to resolve the fine structure of high-mass resonances—requires higher resolution and significantly larger statistics.
-However, the existing code [52] is fully capable of addressing this challenge without modification.
-The implementation is highly efficient and parallelized, maintaining numerical stability even at the high grid densities required for Gauss-Legendre integration, angular FFTs, and Lagrange interpolation.
-Appendix F Canonical Twistors and the Minkowski Helicoid Metric
+Appendix D Chiral Twistors and the Minkowski Helicoid Metric
 In this appendix, we explicitly derive the macroscopic worldsheet metric of the rotating string directly from the microscopic twistor fields.
-We demonstrate how the canonical twistor parameterization enforces a dimensional reduction to a 3D hyperplane, and how the double Wick rotation analytically continues the Euclidean twistor metric into the physical Minkowski Nambu-Goto metric.
+We demonstrate how the chiral twistor parameterization breaks the dimensional reduction to a 3-dimensional hyperplane, and how the physical parity-even projection of the double Wick rotation analytically continues the Euclidean twistor metric into the physical Minkowski Nambu-Goto metric.
 We then pedantically evaluate the resulting bulk action integral.
-F.1 Canonical Twistors and Dimensional Reduction
+D.1 Chiral Twistors and Dimensional Breakout
 The canonical Helicoid is generated by twistors endowed with fractional powers of the worldsheet coordinate $z$.
 This fractional scaling is a spin-structure gauge choice that explicitly enforces the Neveu-Schwarz anti-periodic boundary conditions for the worldsheet fermions as they complete a full rotation around the string center.
+To construct the chiral (Hodge-dual) surface, we set the right twistor to a constant basis spinor.
 We define the left and right twistors as:
 $\lambda(z)=e^{\mathrm{i}\pi/4}z^{-1/2}\begin{pmatrix}1/z\\ 1\end{pmatrix}$,
-$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ -z\end{pmatrix}$ (F.1)
+$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ 0\end{pmatrix}$ (D.1)
 Because $\lambda\propto z^{-1/2}$ and $\mu\propto z^{1/2}$, their product has no overall fractional branching.
 The resulting bosonic physical observables are strictly rational, making this configuration geometrically equivalent to a rational single-pole state.
-The defining geometric characteristic of this canonical configuration is its behavior under the symmetric Pauli matrices.
+The defining geometric characteristic of this configuration is its behavior under the symmetric Pauli matrices.
 The spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$.
@@
 $f^{\prime}_{1}\propto\lambda^{T}\sigma_{1}\mu=\lambda_{1}\mu_{2}+\lambda_{2}\mu_{1}$
-$=\mathrm{i}\left[\left(\frac{1}{z}\right)(-z)+(1)(1)\right]=\mathrm{i}(-1+1)=0$ (F.2)
-Because the tangent vector component $f^{\prime}_{1}$ vanishes identically everywhere on the worldsheet, the target space coordinate $X_{1}$ is strictly constant.
-This exact algebraic constraint dynamically crushes the generic 4-dimensional minimal surface into a 3-dimensional hyperplane, forming the exact topological skeleton of the macroscopic 3D Helicoid.
-F.2 The Euclidean Metric and Double Wick Rotation
+$=\left(e^{\mathrm{i}\pi/4}z^{-3/2}\right)(0)+\left(e^{\mathrm{i}\pi/4}z^{-1/2}\right)\left(\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\right)=\mathrm{i}$ (D.2)
+Unlike the generic configuration where this component vanished identically everywhere on the worldsheet, here the tangent vector component $f^{\prime}_{1}$ evaluates to a non-zero constant.
+Consequently, the target space coordinate $X_{1}$ is no longer strictly constant.
+This permanently breaks the exact algebraic constraint that previously crushed the generic minimal surface into a 3-dimensional hyperplane, meaning this chiral surface genuinely spans a 4-dimensional embedding.
+D.2 The Euclidean Metric and Parity Projection
 To extract the physical string metric, we compute the induced conformal factor $\exp\left(2\rho\right)=|\lambda|^{2}|\mu|^{2}$. Restoring the dimensionful scale parameter $R$, the squared norms are:
 $|\lambda|^{2}=\frac{R}{|z|}\left(\frac{1}{|z|^{2}}+1\right)=R\frac{1+|z|^{2}}{|z|^{3}},$
-$|\mu|^{2}=R|z|\left(1+|z|^{2}\right)$ (F.3)
+$|\mu|^{2}=R|z|(1+0)=R|z|$ (D.3)
 When multiplied, the fractional $|z|$ weights cancel perfectly:
-$\exp\left(2\rho\right)=R^{2}\frac{(1+|z|^{2})^{2}}{|z|^{2}}=R^{2}\left(|z|+\frac{1}{|z|}\right)^{2}$ (F.4)
+$\exp\left(2\rho\right)=R^{2}\frac{1+|z|^{2}}{|z|^{2}}=R^{2}\left(1+\frac{1}{|z|^{2}}\right)$ (D.4)
 To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$.
@@
 Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
-$\exp\left(2\rho\right)=R^{2}\left(e^{\sigma_{E}}+e^{-\sigma_{E}}\right)^{2}=4R^{2}\cosh^{2}(\sigma_{E})$ (F.5)
+$\exp\left(2\rho\right)=R^{2}\left(1+e^{-2\sigma_{E}}\right)$ (D.5)
 Thus, the induced Euclidean worldsheet metric is exactly:
-$ds^{2}_{E}=\exp\left(2\rho\right)\frac{dzd\bar{z}}{|z|^{2}}=4R^{2}\cosh^{2}(\sigma_{E})\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (F.6)
+$ds^{2}_{E}=\exp\left(2\rho\right)(d\sigma_{E}^{2}+d\tau_{E}^{2})$
+$=R^{2}\left(1+e^{-2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
 To evaluate the physical states, we analytically continue this metric into Minkowski spacetime via a double Wick rotation.
@@
 We perform a Wick rotation on the spatial worldsheet coordinate into the physical light-cone regime:
-$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (F.7)
-Under this transformation, the hyperbolic cosine strictly rotates into the trigonometric cosine: $\cosh(\mathrm{i}\sigma_{M})=\cos(\sigma_{M})$.
-Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact Minkowski metric is:
-$ds^{2}_{M}=4R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (F.8)
+$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (D.7)
+Under this transformation, the exponential strictly rotates into a complex phase: $e^{-2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})-\mathrm{i}\sin(2\sigma_{M})$.
+Because the macroscopic QCD string is strictly parity-conserving, the physical state is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations.
+The conjugate right-chiral surface identically provides the inverse complex phase $e^{+2\mathrm{i}\sigma_{M}}$.
+Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the metric:
+$\text{Re}\left[\exp\left(2\rho\right)\right]=R^{2}\big(1+\cos(2\sigma_{M})\big)=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
+Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact physical Minkowski metric is:
+$ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (D.9)
 To explicitly verify that this describes the canonical rotating string, we change from the conformal spatial coordinate $\sigma_{M}$ to the physical target space radial coordinate $r$.
@@
 The differentials transform as:
-$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (F.9)
+$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (D.10)
 Simultaneously, the time component of the metric incorporates the trigonometric identity $\cos^{2}(\sigma_{M})=1-\sin^{2}(\sigma_{M})=1-r^{2}$.
 Substituting these directly into the Minkowski metric, we obtain:
-$ds^{2}_{M}=4R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (F.10)
-This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})dt^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light.
-F.3 Action Evaluation and the Lüscher Shift
-Finally, we pedantically evaluate the total effective action over the fundamental domain of this Minkowski metric.
-The determinant of the conformal Minkowski metric $ds^{2}_{M}=4R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ is $g=16R^{4}\cos^{4}(\sigma_{M})$, yielding the exact proper area measure:
-$\sqrt{-g}=4R^{2}\cos^{2}(\sigma_{M})$ (F.11)
+$ds^{2}_{M}=2R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (D.11)
+This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})d\tau_{E}^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light, but exactly halved in overall magnitude relative to the generic three-dimensional configuration.
+D.3 Action Evaluation and the Halved Lüscher Shift
+Finally, we pedantically evaluate the total effective action over the fundamental domain of this purely real Minkowski metric.
+The determinant of the conformal Minkowski metric $ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ yields the exact proper area measure:
+$\sqrt{-g}=2R^{2}\cos^{2}(\sigma_{M})$ (D.12)
 The center of the rotating string lies at $r=0$ ($\sigma_{M}=0$), and the physical speed-of-light boundary is at $r=1$ ($\sigma_{M}=\pi/2$).
@@
 The exact classical Minkowski area integral is therefore:
-Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}4R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
-$=(2\pi)(4R^{2})\left(\frac{\pi}{4}\right)=2\pi^{2}R^{2}$ (F.12)
+Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}2R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
+$=(2\pi)(2R^{2})\left(\frac{\pi}{4}\right)=\pi^{2}R^{2}$ (D.13)
 Simultaneously, the Liouville action dictates that the worldsheet carries the kinetic anomaly $\frac{1}{12\pi}|\nabla\rho|^{2}$.
-Because the kinetic anomaly density shares the exact same spatial singularity as the conformal area density at the twistor pole, it mathematically merges into the classical area term.
+Because the parity-symmetric conformal factor shares the exact same spatial singularity at the twistor pole as the generic string, it mathematically merges into the classical area term with the identical anomaly coefficient.
 As established in the main text, it acts strictly as a rigid, local shift to the effective string tension:
-$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (F.13)
-Multiplying the classical Minkowski area by this effective tension, we obtain the complete 1-cycle bulk action:
+$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (D.14)
+Multiplying the purely real classical Minkowski area by this effective tension, we obtain the complete one-cycle bulk action:
 $S^{(1)}_{bulk}$ $=-\sigma_{\text{eff}}\times\text{Area}$
-$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(2\pi^{2}R^{2}\right)$
-$=-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$ (F.14)
-This rigorous derivation explicitly demonstrates that the universally mandated $\pi/6$ Lüscher phase is identically generated by the exact geometric mapping of the twistor string to the Minkowski cylinder.
+$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(\pi^{2}R^{2}\right)$
+$=-\pi^{2}\sigma R^{2}-\frac{\pi}{12}$ (D.15)
+This rigorous derivation explicitly demonstrates that for the parity-projected chiral string, the classical string area is exactly halved, which identically matches the universally mandated topological Lüscher phase $\pi/12$, though from a completely different physical and mathematical origins, which we discuss in the text.
+Appendix E The Pinch Singularity and the Topological Barrier
+To mathematically illustrate the appearance of the conjugate pole at $1/\bar{z}_{0}$ and the resulting impenetrable topological barrier at the unit circle, consider the evaluation of the Liouville kinetic anomaly term in the effective action.
+In for a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
+$I=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
+We can rewrite the integrand as a total derivative with respect to $\bar{z}$:
+$I\propto\int_{D}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
+By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $|z|=1$:
+$I\propto\oint_{|z|=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
+Crucially, on the unit circle, the complex conjugate is strictly given by the geometric inversion $\bar{z}=1/z$.
+Substituting this into the logarithm yields the exact boundary representation of the action:
+$I\propto\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
+This exact formula explicitly reveals the analytic structure governing the boundary dynamics.
+The integrand possesses a simple pole at $z=z_{0}$ (originating from the twistor itself) and a logarithmic branch point at $z=1/\bar{z}_{0}$ (originating from the boundary inversion of the conjugate twistor).
+Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror strictly resides outside the disk ($|1/\bar{z}_{0}|>1$).
+However, if any continuous Langevin fluctuation attempts to push the twistor pole across the boundary ($|z_{0}|\to 1$), its conjugate mirror must simultaneously approach the boundary from the outside.
+At the exact moment of crossing, these two singularities violently collide directly on the physical integration contour $|z|=1$, trapping the integration path between them.
+This creates a severe, non-integrable pinch singularity, causing the action integral to macroscopically diverge.
+This geometric divergence acts as an infinite energetic barrier, rigorously protecting the twistor pole from escaping and locking the string into its quantized topological sector.
+Appendix F Algebraic Derivation of Spin Holonomy and Trace T
+To rigorously justify the topological selection rules governing the meson spectrum, we compute the path-ordered Dirac trace $T$ for the one-pole sector ($z_{0}=0$).
+The fundamental link matrix $\mathcal{M}(\theta)$ at the boundary $|z|=1$ is defined as:
+$\mathcal{M}(\theta)=i\mathcal{P}(\theta)=-2\left(e^{i\theta}\eta\xi^{\dagger}-e^{-i\theta}\xi\eta^{\dagger}\right)$ (F.1)
+where $\eta$ is the basis spinor of the frozen $SU(2)_{R}$ sector and $\xi(\theta)$ is the rotating spinor of the active $SU(2)_{L}$ sector.
+In the basis where $\eta=\begin{pmatrix}1\\ 0\end{pmatrix}$ and the twistor $\lambda(z)\propto 1/z$ implies $\xi(\theta)=\begin{pmatrix}0\\ e^{-i\theta}\end{pmatrix}$ at the boundary, the local matrix evaluates to:
+$\mathcal{M}(\theta)=\begin{pmatrix}0&-2e^{2i\theta}\\ 2e^{-2i\theta}&0\end{pmatrix}$ (F.2)
+This matrix represents a local $SU(2)$ rotation that flips the chiral state.
+The propagator $\mathcal{U}$ for one physical cycle is the path-ordered product of these matrices.
+For two adjacent steps $\theta_{1}$ and $\theta_{2}=\theta_{1}+\Delta\theta$, we find:
+$\mathcal{M}(\theta_{2})\mathcal{M}(\theta_{1})=\begin{pmatrix}-4e^{2i\Delta\theta}&0\\ 0&-4e^{-2i\Delta\theta}\end{pmatrix}$ (F.3)
+Assuming even number $2L$ of points on a unit circle in discretized loop, and multiplying these matrices around the entire loop $\oint d\theta=2\pi$ reduces the propagator for one cycle to:
+$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{4\pi i}&0\\ 0&e^{-4\pi i}\end{pmatrix}$ (F.4)
+The physical trace $T$
+[…]
+The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulkates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$.
+which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\mathrm{i}\pi n$ depending on the parity of $L$.

### v3

@@
 By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles.
-The simplest sector with just one pole yields the exact linear Regge spectrum $m^{2}=\frac{\pi\sigma}{4}(n+1/12)$, matching the experimental Regge trajectories for light mesons ($\pi,K$ trajectories with $n=4J$ and $\rho$ trajectory with n=4J2n=4J-2) within 95% confidence intervals.
-The empirically correct intercepts $\alpha_{\pi}(0)=-1/48$ and $\alpha_{\rho}(0)=+23/48$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor pole monodromy, without any ad hoc assumptions of string vibrations.
-Ultimately, this exact geometric quantization explicitly realizes Witten’s Master Field as a critical classical trajectory in twistor space.
+The simplest sector with just one pole yields parametric Regge trajectories governed by trigonometric functions.
+While mathematically non-linear, these geometric trajectories are numerically remarkably close to linear, matching the experimental data for light mesons (the unshifted $\pi,K$ trajectories and the topologically shifted $\rho$ trajectory) within 95% confidence intervals.
+The empirically correct intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=+47/96$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor branch point monodromy, without any ad hoc assumptions of string vibrations.
+Ultimately, this exact geometric quantization explicitly realizes the Master Field as a critical classical trajectory in twistor space.
 1 Introduction
-The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the loop equations.
-At leading order in the $1/N_{c}$ expansion, the Makeenko-Migdal equations  [^1] Referred to hereafter as the MM equations.
-While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible.
-exhibit the characteristic factorization properties of a string theory, yet a concrete, mathematically controlled continuum realization has remained elusive.
+The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the MM loop equations [^1] The Makeenko-Migdal equations, referred to hereafter as the MM equations, exhibit the characteristic factorization properties of a string theory.
+While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible..
+A concrete, mathematically controlled continuum realization has remained elusive.
 Historically, the main obstruction is the ultraviolet singular nature of the coordinate-space Wilson loop functional $W[C]$.
@@
 This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$.
-By analytically continuing the action to the boundary and summing over the exact monodromy phases, we obtain an exact Bohr-Sommerfeld quantization condition.
-For the simplest sector with just one pole, this yields the spectrum:
-$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (1.1)
-where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory.
-Remarkably, the intercepts 1/48,+23/48-1/48,+23/48 emerge directly from the 2D Liouville kinetic anomaly.
-In a profound sense, the culminating result of this construction is the explicit geometric realization of Witten’s original "Master Field" hypothesis for large-$N_{c}$ QCD.
-While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, we discover its exact representation in loop space.
-The true classical limit of Planar QCD is identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory and accessible via modern geometric quantization.
+By analytically continuing the action to the boundary and summing over the infinite number of windings, with $SU(2)$ rotated boundary conditions for twistors, after projecting on the fixed spin state we obtain an exact Bohr-Sommerfeld quantization condition.
+For the simplest sector with just one pole, this yields the unified parametric Regge trajectory:
+$J=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{q}{2}-\frac{a}{48};$ (1.1)
+$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (1.2)
+The topological trace phase $q=0$ identifies the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 rigidly shifts the natural parity ($\rho$) trajectory by $1/2$.
+Remarkably, while mathematically non-linear, these trajectories are numerically very close to straight lines (see Figure 15).
+The empirically correct asymptotic intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=47/96$ emerge directly from the 2D Liouville kinetic anomaly.
+The asymptotic expansion of the Regge trajectory $J(s)$ at $s=m^{2}/\sigma>1$ reads:
+$J(s)$ $=\frac{2s}{\pi^{2}}-\left(\frac{q}{2}+\frac{1}{96}\right)+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+\dots);$ (1.3)
+$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}\approx 0.0000340208;$ (1.4)
+$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}\approx 2.29*10^{-9}$ (1.5)
+We interpret this construction as a geometric realization of the large-$N_{c}$ Master Field [54].
+While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, it admits an explicit formulation in loop space.
+The true classical limit of planar QCD is thus identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory[45].
+This geometric structure is accessible via exact WKB analysis [50, 12], where twistor monodromies dynamically generate flat valleys within the Lefschetz thimbles [55, 3].
 Organization of the paper and navigation points.
@@
 Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
-Section 15 discusses the semiclassical (WKB) limit and the emergence of linear Regge behavior.
 Finally, we highlight (iv) the exact analytic extraction of the mass spectrum.
@@
 with the spinor connection $\omega_{k}$:
-$\nabla_{k}=\partial_{k}+\frac{\mathrm{i}}{2}\omega_{k}\sigma^{3},$ (5.2)
+$\nabla_{k}=\partial_{k}+\frac{\mathrm{i}}{2}\omega_{k}\sigma^{3},\qquad\omega_{k}=\frac{1}{2}e_{kl}\partial_{l}\rho$ (5.2)
 vierbein
@@
 Before studying this phenomenon in general form, we consider the simple solvable limit: the WKB quantization corresponding to the large rotational momentum of our rigid string.
-15 The Linear Regge trajectories in the WKB limit
-15.1 The Chiral Helicoid Saddle Point
-In the semiclassical limit (large effective action), the path integral is dominated by the saddle point of the effective action.
-We look for a solution corresponding to a meson with large angular momentum $J$, which corresponds to a rotating momentum configuration.
-Consider a momentum distribution rotating in the $(1,2)$ plane with frequency $\omega$:
-$P_{1}(\tau)+\mathrm{i}P_{2}(\tau)=P_{0}e^{\mathrm{i}\omega\tau},\qquad P_{3}=P_{4}=0.$ (15.1)
-This source term in the action (11.7) drives the spinor fields.
-We search for a solution that respects this rotational symmetry.
-The natural ansatz corresponds to the helicoid minimal surface described in Appendix B.4.
-In the generic spinor formalism, with rescaling $w(z)=z^{-1/2}$, the conventional Helicoid is generated by $\mu(z)\propto(1,-z)^{T}$.
-However, as established in Appendix F, our geometric prescription requires enforcing the exact chiral (Hodge-dual) configuration.
-We freeze the right-handed twistor to a constant basis spinor:
-$\lambda(z)=\begin{pmatrix}z^{-1}\\ 1\end{pmatrix},\qquad\mu(z)=\mathrm{i}\begin{pmatrix}1\\ 0\end{pmatrix}$ (15.2)
-Substituting this into the null vector definition $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, we find that the spatial tangent vectors no longer identically vanish in the extra dimensions, like they did with conventional helicoid.
-Our Hodge dual minimal surface, with only one of two chiral Gauss maps (9.12), being active, and other being frozen, permanently breaks the 3D embedding constraint and allows the minimal surface to span the 4-dimensional space.
-Integrating this yields the logarithmic cut $f(z)\sim\log z$, which maps the complex plane to the infinite strip of the helicoid.
-The boundary value on the real axis $z=e^{\tau}$ gives a rotating coordinate vector, consistent with the rotating momentum source.
-15.2 The Symmetrized Minkowski Helicoid
-The Minkowski Helicoid is parameterized by time $\tau\in[0,T]$ and radius $r\in[-R,R]$, see [7]. The time $T$ tends to infinity in the end.
-$X(\tau,r)=(\ r\cos\omega\tau,\ r\sin\omega\tau,0,\tau)$ (15.3)
-The induced Minkowski metric for a generic string is $g_{ab}=\text{diag}(1-\omega^{2}r^{2},1)$.
-Crucially, as derived in Appendix F, the conformal factor for the generic surface evaluates to $4\cos^{2}(\sigma_{M})$.
-In stark contrast, our chiral twistor evaluates to an induced Euclidean conformal factor $\exp\left(2\rho\right)\propto(1+e^{-2\sigma_{E}})$.
-Applying our physical prescription—taking the parity-even arithmetic mean of the Left-chiral (SD) and Right-chiral (ASD) configurations—upon double Wick rotation ($\sigma_{E}\to\mathrm{i}\sigma_{M}$), precisely isolates the real part:
-$\text{Re}\left(1+e^{-2\mathrm{i}\sigma_{M}}\right)=1+\cos(2\sigma_{M})=2\cos^{2}(\sigma_{M})$ (15.4)
-Thus, the effective proper area of the symmetrized chiral Minkowski Helicoid is exactly halved relative to the conventional generic string.
-In Minkowski space we are not seeking to minimize the area in the geometric sense, as the area functional is not positive definite and the minimization problem is distinct.
-Instead, our criterion is algebraic: we require a solution to the Loop Equations for the Wilson loop.
-This demands the self-duality of the area derivative and the additivity of the area functional.
-We utilize the fact that the Euclidean Helicoid is a minimal surface for arbitrary angular velocity $\omega$.
-We obtain the Minkowski Helicoid via a double Wick rotation:
-$x_{4}\to\mathrm{i}t,\qquad\omega_{E}\to-\mathrm{i}\omega.$ (15.5)
-Under this transformation, the argument of the trigonometric functions in the coordinate parametrization becomes real ($\omega_{E}x_{4}\to\omega t$), mapping the Euclidean Helicoid to a real surface in Minkowski space.
-Crucially, the condition that $S[C]$ is a zero mode of the Loop Equation depends on algebraic identities (self-duality of the area derivative and additivity).
-By the Identity Theorem, these analytic relations, established for the Euclidean minimal surface, must hold for its analytic continuation into Minkowski space.
-Thus, the symmetrized Minkowski Helicoid remains a valid solution to the QCD Loop Equation, bypassing the ill-posed variational problem of minimizing area in a space with an indefinite metric.
-Figure 14: The helicoid spanned by rotating $\bar{q}q$ pair connected by a rigid stick (string). This minimal surface bounded by a double helix, was discovered by Meusnier in 1785.
-Another important point is that the Helicoid boundary loop (double helix) is infinitely smooth, lacking any cusps or self-intersections.
-This means that no UV divergent renormalization factors could potentially affect our computation of the meson Regge trajectory.
-15.3 Semiclassical Motivation: Twisted Boundary Conditions
-Before solving the spectral equation, we justify the choice of the Helicoid geometry from the perspective of the semiclassical quark path integral.
-Consider the Green’s function for a meson composed of scalar quarks with mass $m_{q}$.
-In the quenched approximation, the effective Minkowski action involves a sum over quark trajectories $C$, weighted by the quark mass term and the area of the confining minimal surface $\Sigma$ bounded by $C$, neglecting the perturbative factor444the perimeter renormalization terms in that factor are included in the physical quark mass..
-This is relevant to our WKB approximation in the Wilson loop at large loops:
-$\mathcal{Z}\sim\int\mathcal{D}C\,\exp\left(-\mathrm{i}m_{q}\oint_{C}ds-\mathrm{i}\sigma S_{\mathrm{Mink}}[\Sigma_{C}]\right)$ (15.6)
-To extract the spectrum of states with angular momentum $J$, we impose twisted boundary conditions.
-We require the quark positions at time $t=T$ to be a rotation of their positions at $t=0$ by an angle $\Theta=\omega T$ in the rotation plane $xy$:
-$(x+\mathrm{i}y)(T)=e^{\mathrm{i}\omega T}(x+\mathrm{i}y)(0)$ (15.7)
-The angular momentum $J$ arises as the thermodynamic conjugate to the twist frequency $\omega$.
-The ground state of this system is determined by the saddle point of the effective action.
-This requires the simultaneous minimization of the boundary length (quark term) and the surface area (gluon term).
+15 Catastrophe Theory and the Geometric Mass Spectrum
+15.1 Twisted Boundary Conditions, Symmetrization, and Spin Projections
+Before solving the exact spectral equations, we justify the choice of the twisted geometry from the perspective of the semiclassical quark path integral.
+In Planar QCD, the effective Minkowski action involves a sum over quark trajectories weighted by the area of the confining minimal surface.
+To extract the spectrum of states with macroscopic angular momentum $J$, we impose twisted boundary conditions.
+We require the quark positions at the ends of the string to undergo a physical spatial rotation by an angle $\phi$ over the period of the motion, with the angular momentum $J$ acting as the thermodynamic conjugate to this twist.
+The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
+The minimal surface spanning two twisted helical boundaries is the helicoid, as illustrated in Figure 14.
+Figure 14: The helicoid spanned by a rotating $\overline{q}q$ pair connected by a rigid stick (string).
+This minimal surface bounded by a double helix, discovered by Meusnier in 1785, provides the macroscopic physical motivation for the twisted boundary conditions.
+Thus, the helicoid is not merely an ansatz; it is the simplest geometric solution mandated by the kinematics of angular momentum in the semiclassical limit.
+To extract the physical spin states of the meson spectrum, we must couple the internal worldsheet geometry to this macroscopic angular momentum $J$ of the target space.
+In our twistor string formulation, we project this twisted kinematics directly into the boundary twistor data.
+This is achieved by introducing the $SU(2)_{L}$ and $SU(2)_{R}$ projections, integrating over the target space rotation angle $\phi$ with the canonical spin weight:
+$\mathcal{P}_{J}=\oint d\phi\exp\left(\mathrm{i}\phi J\right)$ (15.1)
+When the macroscopic boundary of the string rotates by an angle $\phi$, the boundary twistors must geometrically twist.
+To enforce this, we introduce a fractional winding parameter $a$ corresponding to twisted boundary conditions for $\lambda(z)$.
+Furthermore, because macroscopic QCD rigorously conserves parity, the physical effective action must be symmetrized over the Self-Dual (SD) and Anti-Self-Dual (ASD) minimal surfaces.
+In the twistor parameterization, this corresponds to activating the twisted boundary condition for $\lambda(z)$ while keeping the mirror sector $\mu(z)$ frozen to a constant basis spinor, and adding a mirror term with $\mu(z)$ to the exponent.
+The Hodge-duality (or anti-duality) explicitly breaks the 3D embedding constraint, forcing the chiral surface to genuinely span 4-dimensional Minkowski space.
+To analytically generate the perfect resonance leading to the exact WKB saddle point, we must consider the limit of an infinitely large number of geometric windings $k\rightarrow\infty$.
+As the integration variable traverses the flat valley $k$ times, the accumulated action scales strictly linearly with $k$.
+Crucially, this means the transverse Hessian scales as $\mathcal{H}_{k}\propto k$.
+In the asymptotic resonance limit $k\rightarrow\infty$, the transverse restoring force becomes strictly infinite.
+The stochastic variance of the quantum fluctuations scales as $1/k\rightarrow 0$.
+The path integral dynamically localizes to the exact WKB saddle point, completely shielding the topological monodromy from quantum noise.
+These heuristic arguments will be made explicit and quantitative in this section, which will end with exact WKB quantization in a topological sector with only one pole of twistor field.
+15.2 Geometric Quantization via Degenerate Lefschetz Thimbles
+In the semiclassical (WKB) limit, the path integral over the twistor phase space is dominated by the saddle points $\mathcal{V}_{*}(E)$ of the complexified effective action.
+Physical bound states (mesons) correspond to macroscopic divergences in this amplitude.
+Mathematically, these occur at critical energies $E=E_{n}$ where the classical stability matrix (the Hessian $\mathcal{H}$) develops a null space.
+Because the integration variables are coupled complex holomorphic twistors, the vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
+This complex holomorphic geometry natively upgrades the standard square-root branch cut into a strictly simple pole, precisely generating the discrete S-matrix mass poles of the meson spectrum without ad hoc approximations.
+15.3 Catastrophes, flat valleys, and topological barriers
+To dynamically generate an exact simple mass pole $\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat zero-mode that extends infinitely far in configuration space.
+This exact infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields as the boundary parameter winds around the enclosed twistor poles.
+The integrity of this topological winding number is rigorously protected by an infinite topological action barrier.
+If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside.
+This creates a severe pinch singularity, strictly trapping the integration contour and locking the path integral into a conserved topological sector.
+15.4 Poles inside the disk and the monodromies
+For a saddle point with a single pole located strictly inside the unit disk ($|z_{0}|<1$), the total effective action accumulates strictly additively over multiple windings $k$.
+The Liouville bulk action simply duplicates, and the logarithmic boundary primitive picks up a strict topological phase shift at every cycle.
+Together, these terms yield an infinite geometric series $\sum\exp\left(\mathrm{i}k\Delta S_{eff}\right)$.
+The macroscopic divergence of this series generates the simple meson pole precisely when the exact Bohr-Sommerfeld quantization condition $\Delta S_{eff}=2\pi n$ is satisfied.
+15.5 The Exact One-Pole Solution and Minkowski Transition
+For the fundamental meson trajectory, we consider in detail the specific case of a single pole.
+To implement the twisted boundary condition for $\lambda(z)$ while freezing $\mu(z)$ to enforce the symmetrization over SD and ASD surfaces, we utilize the exact chiral 1-pole configuration with fractional power $a$, shifted to a pole position $z_{0}$:
+$\lambda(z)=R(z-z_{0})^{-a}\begin{pmatrix}1\\ 0\end{pmatrix},\qquad\mu(z)=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 1\end{pmatrix}$ (15.2)
+As we trace the boundary loop $z\to z\exp\left(2\pi\mathrm{i}\right)$, the active twistor acquires the fractional monodromy $\lambda\to\exp\left(-2\pi\mathrm{i}a\right)\lambda$.
+Equating this to the physical spin-1/2 transformation $\lambda\to\exp\left(-\mathrm{i}\phi/2\right)\lambda$, we establish the exact geometric identification $\phi=4\pi a$.
+The angular momentum projection operator thereby injects the canonical phase $\exp\left(4\pi\mathrm{i}Ja\right)$.
+To evaluate the bulk action properly, we must transition to the physical Minkowski metric on the symmetrized chiral surface, as established in D.
+We switch to the Minkowski metric on the surface parameterized by $z=\exp\left(\mathrm{i}\sigma_{M}+\tau_{E}\right)$, describing the Nambu-Goto metric $g_{ab}$ via the coordinate transformation $\exp\left(2\rho_{E}\right)d^{2}z=R^{2}\frac{d^{2}z}{(z\bar{z})^{a}}=R^{2}\mathrm{i}d\sigma_{M}d\tau_{E}$.
+Upon double Wick rotation and symmetrization (taking the parity-even arithmetic mean of the Left and Right chiral configurations), the conformal factor of the twisted metric dynamically restricts the causal domain.
+Using the modified radial coordinate $\tilde{\sigma}_{M}=(1-a)\sigma_{M}$, the physical speed-of-light boundary occurs exactly where this conformal metric vanishes, at $\tilde{\sigma}_{M}=\pi/2$.
+The proper Minkowski Area integral factorizes in the light-cone coordinates $z,\bar{z}$:
+$\text{Area}(a)$ $=\frac{\mathrm{i}R^{2}}{2}\int_{z\overline{z}<1}dz\wedge d\overline{z}(z\overline{z})^{-a}$
+$=\frac{-\mathrm{i}R^{2}}{2}\oint\frac{dz}{z}\int_{0}^{1}d(z\overline{z})(z\overline{z})^{-a}$
+$=\frac{-\mathrm{i}R^{2}}{2(1-a)}\oint dzz^{-1}=\frac{\pi R^{2}}{1-a}$ (15.3)
+Multiplying by the physical string tension $\sigma$, the macroscopic Area action evaluates exactly to:
+$S_{M,Area}=\frac{\pi\sigma R^{2}}{1-a}$ (15.4)
+Simultaneously, the Liouville action evaluates the kinetic gradient term.
+Crucially, while the target space metric stretches with the twist, the topological Liouville anomaly resides strictly on the rigidly bounded fundamental domain of the reference worldsheet, $\sigma_{M}\in[0,\pi/2]$:
+$\int|\nabla\rho|^{2}d^{2}z$ $=a^{2}\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}d\sigma_{M}$
+$=a^{2}(2\pi)\left(\frac{\pi}{2}\right)=\pi^{2}a^{2}$ (15.5)
+Multiplying by the universal anomaly coefficient $\frac{1}{12\pi}$, the Minkowski kinetic action yields the strictly positive, rigid Lüscher shift:
+$S_{M,kin}=+\frac{\pi a^{2}}{12}$ (15.6)
+[…]
+$\Delta X_{0}=\int_{0}^{2\pi}\sqrt{2}R\sin\left((1-a)\theta\right)d\theta=\frac{2\sqrt{2}R\sin^{2}(\pi a)}{1-a}$ (15.7)
+We balance the string tension against the boundary energy by minimizing the combined classical action $S(R)=S_{M,Area}-E\Delta X_{0}$ with respect to the geometric scale $R$.
+Since the Area is purely quadratic $(AR^{2})$ and the boundary coupling is strictly linear $(BR)$, the minimization of this quadratic function trivially yields $B^{2}/(4A)$ by the virial theorem.
+This dynamically eliminates the dimensionful size $R$ in one step, leaving a purely intensive, remarkably pristine geometric boundary potential:
+$S_{geom}(a)=-\frac{E^{2}}{\pi\sigma}F(a),\qquad\text{where}\qquad F(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (15.8)
+This establishes the exact, scale-independent geometric potential that governs the twistor branch point degree $a$ at the WKB saddle point.
+15.6 Quark Spin Monodromy, Trace Splitting, and the Unified Spectrum
+The final component of the path integral is the path-ordered Dirac trace $T$ along the quark boundary.
+The global boundary holonomy matrix acting on the twistor space is $U=\text{diag}(\exp\left(-2\pi\mathrm{i}a\right),1)$.
+1. The $\rho$ Trajectory (Vector Current)
+With an extra $\gamma_{\nu}$ projector inserted as the vector quark current, the twistor dyadic projector $P\propto\lambda\mu^{\dagger}$ acts as a strict rank-1 projector, $\left(\begin{smallmatrix}1&1\\ 0&0\end{smallmatrix}\right)$.
+It annihilates the trivial constant mode of the un-twisted right-chiral twistor, leaving exactly:
+$T_{\rho}\propto\text{tr}(UP)$ $=\text{tr}\begin{pmatrix}\exp\left(-2\pi\mathrm{i}a\right)&\exp\left(-2\pi\mathrm{i}a\right)\\ 0&0\end{pmatrix}$
+$=\exp\left(-2\pi\mathrm{i}a\right)$ (15.9)
+This corresponding pure topological phase factor $\exp\left(-2\pi\mathrm{i}a\right)$ adds directly to the rotation weight $\exp\left(4\pi\mathrm{i}Ja\right)$, shifting the angular momentum exactly by $\Delta J=+1/2$.
+2. The $\pi$ and $K$ Trajectories (Pseudoscalar Current)
+With a $\gamma_{5}$ inserted for the pseudoscalar and scalar trajectories, the trace evaluates the sum of the modes:
+$T_{\pi,K}=\text{tr}(U)=1+\exp\left(-2\pi\mathrm{i}a\right)$ (15.10)
+Because the path integral weight is linear in $T$, the amplitude rigorously splits into a sum of two independent path integrals:
   • 1.
-    The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
+    The Light Integral ($T_{1}=1$): Yields a phase $\exp\left(0\right)$. The spin trajectory is completely unshifted. This generates the ultra-light pseudo-Goldstone trajectory (the $\pi$ and $K$).
   • 2.
-    The minimal surface spanning two twisted helical boundaries is the helicoid.
-Thus, the Helicoid is not merely an ansatz; it is the unique geometric solution mandated by the kinematics of angular momentum in the semiclassical limit.
-The balance between the surface tension $\sigma$ (pulling the quarks inward) and the centrifugal barrier (pushing the twisted trajectories outward) determines the physical radius of the meson.
-15.4 Effective Lagrangean
-We construct the effective Lagrangian $L$ by combining the bulk contribution from the Helicoid area (coming from our confining factor) with the worldline terms for the massive quarks.
-The confining factor introduces an effective Minkowski area term.
-We interpret the phase factor for quark propagation as an effective action, adding the worldline terms $2m_{q}\int ds$.
-The total Lagrangian is:
-$L=L_{\text{helicoid}}+L_{\text{quarks}}$ (15.8)
-Because the arithmetic mean of the SD and ASD surfaces rigorously halves the classical area, the bulk contribution, calculated as the area integral in Minkowski space for a ruled surface with angular velocity $\omega$ and radius $R=v/\omega$, is precisely halved relative to the generic string:
-$L_{\text{helicoid}}$ $=-\frac{\sigma}{2}\int_{-R}^{R}dr\sqrt{1-\omega^{2}r^{2}}=-\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}$
-$=-\frac{\sigma}{2\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)$ (15.9)
-Here $\sigma$ is the physical string tension, related to the parameter $\kappa$ in the confining factor $\exp\left(-\kappa S[C]\right)$ by the relation $\sigma=2\sqrt{2}\kappa$, which accounts for the normalization of the Hodge-dual area functional.
-Adding the particle term $2m_{q}\sqrt{1-v^{2}}$ as specified for the massive endpoints, the total effective Lagrangian becomes:
-$L(v,\omega)=-\frac{\sigma}{2\omega}\left(v\sqrt{1-v^{2}}+\arcsin v\right)-2m_{q}\sqrt{1-v^{2}}$ (15.10)
-It is worth noting that this mass term is proportional to the perimeter of the loop in Minkowski space.
-Therefore, the potential perturbative terms proportional to the loop perimeter are included in this physical quark mass $m_{q}$.
-We are assuming that this physical mass (the bare quark mass plus perturbative corrections) adds up to a finite renormalized mass $m_{q}$, which, as the experiment tells us, is much smaller than the QCD mass scale, so that $m_{q}^{2}\ll\sigma$.
-We proceed to compute the conserved quantities via the Legendre transform. The angular momentum $J$ is the conjugate momentum to the rotation frequency:
-$J=\frac{\partial L}{\partial\omega}$ (15.11)
-Taking into account that $v$ depends on $\omega$ via the fixed radial extent ($v=\omega R$), the differentiation yields:
-$J=\frac{\sigma}{2\omega^{2}}(\arcsin v-v\sqrt{1-v^{2}})+\frac{2m_{q}v^{2}}{\omega\sqrt{1-v^{2}}}$ (15.12)
-The Energy $E$ is defined by the Hamiltonian $E=\omega J-L$. Substituting the expressions for $J$ and $L$ and simplifying the algebra, we obtain:
-$E=\frac{\sigma}{\omega}\arcsin v+\frac{2m_{q}}{\sqrt{1-v^{2}}}$ (15.13)
-15.5 Balance of quark forces and boundary condition
-The relationship between the endpoint velocity $v$ and the string length $R$ is not arbitrary; it is determined by the balance of forces at the string endpoint, which depends on the quark mass $m_{q}$.
-We derive this boundary condition by minimizing the action with respect to the geometric radius $R$ of the loop, keeping the angular velocity $\omega$ fixed.
-Recall that the total Lagrangian $L$ is composed of the bulk integral for the halved helicoid and the boundary terms for the quarks:
-$L=-\sigma\int_{0}^{R}dr\sqrt{1-\omega^{2}r^{2}}-2m_{q}\sqrt{1-\omega^{2}R^{2}}$ (15.14)
-The equation of motion for the endpoint radius $R$ is given by the stationarity condition $\frac{\partial L}{\partial R}=0$.
-Differentiating the integral using the Leibniz rule yields the effective tension force, while differentiating the boundary term yields the centrifugal force:
-$\frac{\partial L}{\partial R}=-\sigma\sqrt{1-\omega^{2}R^{2}}-2m_{q}\frac{d}{dR}\left(\sqrt{1-\omega^{2}R^{2}}\right)=0$ (15.15)
-Substituting $v=\omega R$ and computing the exact derivative of the mass term (properly applying the chain rule):
-$-\sigma\sqrt{1-v^{2}}-2m_{q}\left(\frac{-\omega^{2}R}{\sqrt{1-v^{2}}}\right)=0$ (15.16)
-Simplifying the expression, we obtain the exact force balance equation for finite mass:
-$\sigma\sqrt{1-v^{2}}=\frac{2m_{q}v^{2}}{R\sqrt{1-v^{2}}}\implies v^{2}=\frac{\sigma R}{\sigma R+2m_{q}}$ (15.17)
-This equation balances the physically halved confining string tension (LHS) against the relativistic centrifugal inertia of the quark (RHS).
-We now consider the limit $m_{q}\to 0$, and we find the boundary condition
-$v^{2}\to 1$ (15.18)
-Thus, as the quark mass tends to zero, the endpoint velocity $v$ must approach the speed of light ($v\to 1$).
-These equations provide the exact parametric equation for the energy of the classical pair of particles with a given angular momentum and a given radius $R$.
-To recover the Regge trajectories, we consider the limit of massless quarks ($m_{q}\to 0$), where the endpoint velocity approaches the speed of light ($v\to 1$).
-In this limit, the mass terms vanish, and the system is governed purely by the symmetrized confining geometry:
-$E\approx\frac{\sigma}{\omega}\frac{\pi}{2},\qquad J\approx\frac{\sigma}{2\omega^{2}}\frac{\pi}{2}$ (15.19)
-Eliminating $\omega$ between these two relations:
-$J=\frac{\pi\sigma}{4}\frac{1}{\omega^{2}}=\frac{\pi\sigma}{4}\left(\frac{2E}{\pi\sigma}\right)^{2}\implies J=\frac{E^{2}}{\pi\sigma}$ (15.20)
-We thus flawlessly recover the exact linear Regge trajectory for the symmetrized chiral string:
-$E^{2}=\pi\sigma J$ (15.21)
-This result confirms that the linear confinement natively arises directly from the kinematics of the Hodge-dual minimal surface in Minkowski space.
-16 The Catastrophe Theory and the Geometric Mass Spectrum
-16.1 Geometric Quantization via Degenerate Lefschetz Thimbles
-The holomorphic nature of the effective action reveals a profound mathematical pathway to compute the exact meson spectrum analytically, entirely bypassing stochastic simulation.
-This approach relies on the asymptotic evaluation of multiple complex integrals via Picard-Lefschetz theory and the singularity theory of differentiable maps.
-Consider the original, un-reweighted path integral for the quark two-point amplitude $\Pi(E)$ as a function of the Minkowski energy parameter $E$:
-$\Pi(E)=\int_{\Gamma}\mathcal{D}\mathcal{V}_{|z|=1}\,\exp\left(E\,\mathcal{O}(\mathcal{V})-S_{eff}(\mathcal{V})\right)$ (16.1)
-where $\mathcal{V}_{|z|=1}=\left\{z_{2},\left\{\rho,\xi,\eta,\bar{\xi},\bar{\eta},\gamma,\bar{\gamma}\right\}_{k}\right\}$ encompasses the boundary twistor variables extended analytically into the complexified phase space $\mathcal{V}=(\lambda(z),\mu(z))$, and $\Gamma$ is the middle-dimensional integration cycle (the principal Lefschetz thimble) bounding the physical vacuum.
-In the semiclassical (WKB) limit, this integral is dominated by the saddle points $\mathcal{V}_{\star}(E)$ satisfying the classical boundary equations of motion:
-$\nabla S_{eff}(\mathcal{V}_{\star})-E\nabla\mathcal{O}(\mathcal{V}_{\star})=0$ (16.2)
-(which, as shown in Section 13, yields the rigid Helicoid geometry in the special case of twisted boundary conditions with large angular momentum $J$).
-The Gaussian fluctuations around this saddle point yield an amplitude proportional to the inverse square root of the complex Hessian determinant:
-$\Pi(E)\propto\frac{\exp\left(E\,\mathcal{O}(\mathcal{V}_{\star})-S_{eff}(\mathcal{V}_{\star})\right)}{\sqrt{\det\mathcal{H}(E)}}$ (16.3)
-where the fluctuation operator is $\mathcal{H}(E)=\nabla^{2}S_{eff}(\mathcal{V}_{\star})-E\nabla^{2}\mathcal{O}(\mathcal{V}_{\star})$.
-As the continuous energy parameter $E$ is varied, the saddle points drift through the complexified phase space.
-Physical bound states (mesons) correspond to macroscopic divergences in this amplitude.
-Mathematically, these occur at critical energies $E=E_{n}$ where the saddle point becomes strictly shallow or degenerate (a non-Morse critical point).
-At these bifurcation points, the classical stability matrix develops a null space, and the determinant of the Hessian exactly vanishes: $\det\mathcal{H}(E_{n})=0$.
-At this point, the WKB approximation becomes exact, as it dominates the divergent term in the path integral.
-According to the classification of singular integrals in Catastrophe Theory [1], the nature of the resulting divergence depends strictly on the corank of the singularity (the number of simultaneously vanishing eigenvalues).
-If only a single real eigenvalue vanishes (corank 1), the local action along the flat direction scales as $\epsilon x^{2}$ (with $\epsilon\propto E_{n}-E$).
-The integration over this 1D valley yields:
-$\int dx\,e^{-\epsilon x^{2}}\propto\frac{1}{\sqrt{\epsilon}}\propto\frac{1}{\sqrt{E_{n}-E}}$ (16.4)
-Physically, a square-root singularity corresponds to a branch cut, generating a multi-particle continuum threshold, not a bound state.
-However, our twistor-string geometry orchestrates a remarkable mathematical conspiracy.
-Because the integration variables $\mathcal{V}$ are coupled complex holomorphic twistors, $\mathcal{H}(E)$ is a complex operator.
-The vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
-This alters the local topology of the steepest descent path. The Gaussian integral over this 2D flat valley yields:
-$\iint dxdy\exp\left(-\epsilon(x^{2}+y^{2})\right)\propto\frac{1}{\epsilon}\propto\frac{1}{E_{n}-E}$ (16.5)
+    The Heavy Integral ($T_{2}=\exp\left(-2\pi\mathrm{i}a\right)$): Yields the phase $\exp\left(-2\pi\mathrm{i}a\right)$.
+It acts identically to the $\rho$ projector, generating a heavy vector trajectory rigidly shifted by $\Delta J=1/2$ to $J=1$.
+15.7 Asymptotic Linearity and the Chew-Frautschi Plot
+In the macroscopic limit ($a\to 1/2$), the parametric geometric equations generate an exact simple pole, aggressively projecting the system into asymptotic Regge linearity.
+The theoretical intercepts emerge natively from the Liouville anomaly $\frac{\pi a^{2}}{12}$ without a single adjustable parameter:
+  • 1.
+    Pseudoscalar ($q=0$): $\alpha_{\pi}(0)=-1/96$
+  • 2.
+    Vector (q=1q=-1): $\alpha_{\rho}(0)=-1/96+1/2=47/96$
+Figure 15: Exact Twistor String Trajectories $J=\alpha(m^{2})$.
+The plot demonstrates the remarkable geometric linearity of the transcendental equations.
+The $\pi$ and $K$ mesons reside on the unshifted trajectory (with the Kaon explicitly shifted by the perturbative $SU(3)$ flavor breaking $\Delta m_{s}^{2}$), while the vector $\rho$ family is natively shifted upward by $\Delta J=1/2$ due to the rank-1 twistor projection.
+Unweighted least-squares fits of $\sigma$ map the universal curves through the experimental PDG data.
+The physical branch is covered by $a_{\mathrm{min}}<a<1/2$. Here $a_{\mathrm{min}}$ is a solution of the transcendental equation
+$4\pi\cot(\pi a)$ $=\frac{1-2a}{a(1-a)}+\frac{1}{a}$
+$\implies a_{\mathrm{min}}$ $\approx 0.426836\dots$ (15.11)
+Such a narrow range of the parameter $a$ explains the approximate linearity of the Regge trajectories.
+We computed the asymptotic expansion (G.11) of the Regge trajectory at large $s=m^{2}/\sigma$.
+The linear and constant terms dominate, and the higher-order $1/s$ corrections are numerically so small that at $J>1$ the deviations from linearity are invisible.
+But this is, of course, a matter of principle to have a correct formula based on an exact analytic solution instead of heuristic models.
+An interesting property of this analytic formula, discussed in G.2, is that these trajectories do not extend to negative spin or to negative $m^{2}$ (tachyons).
+Instead, they terminate at a geometric singularity characterized by the simultaneous vanishing of the derivatives of our parametric equations, $dJ/da=dm^{2}/da=0$.
+This singularity occurs at a certain critical value of $a=a_{c}<a_{\mathrm{min}}$ that is safely below the end of the physical region $a_{\mathrm{min}}<a<1/2$.
+Therefore, this singularity does not manifest in the physical region, where the trajectories appear virtually linear with the corresponding asymptotic intercepts 1/96-1/96 and $47/96$.
+Thus, this geometric singularity serves as a defender against tachyons.
+The technical details can be found in G.2.
+As shown in Figure 15, the analytic solution maps well onto the experimental data.
+What is truly remarkable is the numerical linearity of these trajectories given the highly transcendental nature of the geometric equations.
+The complex trigonometric terms act as deep-IR regulators, providing the correct chiral symmetry breaking behavior for the anomalously light pion.
+Yet, driven by the exact geometric pole at $a\to 1/2$, the macroscopic strings violently straighten out past $J\approx 1$, validating the strict Regge limit while rigorously housing the microscopic quantum anomalies.
+The unweighted least-squares fits yield effective string tensions $\sigma_{\rho}\approx 0.235$, $\sigma_{\pi}\approx 0.281$, and $\sigma_{K}\approx 0.320\text{ GeV}^{2}$.
+The value for the natural-parity $\rho$ trajectory recovers the canonical macroscopic Regge slope $\alpha^{\prime}=2/(\pi^{2}\sigma_{\rho})\approx 0.86\text{ GeV}^{-2}$, while the slightly higher effective tensions for the $\pi$ and $K$ trajectories physically reflect the dynamical stiffening due to $N_{c}=3$ chiral symmetry breaking and the endpoint inertia of the massive strange quark, respectively.
+15.8 Special case of rational branch points
 […]
-Remark 16.1.
-Corank-2 from a single complex zero mode, and the pinch condition.
-[…]
-[…]
-so that the vanishing of one complex eigenvalue generically flattens two real directions on the thimble (corank =2=2) and produces the pole-like local scaling (16.5).
-In this sense, a “single complex zero mode” is already a corank-2 degeneracy of the real thimble geometry.
-Thus, the complex holomorphic geometry natively upgrades the standard square-root branch cut into a strictly simple pole.
-This rigorously explains how the degenerate saddle points of the Twistor String path integral explicitly generate the discrete mass poles of the meson spectrum.
-(For this pole to manifest macroscopically without being capped by higher-order terms, the steepest descent contour must become infinitely long.
-In the framework of Landau pinch singularities [12], this infinite integration volume is provided dynamically by the Liouville zero-mode stretching the string worldsheet into an infinite cylinder).
-Consequently, the extraction of the exact Planar QCD meson spectrum is formally reduced from a stochastic path integral to a deterministic Generalized Eigenvalue Problem (akin to the Gutzwiller trace formula [19] and the Dashen-Hasslacher-Neveu soliton quantization [6]).
-The physical masses $E_{n}$ are precisely the discrete values for which the classical stability matrix develops a complex null space:
-$\det_{\mathbb{C}}\Big[\nabla^{2}S_{eff}(\mathcal{V}_{\star})-E_{n}\nabla^{2}\mathcal{O}(\mathcal{V}_{\star})\Big]=0$ (16.7)
-Solving this boundary-twistor Fredholm equation in general case provides a direct analytical roadmap for the exact, noise-free computation of the geometric spectrum in future work.
-16.2 Catastrophes, flat valleys, and topological barriers
-In the general theory of functional integration, hadronic resonances emerge when complex saddle points of the effective action become degenerate.
-A vanishing Hessian matrix indicates a local flat valley in the functional landscape, which is the hallmark of Catastrophe Theory.
-However, a local flat valley alone is fundamentally insufficient to produce a true $S$-matrix bound state pole.
-A generic Taylor expansion of the action around a degenerate saddle point will eventually be capped by higher-order terms (e.g., cubic or quartic corrections).
-These terms force the steepest descent contour to curve upwards, trapping the integral and yielding finite fractional powers or branch cuts rather than macroscopic divergences.
-To dynamically generate an exact simple mass pole $\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat zero-mode that extends infinitely far in configuration space without ever curving upwards.
-All higher-order corrections along this specific trajectory must identically vanish, which is a known localization phenomenon in the WKB approximation for path integrals [9].
-In our twistor string framework, this exact infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields.
-The zero-mode trajectory in functional space corresponds to the topological winding of the boundary parameter as it wraps around the enclosed twistor poles.
-Because each winding around the complex pole maps holographically to a uniform translation in target-space Minkowski time, the effective action accumulates strictly additively.
-The total action is exactly linear in the winding number $n$, which serves as the non-compact integration variable along the steepest descent path.
-Consequently, all higher-order nonlinear corrections for this particular field are strictly absent.
-The contour along this monodromy direction remains perfectly flat as $N\to\infty$, allowing the integral to diverge into the exact geometric series that analytically generates the simple meson pole.
-Furthermore, this geometric mechanism reveals that the number of twistor poles enclosed within the unit circle ($|z_{k}|<1$) serves as a strict topological quantum number that classifies the resonance states.
-The integrity of this integer quantum number is rigorously protected by an infinite topological action barrier.
-One might intuitively ask whether continuous fluctuations could cause twistor poles to smoothly drift across the unit circle boundary ($|z|=1$), thereby changing the topological sector.
-However, the holomorphic boundary geometry explicitly forbids this.
-If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside.
-Exactly at the moment of crossing, these two singularities inevitably collide directly on the physical integration contour $|z|=1$.
-This collision creates a classic pinch singularity.
-The integration loop becomes trapped between the colliding poles $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them.
-Consequently, the action strictly diverges at the boundary, creating an impenetrable topological barrier.
-The twistor path integral is therefore dynamically locked into a fixed topological sector, mathematically guaranteeing that the number of enclosed poles is a rigorously conserved topological quantum number.
-16.3 Poles inside the disk and the monodromies
-We are looking for a saddle point with a single pole located at a complex coordinate $z_{0}$ strictly inside the unit disk ($|z_{0}|<1$).
-Why inside?
-Because in this case, there is a non-trivial winding number.
-As the target space surface $X$ winds up in the time direction, it means the quark goes around the loop infinitely many times, creating a bound state of a rotating string.
-This infinite motion is precisely the saddle point we are looking for—it will create the resonance pole through a geometric series over the multiple windings.
-The total effective action evaluated over these multiple windings consists of two parts.
-First, the Liouville bulk action is simply additive because its Lagrangian density is a single-valued, explicitly periodic function of the angle on the boundary circle.
-Therefore, for $k$ windings, its contribution simply duplicates: $S^{(k)}_{bulk}=kS^{(1)}_{bulk}$.
-Second, the boundary energy term in the effective action involves a logarithmic primitive.
-Because the pole $z_{0}$ is enclosed by the contour, analytic continuation around the loop causes this term to pick up a strict topological phase shift at every extra cycle.
-This is the nontrivial monodromy of the effective action.
-Together, these terms lead to an infinite sum over the winding number $k$ of the form $\sum\exp\left(\mathrm{i}k\Delta S_{\text{eff}}\right)$.
-This infinite geometric series is a direct manifestation of the flat valley near the degenerate fixed point.
-The valley is flat near the energy pole because the monodromy of the effective action $\Delta S_{\text{eff}}$ approaches $2\pi n$.
-This makes $\exp\left(\mathrm{i}\Delta S_{\text{eff}}\right)=1$, which leaves the amplitude invariant under winding, so that the sum over windings macroscopically diverges as $\sim 1/(1-\exp\left(\mathrm{i}\Delta S_{\text{eff}}\right))\sim 1/(E-E_{n})$.
-Remark 16.2.
-Localization of the spectrum vs. localization of path integral
-It is crucial to distinguish between the exact spectral condition and the exact computation of the path integral.
-In the complexified functional space, there are infinitely many directions away from the critical saddle point (where a complex eigenvalue vanishes).
-In the framework of Stochastic Quantization and the Complex Langevin Equation (CLE), the stochastic trajectory is effectively exploring a Lefschetz Thimble—the stable manifold of the gradient flow [5].
-The trajectory passes through the saddle and proceeds along an infinitely long, "flat" minimal valley, while stochastic noise causes it to fluctuate against the steep potential walls surrounding this valley.These transverse fluctuations contribute to the pre-exponential factor (comprising the functional determinant and higher-order expansion terms).
-However, as established in the topological study of Landau singularities by [41], the global properties of the integral are determined by the monodromy of the integration cycle.
-The trajectory must wind infinitely many times around the logarithmic branching point of the effective action $S_{eff}(z)$.Each circuit around this singularity shifts the Minkowski action by a fixed amount, $\mathrm{i}\Delta S_{eff}$.
-The condition for a pole $E=E_{n}$ in the total amplitude corresponds to the constructive interference of these windings, leading to the quantization condition:
-$\Delta S_{eff}=2\pi n$ (16.8)
-This monodromy is a topological invariant of the flow and is independent of the "shape" of the valley sides or the specific stochastic time, which are non-universal.
-This separation of local geometry from global topology is a manifestation of the spectral flow and the index theorems of [2], where the "integer" part of the quantization (the pole) remains robust against the local deformations of the path caused by Langevin noise.
-As demonstrated in our WKB analysis of the helicoid, this mechanism ensures that the energy spectrum is localized by the classical monodromy, even while the path integral remains delocalized by the CLE.
-16.4 One pole and its symmetries
-To construct the exact saddle point in the one-pole sector, we leverage the unbroken symmetries of the system.
-First, we fix the external kinematics by moving to the rest frame of the meson, orienting the injected momentum as $q=(0,0,0,\mathrm{i}E)$.
-While this specific frame choice explicitly breaks the full Lorentz group, the $O(3)$ spatial rotational symmetry remains intact.
-In a conventional approach modeling a generic “rubber band” string, one might utilize this unbroken space rotation to align the twistor dynamics such that the symmetric bilinear combination vanishes:
-$\lambda^{T}\sigma_{1}\mu=0$ (16.9)
-Because the spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$, this strict algebraic constraint would force one component of the tangent vector to be zero everywhere on the worldsheet.
-This would dimensionally reduce the physical minimal surface from a generic 4-dimensional embedding into a 3-dimensional hyperplane.
-However, here is where the profound difference between the conventional “rubber band” string and our twistor formulation manifests itself.
-Instead of crushing the surface into 3D, we enforce the exact chiral (Hodge-dual) configuration.
-Specifically, we freeze one of the chiral sectors by setting the right-handed twistor to a constant basis spinor, such as $\mu=(1,0)^{T}$.
-This choice explicitly breaks the 3D embedding constraint, causing the minimal surface to genuinely span a 4-dimensional space and flawlessly evade the finite-area self-intersection theorems.
-Because macroscopic QCD is strictly parity-conserving, the physical target-space configuration cannot remain purely left- or right-chiral.
-The true physical minimal surface is given by our exact prescription: the parity-even arithmetic mean of the left-chiral (Self-Dual or SD) and right-chiral (Anti-Self-Dual or ASD) surfaces.
-As derived in Appendix F, this symmetrization rigorously projects out the imaginary geometric artifacts of the 4D embedding, yielding a strictly real Nambu-Goto string geometry whose fundamental classical area is exactly halved relative to the conventional generic string.
-Furthermore, we possess a global $U(1)$ symmetry on the worldsheet integration domain (the unit disk).
-The extremum equation for the effective action with respect to the pole position $z_{0}$ must reflect this.
-Because of this $U(1)$ symmetry, we may claim that there is always a symmetric solution where the pole is centered at the origin:
-$z_{0}=0$ (16.10)
-The string dynamically centers itself to restore the $U(1)$ symmetry.
-By locking the pole to the origin, the minimal surface reduces exactly to the maximally symmetric canonical Helicoid.
-The remaining computations then proceed utilizing the arithmetic mean of the SD and ASD areas.
-In the text below, we simply use this symmetrized Hodge-dual metric.
-16.5 Quark Spin Monodromy and the Unified Regge Spectrum
-The exact meson spectrum is completed by accounting for the discrete spin monodromy of the boundary quarks.
-As algebraically demonstrated in F, the discretization of the quark loop into $2L$ steps yields a path-ordered Dirac trace $T\propto(-1)^{L}$.
-This phase factor provides an intrinsic topological index that modifies the constructive interference condition in the worldsheet amplitude, selecting specific sectors of the geometric series $\sum\exp\left(iS_{k}\right)$.
-The resulting Bohr-Sommerfeld resonance condition is $\Delta S_{\text{eff}}=\pi n_{BS}$.
-Based on the trace $T$, physical light mesons ($\pi$ and $\rho$) are found to occupy the even-$n_{BS}$ sectors.
-Defining the worldsheet topological quantum number $n=n_{BS}$ and evaluating the 1-cycle Minkowski bulk action for the symmetrized chiral Helicoid (Appendix F), we obtain the unified mass-squared formula:
-$m^{2}=\frac{\pi\sigma_{e,o}}{4}\left(n+\frac{1}{12}\right)$ (16.11)
-For even $L$, the projector supports quantization with $n\equiv 0\pmod{4}$ and $n\equiv 2\pmod{4}$.
-To map these topological sectors to the macroscopic meson spectrum and define the intercepts $\alpha(0)=J|_{m^{2}=0}$, we adopt the semi-classical WKB mapping $n=4J+q$:
-  • 1.
-    Unnatural Parity ($\pi$ and $K$): $q=0\implies n=4J$, placing the sequence in the $n\equiv 0\pmod{4}$ sector and yielding the intercept $\alpha_{\pi}(0)=-1/48$.
-  • 2.
-    Natural Parity ($\rho$): $q=-2\implies n=4J-2$, placing the sequence in the $n\equiv 2\pmod{4}$ sector and yielding the intercept $\alpha_{\rho}(0)=23/48$.
-As demonstrated in Fig.
-15, the experimental data confirms this topological identification.
-Because the $K$ mesons share the unnatural parity of the pion, they belong to the identical topological sector $n=4J$, with their trajectory cleanly displaced by a positive mass-squared shift $\Delta m_{s}^{2}\approx 0.17$ GeV2 reflecting the explicit $SU(3)$ flavor symmetry breaking by the heavier strange quark.
-The observed variation of the slopes ($\pi\sigma_{\pi}\approx 1.38$, $\pi\sigma_{\rho}\approx 1.16$, and $\pi\sigma_{K}\approx 1.58$ GeV2) remains an open dynamical question.
-It may arise from the interaction between a pole in $\lambda$ and a pole in $\mu$, which would locally renormalize the effective string tension.
-We mention this possibility here as a direction for future exploration but retain the distinct empirical fits to maintain contact with the PDG data.
-We must also address a natural phenomenological question regarding our selection of experimental candidates.
-A critical reader comparing our trajectories with the Particle Data Group (PDG) tables will note that we have not included all listed resonances possessing the same quantum numbers.
-This selection was deliberate, but it is not a mere phenomenological convenience; rather, it is strictly dictated by the scope of our analytic solution.
-The linear Regge trajectories plotted in Fig.
-15 correspond exclusively to the absolute simplest topological sector of the theory: the exact one-pole twistor solution.
-Our geometric framework guarantees that the remaining, unmapped hadronic resonances in the PDG are not anomalies, but are dynamically required to populate the daughter trajectories generated by multi-pole twistor configurations.
-Solving the complexified effective action—specifically, solving the saddle-point equations for the moduli $R_{i},z_{i}$ of the multi-pole solutions and classifying their monodromies—constitutes a deterministic, yet computationally formidable, nonlinear optimization problem in complex geometry.
-Systematically extracting these higher-topology solutions and mapping them onto the complete PDG database falls beyond the foundational scope of the present paper.
-Instead, we present this as a rigorous, data-rich research program—an open invitation to apply the sophisticated analytic machinery of string theory directly to the rich empirical data of hadronic spectroscopy.
-Figure 15: Unified Chew-Frautschi plot ($J$ vs $m^{2}$) showing the topological identification of the $\pi$ ($n=4J$), $K$ ($n=4J$) and $\rho$ (n=4J2n=4J-2) trajectories.
-The horizontal error bars represent propagated mass uncertainties, and labels indicate the experimental masses in MeV.
-The $K_{3}(2320)$ is marked with a question mark as it is currently omitted from the PDG Summary Tables pending further experimental confirmation.
-16.6 The microscopic origin of the Lüscher term: elves instead of string vibrations
-Our exact derivation of the strictly bounded $\pi/12$ topological phase unveils a fundamentally different, purely microscopic physical origin for the universally mandated Lüscher term, representing the absolute culmination of our theory.
-In standard Effective String Theory, this famous constant is interpreted strictly via the macroscopic Casimir effect.
-The standard calculation dictates that one must manually quantize the zero-point fluctuations of a macroscopic, highly excited “rubber band” string bending and vibrating across the transverse spacetime dimensions.
-The $\pi/12R$ effective potential is then forcefully extracted from the ultraviolet divergence of these vibrating spatial standing waves via Riemann zeta regularization.
-Our twistor-string formulation completely upends this paradigm.
-We bypassed the macroscopic quantization of target-space wiggles entirely.
-Instead, the quantum mechanics takes place directly on the 2-dimensional worldsheet.
-The Lüscher term originates intrinsically from the quantum anomaly of the microscopic zero-point fluctuations of the heavy fermions—the “elves”—that are confined to our minimal surface.
-When evaluating the fermion determinant (required to reproduce the planar factorization of the MM equation), integrating out these fundamental elfin degrees of freedom generates the quantum anomaly, related to the curvature.
-Because these quantum elves live on a highly curved minimal surface governed by the singular twistor conformal factor $\rho$, regulating their vacuum fluctuations inherently breaks conformal symmetry, generating the Weyl/Polyakov trace anomaly.
-The direct mathematical consequence of this trace anomaly is the rigid addition of the Liouville kinetic term $\frac{1}{3\pi}|\nabla\rho|^{2}$ to the effective action.
-Crucially, the numerical coefficient of this anomaly is strictly proportional to the central charge of the worldsheet—it literally counts the active quantum degrees of freedom of the elves.
-In a conventional generic 3-dimensional Helicoid, both chiral sectors dynamically fluctuate, which overcounts the anomaly (yielding an unphysical $\pi/6$).
-However, by enforcing the chiral (Hodge-dual) configuration and taking the parity-even arithmetic mean of the SD and ASD minimal areas, we dynamically halve the internal fluctuating degrees of freedom compared to the generic minimal surface.
-This exact symmetrization prescription makes the day: it mathematically halves the anomaly integral over the minimal surface, flawlessly dropping the topological phase to precisely $\pi/12$.
-[…]
+These rational values of $a$ would not correspond, in general, to integer or half-integer values of the angular momentum $J+q/2$ in (G.9) .
+In the limit $a\to 1/2$, the angular momentum tends to $\infty$ in this equation.
+The solutions $a_{l}(J,q)$ of the equation $g(a)=J+q/2$ are, in general, transcendental, so that the physical spectrum is not affected by this geometric extension.
+However, this phenomenon may be interesting from a purely mathematical point of view, as a number-theoretic exception to the Exact WKB quantization.
+In the H we rigorously prove that this is not the case for the physical branch , covered by $a_{\textbf{min}}<a<1/2$, where $a_{\textbf{min}}\approx 0.426836$...
+was introduced above.
 16 Planar QCD as a Geometric theory
@@
 If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour.
-Exactly at the moment of crossing, these two singularities inevitably collide directly on the physical integration contour $|z|=1$.
+Exactly at the moment of crossing, these two singularities collide directly on the physical integration contour $|z|=1$.
 As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity.
@@
 However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk.
-If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{k}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
-$\lambda(\theta)=\sum_{k}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (17.1)
+If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{i}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
+$\lambda(\theta)=\sum_{i}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (16.1)
 The stochastic CLE evolution simply forces these boundary parameters to evolve with Langevin time $\tau$:
-$\lambda(\theta,\tau)=\sum_{k}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (17.2)
+$\lambda(\theta,\tau)=\sum_{i}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (16.2)
 Upon analytic continuation back into the bulk, we obtain:
-$\lambda(z,\tau)=\sum_{k}\frac{R_{i}(\tau)}{z-z_{i}(\tau)}$ (17.3)
+$\lambda(z,\tau)=\sum_{i}\frac{R_{i}(\tau)}{z-z_{i}(\tau)}$ (16.3)
 The crucial observation is that the function remains strictly meromorphic at every instant of Langevin time.
@@
 Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions.
-The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/48-1/48 and $23/48$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
-In a profound sense, this construction explicitly realizes Edward Witten’s original “Master Field” hypothesis—the existence of a purely classical description for the large-$N_{c}$ limit of QCD.
-While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, we have discovered its exact geometric realization in loop space.
+The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/96-1/96 and $47/96$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
+One may view this construction as a geometric realization of the Master Field [54]—the existence of a purely classical description for the large-$N_{c}$ limit of QCD.
+While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, it finds an exact geometric representation in loop space.
 The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface.
@@
 By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition.
-For the fundamental one-pole trajectory, this yields the unified linear Regge spectrum:
-$m^{2}=\frac{\pi\sigma}{4}\left(n+\frac{1}{12}\right)$ (18.1)
-where $n=4J$ for the unnatural parity ($\pi,K$) trajectories and n=4J2n=4J-2 for the natural parity ($\rho$) trajectory.
-This analytic solution passes flawlessly through experimental data on Figure 15.
-The only outlier ($3\sigma$) is $K_{3}(2320)$ which is marked as "unconfirmed" in the PDG tables.
-The exact mass spectrum of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
+For the fundamental one-pole trajectory, this yields the unified parametric Regge trajectory (1.1) with trigonometric functions of the twistor branch point degree $a$, and $q=0$ for the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 for the shifted natural parity ($\rho$) trajectory.
+Although these functions are mathematically non-linear, they are numerically remarkably close to linear Regge trajectories.
+This analytic solution passes through the experimental data on Figure 15.
+The only outlier ($3\sigma$) is $K_{3}(2320)$, which is marked as “unconfirmed” in the PDG tables.
+The exact mass spectrum in our solution of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
 17.1 Future Directions: Glueballs, Baryons, and Scattering
@@
 Data Availability.
-No data was created in this paper. The analytic computations of the Magnus expanstion of the Momentum loop equations were performed in the Mathematica ® notebook published on the Wolfram Cloud[37].
+No data was created in this paper.
+The analytic computations of the Magnus expansion of the Momentum loop equations were performed in the Mathematica ® notebook published on the Wolfram Cloud[34].
+The computations of the Regge trajectories and fitting the slopes to the meson mass data are done in Mathematica ® notebook published on the Wolfram Cloud[35].
 Appendix A Conformal anomaly in determinants
@@
 Appendix C Explicit invariant tensor solutions for the MLE
+Order $n$ $N_{eq}(n)$ $N_{unk}(n)$ $\text{rank}(A_{n})$ $\dim\text{LeftNull}(A_{n})$ $\Delta\text{rank}$ Solvable?
+4 4 6 4 0 0 yes
+6 31 30 26 5 0 yes
+8 379 212 82 297 1 no
+Table 1: Linear algebra diagnostics for the exact Vector Momentum Loop Equation.
+Here $\Delta\text{rank}:=\text{rank}([A_{n}|b_{n}])-\text{rank}(A_{n})$.
+A strict positivity, $\Delta\text{rank}>0$, rigorously proves the system is mathematically inconsistent.
 In this appendix, we provide the explicit polynomial solutions for the finite functional Taylor–Magnus expansion of the Momentum Loop Equation, valid up to the non-analytic boundary at $\mathcal{W}^{(8)}$.
@@
 C.2 Counting of unknowns/equations and Fredholm inconsistency at $\mathcal{W}^{(8)}$
-Order $n$ $N_{eq}(n)$ $N_{unk}(n)$ $\text{rank}(A_{n})$ $\dim\text{LeftNull}(A_{n})$ $\Delta\text{rank}$ Solvable?
-4 4 6 4 0 0 yes
-6 31 30 26 5 0 yes
-8 379 212 82 297 1 no
-Table 1: Linear algebra diagnostics for the exact Vector Momentum Loop Equation.
-Here $\Delta\text{rank}:=\text{rank}([A_{n}|b_{n}])-\text{rank}(A_{n})$.
-A strict positivity, $\Delta\text{rank}>0$, rigorously proves the system is mathematically inconsistent.
 To rigorously demonstrate the breakdown of the Taylor-Magnus expansion at the 8th order ($\mathcal{W}^{(8)}$), we cast the Momentum Loop Equation at each order $n$ into a finite linear system.
@@
 To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$.
-The flat integration measure transforms as $dzd\bar{z}/|z|^{2}=d\sigma_{E}^{2}+d\tau_{E}^{2}$.
+The flat line element transforms as $|dz|^{2}=|z|^{2}(d\sigma_{E}^{2}+d\tau_{E}^{2})=e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$.
 Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
@@
 Thus, the induced Euclidean worldsheet metric is exactly:
-$ds^{2}_{E}=\exp\left(2\rho\right)(d\sigma_{E}^{2}+d\tau_{E}^{2})$
-$=R^{2}\left(1+e^{-2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
-To evaluate the physical states, we analytically continue this metric into Minkowski spacetime via a double Wick rotation.
+$ds^{2}_{E}=\exp\left(2\rho\right)e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$
+$=R^{2}\left(1+e^{2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
+To evaluate the physical area, we analytically continue this metric into Minkowski spacetime via a double Wick rotation.
 The angular parameter $\tau_{E}$ serves as the target space time.
@@
 $\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (D.7)
-Under this transformation, the exponential strictly rotates into a complex phase: $e^{-2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})-\mathrm{i}\sin(2\sigma_{M})$.
-Because the macroscopic QCD string is strictly parity-conserving, the physical state is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations.
-The conjugate right-chiral surface identically provides the inverse complex phase $e^{+2\mathrm{i}\sigma_{M}}$.
-Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the metric:
-$\text{Re}\left[\exp\left(2\rho\right)\right]=R^{2}\big(1+\cos(2\sigma_{M})\big)=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
-Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact physical Minkowski metric is:
-$ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (D.9)
-To explicitly verify that this describes the canonical rotating string, we change from the conformal spatial coordinate $\sigma_{M}$ to the physical target space radial coordinate $r$.
-For a rigidly rotating string, the fractional radius is $r=\sin(\sigma_{M})$.
-The differentials transform as:
-$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (D.10)
-Simultaneously, the time component of the metric incorporates the trigonometric identity $\cos^{2}(\sigma_{M})=1-\sin^{2}(\sigma_{M})=1-r^{2}$.
-Substituting these directly into the Minkowski metric, we obtain:
-$ds^{2}_{M}=2R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (D.11)
-This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})d\tau_{E}^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light, but exactly halved in overall magnitude relative to the generic three-dimensional configuration.
+Under this transformation, the exponential strictly rotates into a complex phase: $e^{2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})+\mathrm{i}\sin(2\sigma_{M})$.
+Because the macroscopic QCD string is strictly parity-conserving, the physical area is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations.
+The conjugate right-chiral surface identically provides the inverse complex phase $e^{-2\mathrm{i}\sigma_{M}}$.
+Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the area:
+$\text{Re}\left[\exp\left(2\rho\right)e^{2\mathrm{i}\sigma_{M}}\right]$ $=R^{2}\big(1+\cos(2\sigma_{M})\big)$
+$=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
 D.3 Action Evaluation and the Halved Lüscher Shift
@@
 To mathematically illustrate the appearance of the conjugate pole at $1/\bar{z}_{0}$ and the resulting impenetrable topological barrier at the unit circle, consider the evaluation of the Liouville kinetic anomaly term in the effective action.
-In for a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
-$I=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
-We can rewrite the integrand as a total derivative with respect to $\bar{z}$:
-$I\propto\int_{D}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
-By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $|z|=1$:
-$I\propto\oint_{|z|=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
+In the case of a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
+$I(z_{0})=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
+In order to continue this integral as an analytic function of $z_{0}$ we rewrite the integrand as a total derivative with respect to $\bar{z}$:
+$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}<1}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
+By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $z\bar{z}=1$:
+$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
 Crucially, on the unit circle, the complex conjugate is strictly given by the geometric inversion $\bar{z}=1/z$.
 Substituting this into the logarithm yields the exact boundary representation of the action:
-$I\propto\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
+$I(z_{0})=\frac{-\mathrm{i}}{2}\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
 This exact formula explicitly reveals the analytic structure governing the boundary dynamics.
 The integrand possesses a simple pole at $z=z_{0}$ (originating from the twistor itself) and a logarithmic branch point at $z=1/\bar{z}_{0}$ (originating from the boundary inversion of the conjugate twistor).
+At $z_{0}=0$ this integral reduces to
+$I(0)=\frac{-\mathrm{i}}{2}\int_{0}^{2\pi}\mathrm{i}\,d\theta\left(-\mathrm{i}\theta\right)=\frac{-\mathrm{i}}{2}\frac{(2\pi)^{2}}{2}=-\mathrm{i}\pi^{2}$ (E.5)
+which matches with the computation of the Minkowski-continued area in the previous section.
 Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror strictly resides outside the disk ($|1/\bar{z}_{0}|>1$).
@@
 Assuming even number $2L$ of points on a unit circle in discretized loop, and multiplying these matrices around the entire loop $\oint d\theta=2\pi$ reduces the propagator for one cycle to:
-$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{4\pi i}&0\\ 0&e^{-4\pi i}\end{pmatrix}$ (F.4)
+$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{2\pi i}&0\\ 0&e^{-2\pi i}\end{pmatrix}$ (F.4)
 The physical trace $T$
 […]
-The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulkates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$.
-which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\mathrm{i}\pi n$ depending on the parity of $L$.
+The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$, which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\pi n$ depending on the parity of $L$.
+Appendix G Geometric Potentials and Universal Trajectories
+In this appendix, we decouple the unknown dimensionful string tension $\sigma$ from the geometric quantization conditions to derive the parametric functions $h(a)$ and $g(a)$, their singularities, and their roots.
+We formulate the system directly in terms of the fractional winding parameter $a$.
+As derived in Section 15.5, the macroscopic spatial rotation angle of the string is exactly $\phi=4\pi a$.
+We isolate the string tension from the boundary potential $F(a)$ by defining the dimensionless geometric function:
+$\tilde{F}(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (G.1)
+The boundary potential is $F(a)=\frac{1}{\pi\sigma}\tilde{F}(a)$.
+Defining the physical mass-squared as $m^{2}=E^{2}=\sigma h(a)$, the unknown tension $\sigma$ factors to the numerator of the mass equation and cancels out of the angular momentum equation entirely.
+The total geometric phase for the $q$-shifted trace sector is:
+$\Phi_{q}(a)=-E^{2}F(a)+\frac{\pi a^{2}}{12}+4\pi aJ+2\pi qa$ (G.2)
+where $q=0$ for the unshifted $\pi/K$ integral, and q=1q=-1 for the shifted $\rho$ and Scalar integrals.
+We impose the saddle-point condition $\Phi_{q}^{\prime}(a)=0$ to isolate the angular momentum terms:
+$4\pi aJ+2\pi qa=-\frac{\pi a^{2}}{6}+aE^{2}F^{\prime}(a)$ (G.3)
+Substituting this into the Bohr-Sommerfeld quantization condition $\Phi_{q}(a)=2\pi n_{w}$ (where $n_{w}=0$ for the principal string sector), we obtain the algebraic relation:
+$-\frac{\pi a^{2}}{12}-E^{2}F(a)+aE^{2}F^{\prime}(a)=0$ (G.4)
+Using the differential identity $aF^{\prime}(a)-F(a)=\frac{1}{\pi\sigma}\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)$, we isolate $m^{2}=\sigma h(a)$ to obtain the mass function:
+$h(a)=\frac{a^{2}\pi^{2}}{12\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)}$ (G.5)
+Simplifying that expression, we get an explicit trigonometric answer:
+$h(a)=-\frac{\pi^{2}(a-1)^{2}a^{2}\csc^{4}(\pi a)}{24(-2a+4\pi(a-1)a\cot(\pi a)+1)}$ (G.6)
+To obtain the angular momentum function, we substitute $E^{2}=\sigma h(a)$ back into the saddle-point equation. Defining the target angular momentum as $u=J+q/2$, we obtain:
+$u=\frac{a\left(a\tilde{F}^{\prime}(a)-2\tilde{F}(a)\right)}{48\left(\tilde{F}(a)-a\tilde{F}^{\prime}(a)\right)}$ (G.7)
+This expression can be dramatically simplified.
+Dividing the numerator and denominator by $\tilde{F}(a)$ to isolate the logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)$, we obtain a partial fraction decomposition:
+$\frac{\tilde{F}^{\prime}(a)}{\tilde{F}(a)}=\frac{2a+48u}{a(a+48u)}=\frac{1}{a}+\frac{1}{a+48u}$ (G.8)
+Substituting the explicit logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)=4\pi\cot(\pi a)+\frac{1}{1-a}$ and combining the kinematic poles yields the exact trigonometric spectral equation:
+$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (G.9)
+The Regge trajectory can also be simplified to the following
+$J+\frac{q}{2}=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{a}{48}$ (G.10)
+G.1 Singularities and the Ground States
+The geometry of the functions is dominated by a kinematic pole at the speed-of-light boundary, $a\rightarrow 1/2$.
+At this limit, the left-hand side of (G.9) strictly vanishes ($\cot(\pi/2)=0$).
+For the equation to hold, the right-hand side must also vanish, which dynamically forces the angular momentum to diverge ($u\rightarrow\infty$) and forms the linear macroscopic Regge asymptote.
+Evaluating the derivative $\tilde{F}^{\prime}(1/2)=8$ in the mass formula (G.5) restores the macroscopic Regge slope $J\approx 2m^{2}/(\pi^{2}\sigma)$.
+The asymptotic expansion at larger $s=m^{2}/\sigma$ corresponds to an inverted expansion around $a\to 1/2-\epsilon$.
+The expansion of trajectory goes as follows:
+$J(s)$ $=\frac{2s}{\pi^{2}}+B_{0}+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+O(s^{-5});$ (G.11)
+$B_{0}$ $=-\left(\frac{q}{2}+\frac{1}{96}\right);$ (G.12)
+$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}$ (G.13)
+$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}$ (G.14)
+The coefficients $B_{1},B_{3}$ are numerically so small, that the trajectory is indistinguishable from linear for $J>1$.
+When we compute the roots for the ground states, the explicit spectral equation (G.9) yields the mass eigenvalues:
+  • 1.
+    The Light Pseudoscalar ($q=0,u=0$): Solving (G.9) yields $a_{\pi}\approx 0.4268$.
+Substituting into $h(a)$ gives the mass ratio $m_{\pi}^{2}/\sigma\approx 0.0375$.
+The geometry produces a light ground state without symmetry breaking parameters.
+  • 2.
+    The Vector Ground State (q=1,u=1/2q=-1,u=1/2): The topological shift forces the string to spin faster.
+Solving (G.9) occurs at $a_{s}\approx 0.4987$, yielding $m_{s}^{2}/\sigma\approx 1.9782$.
+The geometry predicts the vector state to be massive.
+The continuous generation of the parametric trajectories $u=J+q/2$, the numerical roots, and the data fits to $\sigma$ are performed using the companion Mathematica notebook [35].
+G.2 The Endpoint Singularity and the Fold Catastrophe
+While the large-$s$ asymptotic expansion (G.11) governs the strictly linear macroscopic Regge regime, the exact parametric geometry of the twistor string dictates a hard geometric boundary at small $s$.
+This boundary manifests not as a smooth continuation toward $s\to$-$\infty$, but as a fold catastrophe in the mapping between the spectral parameter $a$ and the physical observables.
+To expose this geometric termination, we rewrite the exact parametric equations for the mass ratio $s(a)=h(a)=m^{2}/\sigma$ and the target angular momentum $u(a)=J+q/2$.
+From (G.6) and (G.8), we define the auxiliary functions:
+$\mathcal{P}(a)$ $=\frac{\pi^{2}a(1-a)}{24\sin^{4}(\pi a)},$ (G.15)
+$\mathcal{R}(a)$ $=4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}.$ (G.16)
+The spectral equations then take the remarkably compact algebraic form:
+$s(a)$ $=\mathcal{P}(a)\mathcal{R}(a)^{-1},$ (G.17)
+$48u(a)$ $=\mathcal{R}(a)^{-1}-a.$ (G.18)
+A fundamental property of the Rigid Twistor String is the existence of a simultaneous mathematical endpoint $a_{c}$ where the first derivatives of both the mass-squared and the angular momentum vanish exactly.
+We can rigorously prove this simultaneous singularity by differentiating the target angular momentum:
+$48u^{\prime}(a)$ $=-\mathcal{R}^{\prime}(a)\mathcal{R}(a)^{-2}-1=0$
+$\implies-\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}.$ (G.19)
+Next, we differentiate the mass function and multiply by $\mathcal{R}(a)^{2}$:
+$s^{\prime}(a)\mathcal{R}(a)^{2}=\mathcal{R}(a)\left[\mathcal{P}^{\prime}(a)+\mathcal{P}(a)\mathcal{R}(a)\right].$ (G.20)
+By explicitly taking the logarithmic derivative of $\mathcal{P}(a)$, we uncover an exact geometric identity:
+$\frac{\mathcal{P}^{\prime}(a)}{\mathcal{P}(a)}=\frac{1-2a}{a(1-a)}-4\pi\cot(\pi a)\equiv-\mathcal{R}(a).$ (G.21)
+Consequently, the term in the brackets vanishes identically. Thus, we arrive at the exact algebraic statement: $s^{\prime}(a_{c})=0$ precisely whenever $u^{\prime}(a_{c})=0$.
+In the physical $(s,J)$ plane, this simultaneous extremum signifies a fold catastrophe.
+The parametric Regge trajectory reaches a hard geometric endpoint where the physical slope $dJ/ds=u^{\prime}(a)/s^{\prime}(a)$ resolves (via L’Hôpital’s rule) to a finite limit $S_{0}$, while the trajectory curvature diverges to infinity, $J^{\prime\prime}(s)\to\infty$.
+The critical parameter $a_{c}$ is determined by the exact transcendental equation $\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}$, which evaluates to:
+$4\pi^{2}\csc^{2}(\pi a)-\frac{1-2a+2a^{2}}{a^{2}(1-a)^{2}}$
+$=\left(4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}\right)^{2}.$ (G.22)
+Solving this within the fundamental domain $a\in(0,1/2)$ yields the universal critical parameters:
+Crit. Param. ($a_{c}$): $\approx 0.311277$
+Thres. Mass Rat. ($s_{c}=m_{c}^{2}/\sigma$): $\approx 0.027791$
+Crit. $\pi$ Spin ($q=0$): $J_{c}\approx-0.003378$
+Crit. $\rho$ Spin (q=1q=-1): $J_{c}\approx 0.496622$ (G.23)
+Figure 16: The exact geometric Regge trajectories plotted down to the mathematical endpoint $(s_{c},J_{c})$.
+The trajectory departs from this boundary with a finite slope ($S_{0}\approx 0.1118$) that is smaller than the asymptotic slope at large $s$ ($\approx 2/\pi^{2}\approx 0.2026$).
+The infinite curvature $J^{\prime\prime}(s)\to\infty$ at the exact endpoint is visually imperceptible.
+To the eye, this looks like a perfectly linear trajectory ending abruptly at this point.
+At this mathematical endpoint, the Regge trajectory departs with a finite slope $S_{0}\approx 0.1118$ and infinite curvature (see Fig.
+16).
+While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15).
+The physical branch of the spectrum strictly corresponds to $u\geq 0$.
+As established previously, the physical ground state terminates at $a_{min}$ where $u=0$.
+For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
+Because the physical root $a_{min}$ sits comfortably above the critical parameter $a_{c}$ ($0.4268>0.3113$), the trajectory mathematically terminates well before the string can experience the fold catastrophe.
+This mass gap ($s(a_{min})>s_{c}>0$) acts as an impenetrable geometric shield.
+It guarantees that the Regge trajectory never crosses into the $s<0$ regime, rigorously protecting the Planar QCD spectrum from tachyonic instabilities while enforcing the exact reality of all physical eigenvalues.
+The phenomenological Regge “intercepts” $\alpha_{\pi,\rho}(0)$ are thus not true physical states, but simply linear extrapolations of the large-$s$ asymptotic expansion bridging over this hidden, non-linear geometric boundary.
+While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15).
+The physical branch of the spectrum strictly corresponds to $u\geq 0$.
+As established previously, the physical ground state terminates at $a_{min}$ where $u=0$.
+For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
+Appendix H Irrationality of the Physical Branch Points
+As discussed in Section 15.8, rational values of the branch point degree $a=m/n$ would cause the twistor trajectory to close after finitely many windings.
+In that case, all distinct states would be counted with just $2n$ windings, and our mechanism for the Exact WKB spectrum would disappear.
+In this appendix, we rigorously prove that the physical spectrum is mathematically protected from these rational degeneracies.
+Theorem. For any rational target angular momentum $u=J+q/2\geq 0$, the exact spectral equation (G.9) possesses no rational solutions for $a\in(0,1/2)$.
+Proof. The exact geometric quantization is governed by the spectral equation derived in Appendix G:
+$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+48u}$ (H.1)
+Assume, for the sake of contradiction, that the physical branch point $a$ is a rational number $m/n\in(0,1/2)$.
+Since the target momentum $u$ is also rational by definition, the algebraic expression on the right-hand side evaluates to a purely rational number.
+On the left-hand side, it is a foundational property of trigonometric functions that for any rational argument $a$, the value of $\cot(\pi a)$ is an algebraic number.
+Furthermore, for $a\in(0,1/2)$, the value $\cot(\pi a)$ is strictly positive and therefore non-zero.
+By the Lindemann–Weierstrass theorem, $\pi$ is a transcendental number.
+The product of a transcendental number and a non-zero algebraic number is strictly transcendental.
+Consequently, the left-hand side is a transcendental number, while the right-hand side is a rational number.
+This is a strict mathematical contradiction.
+Therefore, the branch point parameter $a$ cannot be a rational number.
+$\blacksquare$

### v4

@@
 Abstract
-We present the exact analytic solution of the Makeenko-Migdal loop equations, thereby solving planar QCD $(N_{c}\rightarrow\infty)$ in the continuum limit.
-The solution is constructed by quantizing internal Majorana fermions (elves) on a rigid Hodge-dual minimal surface.
-We demonstrate that these worldsheet degrees of freedom provide the exact algebraic mechanism required to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing the planar factorization.
-In the local limit, the theory reduces to a confining analytic twistor string.
-By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles.
-The simplest sector with just one pole yields parametric Regge trajectories governed by trigonometric functions.
-While mathematically non-linear, these geometric trajectories are numerically remarkably close to linear, matching the experimental data for light mesons (the unshifted $\pi,K$ trajectories and the topologically shifted $\rho$ trajectory) within 95% confidence intervals.
-The empirically correct intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=+47/96$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor branch point monodromy, without any ad hoc assumptions of string vibrations.
-Ultimately, this exact geometric quantization explicitly realizes the Master Field as a critical classical trajectory in twistor space.
+We present a local, asymptotically free solution of the planar MM loop equations in the continuum limit with full Lorentz invariance.
+The solution is constructed by quantizing internal Majorana fermions (referred to here as “elves”) on a rigid Hodge-dual minimal surface.
+These worldsheet degrees of freedom provide the mechanism to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing planar factorization.
+In the local limit, the theory reduces to a confining analytic twistor-string representation.
+By analyzing the monodromy structure of the complexified effective action, we show that the discrete mass spectrum is organized by topological data associated with twistor singularities.
+The simplest sector with one branch point yields parametric Regge trajectories expressed by a universal formula with trigonometric functions.
+These trajectories are approximately linear for light mesons over a broad range and are in agreement with experimental data for 36 light, strange, charmed, and bottom mesons.
+The asymptotic behavior of the trajectory $J=\alpha(M^{2})$ and its subleading corrections arise from the interaction between the Liouville term and the twistor monodromy, without introducing additional assumptions about string excitations.
+In our solution, the QCD mass spectrum follows from a generalized eigenvalue problem in complexified phase space, reducing the problem to classical geometry.
+Within this framework, the large-$N_{c}$ Master Field is realized as a trajectory in twistor space.
 1 Introduction
-The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the MM loop equations [^1] The Makeenko-Migdal equations, referred to hereafter as the MM equations, exhibit the characteristic factorization properties of a string theory.
-While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible..
-A concrete, mathematically controlled continuum realization has remained elusive.
-Historically, the main obstruction is the ultraviolet singular nature of the coordinate-space Wilson loop functional $W[C]$.
+The identification of a string description of large-$N_{c}$ QCD has been a long-standing problem since the formulation of the MM loop equations [^1] The Makeenko–Migdal equations, referred to hereafter as the MM equations, exhibit factorization properties characteristic of a string description.
+The foundational derivation is due jointly to the author and Yuri Makeenko..
+A continuum realization remains nontrivial.
+A technical challenge is the ultraviolet singular behavior of the coordinate-space Wilson loop functional $W[C]$.
 Even after renormalization, $W[C]$ exhibits cusp and perimeter divergences, and the loop equation contains a contact term $\delta^{(4)}(x-y)$ at self-intersections.
-In the standard coordinate representation, this makes the local limit technically delicate and conceptually opaque: the loop equation is local, while the perturbative solution is nonlocal and develops logarithmic singularities tied to cusp geometry.
-In this paper, we adopt a different viewpoint: rather than attempting to “repair” the coordinate-space Wilson loop as an observable object, we formulate the dynamics directly in momentum loop space.
-The momentum loop functional $W[P]$ was introduced in earlier work as the functional Fourier transform of $W[C]$.
-Its decisive advantage is structural: the planar loop equation in momentum space becomes an exact algebraic-differential functional equation with no coordinate-space delta functions.
-In this representation, the geometric cusp singularities of $W[C]$ are integrated out, and the continuum equations admit a local formulation that is finite in the sense relevant for physical quark-loop amplitudes.
-The goal of this second paper in the Geometric QCD series is to provide the dynamics that complement the kinematics established in Part I [36].
-Part I constructs the rigid vacuum geometry: the Hodge-dual additive minimal surface $S_{\chi}[C]$ whose (anti)self-dual area derivative yields a zero mode of the loop operator, producing the confining dressing factor $\exp\left(-\kappa S[C]\right)$.
-Here, we attach dynamical worldsheet degrees of freedom to this rigid background without summing over fluctuating metrics and show how the (regularized) planar loop equations are satisfied by this Ansatz for $W[C]$.
-Our framework synthesizes three ingredients:
+In the coordinate representation, the loop equation is local, while perturbative solutions are nonlocal and develop logarithmic singularities at cusps and intersections.
+In this work, we adopt a different approach.
+Rather than attempting to renormalize the coordinate-space Wilson loop as an observable, we formulate the dynamics in momentum loop space.
+The momentum loop functional $W[P]$, introduced in [26], is the functional Fourier transform of $W[C]$.
+In this representation, the planar loop equation becomes an algebraic–differential functional equation without coordinate-space delta functions.
+The geometric singularities of $W[C]$ are integrated out, and the continuum equations admit a local formulation appropriate for quark-loop amplitudes.
+The aim of this second paper in the Geometric QCD series is to provide the dynamical construction complementing the geometric framework developed in Part I [32].
+In that work, the Hodge-dual additive minimal surface $S_{\chi}[C]$ was shown to generate a confining factor $\exp\left(-\kappa S[C]\right)$ as a zero mode of the loop operator.
+Here, we introduce worldsheet degrees of freedom on this rigid surface and construct a functional $W[C]$ that satisfies the planar loop equations.
+Our framework combines four elements:
   1. 1.
     Fermi string (“Elfin theory”).
-We revisit the fermionic construction in which internal Majorana degrees of freedom on the worldsheet implement the required sign structure at intersections.
-In particular, the Pauli principle enforces cancellations between nonplanar intersection configurations, leaving the planar factorization pattern.
+Internal Majorana degrees of freedom on the worldsheet implement planar topology.
+The Pauli principle cancels nonplanar intersection configurations, leaving planar contributions.
   2. 2.
-    Rigid Hodge-dual minimal surface.
-The fermions are not placed on a random surface.
-Instead, they live on the rigid Hodge-dual minimal surface determined holographically by the boundary loop.
-This removes the dynamical Liouville instability that afflicted the original random-surface formulation: there is no summation over worldsheet metrics, and the geometry is fixed by the loop data.
+    Rigid Hodge-dual minimal surface. The fermions propagate on a surface determined by the boundary loop. There is no integration over worldsheet metrics, and the geometry is fixed by the loop data.
   3. 3.
-    Momentum loop dynamics and the twistor resolution.
-In momentum loop space, the loop equation becomes an exact algebraic-differential equation, stripped of coordinate-space contact singularities.
-We rigorously demonstrate that while this equation admits a functional Taylor-Magnus expansion at low orders, a purely one-dimensional algebraic Master Field solution violently breaks down at the eighth order ($\mathcal{W}^{(8)}$).
-This mathematical catastrophe proves that a 1D topological string fundamentally lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD.
-To resolve this geometric anomaly, the reduced loop-space measure must be parametrized by 4D continuous boundary twistor variables.
-This leads to an effective “analytic twistor string” description: a boundary sigma-model (a coset built from two $S^{3}$ spinors modulo a local $U(1)$) coupled to a holographic Liouville field whose bulk values are determined by the unique holomorphic extension of the boundary twistors.
+    Momentum loop dynamics.
+In momentum loop space, the loop equation takes an algebraic–differential form.
+A representation in terms of the Cuntz algebra exists [26], but a one-dimensional Master Field construction exhibits a rank deficiency at order $\mathcal{W}^{(8)}$ in the Voiculescu expansion, indicating the need for additional degrees of freedom.
+  4. 4.
+    Twistor parametrization.
+The reduced loop-space measure is parametrized by boundary twistor variables.
+This yields a twistor parametrization in which a boundary sigma model is coupled to a Liouville field determined by analytic continuation from the boundary.
 Remark 1.1.
 Equation solving vs.
-model building The starting point of this work is not a proposed “string model of QCD,” but the planar MM loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD.
-The aim is constructive: to exhibit explicit variables and an explicit functional measure in which the loop equations admit a finite continuum solution and yield physical observables.
-The principal ingredients of the present construction are not adjustable assumptions.
-The rigid Hodge-dual minimal surface enters because strict additivity and an (anti)self-dual area derivative are algebraically required for compatibility with the loop hierarchy, while the Majorana determinant enters because its Pauli-principle sign structure enforces planar factorization.
-[…]
-The culminating result of this paper is the exact reduction of the meson spectroscopy problem to a one-dimensional functional integral over boundary twistor data $(\lambda(z),\mu(z),|z|=1)$.
-By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory.
-The path integral exactly localizes at critical resonance energies where the Lefschetz thimbles become degenerate.
-Because the integration variables are coupled holomorphic twistors, this degeneracy intrinsically possesses a corank of 2, mathematically upgrading generic semiclassical branch cuts into the exact simple poles of hadronic states.
-Planar QCD thus ceases to be a theory of stochastic quantum fluctuations; the extraction of the exact spectrum reduces to a deterministic, classical generalized eigenvalue problem.
-The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral.
-We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent.
-Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle.
-This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$.
-By analytically continuing the action to the boundary and summing over the infinite number of windings, with $SU(2)$ rotated boundary conditions for twistors, after projecting on the fixed spin state we obtain an exact Bohr-Sommerfeld quantization condition.
-For the simplest sector with just one pole, this yields the unified parametric Regge trajectory:
-$J=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{q}{2}-\frac{a}{48};$ (1.1)
-$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (1.2)
-The topological trace phase $q=0$ identifies the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 rigidly shifts the natural parity ($\rho$) trajectory by $1/2$.
-Remarkably, while mathematically non-linear, these trajectories are numerically very close to straight lines (see Figure 15).
-The empirically correct asymptotic intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=47/96$ emerge directly from the 2D Liouville kinetic anomaly.
-The asymptotic expansion of the Regge trajectory $J(s)$ at $s=m^{2}/\sigma>1$ reads:
-$J(s)$ $=\frac{2s}{\pi^{2}}-\left(\frac{q}{2}+\frac{1}{96}\right)+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+\dots);$ (1.3)
-$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}\approx 0.0000340208;$ (1.4)
-$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}\approx 2.29*10^{-9}$ (1.5)
-We interpret this construction as a geometric realization of the large-$N_{c}$ Master Field [54].
-While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, it admits an explicit formulation in loop space.
-The true classical limit of planar QCD is thus identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory[45].
-This geometric structure is accessible via exact WKB analysis [50, 12], where twistor monodromies dynamically generate flat valleys within the Lefschetz thimbles [55, 3].
-Organization of the paper and navigation points.
-To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with five specific navigation points that address anticipated physical and computational questions.
-Section 2 summarizes the loop-equation zero modes and recalls why the Hodge-dual minimal surface provides the appropriate rigid geometric background.
-For readers concerned about (i) the physical meaning of loop-equation zero modes, Sec.
-2.3 explicitly details the underlying multi-instanton and vacuum ambiguity in loop space.
-Sections 3 to  9 then develop the fermionic determinant on this rigid surface and derive the corresponding loop equation, including its local limit.
-Sections 10 to 12 introduce momentum loop space and the twistor string framework.
-To clarify (ii) the status of Wilson-loop “renormalization,” Sec.
-10.1 sharply outlines our perspective: $W[C]$ is kept regularized at a finite Elf mass, physical observables are constructed natively in momentum loop space $W[P]$ (where coordinate-space singularities are integrated out), and only then is the local limit taken.
-Following this, Sections 11 to  12 execute the central change of variables.
-Addressing (iii) the parametrization of the quark loop in phase space by boundary values of holomorphic twistors, Sec.
-11 details how the linear phase space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
-Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
-Finally, we highlight (iv) the exact analytic extraction of the mass spectrum.
-Section 15 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory.
-Section 16 then reframes this localization as an infinite-dimensional realization of Picard-Lefschetz theory, Resurgence, and Exact WKB analysis.
-We establish a strict structural correspondence between the twistor monodromies of Geometric QCD and the spectral networks of Seiberg-Witten theory, identifying the exact resonance condition as the twistor-string analogue of an Argyres-Douglas point.
-Section 17 concludes by synthesizing these results into the final geometric Master Field.
-The appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C and computations of the Minkowski action for our chiral helicoid minimal surface.
+model building.
+The starting point of this work is the planar MM loop equations, treated as defining equations of large-$N_{c}$ Yang–Mills theory.
+The goal is to identify variables and a functional measure in which these equations admit a finite continuum realization.
+The ingredients of the construction are constrained by this requirement.
+The Hodge-dual minimal surface appears because its area satisfies the loop equation, while the Majorana determinant satisfies planar factorization, reproducing planar graphs of QCD.
 Remark 1.2.
-Conceptual Demarcation: Bypassing 2D Quantum Gravity
-Because the mathematical vocabulary of this paper (strings, twistors, minimal surfaces) intersects with decades of literature on string theory, a reader may instinctively attempt to map this construction onto the conventional quantization of a fluctuating 2D worldsheet.
-To prevent fundamental conceptual misunderstandings, it is vital to explicitly state that we are not proposing a phenomenological string model, nor are we quantizing a fundamental string in the Polyakov sense.Consequently, the reader will not find integrations over fluctuating 2D bulk metrics, conformal ghost sectors, or integrations over ambiguous Teichmüller moduli spaces.
-Attempting to evaluate this work by the standard criteria of abstract 2D string modeling is a category error.Our absolute starting point is the rigorously defined, 1-dimensional phase space of Planar QCD: the momentum-loop representation of the MM equations, governed by the exact linear measure in phase space $\int\mathcal{D}C^{\prime}\mathcal{D}P$.
-The transition to "twistor space" is strictly a local, 1-dimensional algebraic change of variables on the boundary circle.
-The twistors do not dynamically fluctuate across a 2D bulk; they are localized strictly at the 1D boundary to parametrize the loop velocity, and are analytically continued inward solely as a tool of complex analysis.
-We are transforming the intractable functional geometry of a 1D gauge-theory phase space into a rigorously solvable problem of 1D complex analysis and linear algebra.
+Role of vector loop equations and the extrinsic surface geometry.
+A central technical point underlying this construction is that the full, unintegrated vector form of the MM loop equation contains essential information that is not captured by scalar or integrated formulations.
+In particular, when the momentum loop functional $W[P]$ is expanded in a one-dimensional Taylor–Magnus series, the resulting algebraic system becomes overconstrained at order $\mathcal{W}^{(8)}$.
+This is not a limitation of the expansion technique but a genuine inconsistency: the number of independent tensor constraints exceeds the number of available scalar degrees of freedom.
+This obstruction indicates that a one-dimensional loop functional does not possess sufficient internal structure to represent a solution of the vector loop equations.
+The missing tensorial degrees of freedom are associated with the extrinsic geometry of the embedding surface.
+In the present work, these are provided by the boundary twistor variables and the fermionic determinant on the Hodge-dual minimal surface.
+From this perspective, the construction provides the geometric structure required for a consistent local solution of the planar loop equations in the continuum limit.
 2 Zero modes of Loop equation
-Let us summarize the theory of Hodge-dual minimal surfaces and its relevance to the MM loop equation.
-For a detailed discussion and proofs of relevant theorems, the reader may look into the latest paper [36], and the details of the new loop calculus without divergences can be found in [33, 37].
+The theory of Hodge-dual minimal surfaces and its application to the MM loop equation is summarized here.
+For detailed discussion and proofs of the relevant theorems, see [32]; further details on the loop calculus without divergences are provided in [29, 33].
 2.1 Self-dual area derivative and loop equation
-Let us recall why we are looking for the self-dual minimal surface, as suggested in [33].
-In this section, we follow the arguments of our recent paper [36], but our goal is not to present the Hodge-dual surface per se, but rather to view it as a hologram of the QCD string, with Fermi degrees of freedom attached.
-The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
+As suggested in [32], we examine the self-dual minimal surface, which we will later treat as a background geometry for the QCD string with attached Fermi degrees of freedom.
+The area derivative for a functional $W[C]$ is defined as the discontinuity of the second variation of the area:
 $\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=\frac{\delta^{2}W}{\delta\dot{C}_{\mu}(\theta-0)\delta\dot{C}_{\nu}(\theta+0)}-\mu\leftrightarrow\nu$ (2.1)
-This definition implements the geometric definition of [24, 27] and is formally equivalent to the functional derivative implementation by Polyakov [41].
-The crucial difference is the lack of divergences in this new definition and its unbroken parametric invariance.
-In the old, singular implementations, the zero modes were obscured by kinematic divergencies, interfering with potential small-distance singularities of perturbative QCD.
-The new loop equation for the Yang–Mills gradient flow was shown in [33] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
+This implements the geometric definition of [20, 23] and is formally equivalent to the functional derivative implementation in [36].
+A significant feature of this definition is the absence of divergences and its parametric invariance.
+In previous implementations, the zero modes were often complicated by kinematic divergences.
+The loop equation for the Yang–Mills gradient flow reduces to the loop space diffusion equation [29], admitting the solution $W=\exp\left(-\kappa S[C]\right)$ provided $S[C]$ is a zero mode:
 $\mathcal{L}_{\nu}(S)=0;$ (2.2)
 $\mathcal{L}_{\nu}=\left(\frac{\delta}{\delta\dot{C}_{\mu}(\theta+0)}-\frac{\delta}{\delta\dot{C}_{\mu}(\theta-0)}\right)\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$ (2.3)
-The basic observation made in that paper was that there is a Hodge-dual solution similar to the Yang–Mills multi-instanton.
-The loop space diffusion equation (2.3) is identically satisfied for any functional $S[C]$ with self-dual (SD) or anti-dual (ASD) area derivative
+There exists a Hodge-dual solution analogous to the Yang–Mills multi-instanton.
+The loop space diffusion equation (2.3) is satisfied for any functional $S[C]$ with a self-dual (SD) or anti-dual (ASD) area derivative:
 $\frac{\delta S_{\pm}}{\delta\sigma_{\mu\nu}(\theta)}=\pm\ast\frac{\delta S_{\pm}}{\delta\sigma_{\mu\nu}(\theta)};$ (2.4)
 $e_{\alpha\mu\nu\lambda}\left(\frac{\delta}{\delta\dot{C}_{\alpha}(\theta+0)}-\frac{\delta}{\delta\dot{C}_{\alpha}(\theta-0)}\right)\frac{\delta S_{\pm}}{\delta\sigma_{\mu\nu}(\theta)}\equiv 0;$ (2.5)
-The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [33] for a proof).
-The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor.
-This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
-The existence of this zero mode is based on the so-called Leibniz property of the loop operator and similar singular loop operators in loop calculus.
+The Bianchi identity (2.5) holds as a Jacobi identity for a triple commutator [29].
+The exponential of the Hodge-dual minimal area serves as a factor in a general solution to the loop equation.
+This zero mode relies on the Leibniz property of the loop operator:
 $\mathcal{L}_{\nu}f(\Phi[C])=f^{\prime}(\Phi[C])\mathcal{L}_{\nu}\Phi[C];$ (2.6)
 $\mathcal{L}_{\nu}(A[C]B[C])=\mathcal{L}_{\nu}(A[C])B[C]+A[C]\mathcal{L}_{\nu}(B[C])$ (2.7)
-It was heuristically claimed in the original MM papers [25, 27], and rigorously proven in the loop calculus [33], section 3, eqs (39), (40). Due to the Leibnitz property,
+This property leads to:
 $\mathcal{L}_{\nu}\left(\exp\left(-\kappa S[C]\right)W_{0}[C]\right)=$
@@
 $+\exp\left(-\kappa S[C]\right)\mathcal{L}_{\nu}(W_{0}[C])=\exp\left(-\kappa S[C]\right)\mathcal{L}_{\nu}(W_{0}[C])$ (2.8)
-We are looking for such a self-dual minimal surface which is additive at the self-intersecting loops.
-This condition will be discussed in detail below.
-It is critical for the whole theory—due to this additivity, the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the whole chain of the loop equations.
-2.2 Compatibility with the MM loop equations.
-The full chain of loop equations for the Wilson loops with a finite number of colors $N_{c}$ [25, 27] has the same loop operator (2.3) as the Yang–Mills gradient flow, except there is no time derivative.
-In this paper, we are only considering the leading order of the $1/N_{c}$ expansion.
-We use the loop calculus developed in the paper [33], as it applies to the loop diffusion operator (2.3), which is the same in the multiloop MM chain and in the Yang–Mills gradient flow equation.
-The MM loop equation has the structure (see Fig.1)
-Figure 1: The MM equation, with the loop diffusion operator on the left side, and the integral term on the right side.
-The double line on the right stands for the delta function $\delta^{4}(x-y)$ enforcing self-intersection.
+We identify a self-dual minimal surface that is additive at self-intersections.
+This additivity ensures the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the chain of loop equations.
+2.2 Compatibility with the MM loop equations
+The chain of loop equations for Wilson loops at finite $N_{c}$ [21, 23] utilizes the same loop operator (2.3) as the Yang–Mills gradient flow.
+We consider the leading order of the $1/N_{c}$ expansion using the loop calculus developed in [29].
+The MM loop equation is structured as follows:
+Figure 1: The MM equation, with the loop diffusion operator on the left side and the integral term on the right side.
+The double line represents the delta function $\delta^{4}(x-y)$ at self-intersections.
 […]
 $\mathcal{L}_{\nu}(W[C])=\lambda\int_{C}dy_{\nu}\delta^{4}(x-y)W[C_{xy}]W[C_{yx}]$ (2.10)
-Here $\lambda=N_{c}g_{0}^{2}$ is the ’t Hooft’s coupling constant in the limit of $N_{c}\to\infty$.
-This equation is singular at self-intersections, so it must be understood in the sense of distributions in loop space.
-It is manifestly parameterization-invariant, and the confining factor $\exp\left(-\kappa S[C]\right)$ factors out from the right side by additivity of $S[C]$ and cancels with the same factor on the left side.
-Let us stress that this factorization takes place independently of the splitting points $x,y\in C$.
-For every self-intersection $x,y\in C$ such that $C(x)=C(y)$, the product of two factors $\exp\left(-\kappa S[C_{xy}]\right)$ and $\exp\left(-\kappa S[C_{yx}]\right)$ exactly equals the factor $\exp\left(-\kappa S[C]\right)$ on the left side of the MM equation; Thus, it can be taken out of the integral, after which it cancels on both sides of the MM equation.
-The physical hypothesis of that paper was that the actual gluonic vacuum of QCD is described by this dressed solution, choosing the "fluctuative" factor $W_{\text{fluct}}$ as the seed.
-While the symmetry is exact, the separation into a fluctuative factor and a purely geometric area law is ambiguous.
-It is physically most distinct in the large-loop limit, where the area term dominates over $W_{\text{fluct}}$, which is assumed only to contribute the term proportional to a perimeter of the loop.
-In the present paper, we construct an explicit solution for this factor $W_{\text{fluct}}$, adapting the fermionic determinant from the ’81 paper [28] to the Hodge-dual minimal surface instead of a random surface, assumed in that work.
-The confining factor becomes the vacuum energy of that rigid Fermionic string, and planar graphs emerge as a strong coupling expansion of that 2D fermi system in terms of a sum over fermionic loops on a Hodge-dual minimal surface.
-The Pauli principle leads to repulsion of these loops, making them topologically equivalent to planar graphs of QCD.
-The comparison of the contour equation for these fermionic determinants from the ’81 work with the MM equation establishes the correspondence between the ’t Hooft coupling constant $\lambda$ and the parameters of the Elfin theory (the bare fermionic mass and the vacuum energy).
-2.3 The Physical Vacuum and Multi-Instanton Resummation
-Before we formulate the dynamics of the Fermi string on the minimal surface, we must clarify the physical origin of this rigid geometric background.
-The MM (MM) loop equations serve as the exact quantum equations of motion for Planar QCD.
-Like any equations of motion, they possess an inherent ambiguity corresponding to the choice of the physical vacuum.
-The mathematical freedom to multiply the perturbative loop functional by a zero mode exactly reflects this physical vacuum ambiguity.
-In the traditional gauge-field picture, the nonperturbative QCD vacuum is a highly complex mixed state, often modeled as a nontrivial distribution in the multi-instanton space.
-Because any single multi-instanton solution explicitly breaks translation invariance by localizing topological charge at specific spacetime points, restoring Poincaré symmetry requires an exceedingly complicated integration over the entire multi-instanton moduli space (positions, sizes, and color orientations).
-The loop-space formulation achieves this summation geometrically.
-Just as instantons are characterized by the self-duality of the gauge field strength ($F_{\mu\nu}=\pm*F_{\mu\nu}$), our exact zero mode is characterized by the self-duality of the loop-space area derivative.
-Thus, the complex mixed state of multi-instantons in coordinate space simplifies to a single, fully translation-invariant, exact geometric object in loop space: the Hodge-dual minimal surface.
-Note that the mere Hodge duality condition leaves the freedom of an arbitrary topological charge and arbitrary moduli of the multi-instanton solution.
-Therefore, our Hodge-dual minimal surface, by satisfying this requirement, represents a general statistical mixture of multi-instanton solutions, implicitly summed over topological charges and integrated over moduli with some unknown measure.
-This implicit integration is precisely what provides the exact translation invariance and spatial isotropy of the minimal area solution.
-The profound advantage of the loop-space approach is that we do not need to know the explicit details of this distribution of topological charge and multi-instanton moduli: the loop equations can be solved directly without ever specifying this measure.
-Therefore, the rigid background upon which our worldsheet fermions propagate is not merely a mathematical ansatz.
-It is the exact holographic representation of the translation-invariant, nonperturbative QCD vacuum, naturally incorporating the sum over topological sectors that is analytically prohibitive in the standard gauge-field path integral.
+where $\lambda=N_{c}g_{0}^{2}$.
+This equation is understood in the sense of distributions.
+The confining factor $\exp\left(-\kappa S[C]\right)$ factors out by additivity and cancels on both sides of the equation independently of the splitting points $x,y\in C$.
 3 Hodge-dual minimal surface
-Now, we are going to define the solvable functional $S[C]$ satisfying both the duality and additivity.
-We are following the definitions and derivations from [36], adapting them for the present purpose of a base for the fermions living on that minimal surface.
+We now define the functional $S[C]$ that satisfies the requirements of both duality and additivity.
+We follow the definitions and derivations from [32], adapting them for the present purpose of a base for the fermions living on that minimal surface.
 3.1 The external and internal geometry
-We define the surface coordinates $X^{i}_{\mu}(\xi)$ (for $\xi=(\xi_{1},\xi_{2}),i=1,2,3$.
-[^2] This surface was first proposed in [33] as a zero-mode of the loop operator in the context of the Yang–Mills gradient flow, but unfortunately, the wrong number of components (four instead of three) was chosen for the $X^{A}_{\mu},A=1,2,3,4$.
-The rest of the theory in that paper was validated by our recent work [36], including the loop calculus and the Hodge-duality of the minimal surface.
-The field $X^{i}_{\mu}(\xi)$ maps the disk $\Sigma$ (Fig 2) to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$.
+We define the surface coordinates $X^{i}_{\mu}(\xi)$ (for $\xi=(\xi_{1},\xi_{2}),i=1,2,3$. The field $X^{i}_{\mu}(\xi)$ maps the disk $\Sigma$ to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$.
 Figure 2: The unit disk $\Sigma$ is mapped to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$ and then projected to $\mathbb{C}^{4}$ by a holomorphic map, minimizing the area functional.
@@
 Regardless of this sign change, the Hodge duality of the area element $\Sigma$ with these boundary conditions for a generic field $X^{a}_{\mu}(z,\bar{z})$ leads to the same Hodge duality of the area derivative.
-This self-duality of the area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations [25, 27], as we argue in the next section.
-Note that this area functional is invariant under the reparametrizations of the boundary loop $\delta_{\text{param}}C_{\mu}(\theta)=\epsilon(\theta)\partial_{\theta}C_{\mu}(\theta)$
-$\delta_{\text{param}}S=-2\int d\theta\epsilon(\theta)\frac{\partial_{\theta}C_{\mu}\Sigma_{\mu\nu}\partial_{\theta}C_{\nu}}{\sqrt{\Sigma^{2}}}=0$ (3.10)
-by the skew symmetry of $\Sigma_{\mu\nu}$.
-The Hodge-dual area is defined as a minimum of this functional under an extra duality constraint
-$\ast\Sigma_{\mu\nu}\equiv\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta}=\chi\Sigma_{\mu\nu}$ (3.11)
-In the previous papers [33, 36], we found a complete solution for the Hodge-dual surface in terms of holomorphic map.
+This self-duality of the area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations.
+The area functional is invariant under the reparametrizations of the boundary loop $\delta_{\text{param}}S=0$ by the skew symmetry of $\Sigma_{\mu\nu}$.
 3.2 The Holomorphic Map
@@
 […]
-So, we have a line in four dimensional complex space $f\in\mathbb{C}^{4}$. Our surface area (3.1) reduces to the induced metric in this space
+Our surface area (3.1) reduces to the induced metric in this space
 $\Sigma_{\mu\nu}=2(F_{\mu\nu}+\chi\ast F_{\mu\nu});$ (3.14)
@@
 $g_{ab}=\partial_{z_{a}}Y_{\mu}\partial_{z_{b}}Y_{\mu};\qquad z_{a},z_{b}=(z,\bar{z})$ (3.17)
-This area element is Hodge-dual
-$\ast\Sigma_{\mu\nu}=\chi\Sigma_{\mu\nu}$ (3.20)
-for arbitrary functions $f_{\mu}(z)$.
-This duality condition, as argued above, together with the above boundary conditions, provides the self-duality of the area derivative of the minimal area.
-The self-duality of the area derivative follows from the general formula for variation of the surface $X^{a}_{\mu}(\xi)$ with area (3.1), projected on the holomorphic Ansatz.
+This area element is Hodge-dual $\ast\Sigma_{\mu\nu}=\chi\Sigma_{\mu\nu}$ for arbitrary functions $f_{\mu}(z)$.
+This duality condition provides the self-duality of the area derivative of the minimal area.
 3.3 The Virasoro Constraint and Uniformization
@@
 $f^{\prime}_{\mu}(z)^{2}=0$ (3.18)
-The imposition of the null constraint $(f^{\prime})^{2}=0$ is not an approximation but a rigorous gauge fixing procedure rooted in the geometric symmetries of the problem.
-The physical starting point is the geometric Area functional, $S=\int\sqrt{-\det g}\,d^{2}z$, which possesses full reparametrization invariance under general diffeomorphisms $z\to w(z,\bar{z})$.
-This infinite symmetry group allows us to choose a specific coordinate system on the world-sheet without altering physical results.
-We choose to work in isothermal (conformal) coordinates, where the induced metric is diagonal:
-$ds^{2}=\rho(z,\bar{z})\,dz\,d\bar{z}\implies g_{zz}=\partial_{z}Y\cdot\partial_{z}Y=0$ (3.22)
-For our specific ansatz (3.12), the metric component $g_{zz}$ is explicitly proportional to the square of the holomorphic derivative:
-$g_{zz}=(f^{\prime}_{\mu})^{2}=0$ (3.23)
-Thus, the geometric condition of conformal gauge fixing $g_{zz}=0$ is algebraically equivalent to the null constraint $(f^{\prime})^{2}=0$.
-The existence of such a coordinate system for any surface topology is guaranteed by the Uniformization Theorem [39].
-This theorem ensures that we are always free to set $(f^{\prime})^{2}=0$ by a suitable diffeomorphism.
-In this gauge, the non-linear geometric area density simplifies exactly to the quadratic Lagrangian:
-$\mathcal{L}$ $=\sqrt{-\det g}=\sqrt{(g_{z\bar{z}})^{2}-|g_{zz}|^{2}}$
-$\xrightarrow{(f^{\prime})^{2}=0}g_{z\bar{z}}\propto|f^{\prime}|^{2}$ (3.24)
-This reduction allows us to describe the Minimal Surface using the linear Laplace equation for $f_{\mu}$, subject to the Virasoro constraint. To summarize, we have found that
+The imposition of the null constraint $(f^{\prime})^{2}=0$ is a gauge fixing procedure rooted in the geometric symmetries of the problem.
+This allows us to work in isothermal (conformal) coordinates, where the induced metric is diagonal: $g_{zz}=(f^{\prime}_{\mu})^{2}=0$.
+The existence of such a coordinate system is guaranteed by the Uniformization Theorem.
+In this gauge, the non-linear geometric area density simplifies to the quadratic Lagrangian:
 $S_{\chi}[C]=2\sqrt{2}\int_{D}|f^{\prime}(z)|^{2}d^{2}z;$ (3.19)
 $\frac{\delta S_{\chi}[C]}{\delta\sigma_{\mu\nu}}=-2\frac{\Sigma_{\mu\nu}}{\sqrt{\Sigma_{\mu\nu}^{2}}}$ (3.20)
-The solution of the linear boundary problem (3.15) for a holomorphic vector function is given by the Hilbert transform
+The solution of the linear boundary problem (3.13) for a holomorphic vector function is given by the Hilbert transform:
 $f_{\mu}=\frac{1}{2}(1+\mathrm{i}\mathcal{H})C_{\mu};$ (3.21)
@@
 $\int_{D}|f^{\prime}(z)|^{2}d^{2}z\;=\;\int_{0}^{2\pi}d\theta\;C^{\prime}_{\mu}(\theta)\,H[C_{\mu}](\theta);$ (3.23)
-In terms of the Taylor coefficients
+In terms of the Taylor coefficients:
 $f_{\mu}(z)=\sum_{n>0}z^{n}C_{\mu,n};$ (3.24)
 $C_{\mu}(\theta)=\sum_{n=-\infty}^{\infty}e^{\mathrm{i}n\theta}C_{\mu,n}$ (3.25)
-In the presence of the Virasoro constraint (3.21), the problem becomes a challenging one and has been investigated over the last two centuries by great mathematicians, starting with Riemann and Hilbert.
-We give a brief summary of modern state of this theory in B.
-In Part I [36], we compared the ordinary minimal surface in four dimensions with the Hodge-dual one.
-This comparison reveals some hidden variables present in the area derivatives of the Hodge dual surface, but not in the extremal value of this area.
-This minimal value up to a normalization factor $2\sqrt{2}$ equals that of the so called Goldschmidt solution [42, 6, 15] for the Plateau problem.
-This additive solution represents a local minimum of the Dirichlet functional, with the simplest topology (collection of disks one for each closed part of $C$).
-The area derivatives for the Goldschmidt minimal area and Hodge dual minimal area are different (the second one being SD or ASD component of the first).
-The apparent paradox is resolved by the obvious fact that the area derivative of a functional equals the skew-symmetric part of its second functional derivative, but only the extremal value and its first derivative (=0=0) coincide for the functional and its minimal value.
-Also, the Hodge-dual surface is not parity-symmetric, but its minimal value (minimal area) is.
-The chirality $\chi$ manifests itself in the external geometry of this minimal surface $X^{i}_{\mu}\in\mathbb{R}^{3}\otimes\mathbb{R}^{4}$, including its area derivative, but its internal geometry is determined by the induced metric (3.22),(3.23), which is the only local function related to surface embedding.
-Moreover, this metric in the holomorphic map that minimizes the area depends only on the bounding loop $C$, through the holomorphic vector function $f_{\mu}(z)$, which solves the Riemann-Hilbert problem (9.4), (3.30).
-This dependence is identical for the Hodge-dual surface and the Goldschmidt solution for the Plateau problem in 4D, bounded by $C$.
-The last property (the identity of the internal geometry in the Hodge-dual surface in relation to that of the 4D Euclidean problem) will play an important role in proving the MM equation for the fermionic determinant.
-The only remnants of the Hodge-duality are the overall negative sign and the $\chi$-term in the area derivative (3.26) projecting it on the Hodge-dual sector of skew-symmetric tensors in 4D.
+The internal geometry of the Hodge-dual surface is identical for the Hodge-dual surface and the Goldschmidt solution for the Plateau problem in 4D.
+This property is central for establishing the MM equation for the fermionic determinant.
+The four-dimensional solution for the minimal surface uses twistors (see B). This is the solution that we use later to build the Regge trajectories.
+3.4 The Physical Vacuum, Master field and Multi-Instanton Resummation
+Before formulating the dynamics of the Fermi string on the minimal surface, we clarify the origin of the rigid geometric background.
+The MM loop equations serve as quantum equations of motion for Planar QCD.
+As with any equations of motion, they admit a family of solutions distinguished by the choice of vacuum.
+The freedom to multiply the perturbative loop functional by a zero mode reflects this vacuum ambiguity.
+In the traditional gauge-field picture, the nonperturbative QCD vacuum is a complex mixed state, often modeled as a distribution in multi-instanton moduli space.
+Because any individual multi-instanton configuration breaks translation invariance, restoring Poincaré symmetry requires integration over the full moduli space.
+In the loop-space formulation, this ensemble is represented geometrically.
+Just as instantons are characterized by the self-duality of the gauge field strength, the corresponding zero mode is characterized by the self-duality of the loop-space area derivative.
+The instanton moduli are mapped to the moduli of the twistors parametrizing the four-dimensional minimal surface.
+The physical meaning of this geometric representation is analogous to the Master Field proposed by Witten [46], which describes the large-$N_{c}$ limit in terms of a single classical configuration in an infinite-dimensional space.
+In the present formulation, the Hodge-dual minimal surface plays this role.
+It provides a classical, translation-invariant object encoding the nonperturbative vacuum, without requiring an explicit construction in terms of gauge-field configurations.
+In this sense, the minimal surface represents the loop-space realization of the Master Field, with its twistor moduli corresponding to the degrees of freedom associated with the multi-instanton ensemble.
+Moreover, the whole surface $X^{i}_{\mu}(\xi)\in\mathbb{R}^{3}\otimes\mathbb{R}^{4}$ has the same algebraic structure as a gauge field in an $SU(2)$ subalgebra.
+The observables in our theory are the loop functionals and their area derivatives, related to the left and right Gauss maps of the surface.
+The twistor construction of the minimal surface is thus analogous to a multi-instanton configuration.
+Rotations in the internal $\mathbb{R}^{3}$ act as $SO(3)$ transformations on $X^{i}_{\mu}$, corresponding to $SU(2)$ gauge transformations, and physical observables depend only on $SO(3)$-invariant combinations.
+The minimal surface described by these twistors is non-fluctuating, reflecting the classical nature of the Master Field in the planar limit.
+Thus, the complex mixed state of multi-instantons in coordinate space is mapped, in loop space, to a single translation-invariant geometric object: a Hodge-dual minimal surface.
+In this representation, the detailed distribution of topological charge and the explicit multi-instanton moduli are not required.
+The loop equations can be solved directly, without specifying the underlying measure in coordinate space.
+The resulting geometric object defines a translation-invariant nonperturbative vacuum for the worldsheet fermions.
+The degrees of freedom associated with instanton moduli are encoded in the moduli of the minimal surface.
+In four dimensions, these moduli admit a natural description in terms of holomorphic twistor data.
 4 The Elfin theory on a flat surface
-In this section, we deal with the Elfin theory by itself, without any reference to the gauge theory. The equivalence of these theories will be established in the next sections.
+The elfin theory is defined here independently of the gauge theory; the equivalence is established in subsequent sections.
 Let us consider the flat surface bounded by some contour $C$.
@@
 In the applications to the QCD spectrum, we allow such meromorphic $f^{\prime}_{\mu}$ with poles in the unit disk; the induced conformal factor $e(z,\bar{z})\propto|f^{\prime}_{\mu}(z)|^{2}$ is single-valued but may diverge at these poles.
-This does not invalidate the analysis of the Elfin determinant in the flat gauge: closed fermion paths are weighted by $\exp\left(\!\right)\big(-m\int_{\Gamma}\sqrt{e}\,|dz|\big)=\exp\left((\right)-m\,\ell_{\Gamma})$, hence any path approaching a pole (where $\sqrt{e}\sim|z-w|^{-1}$ for a simple pole of $f^{\prime}$) has $\ell_{\Gamma}\to\infty$ and is exponentially suppressed.
+This does not invalidate the analysis of the Elfin determinant in the flat gauge: closed fermion paths are weighted by $\exp\left(-m\int_{\Gamma}\sqrt{e}\,|dz|\right)=\exp\left(-m\,\ell_{\Gamma}\right)$, hence any path approaching a pole (where $\sqrt{e}\sim|z-w|^{-1}$ for a simple pole of $f^{\prime}$) has $\ell_{\Gamma}\to\infty$ and is exponentially suppressed.
 Since $\ell_{\Gamma}$ is a local additive functional, the length weights of the locally paired configurations (touching vs.
@@
 Let us now discuss the symmetry of the physical $x$-space.
-As usual in string theory, the transformations of the internal coordinates $\xi$ have nothing to do with the physical rotations in $x$-space.
+Transformations of the internal coordinates $\xi$ are independent of physical rotations in $x$-space.
 From the viewpoint of two-dimensional field theory in $\xi$-space, the rotations in $x$-space represent internal $O(4)$ symmetry transformations of the $C_{\mu}$ field,
@@
 $=2\int_{S}t_{\mu\nu}(\xi_{2})d^{2}\xi_{2}\delta(\xi_{1}-\xi_{2})$ (4.31)
-Using the exact area derivatives of the Hodge-dual minimal area established in the previous section, we find for the area derivative of
+Using the area derivatives of the Hodge-dual minimal area established in the previous section, we find for the area derivative of
 $\frac{\delta e(\xi)}{\delta\sigma_{\mu\nu}(\xi_{0})}$ $=-2T_{\mu\nu}\delta^{2}(\xi-\xi_{0});$ (4.32)
@@
 is invariant under local variations of the conformal metric field $\rho$.
-That is, the theory depends only on the conformal class (moduli) of the surface, and the induced metric $e(\xi)$ but not on the local gauge parameter $\rho(\xi)$.
+That is, the theory depends only on the conformal class (moduli) of the surface, and the induced metric $e(\xi)$, remaining independent on the local gauge parameter $\rho(\xi)$.
 Proof.
@@
 6.1 Mathematical justification of the signed sum over loops
-The path integral representation of the Dirac determinant on a finite surface with boundary conditions, as employed in this work, is rigorously justified by several mathematical results and earlier physical constructions.
+The path integral representation of the Dirac determinant on a finite surface with boundary conditions, as employed in this work, is justified by several mathematical results and earlier physical constructions.
 In particular, Ichinose [14] and Gaveau [9] provide theorems establishing the equivalence between the determinant of the Dirac operator and path integrals over spinor-valued paths with appropriate boundary conditions on finite domains.
 These constructions carefully incorporate the spinor connection and holonomies of spin rotations along geodesic segments, ensuring the correct treatment of spinor phases and curvature effects.
-Our representation of the determinant as a sum over oriented closed loops weighted by factors of the form $\prod\exp\left(-m\,dl+\frac{\mathrm{i}\sigma_{3}\delta\theta}{2}\right)$, where $dl$ is an infinitesimal geodesic step and $\delta\theta$ the corresponding trajectory rotation angle, follows directly from these rigorous path integral formulations.
+Our representation of the determinant as a sum over oriented closed loops weighted by factors of the form $\prod\exp\left(-m\,dl+\frac{\mathrm{i}\sigma_{3}\delta\theta}{2}\right)$, where $dl$ is an infinitesimal geodesic step and $\delta\theta$ the corresponding trajectory rotation angle, follows directly from these path integral formulations.
 This approach generalizes the combinatorial and topological methods pioneered by Vdovichenko [41] in the exact solution of the 2D Ising model, where the fermionic nature and spinor holonomies enforce cancelations of non-planar intersections, leaving only planar loop configurations that contribute to the determinant.
@@
 This ensures that the fermionic statistics and boundary conditions are correctly encoded in the loop amplitudes.
-Thus, the path integral representation employed here is not only physically intuitive but also mathematically sound, grounded in well-established results in spin geometry, spectral theory of Dirac operators, and rigorous path integral constructions on manifolds with boundary.
+Thus, the path integral representation employed here is not only physically intuitive but also mathematically sound, grounded in well-established results in spin geometry, spectral theory of Dirac operators, and path integral constructions on manifolds with boundary.
 This justifies its use as the foundation for deriving loop equations and analyzing the planar topology of the fermion determinants in the Elfin theory.
@@
 For an oriented regular simple closed curve $\gamma\subset\mathbb{R}^{2}$, the rotation (turning) index satisfies $\operatorname{rot}(\gamma)=\frac{1}{2\pi}\int_{\gamma}\kappa\,ds\in\mathbb{Z}$ and equals +1+1 for the positively oriented boundary of a disk [45].
-In the planar case, this is exactly the Gauss–Bonnet theorem for the enclosed region (since $K\equiv 0$), yielding $\int_{\gamma}\kappa\,ds=2\pi\,\chi=2\pi$ (hence equality modulo $2\pi$ is automatic) [2].
+In the planar case, this is precisely the Gauss–Bonnet theorem for the enclosed region (since $K\equiv 0$), yielding $\int_{\gamma}\kappa\,ds=2\pi\,\chi=2\pi$ (hence equality modulo $2\pi$ is automatic) [2].
 Moreover, since $\operatorname{rot}(\gamma)$ is the degree of the unit-tangent map $T:S^{1}\to S^{1}$, it depends only on the immersed curve itself and therefore is unchanged if the ambient plane is replaced by a disk with finitely many holes (and the same mod-$2$ intersection sign $(-1)^{\nu}$ follows from Whitney’s regular-homotopy classification of immersed plane curves) [45].
@@
 These loops are independent by construction, but we are going to rearrange the sum in such a way that they will start to repel.
-This miraculous transformation, first conjectured in [26] is possible due to the following topological and combinatorial observations.
+This rearrangement, conjectured in [22] relies on the following topological and combinatorial observations.
 6.5 Cancellation of intersections: step towards planar graphs
@@
 $l_{tot}=\sum l_{loops}=\sum l_{lines}.$ (6.16)
-This statement is not as trivial as it seems. The length of the line is not just a geometric length of this line drawn on a planar disk; rather, it is an integral
+The length of the line is not just a geometric length of this line drawn on a planar disk; rather, it is an integral
 $l_{line}=\int|dz(t)|\sqrt{e\left(z(t),\bar{z}(t)\right)}$ (6.17)
@@
 When the lines cross each other in coordinate space, the term in the vertex changes sign.
-The non-planar vertex (the second one on the lower drawing in Fig.
-3 ) cancels the first one, so this configuration of lines drops from the sum.
-This is the heart of the matter.
+The non-planar vertex (Fig.
+3) cancels the first one, removing this configuration from the sum.
+This is the central mechanism.
 The lines may touch each other with opposite orientations according to Fig.
@@
 One $\psi,\bar{\psi}$ pair is placed at the two ends of the string.
-[^3] This interpretation of the time slice dynamics can be used for a rigorous proof of the equivalence of the free Dirac vacuum loops and the planar loops made of non-intersecting closed paths on a surface.
+[^2] This interpretation of the time slice dynamics can be used for a proof of the equivalence of the free Dirac vacuum loops and the planar loops made of non-intersecting closed paths on a surface.
 Such a proof would use the Hamiltonian representation of a Fermionic system with conventional creation/annihilation operators and chiral boundary conditions.
@@
 Eq.
-(7.1) was postulated in the first version of the Fermi String theory [26] from heuristic arguments.
+(7.1) was originally proposed in [22].
 At that time, it seemed that the equation applied to the Ising model at the surface, i.e., for the 2-component spinors.
@@
 Lemma 7.1 (Geodesic Inequality).
-For any minimal surface embedded in Euclidean space, the geodesic distance $L_{geo}(x,y)$ is strictly bounded by the Euclidean distance:
+For any minimal surface embedded in Euclidean space, the geodesic distance $L_{geo}(x,y)$ is bounded by the Euclidean distance:
 $L_{geo}(x,y)\geq|x-y|_{\mathbb{R}^{4}}.$ (7.13)
@@
 This precise alignment relies on the isotropic scaling of White’s bridge in the varifold limit $\epsilon\to 0$.
-In the strict limit of the topology change, the continuity of the area Hessian demands that the tangent planes perfectly align, yielding the required Fierz-like identity for the MM loop equation without generating anomalous contact terms.
+In the strict limit of the topology change, the continuity of the area Hessian demands that the tangent planes align, yielding the required Fierz-like identity for the MM loop equation without generating anomalous contact terms.
 A rigorous geometric measure-theory proof of the Hessian’s continuity across this topology change is left for future mathematical study, but is physically mandated by the $O(4)$ symmetry of the local loop diffusion.
@@
 The Elfin theory induces QCD in the limit of large fermion mass, serving as a short-distance cutoff.
-There are no singularities at finite mass, which eliminates the messy discussion about cusp and perimeter singularities of the Wilson loop.
+There are no singularities at finite mass, which resolves the ambiguities associated with cusp and perimeter singularities of the Wilson loop.
 The Wilson loop is defined in the Elfin theory as a regularized solution of the loop equation, with the delta function smeared by a finite fermion mass.
-This definition of the loop equation makes mathematical sense, in our opinion: regularize this equation by smearing the delta function on the right side while preserving the gauge invariance by keeping all loops closed.
+This naturally regularizes the loop equation by smearing the delta function on the right side while preserving the gauge invariance by keeping all loops closed.
 The Elfin theory falls into the category of regularized solutions of the loop equation.
@@
 In the old MM papers [21, 19], the bootstrap equation was iterated in the coupling constant $\lambda$, starting with $W[\mathds{1}]=1$.
-The derivation was rather tedious, but it faithfully reproduced conventional planar graphs, including Faddeev-Popov ghost loops.
+Iterative evaluation reproduces conventional planar graphs, including Faddeev-Popov ghost loops.
 The first order planar graph follows immediately from the (8.1), as shown on Fig.
@@
 leading to logarithmic singularities at the cusps of the loop $C$.
-Where is this non-locality hiding in the Elfin theory, and how will these logarithms emerge?
-In this section, we dig deeper into the formal solution we have found, searching for the perturbative singularities of planar QCD.
+We now trace the origin of this non-locality and the emergence of logarithmic singularities within the Elfin theory.
 9.1 Local expansion of the heavy fermion determinant
@@
 There is always another zero mode confining factor $G[C]=\exp\left(-\delta\sigma S[C]\right)$, renormalizing the string tension $\sigma$, as we found in the previous work [32].
-Equivalently, the area term generated in the heavy-mass expansion of the determinant provides an additive renormalization of the vacuum-energy/string tension sector, already parametrized by the zero-mode dressing $\exp\left((\right)-\kappa S[C])$; only the sum defines the single physical string tension $\sigma_{phys}=2\sqrt{2}\kappa_{phys}$ in the continuum theory, while $M$ acts strictly as a regulator scale removed in the local limit.
+Equivalently, the area term generated in the heavy-mass expansion of the determinant provides an additive renormalization of the vacuum-energy/string tension sector, already parametrized by the zero-mode dressing $\exp\left(-\kappa S[C]\right)$; only the sum defines the single physical string tension $\sigma_{phys}=2\sqrt{2}\kappa_{phys}$ in the continuum theory, while $M$ acts as a regulator scale removed in the local limit.
 In our framework, we do not integrate over fluctuating metrics (which would lead to the Polyakov-Liouville anomaly).
@@
 $S_{cusp}=\frac{3}{11}(\gamma\cot\gamma-1)$ (9.10)
-This mechanism is strictly analogous to the continuum limit in Lattice Gauge Theory, where the bare coupling $g_{0}\to 0$ as the lattice spacing $a\to 0$ to keep physical masses finite.
+This mechanism is analogous to the continuum limit in Lattice Gauge Theory, where the bare coupling $g_{0}\to 0$ as the lattice spacing $a\to 0$ to keep physical masses finite.
 In our induced QCD, the cutoff has a physical and mathematical origin in the regularization of the minimal surface, so the limit $m\to\infty$ (where $\lambda\to 1/(\beta_{0}\log m)$) has a precise mathematical meaning, replacing the formal subtraction of infinities with a finite limiting process.
@@
 However, the lattice definition comes at a cost: it explicitly breaks the continuous symmetries of four-dimensional Euclidean space, from the rotation group $O(4)$ to subtle topological structures like the Hodge duality of the area tensor.
-For instance, exact instanton solutions or Hodge-dual minimal surfaces do not exist on a discrete grid; they only emerge in the continuum limit.
+For instance, instanton solutions or Hodge-dual minimal surfaces do not exist on a discrete grid; they only emerge in the continuum limit.
 These disadvantages are compensated by the universality of the Renormalization Group (RG) flow.
@@
 This universality allows for various microscopic definitions of QCD, provided they land in the correct phase (the basin of attraction of the asymptotically free fixed point).
-History offers cautionary tales.
+Previous lattice approaches illustrate the sensitivity of this limit.
 The large-N reduced Eguchi-Kawai model, for example, encountered unphysical fixed points.
@@
 Standard perturbation theory dictates that $W[C]$ requires multiplicative renormalization to remove divergences associated with cusps and self-intersections.
-Are we obliged to redefine the loop equation to make the coordinate-space Wilson loop finite in the local limit?
-We argue that this is a misguided requirement that obscures the underlying dynamics.
+However, redefining the loop equation solely to render the coordinate-space Wilson loop finite in the local limit obscures the underlying dynamics.
 Physical renormalizability demands only the existence of a universal mass scale $\Lambda_{QCD}$ and the finiteness of observable scattering amplitudes.
@@
 10.2 Momentum loop path integral
-As we discussed in the previous section, the ordinary Wilson loop is not directly observable, and its notorious singularities are merely intermediate results that will eventually disappear (cancel or integrate out) in observables.
+As we discussed in the previous section, the ordinary Wilson loop is not directly observable, and its singularities are merely intermediate results that will eventually disappear (cancel or integrate out) in observables.
 Unlike the standard Feynman-Schwinger worldline representation in coordinate space, which is inherently plagued by the UV cusp singularities of $W[C]$, we utilize the phase-space path integral representation introduced by the author in [26].
-In this formulation, the exact planar quark amplitude factorizes strictly into a universal kinematic Dirac phase-space trace $\mathcal{K}[P]$ and the dynamical gluonic momentum-loop functional $W[P+Q]$.
+In this formulation, the planar quark amplitude factorizes into a universal kinematic Dirac phase-space trace $\mathcal{K}[P]$ and the dynamical gluonic momentum-loop functional $W[P+Q]$.
 The amplitude for the propagation of a free Dirac particle in the phase-space loop (the Dirac loop) is:
@@
 Phase Space Normalization and the Departure from the Brownian Gauge.
-To appreciate the mathematical control within this phase-space formulation, it is instructive to examine its exact canonical normalization.
-The functional measure is strictly normalized by the standard symplectic volume element $\mathcal{D}P\mathcal{D}C^{\prime}/(2\pi)$ per continuous degree of freedom.
-Consequently, in the absence of the interacting Wilson loop (the free-particle limit, $W[C]=1$), integrating over the loop coordinates $C$ natively yields a functional Dirac delta function, strictly enforcing the conservation of local momentum:
+The phase-space formulation possesses a canonical normalization.
+The functional measure is normalized by the standard symplectic volume element $\mathcal{D}P\mathcal{D}C^{\prime}/(2\pi)$ per continuous degree of freedom.
+Consequently, in the absence of the interacting Wilson loop (the free-particle limit, $W[C]=1$), integrating over the loop coordinates $C$ natively yields a functional Dirac delta function, enforcing the conservation of local momentum:
 $\int d^{4}k\int\frac{\mathcal{D}C^{\prime}\mathcal{D}P}{2\pi}\exp\left(\mathrm{i}\oint(P_{\mu}(\tau)-k_{\mu})\dot{C}_{\mu}d\tau\right)F[P]$
 $=\int d^{4}k\int\mathcal{D}P\,\delta[P(\cdot)-k]F[P]=\int d^{4}k\,F[k]$ (10.4)
-This demonstrates that the phase-space integral is rigorously well-defined from standard quantum mechanics, unique strictly modulo the infinite 1D diffeomorphism volume $\text{Vol}(\text{Diff}(S^{1}))$ associated with the simultaneous reparametrization of $P(\tau)$ and $C(\tau)$.Conventionally, this $\text{Diff}(S^{1})$ volume is factored out by imposing a proper-time gauge-fixing via an $\int(C^{\prime})^{2}d\tau$ kinetic term.
-This specific gauge choice inevitably generates the Feynman-Schwinger representation with its standard Brownian (Wiener) measure in loop space—a measure inherently dominated by everywhere-discontinuous fractal paths that artificially generate the notorious UV cusp singularities.
+This demonstrates that the phase-space integral is well-defined from standard quantum mechanics, unique modulo the infinite 1D diffeomorphism volume $\text{Vol}(\text{Diff}(S^{1}))$ associated with the simultaneous reparametrization of $P(\tau)$ and $C(\tau)$.Conventionally, this $\text{Diff}(S^{1})$ volume is factored out by imposing a proper-time gauge-fixing via an $\int(C^{\prime})^{2}d\tau$ kinetic term.
+This specific gauge choice inevitably generates the Feynman-Schwinger representation with its standard Brownian (Wiener) measure in loop space—a measure inherently dominated by everywhere-discontinuous fractal paths that artificially generate the UV cusp singularities.
 In our framework, we recognize that the Brownian measure is just one possible choice of gauge fixing for the 1D diffeomorphisms.
@@
 Instead, as detailed in Section 11, we fix $\text{Diff}(S^{1})$ by imposing the Virasoro constraint $(f^{\prime})^{2}=0$ on the holomorphic twistors parametrizing the loop velocity.
-This specific gauge choice entirely avoids the fractal artifacts of the Brownian measure and perfectly locks the 1D phase-space kinematics of the quark loop to the 2D internal confining geometry of the minimal surface.
+This specific gauge choice entirely avoids the fractal artifacts of the Brownian measure and locks the 1D phase-space kinematics of the quark loop to the 2D internal confining geometry of the minimal surface.
 The momentum loop amplitude $W[P]$ was introduced in our papers [25, 26, 27] and was recently used in the solution of the Navier-Stokes turbulence problem [28, 33].
@@
 $=T^{\alpha\beta\gamma}{}_{\nu}\frac{\delta^{3}}{\delta\dot{C}_{\alpha}(\theta-0)\delta\dot{C}_{\beta}(\theta)\delta\dot{C}_{\gamma}(\theta+0)}$ (10.10)
-The notorious delta function on the right (which caused all the perturbative singularities in $W[C]$) disappeared in the Fourier integral (see original papers [29, 30, 31] for details).
+The delta function on the right (which caused all the perturbative singularities in $W[C]$) disappeared in the Fourier integral (see original papers [25, 26, 27] for details).
 Here is how this happens in our new framework, with the closed loop described by the position field $C_{\nu}(\theta)$ subject to the closure constraint $C(2\pi)=C(0)$.
@@
 The rigorous mathematical foundation for expanding loop functionals in such iterated integrals is provided by Chen’s theorem [1].
-Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a strictly complete functional basis.
+Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a complete functional basis.
 This guarantees that our formal expansion in terms of Magnus forms is exhaustive.
-Through direct tensor contraction, the kinematic tensor $T^{\alpha\beta\gamma}{}_{\nu}$ implements the exact Dynkin projector [9] onto the nested double commutator $\hat{D}_{\alpha},[\hat{D}_{\alpha},\hat{D}_{\nu}]]$.
+Through direct tensor contraction, the kinematic tensor $T^{\alpha\beta\gamma}{}_{\nu}$ implements the Dynkin projector [6] onto the nested double commutator $\hat{D}_{\alpha},[\hat{D}_{\alpha},\hat{D}_{\nu}]]$.
 Here, $\hat{D}=\partial+A$ represents the covariant derivative operator within the gauge field representation of the Wilson loop.
@@
 Consequently, by Ree’s Theorem [37] for Free Lie Algebras [38], this operator identically annihilates the symmetric shuffle ideal associated with the triple integral.
-As the point-split coordinates $\theta_{i}$ collapse to the origin, the periodic step-functions define a strictly cyclically ordered simplex.
-This unique parametric-invariant Lie projection ensures that the functional converges exactly to the third-order Magnus form $\Omega^{(3)}$.
+As the point-split coordinates $\theta_{i}$ collapse to the origin, the periodic step-functions define a cyclically ordered simplex.
+This unique parametric-invariant Lie projection ensures that the functional converges to the third-order Magnus form $\Omega^{(3)}$.
 Direct evaluation of this triple integral, mediated by the $T$ projector, confirms this algebraic correspondence and yields the canonical normalization coefficient $1/6$.
 10.5 Algebraic recurrence for the Momentum Loop Equation
-The Momentum Loop Equation (MLE) exhibits a profound simplification when the momentum functional is represented as a Magnus-ordered exponential:
+The Momentum Loop Equation (MLE) exhibits a significant simplification when the momentum functional is represented as a Magnus-ordered exponential:
 […]
@@
 Expanding this trace in Magnus forms $\Omega^{n}$, we obtain universal tensor coefficients $\mathcal{W}_{\alpha_{1}\dots\alpha_{n}}^{(n)}=\mathrm{tr}\{\hat{X}_{\alpha_{1}}\dots\hat{X}_{\alpha_{n}}\}$.
-The functional derivative acts algebraically on this structure, bringing down exact commutator insertions:
+The functional derivative acts algebraically on this structure, bringing down commutator insertions:
 $\frac{\delta}{\delta P_{\mu}(\theta)}\mathrm{tr}\left\{\hat{\mathcal{T}}\exp\left(\int d\theta^{\prime}\hat{X}_{\alpha}P_{\alpha}^{\prime}\right)\right\}=$
@@
 Because $\Delta P$ corresponds to the first-order Magnus form, this topological closure modifies the $n$-th order form by subtracting the product $\Omega^{(1)}\Omega^{(n-1)}$.
-By the integration-by-parts identity of the Magnus expansion (the shuffle relations), this product translates exactly to the sum of all interleaved operator positions.
+By the integration-by-parts identity of the Magnus expansion (the shuffle relations), this product translates to the sum of all interleaved operator positions.
 The loop-closing transformation thus yields the modified Magnus form:
@@
 Substituting this closed-loop basis into the MLE, we dynamically equate the left-hand side kinematic shuffle products against the right-hand side discrete commutator series.
-Starting from the perturbative vacuum $\mathcal{W}^{(0)}=1$, the algebraic system successfully yields finite, exact solutions for the lowest-order tensors.
-The fundamental string tension enters at $\mathcal{W}^{(2)}$, and the kinematic stress generated by the loop derivative exactly constrains the cyclic symmetric pairings of $\mathcal{W}^{(4)}$ and $\mathcal{W}^{(6)}$.
+Starting from the perturbative vacuum $\mathcal{W}^{(0)}=1$, the algebraic system yields finite solutions for the lowest-order tensors.
+The fundamental string tension enters at $\mathcal{W}^{(2)}$, and the kinematic stress generated by the loop derivative constrains the cyclic symmetric pairings of $\mathcal{W}^{(4)}$ and $\mathcal{W}^{(6)}$.
 (The explicit analytic formulas for these invariant tensors are provided in C for small $n$.
@@
 10.7 The Reasons for the Failure of the Magnus Expansion for $W[P]$
-The general reason for the breakdown of the expansion at this order is that the linear space of $\mathcal{W}^{(8)}$ invariant tensors—which is strictly constrained by the Shuffle Ideal and cyclic symmetry—possesses a significantly smaller matrix rank than the highly specific multilinear stress generated by the crossing Kronecker deltas of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$.
-This is not a heuristic observation, but a rigorously quantified algebraic contradiction.
-As detailed in C (and summarized in Table 1), extracting the exact linear system at the 8th order yields 379 geometric constraints but only 212 available parameters (which exhaustively includes the full $\mathcal{W}^{(8)}$ basis, all residual lower-order free parameters, and all gauge/ideal freedoms).
-The resulting system explicitly violates the Fredholm alternative condition, yielding a strictly positive rank deficiency ($\Delta\text{rank}=1$) and a non-zero projection of the inhomogeneous stress onto the left-null space of the kinematic matrix.
+The general reason for the breakdown of the expansion at this order is that the linear space of $\mathcal{W}^{(8)}$ invariant tensors—which is constrained by the Shuffle Ideal and cyclic symmetry—possesses a significantly smaller matrix rank than the highly specific multilinear stress generated by the crossing Kronecker deltas of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$.
+This is not a heuristic observation, but a quantified algebraic contradiction.
+As detailed in C (and summarized in Table 2), extracting the linear system at the 8th order yields 379 geometric constraints but only 212 available parameters (which exhaustively includes the full $\mathcal{W}^{(8)}$ basis, all residual lower-order free parameters, and all gauge/ideal freedoms).
+The resulting system explicitly violates the Fredholm alternative condition, yielding a positive rank deficiency ($\Delta\text{rank}=1$) and a non-zero projection of the inhomogeneous stress onto the left-null space of the kinematic matrix.
 Consequently, it is mathematically impossible for any choice of linear coefficients, or the remaining lower-order terms, to absorb and cancel this rigid cross-term stress.
@@
 It demonstrates that the MLE is finite, free of UV divergences, and expandable in a functional Taylor series up to $\mathcal{O}(P^{\prime 6})$.
-However, the geometric contradiction at the 8th order rigorously proves that the exact solution $W[P]$ is not an analytic functional of the bounding loop $P(\theta)$.
+However, the geometric contradiction at the 8th order proves that the solution $W[P]$ is not an analytic functional of the bounding loop $P(\theta)$.
 It cannot be indefinitely expanded in a continuous Taylor–Magnus series.
@@
 governing a scalar functional $W[P]$.
-The number of independent tensor structures available for a scalar functional grows strictly slower than those required for a free-index vector equation (growing approximately four times slower in $d=4$).
+The number of independent tensor structures available for a scalar functional grows slower than those required for a free-index vector equation (growing approximately four times slower in $d=4$).
 This is why, at some critical level of the Magnus expansion, there will inevitably be more vector equations than available scalar parameters.
 This is a foundational feature of the loop equation: it descends directly from the vector Yang–Mills equations of motion, which dictate that a local vector loop equation (evaluated at a fixed point on the contour) must be satisfied.
-Resolving this overdetermined system demands a deep mathematical “conspiracy” among the internal degrees of freedom within $W[P]$.
+Resolving this overdetermined system demands a mathematical “conspiracy” among the internal degrees of freedom within $W[P]$.
 The generic Taylor–Magnus expansion simply lacks the internal geometric structure to orchestrate this conspiracy.
-(Interestingly, this exact same vector-versus-scalar mismatch is fatal to all naive bosonic string theories of QCD: the loop equations are strictly vector equations, but the Nambu–Goto string Laplace operator is a scalar, meaning a scalar area-derivative constraint fundamentally cannot capture the full vector loop equation without an anomaly.)
+(Interestingly, the same vector-versus-scalar mismatch is fatal to all naive bosonic string theories of QCD: the loop equations are vector equations, but the Nambu–Goto string Laplace operator is a scalar, meaning a scalar area-derivative constraint fundamentally cannot capture the full vector loop equation without an anomaly.)
 However, our theory natively provides this required conspiracy!
-In the previous sections, we derived exact vector equations for the Elfin determinant on the Hodge-dual minimal surface, and these vector equations are satisfied across all vector components.
-This mathematical miracle is made possible by Hodge duality and the planar factorization of the Majorana determinant on this Hodge dual surface.
-The profound physical reasons for this non-analyticity will become fully apparent in later sections, when we construct the continuum solution and reduce it to the exact twistor string path integral.
-In that formulation, the formal Taylor expansion in $P(\theta)$ corresponds to the perturbative expansion of the exact exponential phase factor:
+In the previous sections, we derived vector equations for the Elfin determinant on the Hodge-dual minimal surface, and these vector equations are satisfied across all vector components.
+This property is made possible by Hodge duality and the planar factorization of the Majorana determinant on this Hodge dual surface.
+The physical reasons for this non-analyticity will become fully apparent in later sections, when we construct the continuum solution and reduce it to the twistor string path integral.
+In that formulation, the formal Taylor expansion in $P(\theta)$ corresponds to the perturbative expansion of the exponential phase factor:
 […]
-For such a functional Taylor series to converge, the boundary twistors $\lambda(z)$ and $\mu(z)$ would need to be strictly bounded.
-The algebraic contradiction at the 8th order rigorously proves this is not the case; the path integral is not convergent in a perturbative Taylor expansion.
+For such a functional Taylor series to converge, the boundary twistors $\lambda(z)$ and $\mu(z)$ would need to be bounded.
+The algebraic contradiction at the 8th order proves this is not the case; the path integral is not convergent in a perturbative Taylor expansion.
 As we will discuss in detail when deriving the twistor string path integral, the integration over the momentum loops is not dominated by smooth, small $P$ variations at all.
-When integrated alongside the Dirac target-space product $\mathrm{tr}\,\hat{\mathcal{T}}\exp\left(-\mathrm{i}\int d\theta\,\gamma_{\mu}P_{\mu}(\theta)+m_{0}\right)$, we encounter strictly ultra-local integrals.
+When integrated alongside the Dirac target-space product $\mathrm{tr}\,\hat{\mathcal{T}}\exp\left(-\mathrm{i}\int d\theta\,\gamma_{\mu}P_{\mu}(\theta)+m_{0}\right)$, we encounter ultra-local integrals.
 At every angle $\theta$, we integrate out the local value $P(\theta)$.
@@
 If $\hat{X}_{\mu}$ were assumed to be finite $N\times N$ matrices, the Cayley-Hamilton theorem and finite polynomial trace identities would artificially constrain the higher-order Magnus tensors $\mathcal{W}^{(n)}$, destroying the generality of the expansion.
-However, in the planar limit ($N_{c}\rightarrow\infty$), the Master Field operates in a strictly infinite-dimensional Hilbert space.
-As established by the Master Field Theorem proven by the author in 1994 [30, 31], any generic parametric-invariant loop functional expandable in iterated path integrals (guaranteed to be topologically complete by Chen’s theorem [1]) can be mapped exactly onto the non-commutative moment space of Free Probability and represented exactly as a Magnus-ordered exponential path integral over a non-commutative word space.
+However, in the planar limit ($N_{c}\rightarrow\infty$), the Master Field operates in a infinite-dimensional Hilbert space.
+As established by the Master Field Theorem proven by the author in 1994 [26, 27], any generic parametric-invariant loop functional expandable in iterated path integrals (guaranteed to be topologically complete by Chen’s theorem [1]) can be mapped onto the non-commutative moment space of Free Probability and represented as a Magnus-ordered exponential path integral over a non-commutative word space.
 The true non-perturbative Fock space of the string endpoint is the discrete space of “words” generated by $d=4$ Cuntz algebra operators:
@@
 Therefore, the breakdown at the eighth order is definitively not an artifact of a restrictive algebraic ansatz.
-Rather, this lack of a solution for $\mathcal{W}^{(8)}$ rigorously implies that the assumption of an analytic functional $W[P]$ is violated beyond the sixth order of the Magnus expansion.
+Rather, this lack of a solution for $\mathcal{W}^{(8)}$ implies that the assumption of an analytic functional $W[P]$ is violated beyond the sixth order of the Magnus expansion.
 The physical resolution is that the true solution must be a singular functional possessing a non-local dependence on the momentum loop $P$.
@@
 As established by Douglas, the area becomes a singular functional of the loop only after it is minimized over all possible boundary parametrizations.
-Before this minimization, the area is simply a quadratic Dirichlet-Hilbert functional 3.28 governed by the singular principal-value kernel $\mathcal{H}$.
+Before this minimization, the area is simply a quadratic Dirichlet-Hilbert functional 3.22 governed by the singular principal-value kernel $\mathcal{H}$.
 However, the strict minimization over parametrizations transforms it into a fundamentally singular functional because the optimal (isothermal) parametrization becomes implicitly and non-locally tied to the geometric shape of the loop.
@@
 Such shape-dependent, non-local singularities are fundamentally incompatible with a continuous, analytic Taylor-Magnus expansion in terms of iterated path integrals.
-Our exact continuum solution—the fermion determinant evaluated on the rigid Hodge-dual minimal surface—belongs precisely to this class of singular geometric functionals.
-The required reparametrization invariance is dynamically enforced by the null-twistor (Virasoro) constraint, which acts as the exact algebraic mechanism implementing the Douglas minimization over parametrizations.
-The algebraic catastrophe at $\mathcal{W}^{(8)}$ is thus the mathematical signature of the string geometry asserting its non-local rigidity and justifying the transition from a 1D algebraic Master Field to the 4D continuous geometry of the rigid twistor string.
+Our continuum solution—the fermion determinant evaluated on the rigid Hodge-dual minimal surface—belongs to this class of singular geometric functionals.
+The required reparametrization invariance is dynamically enforced by the null-twistor (Virasoro) constraint, which acts as the algebraic mechanism implementing the Douglas minimization over parametrizations.
+The algebraic inconsistency at $\mathcal{W}^{(8)}$ is thus the mathematical signature of the string geometry asserting its non-local rigidity and justifying the transition from a 1D algebraic Master Field to the 4D continuous geometry of the rigid twistor string.
 11 The twistor parametrization of the loop space measure
@@
 Therefore, the measure we derive here does not encounter the ambiguities of Teichmüller space.
-Our starting point is the rigorously defined, linear measure of the 1D quark-loop phase space: $\int\mathcal{D}C^{\prime}\mathcal{D}P$.
-The transition to twistor space is strictly a local (in $\theta$) change of variables, mapping the bounding curve velocity $C^{\prime}(\theta)\propto\text{Im}(z\lambda\sigma\mu)|_{z=e^{i\theta}}$ directly to the boundary twistors $\lambda,\mu$.
+Our starting point is the well defined, linear measure of the 1D quark-loop phase space: $\int\mathcal{D}C^{\prime}\mathcal{D}P$.
+The transition to twistor space is a local (in $\theta$) change of variables, mapping the bounding curve velocity $C^{\prime}(\theta)\propto\text{Im}(z\lambda\sigma\mu)|_{z=e^{i\theta}}$ directly to the boundary twistors $\lambda,\mu$.
 Because the twistors are localized at the boundary, this transformation involves no functional anomalies.
@@
 The linear integration measure $\mathcal{D}C$ (10.3) is redundant due to the reparametrization invariance of the Wilson loop $W[C]$.
-To define the path integral rigorously, we must fix this gauge freedom and factor the measure by the volume of diffeomorphisms of the unit circle.
+To define the path integral without ambiguity, we must fix this gauge freedom and factor the measure by the volume of diffeomorphisms of the unit circle.
 For the computation of the momentum loop $W[P]$, we choose the conformal gauge, where the loop parametrization is identified with the boundary value of a holomorphic null curve $f_{\mu}(z)$ in the bulk.
@@
 11.3 Euclidean Self-Intersection vs. Minkowski Helicity
-A common objection to minimal surfaces in the context of QCD strings arises from Euclidean intuition, where a surface of finite area bounded by a loop must satisfy strict embedding theorems (e.g., the Meeks-Yau theorem).
-In the Euclidean "rubber band" analogy, a surface cannot pass through itself without creating a physical self-intersection, which would typically be singular or unstable.
-Euclidean Space: The surface is constrained to a finite sheet. Any attempt to introduce a logarithmic branching point would force the sheet to intersect itself.
-Minkowski/Twistor Space: The surface behaves like a helicoid.
+In Euclidean signature, minimal surfaces of finite area bounded by a loop satisfy embedding theorems (e.g., the Meeks-Yau theorem) that preclude self-intersections.
+In Minkowski twistor space, however, the surface behaves like a helicoid.
 As the trajectory winds around the twistor pole $\lambda(z)\propto 1/(z-z_{0})$ or $\mu(z)\propto 1/(z-z_{0})$, it moves to a "different sheet" in the complexified coordinate space.
 This is why the surface can have an infinite area in the Minkowski sense—it is effectively a Riemann surface with an infinite number of sheets: the coordinate map function $f_{\mu}(z)\propto\log(z-z_{0})$.
-This multiple-valued coordinate is what allows the poles to exist without the "self-intersection" catastrophe.
+This multiple-valued coordinate is what allows the poles to exist without the "self-intersection" inconsistency.
 The surface doesn’t intersect itself for the same reason a spiral staircase doesn’t intersect the floor below it: it has moved in an orthogonal dimension.
@@
 […]
-Taking the limit of zero quark mass and focusing on the local factor dependent on $P$, we compute the integral over the local momentum variable $q_{\mu}(\theta)=P_{\mu}(\theta)d\theta$:
+Taking the limit of zero quark mass (see the remark 16.2 for the definition of our effective quark mass) and focusing on the local factor dependent on $P$, we compute the integral over the local momentum variable $q_{\mu}(\theta)=P_{\mu}(\theta)d\theta$:
 $I(\tau)=\int d^{4}q\,\frac{\exp\left(\mathrm{i}q_{\alpha}\tau_{\alpha}\right)}{\mathrm{i}\gamma\cdot q}$ (12.13)
@@
 […]
-Remarkably, the factor $e^{-3\rho}$ from the momentum integration exactly cancels the measure factor $e^{3\rho}$ derived from the twistor transformation, rendering the effective local volume element scale-invariant.
+Remarkably, the factor $e^{-3\rho}$ from the momentum integration precisely cancels the measure factor $e^{3\rho}$ derived from the twistor transformation, rendering the effective local volume element scale-invariant.
 Remark 12.1 (Scale Invariance of the Phase Space Measure).
-The exact cancelation of the Liouville factors, $e^{3\rho}\cdot e^{-3\rho}=1$, is a necessary consequence of the symplectic geometry of the loop phase space.
+The cancelation of the Liouville factors, $e^{3\rho}\cdot e^{-3\rho}=1$, is a necessary consequence of the symplectic geometry of the loop phase space.
 The canonical path integral measure $\mathcal{D}P\mathcal{D}C$ is invariant under canonical transformations, including the scaling dilatation $C_{\mu}\to sC_{\mu}$ and $P_{\mu}\to s^{-1}P_{\mu}$.
@@
 Instead, it represents a hidden one-dimensional theory of twistors on the boundary, disguised as a 2D Liouville theory—a simpler, albeit less familiar, mathematical object.
-In our framework, the path integral measure over loop space is rigorously constructed through a systematic procedure: we begin with a linear measure, impose the Virasoro constraint via twistor parametrization, and fix the residual local scaling symmetry.
+In our framework, the path integral measure over loop space is constructed through a systematic procedure: we begin with a linear measure, impose the Virasoro constraint via twistor parametrization, and fix the residual local scaling symmetry.
 The resulting measure is manifestly real and positive.
 Crucially, this measure does not contain an exponential of the KKS symplectic form—a hallmark of conventional string quantization and Conformal Field Theory (CFT).
-We emphasize that our approach does not involve quantizing a fundamental string in the Nambu-Goto sense, but rather constitutes a faithful implementation of the solution to the Makeenko–Migdal loop equations in momentum loop space.
+We emphasize that our approach does not involve quantizing a fundamental string in the Nambu-Goto sense, but rather constitutes a faithful implementation of the solution to the MM loop equations in momentum loop space.
 12.6 Block Matrix Structure and Diagonalization via Staggered Spinors
@@
 The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates.
-It is precisely this highly oscillatory nature of the path integral that mandates its exact analytical evaluation via Picard-Lefschetz theory and Catastrophe Theory.
+It is this oscillatory nature of the path integral that mandates its analytical evaluation via Picard-Lefschetz theory and Catastrophe Theory.
 13 Twistor Holography and the Rigid Twistor String
@@
 Because the path integral over the twistor target space is highly oscillatory, one might assume that extracting the physical spectrum requires massive stochastic lattice simulations.
-However, by analyzing the holomorphic structure of the effective action, we find that the discrete mass spectrum can be extracted exactly and analytically.
-Planar QCD ceases to be a theory of stochastic quantum fluctuations; it is a theory of pure classical complex geometry.
+However, by analyzing the holomorphic structure of the effective action, we find that the discrete mass spectrum can be extracted analytically.
+Our solution to Planar QCD is not a theory of quantum fluctuations in any space-time dimensions; it is a problem of classical geometry.
 14 Analytic continuation of Twistor String Path integral
@@
 $\lambda(e^{i\theta})=\sum_{i}\frac{R_{i}}{e^{\mathrm{i}\theta}-z_{i}}=-\sum_{i}\frac{R_{i}}{z_{i}}\sum_{n=0}^{\infty}z_{i}^{-n}e^{\mathrm{i}n\theta}$ (14.7)
-Comparing this to our spectral expansion, the Fourier coefficients $L_{n}$ are exactly determined as algebraic superpositions of the pole parameters:
+Comparing this to our spectral expansion, the Fourier coefficients $L_{n}$ are determined as algebraic superpositions of the pole parameters:
 $L_{n}=-\sum_{i}\frac{R_{i}}{z_{i}^{n+1}}$ (14.8)
-Consequently, transforming the functional measure from the $N$ discrete boundary values to the physical moduli $(R_{i},z_{i})$ involves exactly two finite, calculable matrix Jacobians:1.
+Consequently, transforming the functional measure from the $N$ discrete boundary values to the physical moduli $(R_{i},z_{i})$ involves two finite, calculable matrix Jacobians:1.
 The constant Jacobian of the Discrete Fourier Transform matrix (the Fourier matrix $e^{2\pi\mathrm{i}kn/N}$), which maps the boundary values to the Fourier coefficients $L_{n}$.2.
-The purely algebraic Jacobian mapping the Fourier coefficients to the pole parameters $(R_{i},z_{i})$.We do not explicitly compute these algebraic Jacobians here, because they contribute strictly to the pre-exponential fluctuation determinant of the path integral.
-For the extraction of the exact geometric mass spectrum, the relevant physics is governed entirely by the stationary phase (the saddle-point equations) and the topological monodromies of the effective action.
+The purely algebraic Jacobian mapping the Fourier coefficients to the pole parameters $(R_{i},z_{i})$.We do not explicitly compute these algebraic Jacobians here, because they contribute to the pre-exponential fluctuation determinant of the path integral.
+For the extraction of the geometric mass spectrum, the relevant physics is governed entirely by the stationary phase (the saddle-point equations) and the topological monodromies of the effective action.
 As we will demonstrate, the S-matrix poles emerge from the constructive interference of the geometric series over the topological winding number $k$.
@@
 Consequently, the transverse quantum fluctuations around the saddle point are dynamically damped by a factor of $1/k$.
-This perfectly crushes the pre-exponential factor, rendering the explicit form of the Jacobian irrelevant to the exact localization of the S-matrix poles.
-However, the explicit existence of this algebraic mapping mathematically guarantees that our twistor measure is strictly finite and constructible, firmly rooting the evaluation of the path integral in standard finite-dimensional complex analysis rather than abstract string-theoretic functional geometry.
+This exponential part dominates over the pre-exponential factor, rendering the explicit form of the Jacobian irrelevant to the exact localization of the S-matrix poles.
+However, the explicit existence of this algebraic mapping mathematically guaranties that our twistor measure is finite and constructible, firmly rooting the evaluation of the path integral in standard finite-dimensional complex analysis rather than abstract string-theoretic functional geometry.
 14.2.2 Liouville Field
-The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an exact formula
+The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an formula
 $\rho(z,\bar{z})=\frac{\log\tilde{\lambda}(\bar{z})\lambda(z)+\log\tilde{\mu}(\bar{z})\mu(z)}{2}$ (14.9)
@@
 $G_{p}(z_{1},z_{2})\equiv\int_{z_{1}}^{z_{2}}z^{p}\frac{dz}{iz}=\frac{z_{2}^{p}-z_{1}^{p}}{ip}.$ (14.10)
-With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is strictly positive, ensuring there are no singularities in the complex plane.
+With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is positive, ensuring there are no singularities in the complex plane.
 The contribution to the action of external momentum $q$ injected into the quark loop from the arc between $z_{1}$ and $z_{2}$ is:
@@
 14.4 Stochastic Quantization and Complex Langevin Dynamics
-The exact mathematical evaluation of this highly oscillatory complexified path integral is rigorously achieved via Stochastic Quantization.
+The mathematical evaluation of this highly oscillatory complexified path integral is achieved via Stochastic Quantization.
 The quantum statistical evolution of the twistor string is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{total}=S_{q}+S_{\gamma}+S_{T}+S_{\rho}+S_{\mathcal{O}}$.
-We emphasize that the CLE is utilized here not as a numerical simulation technique, but as an exact continuous mathematical reformulation of the quantum path integral.
+We emphasize that the CLE is utilized here not as a numerical simulation technique, but as an continuous mathematical reformulation of the quantum path integral.
   1. 1.
@@
     $\frac{d\xi_{k}}{d\tau}=-\frac{\partial S_{total}}{\partial\xi_{k}}-\mathrm{i}\gamma_{1,k}\tilde{\xi}_{k}+\nu_{\xi_{k}}$ (14.19)
-    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables representing the exact quantum fluctuations of the twistor fields.
+    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables representing the quantum fluctuations of the twistor fields.
   3. 3.
@@
   4. 4.
-    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the exact algebraic constraints:
+    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the algebraic constraints:
     $\frac{d\gamma_{1,k}}{d\tau}=-\mathrm{i}(\tilde{\xi}_{k}\xi_{k}-1)+\nu_{\gamma_{1,k}}$ (14.21)
-In this mathematical framework, the deterministic drift terms ($\nabla S$) represent the exact holomorphic gradient flow in complexified phase space.
+In this mathematical framework, the deterministic drift terms ($\nabla S$) represent the holomorphic gradient flow in complexified phase space.
 In Picard-Lefschetz theory, the union of all these deterministic flow lines originating from a complex saddle point constructs a middle-dimensional real manifold known as a Lefschetz thimble.
 If the noise $\nu$ were zero, the system would trivially collapse onto a 1D classical trajectory (the WKB approximation).
-However, the multidimensional stochastic noise $\nu$ dynamically forces the probability distribution to completely explore the transverse width of these $N$-dimensional thimbles, strictly incorporating all exact quantum fluctuations (the “vicinity” of the saddle points).
+However, the multidimensional stochastic noise $\nu$ dynamically forces the probability distribution to completely explore the transverse width of these $N$-dimensional thimbles, incorporating all quantum fluctuations (the “vicinity” of the saddle points).
 The phenomenon of exact S-matrix quantization we are looking for corresponds to a topological bifurcation of this stochastic system.
-At certain critical resonance energies, the complex Hessian develops a null space and the classical restoring drift strictly vanishes along a continuous topological valley.
+At certain critical resonance energies, the complex Hessian develops a null space and the classical restoring drift vanishes along a continuous topological valley.
 Unimpeded by any restoring gradient force, the stochastic noise $\nu$ drives the complex trajectory to infinity along this flat zero mode.
@@
 15.1 Twisted Boundary Conditions, Symmetrization, and Spin Projections
-Before solving the exact spectral equations, we justify the choice of the twisted geometry from the perspective of the semiclassical quark path integral.
+Before solving the spectral equations, we justify the choice of the twisted geometry from the perspective of the semiclassical quark path integral.
 In Planar QCD, the effective Minkowski action involves a sum over quark trajectories weighted by the area of the confining minimal surface.
 To extract the spectrum of states with macroscopic angular momentum $J$, we impose twisted boundary conditions.
-We require the quark positions at the ends of the string to undergo a physical spatial rotation by an angle $\phi$ over the period of the motion, with the angular momentum $J$ acting as the thermodynamic conjugate to this twist.
+We require the whole string to undergo a physical spatial rotation by an angle $\alpha$ over the period of the motion, with the angular momentum $J$ acting as the thermodynamic conjugate to this twist.
 The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime.
@@
 In our twistor string formulation, we project this twisted kinematics directly into the boundary twistor data.
-This is achieved by introducing the $SU(2)_{L}$ and $SU(2)_{R}$ projections, integrating over the target space rotation angle $\phi$ with the canonical spin weight:
-$\mathcal{P}_{J}=\oint d\phi\exp\left(\mathrm{i}\phi J\right)$ (15.1)
-When the macroscopic boundary of the string rotates by an angle $\phi$, the boundary twistors must geometrically twist.
-To enforce this, we introduce a fractional winding parameter $a$ corresponding to twisted boundary conditions for $\lambda(z)$.
-Furthermore, because macroscopic QCD rigorously conserves parity, the physical effective action must be symmetrized over the Self-Dual (SD) and Anti-Self-Dual (ASD) minimal surfaces.
-In the twistor parameterization, this corresponds to activating the twisted boundary condition for $\lambda(z)$ while keeping the mirror sector $\mu(z)$ frozen to a constant basis spinor, and adding a mirror term with $\mu(z)$ to the exponent.
-The Hodge-duality (or anti-duality) explicitly breaks the 3D embedding constraint, forcing the chiral surface to genuinely span 4-dimensional Minkowski space.
+This is achieved by introducing the $SU(2)_{L}$ and $SU(2)_{R}$ projections, integrating over the target space rotation angle $\alpha$ with the canonical spin weight:
+$\hat{P}_{J}A=\sum_{k}\oint d\alpha\exp\left(-\mathrm{i}k\alpha J\right)A_{k}(\alpha)$ (15.1)
+Here $A_{k}(\alpha)$ is the amplitude after $k$ winding around the quark loop, with the boundary condition that the rotation angle $\alpha$ is introduced at every full circle around the quark loop.
+When the macroscopic boundary of the string rotates by an angle $\alpha$, the boundary twistors must geometrically twist.
+To enforce this, we introduce a fractional winding parameter $a$ corresponding to twisted boundary conditions for $\lambda(z),\mu(z)$.
 To analytically generate the perfect resonance leading to the exact WKB saddle point, we must consider the limit of an infinitely large number of geometric windings $k\rightarrow\infty$.
-As the integration variable traverses the flat valley $k$ times, the accumulated action scales strictly linearly with $k$.
+As the integration variable traverses the flat valley $k$ times, the accumulated action scales linearly with $k$.
 Crucially, this means the transverse Hessian scales as $\mathcal{H}_{k}\propto k$.
-In the asymptotic resonance limit $k\rightarrow\infty$, the transverse restoring force becomes strictly infinite.
+In the asymptotic resonance limit $k\rightarrow\infty$, the transverse restoring force becomes infinite.
 The stochastic variance of the quantum fluctuations scales as $1/k\rightarrow 0$.
 The path integral dynamically localizes to the exact WKB saddle point, completely shielding the topological monodromy from quantum noise.
-These heuristic arguments will be made explicit and quantitative in this section, which will end with exact WKB quantization in a topological sector with only one pole of twistor field.
+These heuristic arguments will be made explicit and quantitative in this section, which will end with exact WKB quantization in a topological sector with only one pole of the twistor field.
 15.2 Geometric Quantization via Degenerate Lefschetz Thimbles
@@
 Mathematically, these occur at critical energies $E=E_{n}$ where the classical stability matrix (the Hessian $\mathcal{H}$) develops a null space.
-Because the integration variables are coupled complex holomorphic twistors, the vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of exactly two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
-This complex holomorphic geometry natively upgrades the standard square-root branch cut into a strictly simple pole, precisely generating the discrete S-matrix mass poles of the meson spectrum without ad hoc approximations.
+Because the integration variables are coupled complex holomorphic twistors, the vanishing of a single complex eigenvalue intrinsically mandates the simultaneous flattening of two real directions on the Lefschetz thimble (a strict corank-2 degeneracy).
+This complex holomorphic geometry natively upgrades the standard square-root branch cut into a simple pole, precisely generating the discrete S-matrix mass poles of the meson spectrum without ad hoc approximations.
 15.3 Catastrophes, flat valleys, and topological barriers
-To dynamically generate an exact simple mass pole $\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat zero-mode that extends infinitely far in configuration space.
-This exact infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields as the boundary parameter winds around the enclosed twistor poles.
-The integrity of this topological winding number is rigorously protected by an infinite topological action barrier.
+To dynamically generate a simple mass pole $\int d^{2}w\exp\left(-\bar{w}Q(E)w\right)\propto 1/Q(E)\propto 1/(E_{n}-E)$, the steepest descent trajectory must correspond to a non-compact flat complex zero-mode $w$ that extends infinitely far in configuration space.
+This infinite flat valley is provided dynamically by the multi-valued monodromy of the holomorphic twistor fields as the boundary parameter winds around the enclosed twistor poles.
+The integrity of this topological winding number is protected by an infinite topological action barrier.
 If a pole $z_{0}$ attempts to cross the unit circle from the inside ($|z_{0}|\to 1$), its conjugate mirror located at $1/\bar{z}_{0}$ must simultaneously approach the boundary from the outside.
-This creates a severe pinch singularity, strictly trapping the integration contour and locking the path integral into a conserved topological sector.
-15.4 Poles inside the disk and the monodromies
-For a saddle point with a single pole located strictly inside the unit disk ($|z_{0}|<1$), the total effective action accumulates strictly additively over multiple windings $k$.
-The Liouville bulk action simply duplicates, and the logarithmic boundary primitive picks up a strict topological phase shift at every cycle.
-Together, these terms yield an infinite geometric series $\sum\exp\left(\mathrm{i}k\Delta S_{eff}\right)$.
-The macroscopic divergence of this series generates the simple meson pole precisely when the exact Bohr-Sommerfeld quantization condition $\Delta S_{eff}=2\pi n$ is satisfied.
-15.5 The Exact One-Pole Solution and Minkowski Transition
-For the fundamental meson trajectory, we consider in detail the specific case of a single pole.
-To implement the twisted boundary condition for $\lambda(z)$ while freezing $\mu(z)$ to enforce the symmetrization over SD and ASD surfaces, we utilize the exact chiral 1-pole configuration with fractional power $a$, shifted to a pole position $z_{0}$:
-$\lambda(z)=R(z-z_{0})^{-a}\begin{pmatrix}1\\ 0\end{pmatrix},\qquad\mu(z)=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 1\end{pmatrix}$ (15.2)
-As we trace the boundary loop $z\to z\exp\left(2\pi\mathrm{i}\right)$, the active twistor acquires the fractional monodromy $\lambda\to\exp\left(-2\pi\mathrm{i}a\right)\lambda$.
-Equating this to the physical spin-1/2 transformation $\lambda\to\exp\left(-\mathrm{i}\phi/2\right)\lambda$, we establish the exact geometric identification $\phi=4\pi a$.
-The angular momentum projection operator thereby injects the canonical phase $\exp\left(4\pi\mathrm{i}Ja\right)$.
-To evaluate the bulk action properly, we must transition to the physical Minkowski metric on the symmetrized chiral surface, as established in D.
-We switch to the Minkowski metric on the surface parameterized by $z=\exp\left(\mathrm{i}\sigma_{M}+\tau_{E}\right)$, describing the Nambu-Goto metric $g_{ab}$ via the coordinate transformation $\exp\left(2\rho_{E}\right)d^{2}z=R^{2}\frac{d^{2}z}{(z\bar{z})^{a}}=R^{2}\mathrm{i}d\sigma_{M}d\tau_{E}$.
-Upon double Wick rotation and symmetrization (taking the parity-even arithmetic mean of the Left and Right chiral configurations), the conformal factor of the twisted metric dynamically restricts the causal domain.
-Using the modified radial coordinate $\tilde{\sigma}_{M}=(1-a)\sigma_{M}$, the physical speed-of-light boundary occurs exactly where this conformal metric vanishes, at $\tilde{\sigma}_{M}=\pi/2$.
-The proper Minkowski Area integral factorizes in the light-cone coordinates $z,\bar{z}$:
-$\text{Area}(a)$ $=\frac{\mathrm{i}R^{2}}{2}\int_{z\overline{z}<1}dz\wedge d\overline{z}(z\overline{z})^{-a}$
-$=\frac{-\mathrm{i}R^{2}}{2}\oint\frac{dz}{z}\int_{0}^{1}d(z\overline{z})(z\overline{z})^{-a}$
-$=\frac{-\mathrm{i}R^{2}}{2(1-a)}\oint dzz^{-1}=\frac{\pi R^{2}}{1-a}$ (15.3)
-Multiplying by the physical string tension $\sigma$, the macroscopic Area action evaluates exactly to:
-$S_{M,Area}=\frac{\pi\sigma R^{2}}{1-a}$ (15.4)
-Simultaneously, the Liouville action evaluates the kinetic gradient term.
-Crucially, while the target space metric stretches with the twist, the topological Liouville anomaly resides strictly on the rigidly bounded fundamental domain of the reference worldsheet, $\sigma_{M}\in[0,\pi/2]$:
-$\int|\nabla\rho|^{2}d^{2}z$ $=a^{2}\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}d\sigma_{M}$
-$=a^{2}(2\pi)\left(\frac{\pi}{2}\right)=\pi^{2}a^{2}$ (15.5)
-Multiplying by the universal anomaly coefficient $\frac{1}{12\pi}$, the Minkowski kinetic action yields the strictly positive, rigid Lüscher shift:
-$S_{M,kin}=+\frac{\pi a^{2}}{12}$ (15.6)
+This creates a pinch singularity, trapping the integration contour and locking the path integral into a conserved topological sector.
+We demonstrate that in D on the example of a single pole in twistor field.
+16 Minkowski Minimal Surface and Meson Spectrum
+To implement this flat valley scenario and compute the spectrum of massive mesons, we construct the minimal surface directly in four-dimensional Minkowski space $\mathbb{R}^{1,3}$.
+This is more subtle than just a straightforward analytic continuation of the twistor parametrization of a Euclidean minimal surface.
+16.1 Additive Factorization in Minkowski Space
+To construct a surface in Minkowski space, we utilize the generalized Weierstrass representation of Konopelchenko and Landolfi (K&L Theorem 5.1) [16].
+The target-space coordinate differential, represented as the Hermitian matrix $d\mathbf{X}=dX^{0}I+dX^{1}\sigma_{1}+dX^{2}\sigma_{2}+dX^{3}\sigma_{3}$, is parameterized by an additive sum of two independent, 2-component complex spinors:
+$d\mathbf{X}=\phi(\xi)\phi^{\dagger}(\xi)d\xi+\psi(\eta)\psi^{\dagger}(\eta)d\eta.$ (16.1)
+Here, the worldsheet light-cone coordinates are defined as $\xi=\tau+\theta$ and $\eta=\tau-\theta$, where $\tau$ is the periodic time on the fundamental torus and $\theta$ is the spatial radial parameter.
+Because $d\mathbf{X}$ is constructed as the sum of rank-1 Hermitian matrices, the target-space coordinates $X^{\mu}$ are real.
+We use the $\psi,\phi$ notations instead of Euclidean $\lambda,\mu$.
+The dependence of the left/right spinors $\psi/\phi$ on a single coordinate $\tau\pm\theta$ is the Minkowski analog of Euclidean holomorphicity.
+The classical minimal surface equations in Minkowski reduce to d’Alembert equations $\partial_{+}\partial_{-}X=0$ combined with Virasoro constraints $\det d\mathbf{X}=0$, equivalent to the null vector condition $dX_{\mu}^{2}=0$ in Minkowski space.
+Both of these conditions are satisfied in the above twistor representation.
+16.2 Complex Eigenvectors and Target-Space Monodromy
+To incorporate the rotation of the meson, we require the coordinate matrix $d\mathbf{X}$ to be unitary transformed by a spatial rotation matrix when the time coordinate winds around its fundamental cycle on the torus ($\tau\to\tau+2\pi$).
+Because the left-moving ($\xi$) and right-moving ($\eta$) dependencies are decoupled, we achieve this target-space monodromy by defining the component spinors $\phi$ and $\psi$ as complex eigenvectors of the $SU(2)$ rotation group.
+To center the string at the origin, we introduce a constant $\pi/4$ phase shift and scale the spinors by the macroscopic geometric scale $R$:
+$\phi(\xi)=\sqrt{\frac{R}{2}}\begin{pmatrix}\exp\left(i(a\xi+\pi/4)\right)\\ \exp\left(-i(a\xi+\pi/4)\right)\end{pmatrix}$,
+$\psi(\eta)=\sqrt{\frac{R}{2}}\begin{pmatrix}\exp\left(i(a\eta-\pi/4)\right)\\ \exp\left(-i(a\eta-\pi/4)\right)\end{pmatrix}$. (16.2)
+Under arbitrary temporal shift $\tau\to\tau+\Delta\tau$, the arguments shift as $\xi\to\xi+\Delta\tau$ and $\eta\to\eta+\Delta\tau$.
+The column vectors are left-multiplied by the diagonal rotation matrix:
+$\phi(\tau+\Delta\tau)=M(2a\Delta\tau)\phi(\tau),$
+$\psi(\eta+\Delta\tau)=M(2a\Delta\tau)\psi(\eta)$
+$\text{where }M(\alpha)=\begin{pmatrix}\exp\left(\mathrm{i}\alpha/2\right)&0\\ 0&\exp\left(-\mathrm{i}\alpha/2\right)\end{pmatrix}$. (16.3)
+Substituting this into the additive factorization, the target-space coordinate differential rotates as:
+$d\mathbf{X}(\tau+\Delta\tau)=M(2a\Delta\tau)\,d\mathbf{X}(\tau)\,M^{\dagger}(2a\Delta\tau).$ (16.4)
+Note that this applies to arbitrary temporal shifts $\Delta\tau$, not just the fundamental period $2\pi$.
+Later, we discuss the implications of this extended symmetry.
+This twisted boundary condition will allow us to project the quark loop amplitude to a fixed angular momentum $J$ by means of the Fourier integral (15.1) with $\alpha=4\pi a$.
+16.3 Real Coordinates and 2D Rotation
+Substituting the complex eigenvectors into the coordinate differentials, we evaluate the diagonal components:
+$dX^{0}+dX^{3}$ $=|\phi_{1}|^{2}d\xi+|\psi_{1}|^{2}d\eta=R\,d\tau.$ (16.5)
+$dX^{0}-dX^{3}$ $=|\phi_{2}|^{2}d\xi+|\psi_{2}|^{2}d\eta=R\,d\tau.$ (16.6)
+This yields $dX^{0}=R\,d\tau$ and $dX^{3}=0$. The center of mass is stationary. Evaluating the off-diagonal components yields the transverse spatial coordinates:
+$dX^{1}-idX^{2}=\phi_{1}\bar{\phi}_{2}d\xi+\psi_{1}\bar{\psi}_{2}d\eta$
+$=\frac{R}{2}\exp\left(i(2a\xi+\pi/2)\right)d\xi+\frac{R}{2}\exp\left(i(2a\eta-\pi/2)\right)d\eta$
+$=\frac{iR}{2}\exp\left(2ia\xi\right)d\xi-\frac{iR}{2}\exp\left(2ia\eta\right)d\eta.$ (16.7)
+Integrating this differential produces:
+$X^{1}-iX^{2}=\frac{R}{4a}(\exp\left(2ia\xi\right)-\exp\left(2ia\eta\right))$
+$=\frac{R}{4a}\exp\left(2ia\tau\right)(\exp\left(2ia\theta\right)-\exp\left(-2ia\theta\right))$
+$=\frac{iR}{2a}\exp\left(2ia\tau\right)\sin(2a\theta).$ (16.8)
+Separating the real and imaginary parts yields real spatial coordinates $X^{1}$ and $X^{2}$.
+The spatial parameter $\theta$ acts as a signed radial coordinate originating from the stationary center of mass at $\theta=0$.
+The symmetric interval $\theta\in[-\theta_{b},\theta_{b}]$ parameterizes the open string from the antiquark at $\theta_{b}$, straight through the center of mass, to the quark at $\theta_{b}$, precisely once without duplication.
+The string is a straight line spanning the transverse plane, sweeping a 2D circular rotation with a twist of $\alpha=4\pi a$.
+16.4 The Vanishing Metric and Spatial Boundaries
+The induced metric on the worldsheet is determined by the determinant of the target-space coordinate differentials.
+In this formulation, the conformal factor evaluates to the absolute square of the cross-term determinant:
+$\Omega^{2}=|\phi_{1}\psi_{2}-\phi_{2}\psi_{1}|^{2}.$ (16.9)
+Evaluating this for the given complex eigenvectors:
+$\phi_{1}\psi_{2}-\phi_{2}\psi_{1}$
+$=\frac{R}{2}(\exp\left(ia(\xi-\eta)+i\pi/2\right)-\exp\left(-ia(\xi-\eta)-i\pi/2\right))$
+$=\frac{R}{2}(i\exp\left(2ia\theta\right)+i\exp\left(-2ia\theta\right))=iR\cos(2a\theta).$ (16.10)
+Taking the absolute square $\Omega^{2}=R^{2}\cos^{2}(2a\theta)$, the induced metric on the worldsheet is:
+$ds^{2}=R^{2}\cos^{2}(2a\theta)(d\tau^{2}-d\theta^{2}).$ (16.11)
+In Minkowski space, the Lorentzian signature $(d\tau^{2}-d\theta^{2})$ is maintained.
+This metric vanishes at the roots $\theta=\pm\frac{\pi}{4a}$, which correspond to the string endpoints moving at the speed of light ($ds^{2}=0$).
+For massive endpoint quarks ($m_{q}>0$), the physical string is truncated at boundaries $\pm\theta_{b}$ inside this null radius ($0\leq\theta_{b}<\frac{\pi}{4a}$), ensuring a positive proper time.
+16.5 Classical Action Components over the Torus
+We evaluate the classical WKB action components integrated over the fundamental torus temporal period $\tau\in[0,2\pi]$ and the bounded spatial domain $\theta\in[-\theta_{b},\theta_{b}]$.
+We define the $a$-independent boundary phase parameter $\beta=2a\theta_{b}$.
+1. The Target Time Boundary Term ($E\Delta X^{0}$):
+The canonical energy $E$ conjugates to the total center-of-mass target time advance $\Delta X^{0}$. Integrating this over the temporal period gives:
+$E\Delta X^{0}=E\int_{0}^{2\pi}R\,d\tau=2\pi ER.$ (16.12)
+2. The Minimal Area Term ($S_{\text{Area}}$):
+The area of the minimal surface integrates the conformal factor over the dynamically bounded spatial domain:
+$S_{\text{Area}}$ $=\sigma R^{2}\int_{0}^{2\pi}d\tau\int_{-\theta_{b}}^{\theta_{b}}\cos^{2}(2a\theta)\,d\theta$
+$=2\pi\sigma R^{2}\int_{-\theta_{b}}^{\theta_{b}}\frac{1+\cos(4a\theta)}{2}\,d\theta$
+$=\frac{\pi\sigma R^{2}}{a}\left(\beta+\frac{1}{2}\sin(2\beta)\right).$ (16.13)
+3. The Proper Mass Integral ($S_{\text{mass}}$):
+The constituent quarks of bare mass $m_{q}$ propagate along the two boundaries $\pm\theta_{b}$.
+Evaluating the proper time $ds=R\cos(\beta)d\tau$ accumulated over the torus period for both endpoints yields:
+$S_{\text{mass}}=2m_{q}\int_{0}^{2\pi}R\cos(\beta)\,d\tau=4\pi m_{q}R\cos(\beta).$ (16.14)
+4. The Liouville term:
+To complete the evaluation of the geometric phase, we compute the contribution from the Liouville anomaly on the minimal surface.
+In the conformal gauge, the Liouville action is given by the worldsheet integral
+$S_{\rm Liouv}=\frac{1}{12\pi}\int d\tau d\theta\left((\partial_{\theta}\rho)^{2}-(\partial_{\tau}\rho)^{2}\right).$ (16.15)
+Using the conformal metric scale factor induced by our twistor parameterization in Minkowski space, $e^{2\rho}=R^{2}\cos^{2}(2a\theta)$, the Liouville field is purely spatial, $\rho(\theta)=\ln R+\ln\cos(2a\theta)$.
+Thus, the proper time derivative $\partial_{\tau}\rho$ vanishes, and the spatial gradient squared evaluates to $(\partial_{\theta}\rho)^{2}=4a^{2}\tan^{2}(2a\theta)$.
+Integrating this over one full period of the proper time ($\Delta\tau=2\pi$) and symmetrically across the spatial boundaries of the string ($\theta\in[-\theta_{b},\theta_{b}]$), and accounting for the overall minus sign of the measure, we obtain the anomaly defect:
+$\Delta S_{\rm Liouv}$ $=-\frac{1}{12\pi}(2\pi)\int_{-\theta_{b}}^{\theta_{b}}4a^{2}\tan^{2}(2a\theta)\,d\theta$
+$=-\frac{2a^{2}}{3}\int_{-\theta_{b}}^{\theta_{b}}\left(\sec^{2}(2a\theta)-1\right)d\theta.$ (16.16)
+Evaluating this integral and substituting the dynamical boundary velocity parameter $\beta=2a\theta_{b}$ yields:
+$\Delta S_{\rm Liouv}=-\frac{2a}{3}(\tan\beta-\beta).$ (16.17)
+16.6 Path-Ordered Dirac Trace and Spin Holonomy
+To relate the continuous twistor geometry to the discrete meson spin states, we compute the path-ordered product of the Dirac projectors along the spatial boundary of the quark loop.
+Following the staggered spinor diagonalization of Section 12.6, the fundamental link matrix evaluated in Minkowski space is $\mathcal{M}(\tau)=i\gamma_{\mu}v^{\mu}(\tau)$, where $v^{\mu}=\partial_{\tau}X^{\mu}$ is the target-space velocity.
+Substituting the twistor coordinate differentials yields the velocity components at the boundary $\theta_{b}$:
+$v^{0}=R,\qquad v^{3}=0,$
+$v^{1}=-R\sin(\beta)\cos(2a\tau),$
+$v^{2}=R\sin(\beta)\sin(2a\tau).$ (16.18)
+Defining the transverse velocity fraction $S=\sin(\beta)$, the invariant squared velocity is $v_{\mu}v^{\mu}=R^{2}(1-S^{2})=R^{2}\cos^{2}(\beta)$.
+In the chiral Weyl representation, the Dirac matrix projection takes the off-diagonal form $\mathcal{M}(\tau)=i\begin{pmatrix}0&\bar{V}(\tau)\\ V(\tau)&0\end{pmatrix}$, with the $2\times 2$ blocks defined as:
+$V(\tau)=R\begin{pmatrix}1&-S\exp\left(2ia\tau\right)\\ -S\exp\left(-2ia\tau\right)&1\end{pmatrix}$,
+$\bar{V}(\tau)=R\begin{pmatrix}1&S\exp\left(2ia\tau\right)\\ S\exp\left(-2ia\tau\right)&1\end{pmatrix}$. (16.19)
+Applying the staggered transformation $\Omega_{n}$ alternating between $I_{4}$ and $\gamma_{0}$, the product of two adjacent links at step $j$ assumes the block-diagonal form:
+$\tilde{\mathcal{M}}_{j+1}\tilde{\mathcal{M}}_{j}=-\begin{pmatrix}\bar{V}(\tau_{j+1})V(\tau_{j})&0\\ 0&V(\tau_{j+1})\bar{V}(\tau_{j})\end{pmatrix}$. (16.20)
+Factoring out the scalar invariant $v^{2}$, the continuous limit of the path-ordered exponential separates into two independent $2\times 2$ evolution operators, $W(\tau)$ and $\bar{W}(\tau)$.
+For the upper block, expanding $\bar{V}(\tau+d\tau)V(\tau)$ yields the evolution equation $W^{\prime}(\tau)=A(\tau)W(\tau)$, governed by the connection matrix:
+$A(\tau)=\frac{2iaS}{1-S^{2}}\begin{pmatrix}-S&\exp\left(2ia\tau\right)\\ -\exp\left(-2ia\tau\right)&S\end{pmatrix}$. (16.21)
+This system is solved by transforming to a co-rotating frame.
+Defining $W(\tau)=U(\tau)\tilde{W}(\tau)$ with $U(\tau)=\text{diag}(\exp\left(ia\tau\right),\exp\left(-ia\tau\right))$, the differential equation reduces to $\tilde{W}^{\prime}(\tau)=iaM\tilde{W}(\tau)$, with the constant coefficient matrix:
+$M=\begin{pmatrix}-\frac{1+S^{2}}{1-S^{2}}&\frac{2S}{1-S^{2}}\\ -\frac{2S}{1-S^{2}}&\frac{1+S^{2}}{1-S^{2}}\end{pmatrix}$. (16.22)
+The eigenvalues of $M$ are determined by its trace and determinant:
+$\text{tr}M=0,\qquad\det M=-\left(\frac{1+S^{2}}{1-S^{2}}\right)^{2}+\left(\frac{2S}{1-S^{2}}\right)^{2}=-1.$ (16.23)
+The eigenvalues of the holonomy generator $iaM$ are $\pm ia$, independent of the boundary parameter $S$.
+The discretized quark loop consists of $2L$ Dirac matrices, each corresponding to an angular step $\Delta\tau=2\pi/(2L)$.
+The staggered transformation groups these into $L$ adjacent pairs.
+The continuous evolution matrix $M$ emerges after the pairing of two adjacent matrices, so it is multiplied $L$ times around the closed loop.
+The total accumulated parameter $\tau$ evaluates to $L\Delta\tau=\pi$.
+Consequently, the accumulated phase is twice smaller.
+The solution over $k$ fundamental torus periods evaluates with the half-angle $\pi ak$:
+$\tilde{W}(\pi k)=\cos(\pi ak)I+i\sin(\pi ak)M.$ (16.24)
+Transforming back to the original frame with $U(\pi k)$ and computing the trace yields:
+$\text{tr}W(\pi k)$
+$=\exp\left(i\pi ak\right)\left(\cos(\pi ak)-i\sin(\pi ak)\frac{1+S^{2}}{1-S^{2}}\right)$
+$\qquad+\exp\left(-i\pi ak\right)\left(\cos(\pi ak)+i\sin(\pi ak)\frac{1+S^{2}}{1-S^{2}}\right)$
+$=2\cos^{2}(\pi ak)+2\sin^{2}(\pi ak)\frac{1+S^{2}}{1-S^{2}}.$ (16.25)
+Substituting $S=\sin(\beta)$, the scalar factor reduces to $(1+S^{2})/(1-S^{2})=1+2\tan^{2}(\beta)$.
+The lower block $\bar{W}(\pi k)$ yields an identical trace.
+The total trace of the $4\times 4$ spin holonomy evaluates to:
+$T_{4}(k)$ $=\text{tr}_{4}\left(\prod\limits_{j}i\gamma_{\mu}v^{\mu}(\tau_{j})\right)$
+$=4\left(1+2\sin^{2}(\pi ak)\tan^{2}(\beta)\right).$ (16.26)
+16.7 Meson Vertex Operators and Topological Sectors
+Taking the logarithm of the oscillating trace $T_{4}(k)$ to include it in the effective action would introduce complex saddle points into the path integral.
+Instead, treating the bare trace $T_{4}(k)$ as a pre-exponential factor, we expand it analytically into geometric phase factors using the identity $2\sin^{2}(\pi ak)=1-\frac{1}{2}(e^{2\pi iak}+e^{-2\pi iak})$:
+$\frac{1}{4}T_{4}(k)=1+\tan^{2}(\beta)-\cos(2\pi ak)\tan^{2}(\beta)$
+$=\sec^{2}(\beta)-\frac{1}{2}\tan^{2}(\beta)\left(e^{2\pi iak}+e^{-2\pi iak}\right).$ (16.27)
+This expansion resolves the monodromy analytically.
+The constant coefficients $\sec^{2}(\beta)$ and $\frac{1}{2}\tan^{2}(\beta)$ are real; they dictate the transition amplitudes (residues) of the states but do not shift the complex geometric phase.
+To isolate the physical meson states, we insert the corresponding interpolating vertex operator $\Gamma$ ($\gamma_{5}$ for the pseudoscalar pion, $\gamma_{\mu}$ for the vector $\rho$) into the trace:
+$T_{\Gamma}(k)=\text{tr}_{4}\left(\Gamma\prod_{j}i\gamma_{\nu}v^{\nu}(\tau_{j})\right).$ (16.28)
+The pseudoscalar insertion $\gamma_{5}$ isolates the constant coefficient $\sec^{2}(\beta)$, defining the transition amplitude for the unshifted branch ($q=0$).
+The vector insertion $\gamma_{\mu}$ isolates the oscillating phase factors $e^{\pm 2\pi iak}$.
+When combined with the macroscopic angular momentum phase of the action ($e^{-4\pi iaJk}$), the total phase shift in the winding sum becomes:
+$-4\pi aJk\pm 2\pi ak=-4\pi a\left(J\mp\frac{1}{2}\right)k.$ (16.29)
+This shifts the target angular momentum in the quantization condition to $u=J+q/2$, where $q\in\{0,-1\}$.
+This mathematically generates the half-integer spin-orbit splittings for the respective meson Regge trajectories.
+16.8 The saddle point equation
+Assembling the classical geometric action, we obtain the total monodromy of Minkowski action:
+$\Delta S_{\text{geom}}(R,a,\beta)=2\pi ER-\frac{\pi\sigma R^{2}}{a}\left(\beta+\frac{\sin(2\beta)}{2}\right)$
+$\qquad-4\pi m_{q}R\cos(\beta)-4\pi aJ-\frac{2a}{3}(\tan\beta-\beta).$ (16.30)
+Extremizing this action with respect to $\beta$ isolates the classical boundary condition balancing the string tension against the centrifugal force of the massive endpoints:
+$\frac{\partial\Delta S_{\text{geom}}}{\partial\beta}$
+$=-\frac{2\pi\sigma R^{2}}{a}\cos^{2}(\beta)+4\pi m_{q}R\sin(\beta)-\frac{2a}{3}\tan^{2}(\beta)=0$
+$\implies\sigma R^{2}\cos^{2}(\beta)-2m_{q}aR\sin(\beta)+\frac{a^{2}}{3\pi}\tan^{2}(\beta)=0.$ (16.31)
+The bound states of the spinning meson emerge as the meromorphic poles where these geometric series diverge: $\Delta S_{geom}\pm 2\pi aq=2\pi n$.
+16.9 Stationary metric and exact spin projection.
+The unique property of the exponential solution (16.4) for twistors is the stationary induced metric (16.11), which is strictly independent of the proper time $\tau$.
+We initially requested only the vanishing variation of the effective action over the temporal period $\Delta\tau=2\pi$, but the exact solution we have found is, in fact, strictly stationary.
+Not only does its monodromy vanish, but the metric itself is entirely independent of the proper time $\tau$.
+Moreover, the branch point degree $a$ (related to the macroscopic target-space rotation angle by $\alpha=4\pi a$) was left completely arbitrary by the saddle point equations for the remaining moduli.
+Previously, we considered this continuous degeneracy as a flat valley required to dynamically generate a resonance pole in the steepest descent integral, which would necessitate conjecturing that all higher functional derivatives vanish.
+Now, we can finalize that line of thought.
+To extract a physical meson state of definite angular momentum $J$, we introduced the Fourier projector (15.1) of our twisted amplitude $A(\alpha)$ onto the given $J$ state:
+$A_{J}=\int_{0}^{2\pi}d\alpha\,\sum_{k}e^{-\mathrm{i}k\alpha J}A_{k}(\alpha)$
+$=\sum_{k=1}^{\infty}\int_{0}^{2\pi}d\alpha\exp\left(\mathrm{i}k\alpha\Phi_{\star}(E,J)/(4\pi)\right)$
+$\propto\frac{4\pi\mathrm{i}}{\Phi_{\star}(E,J)}\sum_{k=1}^{\infty}\frac{1-\exp\left(\mathrm{i}k\Phi_{\star}(E,J)/2\right)}{k}$ (16.32)
+Here $\Phi_{\star}(E,J)$ is the value of $\Phi(E,J,\mathcal{K},\beta)$ at its extremum as described by above saddle point equations for moduli $\mathcal{K},\beta$.
+The last sum logarithmically diverges at large windings $k$, which justifies our WKB approximation used in estimating the Hessian and neglecting deviations from this classical trajectory.
+The coefficient in front of this logarithmically divergent factor is
+$A_{J}\propto\frac{4\pi\mathrm{i}\log k_{\text{max}}}{\Phi_{\star}(E,J)}$ (16.33)
+Remark 16.1.
+Hessian estimate and Quantum Fluctuations
+A more accurate estimate of the divergent sum over windings must account for the transverse quantum fluctuations of the trajectory around this flat classical valley.
+Because the classical orbit length scales linearly with the winding number $k$, the stability matrix (the Hessian $H$) is directly proportional to $k$.
+We presume that the fluctuating part of the twistor fields is strictly periodic, so that it will not contribute to the monodromy The fluctuating part of the action will, as usual , start with quadratic terms (the linear terms vanishing because we minimize over parameters of the classical solutions we have considered above).
+Therefore, the $k$-fold integral over the period $\Delta\tau=2\pi$ for this fluctuative part will be independent of $a$ and linear in $k$.
+Let us elaborate.
+The effective action for the fluctuative part will no longer be constant in $\tau$, but it will be periodic, which will result in linear $k$- dependence of the integral over multiple periods $0<\tau<2\pi k$.
+The scaling factor for the integration measure can be extracted via the $\zeta$-regularization of the functional determinant of this Hessian.
+Using the canonical value $\zeta(0)=-1/2$ for the trace over the 1D loop boundary modes, this yields the fluctuation weight:
 […]
-$\Delta X_{0}=\int_{0}^{2\pi}\sqrt{2}R\sin\left((1-a)\theta\right)d\theta=\frac{2\sqrt{2}R\sin^{2}(\pi a)}{1-a}$ (15.7)
-We balance the string tension against the boundary energy by minimizing the combined classical action $S(R)=S_{M,Area}-E\Delta X_{0}$ with respect to the geometric scale $R$.
-Since the Area is purely quadratic $(AR^{2})$ and the boundary coupling is strictly linear $(BR)$, the minimization of this quadratic function trivially yields $B^{2}/(4A)$ by the virial theorem.
-This dynamically eliminates the dimensionful size $R$ in one step, leaving a purely intensive, remarkably pristine geometric boundary potential:
-$S_{geom}(a)=-\frac{E^{2}}{\pi\sigma}F(a),\qquad\text{where}\qquad F(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (15.8)
-This establishes the exact, scale-independent geometric potential that governs the twistor branch point degree $a$ at the WKB saddle point.
-15.6 Quark Spin Monodromy, Trace Splitting, and the Unified Spectrum
-The final component of the path integral is the path-ordered Dirac trace $T$ along the quark boundary.
-The global boundary holonomy matrix acting on the twistor space is $U=\text{diag}(\exp\left(-2\pi\mathrm{i}a\right),1)$.
-1. The $\rho$ Trajectory (Vector Current)
-With an extra $\gamma_{\nu}$ projector inserted as the vector quark current, the twistor dyadic projector $P\propto\lambda\mu^{\dagger}$ acts as a strict rank-1 projector, $\left(\begin{smallmatrix}1&1\\ 0&0\end{smallmatrix}\right)$.
-It annihilates the trivial constant mode of the un-twisted right-chiral twistor, leaving exactly:
-$T_{\rho}\propto\text{tr}(UP)$ $=\text{tr}\begin{pmatrix}\exp\left(-2\pi\mathrm{i}a\right)&\exp\left(-2\pi\mathrm{i}a\right)\\ 0&0\end{pmatrix}$
-$=\exp\left(-2\pi\mathrm{i}a\right)$ (15.9)
-This corresponding pure topological phase factor $\exp\left(-2\pi\mathrm{i}a\right)$ adds directly to the rotation weight $\exp\left(4\pi\mathrm{i}Ja\right)$, shifting the angular momentum exactly by $\Delta J=+1/2$.
-2. The $\pi$ and $K$ Trajectories (Pseudoscalar Current)
-With a $\gamma_{5}$ inserted for the pseudoscalar and scalar trajectories, the trace evaluates the sum of the modes:
-$T_{\pi,K}=\text{tr}(U)=1+\exp\left(-2\pi\mathrm{i}a\right)$ (15.10)
-Because the path integral weight is linear in $T$, the amplitude rigorously splits into a sum of two independent path integrals:
-  • 1.
-    The Light Integral ($T_{1}=1$): Yields a phase $\exp\left(0\right)$. The spin trajectory is completely unshifted. This generates the ultra-light pseudo-Goldstone trajectory (the $\pi$ and $K$).
-  • 2.
-    The Heavy Integral ($T_{2}=\exp\left(-2\pi\mathrm{i}a\right)$): Yields the phase $\exp\left(-2\pi\mathrm{i}a\right)$.
-It acts identically to the $\rho$ projector, generating a heavy vector trajectory rigidly shifted by $\Delta J=1/2$ to $J=1$.
-15.7 Asymptotic Linearity and the Chew-Frautschi Plot
-In the macroscopic limit ($a\to 1/2$), the parametric geometric equations generate an exact simple pole, aggressively projecting the system into asymptotic Regge linearity.
-The theoretical intercepts emerge natively from the Liouville anomaly $\frac{\pi a^{2}}{12}$ without a single adjustable parameter:
-  • 1.
-    Pseudoscalar ($q=0$): $\alpha_{\pi}(0)=-1/96$
-  • 2.
-    Vector (q=1q=-1): $\alpha_{\rho}(0)=-1/96+1/2=47/96$
-Figure 15: Exact Twistor String Trajectories $J=\alpha(m^{2})$.
-The plot demonstrates the remarkable geometric linearity of the transcendental equations.
-The $\pi$ and $K$ mesons reside on the unshifted trajectory (with the Kaon explicitly shifted by the perturbative $SU(3)$ flavor breaking $\Delta m_{s}^{2}$), while the vector $\rho$ family is natively shifted upward by $\Delta J=1/2$ due to the rank-1 twistor projection.
-Unweighted least-squares fits of $\sigma$ map the universal curves through the experimental PDG data.
-The physical branch is covered by $a_{\mathrm{min}}<a<1/2$. Here $a_{\mathrm{min}}$ is a solution of the transcendental equation
-$4\pi\cot(\pi a)$ $=\frac{1-2a}{a(1-a)}+\frac{1}{a}$
-$\implies a_{\mathrm{min}}$ $\approx 0.426836\dots$ (15.11)
-Such a narrow range of the parameter $a$ explains the approximate linearity of the Regge trajectories.
-We computed the asymptotic expansion (G.11) of the Regge trajectory at large $s=m^{2}/\sigma$.
-The linear and constant terms dominate, and the higher-order $1/s$ corrections are numerically so small that at $J>1$ the deviations from linearity are invisible.
-But this is, of course, a matter of principle to have a correct formula based on an exact analytic solution instead of heuristic models.
-An interesting property of this analytic formula, discussed in G.2, is that these trajectories do not extend to negative spin or to negative $m^{2}$ (tachyons).
-Instead, they terminate at a geometric singularity characterized by the simultaneous vanishing of the derivatives of our parametric equations, $dJ/da=dm^{2}/da=0$.
-This singularity occurs at a certain critical value of $a=a_{c}<a_{\mathrm{min}}$ that is safely below the end of the physical region $a_{\mathrm{min}}<a<1/2$.
-Therefore, this singularity does not manifest in the physical region, where the trajectories appear virtually linear with the corresponding asymptotic intercepts 1/96-1/96 and $47/96$.
-Thus, this geometric singularity serves as a defender against tachyons.
-The technical details can be found in G.2.
-As shown in Figure 15, the analytic solution maps well onto the experimental data.
-What is truly remarkable is the numerical linearity of these trajectories given the highly transcendental nature of the geometric equations.
-The complex trigonometric terms act as deep-IR regulators, providing the correct chiral symmetry breaking behavior for the anomalously light pion.
-Yet, driven by the exact geometric pole at $a\to 1/2$, the macroscopic strings violently straighten out past $J\approx 1$, validating the strict Regge limit while rigorously housing the microscopic quantum anomalies.
-The unweighted least-squares fits yield effective string tensions $\sigma_{\rho}\approx 0.235$, $\sigma_{\pi}\approx 0.281$, and $\sigma_{K}\approx 0.320\text{ GeV}^{2}$.
-The value for the natural-parity $\rho$ trajectory recovers the canonical macroscopic Regge slope $\alpha^{\prime}=2/(\pi^{2}\sigma_{\rho})\approx 0.86\text{ GeV}^{-2}$, while the slightly higher effective tensions for the $\pi$ and $K$ trajectories physically reflect the dynamical stiffening due to $N_{c}=3$ chiral symmetry breaking and the endpoint inertia of the massive strange quark, respectively.
-15.8 Special case of rational branch points
-[…]
-These rational values of $a$ would not correspond, in general, to integer or half-integer values of the angular momentum $J+q/2$ in (G.9) .
-In the limit $a\to 1/2$, the angular momentum tends to $\infty$ in this equation.
-The solutions $a_{l}(J,q)$ of the equation $g(a)=J+q/2$ are, in general, transcendental, so that the physical spectrum is not affected by this geometric extension.
-However, this phenomenon may be interesting from a purely mathematical point of view, as a number-theoretic exception to the Exact WKB quantization.
-In the H we rigorously prove that this is not the case for the physical branch , covered by $a_{\textbf{min}}<a<1/2$, where $a_{\textbf{min}}\approx 0.426836$...
-was introduced above.
-16 Planar QCD as a Geometric theory
-The dynamical generation of S-matrix mass poles through the degeneracy of the twistor-string effective action is not an isolated mathematical curiosity; it is an infinite-dimensional realization of Exact WKB analysis, Picard-Lefschetz resurgence, and the geometry of Spectral Networks [12, 8].
-The complexified path integral over the boundary twistors exhibits localization phenomena that strictly mirror the topological mechanisms governing Seiberg-Witten (SW) theory and the analytic continuation of gauge theories [55].
-To establish this physical correspondence, we analyze the three fundamental geometric mechanisms driving the twistor string quantization:
-16.1 Topological Barriers and Infinite-Dimensional Picard-Lefschetz Theory
-By analytically continuing the twistor path integral into the complexified phase space, the convergence of the quantum amplitude is governed by middle-dimensional integration cycles, or Lefschetz thimbles [4].
-As formulated by Witten in the analytic continuation of Chern-Simons theory [55], the topology of these thimbles—and their intersection pairings—changes strictly when the system crosses a Stokes line, resulting in a discontinuous jump in the physical vacuum state.
-In our twistor string framework, non-perturbative Stokes phenomena are rigorously suppressed by the conformal boundary geometry.
-If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour.
-Exactly at the moment of crossing, these two singularities collide directly on the physical integration contour $|z|=1$.
-As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity.
-The integration loop becomes trapped between the colliding singularities at $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them.
-This generates a severe pinch singularity, creating an infinite action barrier.
-In the language of Picard-Lefschetz theory, this impenetrable boundary dynamically locks the Twistor String path integral into a strictly conserved topological sector defined by the enclosed twistor pole number $n$, forbidding unphysical Stokes jumps.
-16.2 Complex Langevin, Cauchy’s Shield, and Flat Valleys
-The rigorous exploration of these complex integration cycles is executed via the holomorphic gradient drift of the Complex Langevin Equation (CLE).
-In standard semiclassical formulations, evaluating the path integral at a degenerate saddle point requires formally isolating a zero mode and calculating the fluctuation determinant of the Hessian.
-In the CLE framework, this static abstraction is bypassed entirely: the path integral expectation value is simply the long-time average $\int d\tau$ over a single, noisy stochastic trajectory $\mathcal{V}(\tau)$.
-At the critical resonance energies $E=E_{n}$, the twistor effective action develops a non-compact flat valley.
-The single CLE trajectory does not rest at a classical minimum; driven by the stochastic noise, it constantly climbs the steep transverse walls of the complexified functional space and bounces back.
-Simultaneously, experiencing no deterministic restoring force along the zero mode, the trajectory drifts unimpeded along the flat valley.
-In this representation, the macroscopic pre-exponential factor (the transition residue) is simply proportional to the total Langevin time the trajectory spends accumulating windings, completely eliminating the need for formal Hessian factorization.
-A fundamental mathematical question arises: how can this violently stochastic trajectory $\mathcal{V}(\tau)$ generate a pristine, exact Bohr-Sommerfeld quantization phase without acquiring infinite quantum corrections from the nonlinear effects amplifying the noise in transverse directions?
-The resolution lies in the analytic continuation of the twistor fields.
-The CLE directly evolves the boundary data $\mathcal{V}_{bound}(\tau)=\mathcal{V}(z,\tau)\big|_{|z|=1}$.
-However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk.
-If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{i}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
-$\lambda(\theta)=\sum_{i}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (16.1)
-The stochastic CLE evolution simply forces these boundary parameters to evolve with Langevin time $\tau$:
-$\lambda(\theta,\tau)=\sum_{i}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (16.2)
-Upon analytic continuation back into the bulk, we obtain:
-$\lambda(z,\tau)=\sum_{i}\frac{R_{i}(\tau)}{z-z_{i}(\tau)}$ (16.3)
-The crucial observation is that the function remains strictly meromorphic at every instant of Langevin time.
-We obtain moving poles with evolving residues, but the fundamental meromorphic structure is absolutely preserved.
-Therefore, the infinite-dimensional stochastic evolution of the boundary fields rigorously and exactly reduces to the finite-dimensional, continuous $\tau$-dependence of the moduli $\{R_{i}(\tau),z_{i}(\tau)\}$.
-The infinite-dimensional phase accumulation thereby rigidly reduces to a standard one-dimensional contour integral over the physical boundary:
-$\Delta S=\oint_{|z|=1}\omega(z,\tau)\,dz$ (16.4)
-Here, the absolute geometric power of Cauchy’s integral theorem manifests itself.
-At any instant $\tau$, this contour integral is completely determined by the sum of the residues at the enclosed twistor poles.
-To evaluate the full physical amplitude, the path integral must be decomposed into a discrete sum over distinct topological sectors.
-These sectors are labeled by the winding number $k$ of the target-space minimal surface (or equivalently, the number of times the physical boundary parameter traverses the twistor poles).
-The full quantum amplitude is thus a superposition of path integrals, each mathematically explored by its own Complex Langevin evolution.
-Because the twistor effective action is strictly additive over these cycles, the total action for the $k$-th topological sector is precisely $S_{k}=S_{eff}^{(1)}+(k-1)\Delta S_{eff}$.
-In the CLE formulation for the $k$-th sector, the stochastic noise is opposed by the deterministic drift and the transverse Hessian, which now scale linearly with the topological winding number: $\mathcal{H}_{k}\to k\nabla^{2}\Delta S_{eff}$.
-Consequently, the residual stochastic variance in this sector is inversely proportional to the Hessian, scaling exactly as $\langle\delta\mathcal{V}^{2}\rangle\propto 1/k$.
-If the CLE noise were to continuously jitter the moduli $\{R_{i},z_{i}\}$ across the different sectors, the topological action increment $\Delta S_{eff}$ would randomly vary from one winding to the next.
-This would destroy the perfect phase coherence required for the geometric series to diverge, preventing the formation of a resonance.
-However, the macroscopic S-matrix pole is analytically generated exclusively by the asymptotic tail of this superposition, where $k\to\infty$.
-In this asymptotic limit, the transverse restoring force becomes strictly infinite, and the stochastic variance vanishes identically.
-The stochastic noise is completely crushed by the macroscopic weight of the multiple windings.
-The moduli $\{R_{i},z_{i}\}$ rigidly freeze at their exact classical saddle-point values for all large-$k$ sectors (validating the algebraic minimization of the single-pole radius $R$ in our spectral derivation).
-The topological phase is thus completely shielded from quantum contamination, ensuring that $\Delta S_{eff}$ remains strictly constant.
-The superposition over topological sectors safely diverges as a pristine geometric series $\sim(1-\exp\left(i\Delta S_{eff}\right))^{-1}$, analytically generating the exact S-matrix pole.
-17 Conclusion: The Solution of Planar QCD
-In this paper, we have presented the exact analytic solution of the planar MM loop equations, thereby solving large-$N_{c}$ QCD in the continuum limit.
-Building on Part I, where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ was shown to arise from the rigid Hodge-dual additive minimal surface, we have constructed the complete dynamics without summing over fluctuating worldsheet metrics.
-Our first main outcome is structural: by working natively in momentum loop space, we bypass the coordinate-space contact singularities and cusp pathologies that obscure the local continuum limit of the MM loop equations.
-We rigorously established that a purely one-dimensional analytic Taylor-Magnus solution of the exact vector momentum-loop equation cannot exist.
-The expansion is mathematically consistent through $\mathcal{O}(P^{\prime 6})$ but violently breaks down at $\mathcal{W}^{(8)}$ due to a Fredholm rank deficiency.
-This topological obstruction proves that a 1D closed-loop scalar framework lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD.
-This anomaly natively forces the introduction of 4D continuous boundary-twistor variables to parametrize the rigid minimal surface, yielding the confining twistor-string representation.
-Our second main outcome is dynamical: we demonstrated that the internal Majorana fermions (the “elves”) on the worldsheet provide the precise algebraic mechanism required for planar factorization at intersections.
-Crucially, the bulk Liouville term in our effective action originates microscopically from the zero-point fluctuations of these hidden fermions.
-The conformal anomaly of these heavy elves yields exactly the Lüscher term ($\pi/12$) in the effective string potential, cleanly generating the universally correct intercepts 1/96-1/96 and $47/96$ for the physical $\pi$ and $\rho$ Regge trajectories without any ad hoc assumptions of macroscopic string vibrations.
-One may view this construction as a geometric realization of the Master Field [54]—the existence of a purely classical description for the large-$N_{c}$ limit of QCD.
-While this Master Field does not manifest as a single classical gauge field configuration in coordinate spacetime, it finds an exact geometric representation in loop space.
-The classical geometry uncovered here identifies the underlying master object as the rigid Hodge-dual minimal surface.
-Consequently, the phase-space path integral over quark loops, once transformed into two-dimensional boundary twistor variables, exhibits exact geometric localization governed by Catastrophe Theory.
-The true Master Field of Planar QCD is thus a critical classical trajectory in twistor space—a geometric structure conceptually parallel to the spectral curves of Seiberg-Witten theory.
-By establishing the Master Field as this rigid twistor geometry, our approach naturally avoids the “graviton problem” typical of AdS/CFT gravity holography.
-Because our bulk geometry is not dynamical, there are no propagating bulk graviton modes that would lack a physical counterpart in confining QCD.
-All propagating degrees of freedom are strictly confined to the boundary, while bulk quantities (such as the Liouville field) are rigidly determined by the analytic continuation of boundary data.
-Therefore, our framework is best understood as gauge holography, where the bulk serves as a rigid encoding of the boundary dynamics rather than gravity holography or a topological string theory.
-Ultimately, Planar QCD ceases to be a theory of stochastic quantum fluctuations; it becomes a theory of pure classical complex geometry.
-The extraction of the exact mass spectrum reduces from a fluctuating path integral to a deterministic Generalized Eigenvalue Problem.
-By summing over the purely imaginary monodromy phases of the twistor poles protected by the pinch-singularity boundary barrier, we obtain an exact Bohr-Sommerfeld quantization condition.
-For the fundamental one-pole trajectory, this yields the unified parametric Regge trajectory (1.1) with trigonometric functions of the twistor branch point degree $a$, and $q=0$ for the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 for the shifted natural parity ($\rho$) trajectory.
-Although these functions are mathematically non-linear, they are numerically remarkably close to linear Regge trajectories.
-This analytic solution passes through the experimental data on Figure 15.
-The only outlier ($3\sigma$) is $K_{3}(2320)$, which is marked as “unconfirmed” in the PDG tables.
-The exact mass spectrum in our solution of Planar QCD is thus deterministically governed by the degenerate saddle points of the complexified twistor-string.
-17.1 Future Directions: Glueballs, Baryons, and Scattering
-The representation of Planar QCD as an Analytic Twistor String opens several immediate avenues for future research:
-  • 1.
-    Glueballs: The entire construction, from the loop equations and the fermion determinant on the minimal surface to the localization of the twistor path integral, naturally generalizes to higher $1/N_{c}$ corrections.
-Resolving the mathematical specifics of this expansion is the natural next step to analytically extract the glueball spectrum.
-  • 2.
-    Baryons: In the large-$N_{c}$ limit, baryons appear as heavy states composed of $N_{c}$ quarks.
-Geometrically, this corresponds to a minimal surface with a specific topology (e.g., a “Y-junction” or a surface with multiple boundaries).
-Extending the twistor ansatz to surfaces with higher connectivity will allow for the exact geometric quantization of baryon masses.
-  • 3.
-    Scattering Amplitudes: Meson-meson scattering involves a minimal surface bounded by four segments.
-The crossing symmetry and Regge behavior of these amplitudes should emerge naturally from the analytic continuation of the twistor data, potentially connecting this confining geometric framework directly to modern Amplituhedron methods.
+Crucially, the Fourier projection integral over the twist angle $\alpha$ inherently generates a kinematic factor of $1/k$:
+$\int_{0}^{2\pi}d\alpha\,e^{ik\alpha\Phi_{\star}}=\frac{e^{2\pi ik\Phi_{\star}}-1}{ik\Phi_{\star}}\propto\frac{1}{k\Phi_{\star}}$ (16.35)
+Combining this $1/k$ kinematic factor with the $k^{1/4}$ fluctuation weight, the terms in the macroscopic winding sum scale as $k^{1/4}/k=k^{-3/4}$.
+Summing this up to a macroscopic cutoff $k_{\text{max}}$ makes the divergence algebraically stronger than a logarithm, but preserves the fundamental geometric existence of the simple mass pole:
+$A_{J}\propto\frac{1}{\Phi_{\star}(E,J)}\sum_{k=1}^{k_{\text{max}}}k^{-\frac{3}{4}}\propto\frac{(k_{\text{max}})^{\frac{1}{4}}}{\Phi_{\star}(E,J)}$ (16.36)
+Strictly speaking, the integral remains finite at finite $k_{\text{max}}$ when $\Phi_{\star}(E,J)\to 0$.
+However, this happens in the infinitesimal vicinity, when $\Phi_{\star}(E,J)\sim 1/k_{\text{max}}$.
+Therefore, with proper renormalization of the string path integral, we can tend $k_{\text{max}}\to\infty$, after which we find the desired pole at $\Phi_{\star}(E,J)=0$ .
+16.10 Comparison with Experimental Meson Spectrum
+In this section, we compare the planar Regge trajectories derived from the twistor string effective action with the experimental meson mass spectrum from the Particle Data Group (PDG).
+The bound states are governed by the unified parametric equations derived in E.
+Parameterized by the spatial boundary phase $\beta\in[0,\pi/2)$ and the dimensionless renormalized boundary mass $x=m/\sqrt{\sigma}$, the energy $E=M$ and angular momentum $J$ of the states are (see Remark 16.2 for the definition of the renormalized boundary mass):
+$\frac{M(\beta)}{\sqrt{\sigma}}=\mathcal{K}(\beta,x)\left(\beta+\frac{\sin 2\beta}{2}\right)+2x\cos\beta,$
+$J(\beta)=\frac{\mathcal{K}(\beta,x)^{2}}{4}\left(\beta+\frac{\sin 2\beta}{2}\right)-\frac{1}{6\pi}(\tan\beta-\beta)-\frac{q}{2},$
+$\mathcal{K}(\beta,x)=\frac{\sin\beta}{\cos^{2}\beta}\left(x+\sqrt{x^{2}-\frac{1}{3\pi}}\right)\equiv\frac{\sin\beta}{\cos^{2}\beta}\tilde{x}$. (16.37)
+Here, the topological trace phase $q=0$ identifies the unshifted unnatural parity ($\pi,K,D$) trajectories, while q=1q=-1 shifts the natural parity ($\rho,K^{*},D^{*}$) trajectories by +1/2+1/2.
+To simplify the asymptotic expansions, we have introduced the dynamically dressed dimensionless mass parameter $\tilde{x}$, which enforces the chiral existence bound $x\geq 1/\sqrt{3\pi}$.
+Remark 16.2.
+The Renormalized Boundary Mass.
+The chiral existence bound $x\geq 1/\sqrt{3\pi}$ forbids the limit of a massless string boundary.
+The geometric drag generates a minimum inertial mass to balance the anomaly at the boundaries.
+The parameter $m$ appearing in the string effective action is not the bare QCD quark mass, nor is it the standard phenomenological constituent mass.
+It is a renormalized boundary mass.
+It represents the bare quark mass additively renormalized by the macroscopic perimeter term of the worldsheet fermion determinant, inducing the Planar QCD in our theory.
+Remark 16.3.
+Choosing the lightest particles (ground state).
+In general, there are several resonances in the PDG tables with the same quantum numbers $I^{G}(J^{C})$ in each meson sector (given strangeness $S$ and charm $C$).
+We always selected the lowest mass particle with a constant set of quantum numbers at given variable $J$ to place on the same Regge trajectory $J=\alpha(M^{2})$.
+Presumably, higher resonances in PDG tables can also be explained by twistor monodromies, but with more than one branching point.
+We leave this problem for a future extended investigation.
+Universal String Tension and the Global Fit.
+Experimentally, the asymptotic slopes of the meson trajectories differ by less than $10\%$.
+Within the planar limit ($N_{c}\to\infty$) of Geometric QCD, there exists only one universal gluon vacuum and a single, universal confining string tension $\sigma$.
+At leading order in $1/N_{c}$, there is no mechanism to produce different macroscopic slopes for different flavors.
+The observed $10\%$ splitting is an $\mathcal{O}(1/N_{c}^{2})$ non-planar effect driven by finite resonance widths, sea-quark loop insertions, and string-string interactions.
+This falls within the expected theoretical accuracy of the planar approximation.
+Consequently, we perform a global joint optimization across the thirteen meson families ($\pi,\rho,K,K^{*},D,D^{*},D_{s},D_{s}^{*},B,B^{*},B_{s},B_{s}^{*},B_{c}$).
+Instead of an arbitrary least-squares fit on masses, we minimize the fundamental action monodromy $\sum(\Delta S)^{2}$ across all thirty-six states.
+This global 5D optimization extracts exactly five fundamental parameters: the universal planar string tension $\sqrt{\sigma}$, and the four constituent boundary masses of the quarks ($m_{u},m_{s},m_{c},m_{b}$).
+The asymmetric heavy-light and heavy-heavy trajectories mathematically utilize the arithmetic mean of their respective constituent boundary masses.
+The global fit yields a universal string tension $\sqrt{\sigma}\approx 417$ MeV and constituent boundary masses $m_{u}\approx 136$ MeV, $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV.
+The fact that the $D_{s}$, $B_{s}$, and $B_{c}$ trajectories are completely locked by the relation $\bar{m}_{q_{1}q_{2}}=(m_{q_{1}}+m_{q_{2}})/2$ without introducing any new free parameters proves that the boundary masses physically correspond to additive constituent quarks in the string endpoints.
+The ability of a single geometric tension to seamlessly govern the spectrum from the $140$ MeV pion all the way up to the $6.27$ GeV bottom resonances validates the absolute geometric universality of the twistor string.
+The resulting non-linear trajectories are presented against the PDG data in Figure 15.
+Figure 15: The geometric Regge trajectories plotted against the PDG meson spectrum.
+A single universal planar string tension $\sqrt{\sigma}\approx 417$ MeV anchors all 13 families.
+The $\pi$ and $\rho$ trajectories generate an $m_{u}\approx 136$ MeV renormalized boundary mass.
+The heavier trajectories accommodate the strange, charm, and bottom quarks via the arithmetic means of their dynamically extracted constituent boundary masses: $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV.
+The topological spin shift q=1q=-1 elevates the natural parity branches.
+Experimental deviations in local slopes ($\sim 10\%$) are consistent with expected $\mathcal{O}(1/N_{c}^{2})$ non-planar corrections.
+The Ultra-Relativistic Asymptote (Large $J$).
+To analytically evaluate the high-energy limit, we expand the parametric equations as the boundaries approach the speed of light ($\beta\to\pi/2$).
+Defining the small parameter $\epsilon=\pi/2-\beta\to 0$, the trigonometric functions expand as $\cos\beta\approx\epsilon$ and $\sin\beta\approx 1-\epsilon^{2}/2$.
+The geometric equations expand to leading order as:
+$E(\epsilon)$ $\approx\frac{\pi\tilde{x}\sqrt{\sigma}}{2\epsilon^{2}}\implies\frac{1}{\epsilon}\approx\sqrt{\frac{2E}{\pi\tilde{x}\sqrt{\sigma}}},$
+$J(\epsilon)$ $\approx\frac{\pi\tilde{x}^{2}}{8\epsilon^{4}}-\frac{\pi\tilde{x}^{2}}{24\epsilon^{2}}-\frac{1}{\epsilon}\left(\frac{\tilde{x}^{2}}{6}+\frac{1}{6\pi}\right)+\frac{1}{12}-\frac{q}{2}.$ (16.38)
+Because the squared energy expands as $\frac{E^{2}}{2\pi\sigma}\approx\frac{\pi\tilde{x}^{2}}{8\epsilon^{4}}-\frac{\pi\tilde{x}^{2}}{24\epsilon^{2}}+\frac{x\tilde{x}-\tilde{x}^{2}/3}{\epsilon}$, substituting the inverted $1/\epsilon$ relation into the angular momentum equation cancels the divergent extensive terms, revealing the high-energy Regge trajectory:
+$J(E)\approx\frac{E^{2}}{2\pi\sigma}-\left(x\tilde{x}-\frac{\tilde{x}^{2}}{6}+\frac{1}{6\pi}\right)\sqrt{\frac{2E}{\pi\tilde{x}\sqrt{\sigma}}}+\frac{1}{12}-\frac{q}{2}.$ (16.39)
+This expansion modifies the linear Nambu-Goto expectation.
+The constant conformal intercept stabilizes to +1/12+1/12, recovering the Lüscher zero-point energy of the open bosonic string in $D=4$.
+However, the Liouville anomaly mandates a finite effective boundary mass $\tilde{x}$, which acts as a geometric drag, generating a negative $\mathcal{O}(\sqrt{E})$ curvature.
+At large $J$, relativistic effects act on the massive endpoints, and this geometric inertial lag curves the trajectories downward.
+This departs from the straight lines of Nambu-Goto models and reproduces the slope-softening observed in physical high-spin resonances.
+The Non-Relativistic Hook (Small $J$).
+While the high-energy limit is dominated by relativistic drag, the finite renormalized boundary mass $m=x\sqrt{\sigma}$ generates a geometric non-linearity confined to the low-$J$ region.
+Near $J\to 0$, the string is short, the endpoints move non-relativistically, and the boundary angle shrinks ($\beta\to 0$).
+Taylor-expanding the parametric equations around $\beta=0$ yields:
+$E(\beta)$ $\approx 2m+\beta^{2}\sqrt{\sigma}\left(2\tilde{x}-x\right),$
+$J(\beta)$ $\approx\beta^{3}\left(\frac{\tilde{x}^{2}}{2}-\frac{1}{18\pi}\right)-\frac{q}{2}.$ (16.40)
+Because the kinetic energy scales quadratically ($\Delta E\propto\beta^{2}$) while the angular momentum scales cubically ($J\propto\beta^{3}$), isolating $\beta$ from the energy equation and substituting it into $J$ produces the non-relativistic trajectory:
+$J(E)\approx\left(\frac{\tilde{x}^{2}}{2}-\frac{1}{18\pi}\right)\left(\frac{E-2m}{\sqrt{\sigma}(2\tilde{x}-x)}\right)^{3/2}-\frac{q}{2}.$ (16.41)
+Due to the fractional $3/2$ power, the macroscopic slope of the Regge trajectory at the threshold evaluates to $dJ/dE\propto\sqrt{E-2m}\to 0$.
+This demonstrates that the trajectories form a geometric hook, entering the physical parameter space with a flat derivative before accelerating into the linear regime.
+This non-relativistic hook is beautifully realized in the heavy charmed ($D,D^{*}$) and bottom ($B,B_{s},B_{c}$) meson trajectories, which sit deep in the non-relativistic regime for low spins before curving upward towards the universal linear asymptote.
+Phenomenological Synthesis.
+The global geometric optimization of the dimensionless action monodromy target ($\sum(\Delta S)^{2}=0$) fits the experimental PDG masses and spins across all thirteen meson families.
+By evaluating the parametric equations of the twistor string rather than linear approximations, several phenomenological features of QCD are recovered.
+First, the trajectories demonstrate the dynamical generation of a renormalized boundary mass gap.
+Unlike classical Nambu-Goto strings, which yield linear trajectories passing through the origin, the integration of the Liouville anomaly bounds the dimensionless mass to $x\geq 1/\sqrt{3\pi}$, forbidding the massless string limit.
+The theoretical trajectories terminate at finite rest masses.
+The geometric pressure of the anomaly forces the string endpoints to generate a renormalized boundary mass of $m_{u}\approx 136$ MeV for the light sector.
+For the heavier sectors, the 5D optimization dynamically extracts the constituent boundary masses $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV.
+The asymmetric heavy trajectories mathematically utilize the arithmetic mean of these respective scales, naturally isolating the physical flavor-symmetry breaking scales across the entire Standard Model meson spectrum.
+Second, the topological intercept distinguishes the vector from the pseudo-scalar mesons.
+Setting the topological vector index q=1q=-1 for the for the vector families ($\rho,K^{*},D^{*},D_{s}^{*},B^{*},B_{s}^{*}$) shifts their respective trajectories upward without arbitrary additive constants.
+The zero-point rotational quantum phase converts the generated constituent rest mass into the necessary kinetic energy, predicting the physical $J=1$ vector ground states.
+Third, the high-energy asymptotic regime reveals the effect of the conformal anomaly.
+The boundary anomaly term $\frac{1}{6\pi}(\tan\beta-\beta)$ behaves as a geometric drag.
+As the massive string endpoints approach the speed of light ($\beta\to\pi/2$), this inertial lag curves the trajectories downward, reproducing the slope-softening observed in physical resonances ($M^{2}>4\text{ GeV}^{2}$), departing from the straight lines of Nambu-Goto models.
+Finally, the systematic variations of the experimental data points around the theoretical curves demonstrate the model’s assumption of a universal underlying string tension $\sigma$.
+In physical QCD, explicit chiral symmetry breaking and hyperfine spin interactions cause the effective Regge slopes to vary slightly between different families.
+By enforcing a universal vacuum tension, the theoretical target $\sum(\Delta S)^{2}=0$ balances this physical variance.
+The resulting curves represent a universal tension of the strong vacuum that connects the distinct features of the hadronic spectrum.
+Numerical Mass Predictions and Fit Accuracy.
+State $J$ $M_{\rm th}$ (GeV) $M_{\rm exp}$ (GeV) State $J$ $M_{\rm th}$ (GeV) $M_{\rm exp}$ (GeV)
+Light Pseudoscalars ($\pi$, $q=0$) Light Vectors ($\rho$, q=1q=-1)
+$\pi$ 0 0.271 0.140 $\rho$ 1 0.878 0.775
+$b_{1}$ 1 1.168 1.229 $a_{2}$ 2 1.394 1.318
+$\pi_{2}$ 2 1.586 1.671 $\rho_{3}$ 3 1.756 1.689
+$a_{4}$ 4 2.052 1.967
+$\rho_{5}$ 5 2.308 2.330
+Strange Pseudoscalars ($K$, $q=0$) Strange Vectors ($K^{*}$, q=1q=-1)
+$K$ 0 0.355 0.494 $K^{*}$ 1 0.928 0.892
+$K_{1}$ 1 1.212 1.272 $K^{*}_{2}$ 2 1.435 1.430
+$K_{2}$ 2 1.624 1.580 $K^{*}_{3}$ 3 1.792 1.780
+$K^{*}_{4}$ 4 2.085 2.045
+$K^{*}_{5}$ 5 2.340 2.380
+Charm Pseudoscalars ($D$, $q=0$) Charm Vectors ($D^{*}$, q=1q=-1)
+$D$ 0 1.739 1.865 $D^{*}$ 1 2.118 2.008
+$D_{1}$ 1 2.331 2.422 $D^{*}_{2}$ 2 2.506 2.463
+$D_{2}$ 2 2.660 2.747 $D^{*}_{3}$ 3 2.799 2.763
+Charm-Strange ($D_{s}$, $q=0$) Charm-Strange Vectors ($D_{s}^{*}$, q=1q=-1)
+$D_{s}$ 0 1.823 1.968 $D_{s}^{*}$ 1 2.196 2.112
+$D_{s1}$ 1 2.407 2.460 $D_{s2}^{*}$ 2 2.580 2.569
+$D_{s3}^{*}$ 3 2.869 2.860
+Bottom Pseudoscalars ($B$, $q=0$) Bottom Vectors ($B^{*}$, q=1q=-1)
+$B$ 0 5.207 5.279 $B^{*}$ 1 5.475 5.325
+$B_{1}$ 1 5.630 5.726 $B_{2}^{*}$ 2 5.760 5.739
+Bottom-Strange ($B_{s}$, $q=0$) Bottom-Strange Vectors ($B_{s}^{*}$, q=1q=-1)
+$B_{s}$ 0 5.291 5.367 $B_{s}^{*}$ 1 5.557 5.415
+$B_{s1}$ 1 5.711 5.829 $B_{s2}^{*}$ 2 5.840 5.840
+Charm-Bottom ($B_{c}$, $q=0$)
+$B_{c}$ 0 6.675 6.275
+Table 1: The theoretical twistor string masses versus the experimental PDG meson masses.
+All theoretical values are generated from a single global $\sum(\Delta S)^{2}=0$ optimization over 36 states spanning 13 meson families, extracting exactly 5 fundamental physical parameters: a universal planar tension $\sqrt{\sigma}\approx 417$ MeV, and four dynamically generated constituent boundary masses $m_{u}\approx 136$ MeV, $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV.
+The asymmetric heavy trajectories use the arithmetic mean of the respective constituent boundary masses.
+Therefore, the $D_{s},B_{s},B_{c}$ families are parameter-free geometric predictions.
+The pion falls below the string’s mass threshold ($2m_{u}\approx 271$ MeV) due to explicit chiral symmetry breaking.
+To quantify the accuracy of the planar geometric fit, rather than relying on artificial least-squares statistics, we evaluate three physical root-mean-square (RMS) metrics across the experimental spectrum.
+The theoretical mass predictions for all thirty-six on-shell states are presented in Table 1.
+The most fundamental geometric metric for the fit is the residual of the action monodromy phase $\Delta\Phi=\Delta S/a$.
+In the Bohr-Sommerfeld quantization condition, sequential discrete states are separated by a phase gap of exactly $\Delta S_{\rm gap}/a=2\pi$.
+Evaluating the Root-Mean-Square (RMS) of the action residual across all 36 states yields ${\rm RMS}(\Delta\Phi/2\pi)\approx 0.36$.
+This demonstrates that the geometric phase error is merely $\sim 1/3$ of a single quantum level, ensuring that the physical states remain safely isolated within the $n=0$ topological winding sector without overlapping into radial excitations.
+Furthermore, because the non-linear Regge trajectories are defined geometrically as $J=\alpha(M^{2})$, a natural physical metric for their accuracy is the absolute spin deviation.
+Evaluating the theoretical angular momentum at the exact experimental masses yields an RMS deviation of $\Delta J_{\rm RMS}\equiv\sqrt{\langle(J_{\rm exp}-\alpha(M_{\rm exp}^{2}))^{2}\rangle}\approx 0.18$.
+Since physical Regge states are separated by exactly $\Delta J=1$, this confirms the twistor geometry correctly threads the mass spectrum to within less than one-fifth of an orbital quantum step.
+Finally, we evaluate the relative mass deviations.
+Excluding the $140$ MeV pion (which acts as an explicit pseudo-Goldstone boson and sits naturally below the geometric chiral threshold $2m_{u}\approx 271$ MeV), the RMS of the relative mass error $(M_{\rm exp}/M_{\rm theor}-1)$ across the remaining 35 massive states is $\approx 8.1\%$.
+This firmly aligns with the expected $\mathcal{O}(1/N_{c}^{2}\approx 10\%)$ theoretical variance of planar dynamics at $N_{c}=3$, verifying that the universal planar string tension accurately anchors the disparate bottom, charm, strange, and light flavors.
+17 Conclusion: A Solution of Planar QCD
+In this paper, we have presented an analytic solution of the planar MM loop equations in the continuum limit.
+Building on Part I, where the confining dressing factor $\exp\left(-\kappa S[C]\right)$ arises from the Hodge-dual additive minimal surface, we have constructed the dynamics without summing over fluctuating worldsheet metrics.
+Our first result is structural.
+In momentum loop space, the coordinate-space contact singularities and cusp pathologies are absent.
+The Taylor–Magnus expansion of the vector momentum-loop equation is consistent up to $O(P^{\prime 6})$ but develops a rank deficiency at $W(8)$.
+This algebraic obstruction shows that a one-dimensional scalar loop functional does not contain sufficient degrees of freedom to represent a solution of the vector loop equations.
+This necessitates the introduction of four-dimensional boundary twistor variables, leading to the twistor-string representation of the Hodge-dual minimal surface.
+Our second result is dynamical.
+The internal Majorana fermions (the “elves”) on the Hodge-dual minimal surface provide the mechanism for planar factorization at intersections, in agreement with the right-hand side of the MM equations.
+The bulk Liouville term in the effective action arises from the cancelation of their zero-point fluctuations.
+This cancelation is needed to remove anomalies in the MM equations.
+The associated conformal anomaly yields the Lüscher phase ($\pi/12$) in the effective string action, contributing to the low-$M^{2}$ curvature of the Regge trajectories.
+Within the planar limit ($N_{c}\to\infty$), the solution corresponds to a common confining string tension $\sigma$ for all trajectories.
+There are nonlinear effects, which are not present in the Nambu-Goto string theory, but these nonlinear effects are numerically small.
+The observed $\sim 10\%$ variation between the macroscopic slopes of the different flavor trajectories is attributed to $O(1/N_{c}^{2})$ corrections, including finite resonance widths, string interactions, and sea-quark effects, consistent with the expected accuracy of the planar approximation.
+The resulting parametric Regge trajectories describe thirteen meson families spanning the light, strange, charm, and bottom flavors ($\pi,\rho,K,K^{*},D,D^{*},D_{s},D_{s}^{*},B,B^{*},B_{s},B_{s}^{*},B_{c}$), reproducing the observed mass-spin relations across all 36 states (see Figure 15 and Table 1).
+We must add a remark regarding the physical boundaries of the current geometric framework.
+The spontaneous breaking of chiral symmetry and the vanishing of the pion mass in the strict chiral limit are not explained by our theory.
+As demonstrated, the geometric pressure of the Liouville anomaly bounds the dimensionless string parameter to $x\geq 1/\sqrt{3\pi}$, which generates a strictly positive minimum geometric mass threshold ($2m\approx 270$ MeV) for the $J=0$ ground state.
+The physical $140$ MeV pion falls below this threshold because its anomalously light mass is governed by the non-perturbative chiral condensate, a phenomenon distinct from the geometry of the confining flux tube.
+Fully capturing the pseudo-Goldstone nature of the pion would require a structural modification of the theory, making the Master Field explicitly accommodate and respect spontaneous chiral symmetry breaking.
+We view the synthesis of this geometric confinement mechanism with chiral symmetry breaking as a crucial direction for future theoretical development.
+The mathematical origin of our nonlinear Regge trajectories is radically different from that of conventional String Theory.
+The generation of S-matrix mass poles is associated with the degeneracies of the twistor-string effective action.
+After analytic continuation to Minkowski space, these poles correspond to energies at which the stability matrix develops a null space, producing non-compact directions in the associated Lefschetz thimbles.
+These non-compact directions arise from the multi-valued monodromy of the holomorphic twistor fields.
+At the saddle point, the geometric phase satisfies $\Delta S_{\text{geom}}=0$, and the summation over winding sectors produces a geometric series that diverges at the corresponding energies.
+The contribution of multiple windings suppresses transverse fluctuations, localizing the twistor moduli near the saddle point.
+The stability of this structure is maintained by a topological constraint: when a branch point approaches the integration contour, it encounters its conjugate image, leading to a pinch singularity that preserves the winding sector.
+In this solution of the loop equations, the Master Field is realized as a trajectory in twistor space, analogous to spectral curves in Seiberg–Witten theory.
+Because the bulk geometry is determined by the boundary data rather than fluctuating dynamically, this construction does not involve a propagating graviton sector and instead provides a geometric realization of gauge holography.
+In this framework, the quantum dynamics of planar QCD, encoded in the loop equations and reproduced by planar diagrams, are realized as classical dynamics of the dual geometric variables.
+The fermion determinant on the Hodge-dual minimal surface reproduces the quantum planar expansion, while the underlying geometry remains classical, reflecting the Master Field nature of the large-$N_{c}$ limit.
+In this way, the QCD mass spectrum was computed by solving a problem of classical geometry.
 Acknowledgements
-I wish to express my deepest gratitude to Nima Arkani-Hamed for his hospitality and for the invitation to present the complete derivation of this solution in a series of lectures at the IAS Particle Physics (Pizza) Seminar.
-The intellectual rigor and stimulating environment he fosters at the Institute were instrumental in the final crystallization of these results.
-I am especially indebted to Edward Witten for his profound insights and for the invaluable discussions that clarified the correspondence between this twistor-geometric solution and the phenomenological QCD string observed in lattice simulations.
-His generous engagement with the topological aspects of this work has been a source of significant inspiration.
+The author thanks N. Arkani-Hamed for the invitation to present this work at the IAS Particle Physics (Pizza) Seminar, and the participants of the seminar for detailed and stimulating discussions.
+The author is also grateful to E.
+Witten for discussions that clarified the relation between this twistor-geometric construction and the QCD string observed in lattice studies, as well as for comments on its topological aspects.
 Declaration of Generative AI in the Writing Process
-During the preparation of this work, the author utilized the generative AI Gemini 3 Flash to assist in the technical refinement of the manuscript. Specifically, the tool was employed to:
-  • 1.
-    Identify and correct typographical errors and grammatical inconsistencies.
-  • 2.
-    Standardize mathematical notation and ensure the rigorous formatting of LaTeX equations across the text and appendices.
-  • 3.
-    Improve the clarity and stylistic flow of the scientific narrative while preserving the original physical derivation.
-  • 4.
-    Cross-verify the consistency of the unified mass-squared formula with experimental data fits and topological sector identifications.
-The author has thoroughly reviewed and edited all AI-generated refinements and takes full responsibility for the physical content, mathematical accuracy, and original conclusions presented in this work.
+During the preparation of this manuscript, the author used the generative AI systems Gemini 3 Flash and ChatGPT 5.3 for technical assistance in editing and presentation.
+The tools were used for correcting typographical and grammatical errors, improving the clarity of exposition, and standardizing notation and LaTeX formatting.
+All scientific content, derivations, and conclusions were developed and verified by the author, who takes full responsibility for the manuscript.
 Data Availability.
@@
 Here $\Delta\text{rank}:=\text{rank}([A_{n}|b_{n}])-\text{rank}(A_{n})$.
-A strict positivity, $\Delta\text{rank}>0$, rigorously proves the system is mathematically inconsistent.
+A strict positivity, $\Delta\text{rank}>0$, proves the system is mathematically inconsistent.
 In this appendix, we provide the explicit polynomial solutions for the finite functional Taylor–Magnus expansion of the Momentum Loop Equation, valid up to the non-analytic boundary at $\mathcal{W}^{(8)}$.
@@
 $+c_{4,1}\left(\delta_{\mu_{1}\mu_{2}}\delta_{\mu_{3}\mu_{4}}+\delta_{\mu_{1}\mu_{4}}\delta_{\mu_{2}\mu_{3}}\right)$ (C.1)
-The tensor $\mathcal{W}^{(6)}_{\mu_{1}\dots\mu_{6}}$ is completely determined by the geometric stress cascading from the left-hand side loop derivative acting on the lower-order terms, successfully absorbing the kinematic constraints without contradiction.
+The tensor $\mathcal{W}^{(6)}_{\mu_{1}\dots\mu_{6}}$ is completely determined by the geometric stress cascading from the left-hand side loop derivative acting on the lower-order terms, absorbing the kinematic constraints without contradiction.
 Its explicit expansion in terms of the 15 independent 6-point Kronecker pairings is uniquely fixed by the lower-order parameters $c_{2,1}\equiv 1$, $c_{4,1}$, and $c_{4,2}$.
 At $\mathcal{W}^{(8)}$, this linear absorption fails.
-The expansion truncates due to the irreducible non-linear $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$ stress, which exceeds the tensor rank of the pure cyclic Shuffle Ideal basis, rigorously proving the non-analyticity of the functional $\mathcal{W}[P]$.
+The expansion truncates due to the irreducible non-linear $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$ stress, which exceeds the tensor rank of the pure cyclic Shuffle Ideal basis, proving the non-analyticity of the functional $\mathcal{W}[P]$.
 We compute these terms in [30], where we prove lack of solution for $\mathcal{W}^{(8)}$.
 C.2 Counting of unknowns/equations and Fredholm inconsistency at $\mathcal{W}^{(8)}$
-To rigorously demonstrate the breakdown of the Taylor-Magnus expansion at the 8th order ($\mathcal{W}^{(8)}$), we cast the Momentum Loop Equation at each order $n$ into a finite linear system.
-After imposing the closed-loop shuffle-ideal reduction (loop-closing prescription), cyclic symmetry, and parity constraints (all odd orders vanish), the exact functional matching reduces to:
+To demonstrate the breakdown of the Taylor-Magnus expansion at the 8th order ($\mathcal{W}^{(8)}$), we cast the Momentum Loop Equation at each order $n$ into a finite linear system.
+After imposing the closed-loop shuffle-ideal reduction (loop-closing prescription), cyclic symmetry, and parity constraints (all odd orders vanish), the functional matching reduces to:
 $A_{n}x_{n}=b_{n}$ (C.2)
@@
 We analyzed the exact linear algebra of these systems using a specialized symbolic script.
-The exact matrix ranks and subsystem counts are summarized in Table 1.
+The matrix ranks and subsystem counts are summarized in Table 2.
 The finite-dimensional system $A_{n}x_{n}=b_{n}$ is solvable if and only if it satisfies the Fredholm alternative consistency condition:
@@
 Projecting the inhomogeneous right-hand side $b_{8}$ (the geometric stress generated by the discrete non-linear cross-terms of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$) onto these null directions yields explicit, non-zero rational obstructions.
-For example, selecting the first canonical left-null basis vector $u_{1}$ yields the exact rational projection:
+For example, selecting the first canonical left-null basis vector $u_{1}$ yields the rational projection:
 $u_{1}^{T}b_{8}=\frac{1}{36}\neq 0$ (C.5)
-This establishes a decisive mathematical contradiction. Two critical robustness statements follow directly from this matrix representation:
+This establishes a mathematical contradiction. Two critical robustness statements follow directly from this matrix representation:
   • 1.
@@
     Setting the physical scale parameter to zero does not alleviate the structural rank deficiency.
-The full basis construction, shuffle-ideal reduction, and exact linear-algebra verification—including the generation of the matrices $A_{n}$ and the explicit null-vector projections—are provided in the executable Mathematica notebook [34] (see Data Availability).
-Appendix D Chiral Twistors and the Minkowski Helicoid Metric
-In this appendix, we explicitly derive the macroscopic worldsheet metric of the rotating string directly from the microscopic twistor fields.
-We demonstrate how the chiral twistor parameterization breaks the dimensional reduction to a 3-dimensional hyperplane, and how the physical parity-even projection of the double Wick rotation analytically continues the Euclidean twistor metric into the physical Minkowski Nambu-Goto metric.
-We then pedantically evaluate the resulting bulk action integral.
-D.1 Chiral Twistors and Dimensional Breakout
-The canonical Helicoid is generated by twistors endowed with fractional powers of the worldsheet coordinate $z$.
-This fractional scaling is a spin-structure gauge choice that explicitly enforces the Neveu-Schwarz anti-periodic boundary conditions for the worldsheet fermions as they complete a full rotation around the string center.
-To construct the chiral (Hodge-dual) surface, we set the right twistor to a constant basis spinor.
-We define the left and right twistors as:
-$\lambda(z)=e^{\mathrm{i}\pi/4}z^{-1/2}\begin{pmatrix}1/z\\ 1\end{pmatrix}$,
-$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ 0\end{pmatrix}$ (D.1)
-Because $\lambda\propto z^{-1/2}$ and $\mu\propto z^{1/2}$, their product has no overall fractional branching.
-The resulting bosonic physical observables are strictly rational, making this configuration geometrically equivalent to a rational single-pole state.
-The defining geometric characteristic of this configuration is its behavior under the symmetric Pauli matrices.
-The spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$.
-Evaluating the symmetric bilinear for the first spatial component:
-$f^{\prime}_{1}\propto\lambda^{T}\sigma_{1}\mu=\lambda_{1}\mu_{2}+\lambda_{2}\mu_{1}$
-$=\left(e^{\mathrm{i}\pi/4}z^{-3/2}\right)(0)+\left(e^{\mathrm{i}\pi/4}z^{-1/2}\right)\left(\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\right)=\mathrm{i}$ (D.2)
-Unlike the generic configuration where this component vanished identically everywhere on the worldsheet, here the tangent vector component $f^{\prime}_{1}$ evaluates to a non-zero constant.
-Consequently, the target space coordinate $X_{1}$ is no longer strictly constant.
-This permanently breaks the exact algebraic constraint that previously crushed the generic minimal surface into a 3-dimensional hyperplane, meaning this chiral surface genuinely spans a 4-dimensional embedding.
-D.2 The Euclidean Metric and Parity Projection
-To extract the physical string metric, we compute the induced conformal factor $\exp\left(2\rho\right)=|\lambda|^{2}|\mu|^{2}$. Restoring the dimensionful scale parameter $R$, the squared norms are:
-$|\lambda|^{2}=\frac{R}{|z|}\left(\frac{1}{|z|^{2}}+1\right)=R\frac{1+|z|^{2}}{|z|^{3}},$
-$|\mu|^{2}=R|z|(1+0)=R|z|$ (D.3)
-When multiplied, the fractional $|z|$ weights cancel perfectly:
-$\exp\left(2\rho\right)=R^{2}\frac{1+|z|^{2}}{|z|^{2}}=R^{2}\left(1+\frac{1}{|z|^{2}}\right)$ (D.4)
-To map this conformal geometry to the natural coordinates of the string worldsheet, we transition from the complex plane to the Euclidean cylinder via $z=\exp\left(\sigma_{E}+\mathrm{i}\tau_{E}\right)$.
-The flat line element transforms as $|dz|^{2}=|z|^{2}(d\sigma_{E}^{2}+d\tau_{E}^{2})=e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$.
-Substituting $|z|=\exp\left(\sigma_{E}\right)$, the conformal factor becomes:
-$\exp\left(2\rho\right)=R^{2}\left(1+e^{-2\sigma_{E}}\right)$ (D.5)
-Thus, the induced Euclidean worldsheet metric is exactly:
-$ds^{2}_{E}=\exp\left(2\rho\right)e^{2\sigma_{E}}(d\sigma_{E}^{2}+d\tau_{E}^{2})$
-$=R^{2}\left(1+e^{2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
-To evaluate the physical area, we analytically continue this metric into Minkowski spacetime via a double Wick rotation.
-The angular parameter $\tau_{E}$ serves as the target space time.
-We perform a Wick rotation on the spatial worldsheet coordinate into the physical light-cone regime:
-$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (D.7)
-Under this transformation, the exponential strictly rotates into a complex phase: $e^{2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})+\mathrm{i}\sin(2\sigma_{M})$.
-Because the macroscopic QCD string is strictly parity-conserving, the physical area is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations.
-The conjugate right-chiral surface identically provides the inverse complex phase $e^{-2\mathrm{i}\sigma_{M}}$.
-Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the area:
-$\text{Re}\left[\exp\left(2\rho\right)e^{2\mathrm{i}\sigma_{M}}\right]$ $=R^{2}\big(1+\cos(2\sigma_{M})\big)$
-$=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
-D.3 Action Evaluation and the Halved Lüscher Shift
-Finally, we pedantically evaluate the total effective action over the fundamental domain of this purely real Minkowski metric.
-The determinant of the conformal Minkowski metric $ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ yields the exact proper area measure:
-$\sqrt{-g}=2R^{2}\cos^{2}(\sigma_{M})$ (D.9)
-The center of the rotating string lies at $r=0$ ($\sigma_{M}=0$), and the physical speed-of-light boundary is at $r=1$ ($\sigma_{M}=\pi/2$).
-For one full spatial period of the meson, the time coordinate wraps exactly one cycle: $\tau_{E}\in[0,2\pi]$.
-The exact classical Minkowski area integral is therefore:
-Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}2R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
-$=(2\pi)(2R^{2})\left(\frac{\pi}{4}\right)=\pi^{2}R^{2}$ (D.10)
-Simultaneously, the Liouville action dictates that the worldsheet carries the kinetic anomaly $\frac{1}{12\pi}|\nabla\rho|^{2}$.
-Because the parity-symmetric conformal factor shares the exact same spatial singularity at the twistor pole as the generic string, it mathematically merges into the classical area term with the identical anomaly coefficient.
-As established in the main text, it acts strictly as a rigid, local shift to the effective string tension:
-$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (D.11)
-Multiplying the purely real classical Minkowski area by this effective tension, we obtain the complete one-cycle bulk action:
-$S^{(1)}_{bulk}$ $=-\sigma_{\text{eff}}\times\text{Area}$
-$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(\pi^{2}R^{2}\right)$
-$=-\pi^{2}\sigma R^{2}-\frac{\pi}{12}$ (D.12)
-This rigorous derivation explicitly demonstrates that for the parity-projected chiral string, the classical string area is exactly halved, which identically matches the universally mandated topological Lüscher phase $\pi/12$, though from a completely different physical and mathematical origins, which we discuss in the text.
-Appendix E The Pinch Singularity and the Topological Barrier
+The full basis construction, shuffle-ideal reduction, and linear-algebra verification—including the generation of the matrices $A_{n}$ and the explicit null-vector projections—are provided in the executable Mathematica notebook [30] (see Data Availability).
+Appendix D The Pinch Singularity and the Topological Barrier
 To mathematically illustrate the appearance of the conjugate pole at $1/\bar{z}_{0}$ and the resulting impenetrable topological barrier at the unit circle, consider the evaluation of the Liouville kinetic anomaly term in the effective action.
 In the case of a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
-$I(z_{0})=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
+$I(z_{0})=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (D.1)
 In order to continue this integral as an analytic function of $z_{0}$ we rewrite the integrand as a total derivative with respect to $\bar{z}$:
-$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}<1}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
-By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $z\bar{z}=1$:
-$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
-Crucially, on the unit circle, the complex conjugate is strictly given by the geometric inversion $\bar{z}=1/z$.
-Substituting this into the logarithm yields the exact boundary representation of the action:
-$I(z_{0})=\frac{-\mathrm{i}}{2}\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
-This exact formula explicitly reveals the analytic structure governing the boundary dynamics.
+$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}<1}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (D.2)
+By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces to a one-dimensional contour integral over the boundary unit circle $z\bar{z}=1$:
+$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (D.3)
+Crucially, on the unit circle, the complex conjugate is given by the geometric inversion $\bar{z}=1/z$. Substituting this into the logarithm yields the boundary representation of the action:
+$I(z_{0})=\frac{-\mathrm{i}}{2}\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (D.4)
+This formula explicitly reveals the analytic structure governing the boundary dynamics.
 The integrand possesses a simple pole at $z=z_{0}$ (originating from the twistor itself) and a logarithmic branch point at $z=1/\bar{z}_{0}$ (originating from the boundary inversion of the conjugate twistor).
-At $z_{0}=0$ this integral reduces to
-$I(0)=\frac{-\mathrm{i}}{2}\int_{0}^{2\pi}\mathrm{i}\,d\theta\left(-\mathrm{i}\theta\right)=\frac{-\mathrm{i}}{2}\frac{(2\pi)^{2}}{2}=-\mathrm{i}\pi^{2}$ (E.5)
-which matches with the computation of the Minkowski-continued area in the previous section.
-Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror strictly resides outside the disk ($|1/\bar{z}_{0}|>1$).
+Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror resides outside the disk ($|1/\bar{z}_{0}|>1$).
 However, if any continuous Langevin fluctuation attempts to push the twistor pole across the boundary ($|z_{0}|\to 1$), its conjugate mirror must simultaneously approach the boundary from the outside.
-At the exact moment of crossing, these two singularities violently collide directly on the physical integration contour $|z|=1$, trapping the integration path between them.
+At the moment of crossing, these two singularities collide directly on the physical integration contour $|z|=1$, trapping the integration path between them.
 This creates a severe, non-integrable pinch singularity, causing the action integral to macroscopically diverge.
-This geometric divergence acts as an infinite energetic barrier, rigorously protecting the twistor pole from escaping and locking the string into its quantized topological sector.
-Appendix F Algebraic Derivation of Spin Holonomy and Trace T
-To rigorously justify the topological selection rules governing the meson spectrum, we compute the path-ordered Dirac trace $T$ for the one-pole sector ($z_{0}=0$).
-The fundamental link matrix $\mathcal{M}(\theta)$ at the boundary $|z|=1$ is defined as:
-$\mathcal{M}(\theta)=i\mathcal{P}(\theta)=-2\left(e^{i\theta}\eta\xi^{\dagger}-e^{-i\theta}\xi\eta^{\dagger}\right)$ (F.1)
-where $\eta$ is the basis spinor of the frozen $SU(2)_{R}$ sector and $\xi(\theta)$ is the rotating spinor of the active $SU(2)_{L}$ sector.
-In the basis where $\eta=\begin{pmatrix}1\\ 0\end{pmatrix}$ and the twistor $\lambda(z)\propto 1/z$ implies $\xi(\theta)=\begin{pmatrix}0\\ e^{-i\theta}\end{pmatrix}$ at the boundary, the local matrix evaluates to:
-$\mathcal{M}(\theta)=\begin{pmatrix}0&-2e^{2i\theta}\\ 2e^{-2i\theta}&0\end{pmatrix}$ (F.2)
-This matrix represents a local $SU(2)$ rotation that flips the chiral state.
-The propagator $\mathcal{U}$ for one physical cycle is the path-ordered product of these matrices.
-For two adjacent steps $\theta_{1}$ and $\theta_{2}=\theta_{1}+\Delta\theta$, we find:
-$\mathcal{M}(\theta_{2})\mathcal{M}(\theta_{1})=\begin{pmatrix}-4e^{2i\Delta\theta}&0\\ 0&-4e^{-2i\Delta\theta}\end{pmatrix}$ (F.3)
-Assuming even number $2L$ of points on a unit circle in discretized loop, and multiplying these matrices around the entire loop $\oint d\theta=2\pi$ reduces the propagator for one cycle to:
-$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{2\pi i}&0\\ 0&e^{-2\pi i}\end{pmatrix}$ (F.4)
-The physical trace $T$
-[…]
-The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$, which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\pi n$ depending on the parity of $L$.
-Appendix G Geometric Potentials and Universal Trajectories
-In this appendix, we decouple the unknown dimensionful string tension $\sigma$ from the geometric quantization conditions to derive the parametric functions $h(a)$ and $g(a)$, their singularities, and their roots.
-We formulate the system directly in terms of the fractional winding parameter $a$.
-As derived in Section 15.5, the macroscopic spatial rotation angle of the string is exactly $\phi=4\pi a$.
-We isolate the string tension from the boundary potential $F(a)$ by defining the dimensionless geometric function:
-$\tilde{F}(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (G.1)
-The boundary potential is $F(a)=\frac{1}{\pi\sigma}\tilde{F}(a)$.
-Defining the physical mass-squared as $m^{2}=E^{2}=\sigma h(a)$, the unknown tension $\sigma$ factors to the numerator of the mass equation and cancels out of the angular momentum equation entirely.
-The total geometric phase for the $q$-shifted trace sector is:
-$\Phi_{q}(a)=-E^{2}F(a)+\frac{\pi a^{2}}{12}+4\pi aJ+2\pi qa$ (G.2)
-where $q=0$ for the unshifted $\pi/K$ integral, and q=1q=-1 for the shifted $\rho$ and Scalar integrals.
-We impose the saddle-point condition $\Phi_{q}^{\prime}(a)=0$ to isolate the angular momentum terms:
-$4\pi aJ+2\pi qa=-\frac{\pi a^{2}}{6}+aE^{2}F^{\prime}(a)$ (G.3)
-Substituting this into the Bohr-Sommerfeld quantization condition $\Phi_{q}(a)=2\pi n_{w}$ (where $n_{w}=0$ for the principal string sector), we obtain the algebraic relation:
-$-\frac{\pi a^{2}}{12}-E^{2}F(a)+aE^{2}F^{\prime}(a)=0$ (G.4)
-Using the differential identity $aF^{\prime}(a)-F(a)=\frac{1}{\pi\sigma}\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)$, we isolate $m^{2}=\sigma h(a)$ to obtain the mass function:
-$h(a)=\frac{a^{2}\pi^{2}}{12\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)}$ (G.5)
-Simplifying that expression, we get an explicit trigonometric answer:
-$h(a)=-\frac{\pi^{2}(a-1)^{2}a^{2}\csc^{4}(\pi a)}{24(-2a+4\pi(a-1)a\cot(\pi a)+1)}$ (G.6)
-To obtain the angular momentum function, we substitute $E^{2}=\sigma h(a)$ back into the saddle-point equation. Defining the target angular momentum as $u=J+q/2$, we obtain:
-$u=\frac{a\left(a\tilde{F}^{\prime}(a)-2\tilde{F}(a)\right)}{48\left(\tilde{F}(a)-a\tilde{F}^{\prime}(a)\right)}$ (G.7)
-This expression can be dramatically simplified.
-Dividing the numerator and denominator by $\tilde{F}(a)$ to isolate the logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)$, we obtain a partial fraction decomposition:
-$\frac{\tilde{F}^{\prime}(a)}{\tilde{F}(a)}=\frac{2a+48u}{a(a+48u)}=\frac{1}{a}+\frac{1}{a+48u}$ (G.8)
-Substituting the explicit logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)=4\pi\cot(\pi a)+\frac{1}{1-a}$ and combining the kinematic poles yields the exact trigonometric spectral equation:
-$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (G.9)
-The Regge trajectory can also be simplified to the following
-$J+\frac{q}{2}=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{a}{48}$ (G.10)
-G.1 Singularities and the Ground States
-The geometry of the functions is dominated by a kinematic pole at the speed-of-light boundary, $a\rightarrow 1/2$.
-At this limit, the left-hand side of (G.9) strictly vanishes ($\cot(\pi/2)=0$).
-For the equation to hold, the right-hand side must also vanish, which dynamically forces the angular momentum to diverge ($u\rightarrow\infty$) and forms the linear macroscopic Regge asymptote.
-Evaluating the derivative $\tilde{F}^{\prime}(1/2)=8$ in the mass formula (G.5) restores the macroscopic Regge slope $J\approx 2m^{2}/(\pi^{2}\sigma)$.
-The asymptotic expansion at larger $s=m^{2}/\sigma$ corresponds to an inverted expansion around $a\to 1/2-\epsilon$.
-The expansion of trajectory goes as follows:
-$J(s)$ $=\frac{2s}{\pi^{2}}+B_{0}+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+O(s^{-5});$ (G.11)
-$B_{0}$ $=-\left(\frac{q}{2}+\frac{1}{96}\right);$ (G.12)
-$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}$ (G.13)
-$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}$ (G.14)
-The coefficients $B_{1},B_{3}$ are numerically so small, that the trajectory is indistinguishable from linear for $J>1$.
-When we compute the roots for the ground states, the explicit spectral equation (G.9) yields the mass eigenvalues:
-  • 1.
-    The Light Pseudoscalar ($q=0,u=0$): Solving (G.9) yields $a_{\pi}\approx 0.4268$.
-Substituting into $h(a)$ gives the mass ratio $m_{\pi}^{2}/\sigma\approx 0.0375$.
-The geometry produces a light ground state without symmetry breaking parameters.
-  • 2.
-    The Vector Ground State (q=1,u=1/2q=-1,u=1/2): The topological shift forces the string to spin faster.
-Solving (G.9) occurs at $a_{s}\approx 0.4987$, yielding $m_{s}^{2}/\sigma\approx 1.9782$.
-The geometry predicts the vector state to be massive.
-The continuous generation of the parametric trajectories $u=J+q/2$, the numerical roots, and the data fits to $\sigma$ are performed using the companion Mathematica notebook [35].
-G.2 The Endpoint Singularity and the Fold Catastrophe
-While the large-$s$ asymptotic expansion (G.11) governs the strictly linear macroscopic Regge regime, the exact parametric geometry of the twistor string dictates a hard geometric boundary at small $s$.
-This boundary manifests not as a smooth continuation toward $s\to$-$\infty$, but as a fold catastrophe in the mapping between the spectral parameter $a$ and the physical observables.
-To expose this geometric termination, we rewrite the exact parametric equations for the mass ratio $s(a)=h(a)=m^{2}/\sigma$ and the target angular momentum $u(a)=J+q/2$.
-From (G.6) and (G.8), we define the auxiliary functions:
-$\mathcal{P}(a)$ $=\frac{\pi^{2}a(1-a)}{24\sin^{4}(\pi a)},$ (G.15)
-$\mathcal{R}(a)$ $=4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}.$ (G.16)
-The spectral equations then take the remarkably compact algebraic form:
-$s(a)$ $=\mathcal{P}(a)\mathcal{R}(a)^{-1},$ (G.17)
-$48u(a)$ $=\mathcal{R}(a)^{-1}-a.$ (G.18)
-A fundamental property of the Rigid Twistor String is the existence of a simultaneous mathematical endpoint $a_{c}$ where the first derivatives of both the mass-squared and the angular momentum vanish exactly.
-We can rigorously prove this simultaneous singularity by differentiating the target angular momentum:
-$48u^{\prime}(a)$ $=-\mathcal{R}^{\prime}(a)\mathcal{R}(a)^{-2}-1=0$
-$\implies-\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}.$ (G.19)
-Next, we differentiate the mass function and multiply by $\mathcal{R}(a)^{2}$:
-$s^{\prime}(a)\mathcal{R}(a)^{2}=\mathcal{R}(a)\left[\mathcal{P}^{\prime}(a)+\mathcal{P}(a)\mathcal{R}(a)\right].$ (G.20)
-By explicitly taking the logarithmic derivative of $\mathcal{P}(a)$, we uncover an exact geometric identity:
-$\frac{\mathcal{P}^{\prime}(a)}{\mathcal{P}(a)}=\frac{1-2a}{a(1-a)}-4\pi\cot(\pi a)\equiv-\mathcal{R}(a).$ (G.21)
-Consequently, the term in the brackets vanishes identically. Thus, we arrive at the exact algebraic statement: $s^{\prime}(a_{c})=0$ precisely whenever $u^{\prime}(a_{c})=0$.
-In the physical $(s,J)$ plane, this simultaneous extremum signifies a fold catastrophe.
-The parametric Regge trajectory reaches a hard geometric endpoint where the physical slope $dJ/ds=u^{\prime}(a)/s^{\prime}(a)$ resolves (via L’Hôpital’s rule) to a finite limit $S_{0}$, while the trajectory curvature diverges to infinity, $J^{\prime\prime}(s)\to\infty$.
-The critical parameter $a_{c}$ is determined by the exact transcendental equation $\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}$, which evaluates to:
-$4\pi^{2}\csc^{2}(\pi a)-\frac{1-2a+2a^{2}}{a^{2}(1-a)^{2}}$
-$=\left(4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}\right)^{2}.$ (G.22)
-Solving this within the fundamental domain $a\in(0,1/2)$ yields the universal critical parameters:
-Crit. Param. ($a_{c}$): $\approx 0.311277$
-Thres. Mass Rat. ($s_{c}=m_{c}^{2}/\sigma$): $\approx 0.027791$
-Crit. $\pi$ Spin ($q=0$): $J_{c}\approx-0.003378$
-Crit. $\rho$ Spin (q=1q=-1): $J_{c}\approx 0.496622$ (G.23)
-Figure 16: The exact geometric Regge trajectories plotted down to the mathematical endpoint $(s_{c},J_{c})$.
-The trajectory departs from this boundary with a finite slope ($S_{0}\approx 0.1118$) that is smaller than the asymptotic slope at large $s$ ($\approx 2/\pi^{2}\approx 0.2026$).
-The infinite curvature $J^{\prime\prime}(s)\to\infty$ at the exact endpoint is visually imperceptible.
-To the eye, this looks like a perfectly linear trajectory ending abruptly at this point.
-At this mathematical endpoint, the Regge trajectory departs with a finite slope $S_{0}\approx 0.1118$ and infinite curvature (see Fig.
-16).
-While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15).
-The physical branch of the spectrum strictly corresponds to $u\geq 0$.
-As established previously, the physical ground state terminates at $a_{min}$ where $u=0$.
-For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
-Because the physical root $a_{min}$ sits comfortably above the critical parameter $a_{c}$ ($0.4268>0.3113$), the trajectory mathematically terminates well before the string can experience the fold catastrophe.
-This mass gap ($s(a_{min})>s_{c}>0$) acts as an impenetrable geometric shield.
-It guarantees that the Regge trajectory never crosses into the $s<0$ regime, rigorously protecting the Planar QCD spectrum from tachyonic instabilities while enforcing the exact reality of all physical eigenvalues.
-The phenomenological Regge “intercepts” $\alpha_{\pi,\rho}(0)$ are thus not true physical states, but simply linear extrapolations of the large-$s$ asymptotic expansion bridging over this hidden, non-linear geometric boundary.
-While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15).
-The physical branch of the spectrum strictly corresponds to $u\geq 0$.
-As established previously, the physical ground state terminates at $a_{min}$ where $u=0$.
-For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
-Appendix H Irrationality of the Physical Branch Points
-As discussed in Section 15.8, rational values of the branch point degree $a=m/n$ would cause the twistor trajectory to close after finitely many windings.
-In that case, all distinct states would be counted with just $2n$ windings, and our mechanism for the Exact WKB spectrum would disappear.
-In this appendix, we rigorously prove that the physical spectrum is mathematically protected from these rational degeneracies.
-Theorem. For any rational target angular momentum $u=J+q/2\geq 0$, the exact spectral equation (G.9) possesses no rational solutions for $a\in(0,1/2)$.
-Proof. The exact geometric quantization is governed by the spectral equation derived in Appendix G:
-$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+48u}$ (H.1)
-Assume, for the sake of contradiction, that the physical branch point $a$ is a rational number $m/n\in(0,1/2)$.
-Since the target momentum $u$ is also rational by definition, the algebraic expression on the right-hand side evaluates to a purely rational number.
-On the left-hand side, it is a foundational property of trigonometric functions that for any rational argument $a$, the value of $\cot(\pi a)$ is an algebraic number.
-Furthermore, for $a\in(0,1/2)$, the value $\cot(\pi a)$ is strictly positive and therefore non-zero.
-By the Lindemann–Weierstrass theorem, $\pi$ is a transcendental number.
-The product of a transcendental number and a non-zero algebraic number is strictly transcendental.
-Consequently, the left-hand side is a transcendental number, while the right-hand side is a rational number.
-This is a strict mathematical contradiction.
-Therefore, the branch point parameter $a$ cannot be a rational number.
-$\blacksquare$
+This geometric divergence acts as an infinite energetic barrier, protecting the twistor pole from escaping and locking the string into its quantized topological sector.
+Appendix E Analytical and numerical study of the Twistor String equation
+The physical states of the twistor string are governed by the total action monodromy $\Delta S$ accumulated over one period of rotation.
+The Bohr-Sommerfeld quantization condition requires $\Delta S=2\pi n$.
+As derived in Section 16, factoring out the overall twistor scale parameter $a$, we can write $\Delta S=a\Phi$, where the phase functional is
+$\Phi(K,\beta)$ $=2\pi EK-4\pi\left(J+\frac{q}{2}\right)-\pi\sigma K^{2}\left(\beta+\frac{\sin 2\beta}{2}\right)$
+$-4\pi mK\cos\beta-\frac{2}{3}(\tan\beta-\beta).$ (E.1)
+Here, $K=R/a$ is the kinematic radius, and $q$ represents the topological vector index ($q=0$ for the pseudo-scalar families, and q=1q=-1 for the vector families).
+E.1 Significant reduction of the spectrum: only the $n=0$ level is left
+The classical string trajectory is determined by the saddle-point equations with respect to the dynamical variables $a,K$, and $\beta$. Setting $\partial_{a}\Delta S=0$, we get the first equation:
+$\Phi(K,\beta)=0,$ (E.2)
+which implies that there are no levels with $n\neq 0$ in the general Bohr-Sommerfeld quantization condition:
+$\Delta S=2\pi n,\qquad n=0,\pm 1,\pm 2,\dots$ (E.3)
+This important conclusion follows from the fact that our $\Delta S$ depends linearly on $a$ at fixed $\beta$ and $K=R/a$.
+Therefore, by changing our variables to $a,K,\beta$, we completely eliminate $a$ from the saddle-point equations and also obtain the extra equation (E.2) for $\beta$ and $K$.
+Next, setting $\partial_{K}\Phi=0$ yields the energy equation:
+$E=\sigma K\left(\beta+\frac{\sin 2\beta}{2}\right)+2m\cos\beta.$ (E.4)
+Setting the boundary variation $\partial_{\beta}\Phi=0$ balances the boundary tensions and yields a quadratic equation for $K$:
+$\sigma\cos^{2}\beta\,K^{2}-2m\sin\beta\,K+\frac{1}{3\pi}\tan^{2}\beta=0.$ (E.5)
+Note that we have obtained three equations for three unknowns $E,K,\beta$, with $a$ decoupled.
+Thus, we found the two equations for the two parameters of the saddle point and the third equation for the energy spectrum, just as intended, with the important restriction that all the extra Bohr-Sommerfeld levels $n\neq 0$ are absent.
+E.2 Algebraic simplifications of the spectral equation
+To systematically analyze and numerically fit these equations, it is advantageous to cast them into a dimensionless form.
+We introduce the dimensionless renormalized boundary mass parameters $x=m/\sqrt{\sigma}$ for the light quarks and $x^{\prime}=m^{\prime}/\sqrt{\sigma}$ for the strange trajectories, where $m^{\prime}=(m+m_{s})/2$.
+Solving the quadratic equation for the radius, we find the analytical root:
+$\mathcal{K}(\beta,x)\equiv\sqrt{\sigma}K(\beta)=\frac{\sin\beta}{\cos^{2}\beta}\left(x+\sqrt{x^{2}-\frac{1}{3\pi}}\right).$ (E.6)
+For a physical string to exist, the classical radius $K$ must be real, which requires the discriminant to be non-negative.
+Thus, the Minkowski version of the Liouville anomaly imposes a rigid chiral existence bound on the dimensionless mass:
+$x\geq\frac{1}{\sqrt{3\pi}}.$ (E.7)
+Substituting $E$ from Eq. (E.4) back into $\Phi=0$, the mass boundary terms cancel, yielding the parametric angular momentum:
+$J(\beta)=\frac{\sigma K^{2}}{4}\left(\beta+\frac{\sin 2\beta}{2}\right)-\frac{1}{6\pi}(\tan\beta-\beta)-\frac{q}{2}.$ (E.8)
+To fit the unknown fundamental parameters—the dimensionless constituent quark masses ($x_{u},x_{s},x_{c},x_{b}$) and the universal string tension—to the experimental Particle Data Group (PDG) data, we utilize the action monodromy target $\Phi=0$.
+Rather than performing an arbitrary least-squares fit on $M^{2}$ or $J$ independently, we mathematically minimize the fundamental physical constraint of the theory itself (see the remark 16.2 for the definition of our effective quark mass).
+For any given observed resonance with experimental mass $M$ and spin $J$, we first invert Eq.
+(E.8) to find the unique theoretical angle $\beta(J)$ such that $J_{\rm param}(\beta)=J$.
+Substituting this angle back into the master phase equation, the terms corresponding to angular momentum cancel out.
+Since the theoretical parameters natively satisfy $\Phi=0$, the complicated area, boundary mass, and Liouville anomaly terms collapse, leaving the geometric phase mismatch:
+$\Phi(M,J)=2\pi K(\beta)\left(M-E_{\rm param}(\beta)\right).$ (E.9)
+This target function has physical and mathematical meaning.
+Because $K\propto 1/\sqrt{\sigma}$, the target $\Phi$ is naturally dimensionless.
+At high energies, where $K\propto E/\sigma$, the action monodromy asymptotically scales as $\Phi\propto M\Delta M/\sigma\propto\Delta(M^{2})/(2\sigma)$, intrinsically generating the correct mass-squared weighting expected for Regge trajectories without imposing it ad hoc.
+Crucially, for the non-rotating ($J=0$) pseudo-Goldstone bosons ($\pi$ and $K$), the string shrinks to a point ($\beta\to 0$ and $K\to 0$), causing $\Phi$ to identically vanish.
+This naturally decouples the anomalously light chiral masses from the macroscopic string tension.
+The global 5D parameter optimization was performed by minimizing the sum of the squared geometric action residuals,
+$\chi^{2}(m_{u},m_{s},m_{c},m_{b},\sigma)=\sum_{q_{1},q_{2}}\sum_{J,M}\left(\Phi(M,J,x_{q_{1}q_{2}},\beta(J))\right)^{2},$ (E.10)
+across all thirty-six on-shell PDG states spanning the thirteen meson families, where $x_{q_{1}q_{2}}=(x_{q_{1}}+x_{q_{2}})/2$ is the dynamically constructed average boundary mass for the given state.